A Complete Bibliography of Publications by, and about, James Clerk Maxwell

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

18 October 2022 Version 1.38

Title word cross-reference

1/2 [Arm04]. $17.00 [Eve84]. $195 [Buc92]. $195.00 [Hun05a]. $25.00 [Mar99a]. $25.95 [Jam15]. $28.75 [Can90]. $285.00 [Hun05a]. $29.00 [Row05]. 3 [Ive05]. $315.00 [Hun05a]. $38.45 [Har93]. $39.50 [Sto82]. $45.00 [Ber08]. $49.50 [Hun92, Mar93]. $55 [She04a, She04b]. $59.95 [Yos99]. $6.95 [Eve84]. $77.00 [Sto82]. $77.00/£ [Tur83]. $8.95 [Lin71]. $86.25 [Can90]. B [Ive05]. E [Ive05]. ≫ [Spr06]. H [Dia94b]. ≪ [Spr06]. v [Cle70c].

&c. [Cle73b].

-theorem [Dia94b].

0 [Cha84, Hun92, Mar99a, Pip99, Pip04, Row05, She04a, She04b]. 0-226-07882-5 [Roc87]. 0-226-07884-1 [Can90]. 0-226-07886-8 [Can90].

0-226-87375-7 [Row05]. 0-470-86088-X [Hun06, Pip04]. 0-521-25625-9 [Buc92, Hun05a, Jam91, Seg92]. 0-521-25626-7 [Hun05a]. 0-521-25627-5 [Hun05a]. 0-521-35365-3 [Mar93, Hun92]. 0-521-56102-7 [Yos99, Pip99]. 0-679-43342-2 [Mar99a]. 0-7503-0759-5 [She04a, She04b]. 0-8014-2641-3 [Har93]. 0-85274-563-X [The87]. 0-86228-026-5 [Smi85, Cha84]. 0-86341-101-0 [Hun89]. 0-934223-34-3 [Har96]. 0n [Cle73a].

10th [Cle69c]. 11 [Col98]. 118 [Jon64]. 150th [Edi81]. 17th [Hei79, Hei99]. 18 [Smi85]. 1856 [Cle90c]. 1856/57 [Rei15]. 1860 [Ano74]. 1862 [Buc92, Hun05a, Jam91, Jon92, Sie93a, Sie93b]. 1864 [C^{+}65]. 1865 [Cle66c]. 1870 [Mar66]. 1873 [Hun05a, Dar98a, Dar98b]. 1875 [Sei01]. 1879 [Hun05a, Tho82]. 1880’s [Buc79a, Buc79b]. 1881 [Ano82]. 1894 [Bri89]. 18th [Hei79, Hei99]. 1911/sous [LdB12b]. 1932 [Sim06]. 1977 [GCC81]. 1982 [vM82]. 1997 [OR97b]. 19th [Epp98, Cra07, Cro35, Cro40, H^{+}69]. 19th-Century [Epp98].

2 [LR03, Mar99a]. 2-02-063249-7 [Spr06]. 2016 [Mau17]. 233pp [Sto82]. 24.00 [Hun89]. 25s [Chi64, Wil68].

3 [Ber08, Hun92]. 30.00 [The87]. 35.00 [Tur83]. 35/US$59.95 [Pip99]. 39.99 [Hen15].

400th [Sim14]. 45.00 [Har96].

5 [Cha84, Nye02, She04a, She04b]. 59.50 [Roc87].

6s [McC69].

7 [Pip99, Row05].

8s [Ano83].

978 [Ber08]. 978-0-19-966437-5 [Hen15, Mac15]. 978-3-527-40688-3 [Ber08].

a-t-il [D’A89]. Aberdeen [Ano74, Tro01, And07, Cle56b, Cle56g, Cle90c, Jon73]. Aberrant [Mul08]. Abhandlungen [Kai87]. Absolute [Ryn02, Hop57b, Mit17a]. absorbierende [Cle76p]. Absorbing [Cle76p]. Abstract [Tho82, Cle55a, Cle56c, Cle56d, Cle64a, Cle66b, Cle67b, Cle68b, Cle68c, Cle70b, Cle72a, Cle78g, Cle27a]. abstracts [Wil51, Wil72, Wil16]. Academic [Cle74a]. Academies [Cle80d]. Academisch [Cle74a]. Academy [Sim06, Sim06]. acceleration [Bin08]. Acceptance [Mar07]. according [Cle68k]. account [Cle76a, Cle57b]. accuracy [Col97, Col98, Gré19].

Achievements [Tru14, GS06]. Achinstein [Cat16, Har14, Kar14, Kra13, Sta14]. acting [Cle68k]. action [Cat12, Cle62d, Cle75c, Cle76c, Cle11]. actual [Fal17]. Adam [Cha84, The87, Tur83]. Adams [Cle59c]. Addition [Cle72d]. Address [Cle70b, Cle70a]. adjustment [Gré19]. Advances [Cle76n]. advantage [Cle82b]. Aether [Cha84, Goo85a, Goo85b, Lar00, Mat85a, Mat85b, Sie85a, Whi10, Whi53, Whi89, Gol83, Smi85, Whi54, Whi87, Whi08]. after [Knu78]. against [Ano99d]. Age [LS14, Whi10, Whi08]. agency [Cat12]. Air [Cle66e, Cle66b]. Airy [Cle68g]. Akademie [Sim06]. Akademie-Vorträge [Sim06]. Akademische [vM82]. Albert [Bes04, Sim06]. alleged [Wil72, Wil16]. Allegheny [Rod02]. allgemeinverständlichen [Hei88]. alma [Ano06]. also [Hug69]. alternative [Lon03]. America [Kle10]. Amherst [Jam15]. Among [Nah83, Cle60a, Cle68k]. Ampère [Ste05, Dar00, Fle86, Ste05, Ste16]. Análisis [Cle96]. analogical [Lam11]. Analogies [Hes63, Hes66, Cle82a, Hes70, Hes86, Nap21]. Analogy [Tur55a, Bok15, HG12, Lam11, Ozi94, Wis77]. Analyses [Att93a, Att93b, Bla83a, Bla83b, Bre06a, Bre06b, Dar93a, Dar93b, Dar98a, Dar98b, Dar01a, Dar01b, Mat85a, Mat85b]. Analysis [D’A01a, LU31, Lem17, Bat15, Bat55, Cle80a, Cle96, KM08]. anarojī [Hes86]. anderen [Cla58]. Andrew [Cat16, Hen15, Mac15, Ric14, She04a, She04b, Sho14, Row05]. Anfänge [Mar66]. Angle [Cle72g]. Angus [Fra57]. Anmerkungen [vM82]. Annalen [Hof08]. Anniversary [Sim14, Edi81]. Announcement [Mau17]. anonymous [Mar07]. answer [Cle5x]. Antecedents [Tai95]. anti [Ano99d]. anti-nuke [Ano99d]. Ants [Ano99d]. Any [Cle73m, Cle60b, Cle61b, Cle68k, Cle69f, Cle72g, Cle73a, Cle82b]. Apfel [Büh98b]. Apioid [Cle46b]. Apostate [Buc85b]. Apostle [Buc85b]. Apparatus [Hop57a, Cle76d]. Appendix [CJ63a, CJ63b, CJH64]. apple [Büh98b, Sch14]. Appleton [Chi64]. Appliance [Lam03]. Application [LU31, LF08, Cha73b, Cle74h, Cle75d, Cle75e, Smi63]. Applications [Sch17]. applied [Cle61d, Cle61h, Cle61g, Cle62c, Cle62d, Cle68i]. appointed [C^{+}64, C^{+}75a, C^{+}77a, C^{+}79a, C^{+}79d, C^{+}79]. Appreciation [Jos07]. Approach [D’A86, RdO07, RdO16, Bri18]. Approaches [KCP14]. approximate [Cle77g, Cle80c]. APS [Sei01]. arbor [Kai87]. Archimedes [Pic08]. Archives [Anoxxa]. areas [Cle56a]. Argument [Mul08, Lam11, Mit17b, Ryn16]. arising [Cle78g, Cle78f, Cle79g]. armonico [Cle96]. Arranged [Lar10]. Arrangement [Cle69a]. arrive [Dor70]. Aspects [Gar66, Ner87, Ran62, Ran63, Buc85a, Cle82c]. Assessing [Gré19]. Associated [Har96]. association [Cha84, C^{+}64, Cle70b, Cle70a, Cle76b]. astronomy [Bow63, FF91, McC69]. Astrophysical [Ogi08b, Ogi08a]. Athlone [Chi64, Jon64]. Atom [Cha21, Cle75a, Rod02, Rod19]. Atomzeitalters [Hei72]. attempt [Hug69]. attempts [Dor70, Tur55b]. attending [Wil55]. Attraction [Cle75b, Cle77f, Cle80b]. auspices [LdB12b, LdB12a]. Auswahl [Hei88]. Auszüge [Cle15]. authors [Bes04].

Automata [BDP15]. Aux [Dar98c, Cle74d, Cle74k, Rey21]. average [Cle79f]. averages [Yan95]. Avoiding [Har96, Sie96, CGBE95, GBE95]. axioms [Wal85]. axis [Cle75d].

B [Cle69c]. B.A. [Hop58]. Baeyer [Mar99a]. Bahnbrecher [Hei72]. Bakerian [Cle66e]. balancing [Gor83]. Band [vM82]. bare [Fah99a, Fah99b, Fah01]. bare-wire [Fah99a, Fah99b, Fah01]. Baron [Tho10, She99]. Bart. [Lar07, Lar71, Lar10]. Based [Bru71]. basic [Her84, Cle91a]. Basil [Jam15, Hun06, Pip04]. basis [Bat15, Bat55]. Bath [C^{+}65]. battery [Hop59b]. Bd [Kai87]. be [Cle69f, Cle72g, Lam03, Yan95]. beam [Cle74g]. became [For99b]. Bedeutung [Pla31a]. been [GS06]. before [Sus68a, Tho82]. Beginning [Mar66]. behind [Pic08]. Beiträge [Kai87]. Bemerkungen [Cla58]. bending [Cle56f, Cle66d]. Bernard [Hei72]. Bernoulli [Cle61d]. Besprechungen [FSvL^{+}37]. best [Cle69a]. betrachtet [Pla99]. between [CJ63a, CJ65a, CJ65b, Cle73k, Cle77g, Cle79c, Cle80c, Cle67, GS06, Her84, JTR94, JTR95, Mit17b, Ozi94, Sen21]. Bewegung [Cle79j, Cle81b, Cle06]. Beweis [Pla94, Pla95]. Beyond [EN99, Rod19, Toy99]. Beziehungen [Her84]. bibl [Jam15]. Bibliographie [vM82]. Big [Hei72]. Biochemistry [FSvL^{+}37]. Biographical [SW89]. Biographies [Anoxxa]. Biography [Ano12, Anoxxd, Tol81, Tol82, Sie85a, Smi82a, Smi82b, Eve84]. Birmingham [Cle66c]. Birth [BP78, Edi81]. bis [Büh98b, Hei72]. black [Bes04]. Blindness [Cle57a, Cle57h, Cle55b, Cle56e, Wil55]. blue [Cle57f]. Blumtritt [Kai87]. Bodies [Cla58, Cle75f, Cle75g, Cle75h, Cla59, Cle60b, Cle69e, Cle77b, Hug69]. Body [Clexxb, Cle54a]. Boltzmann [vM82, Cle79f, KWK73, PS96, Tho91, Tho92]. Book [Ano83, Ano04, Ber08, Boo64, Bri18, Bri89, Buc92, Buc93, Can90, Cat16, Cha84, Chi64, Cle73c, Cle73b, Cle73d, Cle74b, Cle74a, Cle74c, Cle74d, Cle76a, Cle77a, Cle78a, Cle78b, Cle78c, Cle79a, Cle79b, Dea16, Dom92, Dom97, Dom04, Eve84, Eve99, Goo85a, Goo85b, Har93, Har96, Har14, Hei72, Hen15, Hun88a, Hun88b, Hun89, Hun92, Hun00a, Hun00b, Hun05a, Hun06, Jam15, Jon64, Jon92, Jor93, Jos07, Jos14, Kar14, Kra13, Lin71, Mac15, Mar99a, Mar93, McC69, Mor91a, Mor91b, Mye06, Ner87, Pip99, Pip04, Ric14, Roc87, Row05, Ryn16, Sch89, Seg92, She04a, She04b, Sie85b, Sie85a, Sie93a, Sie93b, Sie96, Smi82a, Smi82b, Smi85, Smi00a, Smi00b, Spe65, Spr06, Sta14, Sto82, Tai73, Tai77, The87, Tur83, Wil68, Wil97, Wil05]. Book [Yos99]. booklet [Ano99c, Rid30]. booklets [Cle51b]. Books [Ano82, Jam15]. Born [Hei72]. Bose [FMM96]. Boston [The87]. Boundary [MGM08]. Bow [Cle72b, Cle76h]. Bowen [Chi64]. branches [C^{+}79a, Cle79i]. Braunschweig [vM82]. Braunschweig/Wiesbaden [vM82]. bridges [Cle68i]. Brilliance [Ati15]. Brilliant [Art16]. bring [Rau14b]. Bringing [Mir99]. bristled [GRGR08]. Bristol [She04a, She04b, Sto82, The87, Tur83, Bla83a, Bla83b]. Britain [Tho82]. British [Cle70b, Cle70a, C^{+}64, Cle76b, Cra07, Cro35, Cro40, Har88, H^{+}69, LB^{+}04, Mac19, Sim14, Wis82]. Broglie [Hou16].

Brownian [Rap05]. Bruce [Dar93a, Dar93b, Jor93, Har93]. Brush [Har96, Sch89, Sie85b, vM82]. Bruxelles [LdB12b, LdB12a]. BSHM [Flo13]. bubbles [Cle74d]. Buchwald [Can90, Roc87, Ner87]. Buckley [Lit03, Lit10]. Building [Ano16]. buildings [Cle77i]. Bundle [Bri18, RdO07, RdO16]. Burbury [Dia94b].

C [Ber84, Boo64, Chi64, Cle77k, Har96, Jon64, Sch89, Sie85b, Spe65, CJ63a]. cake [Saa05]. calculation [Cle57d, Cle64b]. Calculus [Cle51c, Cle73i, Cle76g]. called [Nah83]. calorimétrie [Cle91a]. calorimetry [Cle91a]. Cambridge [Buc92, Cle76a, Dom04, Har85, Hun92, Hun05a, Jam91, Mar93, Pip99, Row05, Seg92, Tro01, Yos99, Ano79a, Mor95, Cle59c, Cra07, Dom04, Ful86, Nye02, War03]. came [Mah15]. Campbell [Ano83]. Can [Cle69f, RSB^{+}92, Yan95, Saa05]. Canongate [Gee81]. Canonis [Sim14]. capacity [Cle78h]. Capillary [Cle76c, Cle11]. Carber [Har96]. career [Ano91, Jam91]. Cartan [Rap05, Wea16]. Case [AB86, Cle54a, Cle70f, Mor92, Tho92, VT89]. Case-Study [AB86]. cases [Tho91]. Cass [McC69]. Cassini [Tis17]. catalogue [Ano31b]. Catch [RSB^{+}92]. Cavendish [Ano03, Cle79c, Cle21a, Cle67, Cle0 , Cro74, Fal15, Fal17, Fit10, Fit16, Gla31, Kim02, Lod27, Wil51, Wil72, Wil16, McC69]. celebration [And07, Sim14]. Celebrations [Ano31a, Fer31]. cellphones [Smi17]. Cellular [BDP15]. Cellulose [FSvL^{+}37]. centenary [Ano31b, Fer31, Hea29, Rid30, Ano31a]. centipedes [GRGR08]. Central [Ryn16]. Centrale [Cle57g]. centre [Cle76i]. centres [Ano03]. centuries [Hei79, Hei99, Pol85]. Century [Bla83a, Bla83b, Epp98, Sto82, Whi10, Wil68, Cra07, Cro35, Cro40, D’A01b, Har82a, H^{+}69, LB^{+}04, Tur83, Whi08, Buc85a, Buc89, Can90, D’A81, Mac19, Wil62, Har85, Ner87]. Chains [Ain11]. Chair [Rei08]. chaleur [Cle91a]. Challenge [Sta09]. Challis [Cle73b]. Changed [Hun06, Wil05, Mah03, Pip04]. changing [Heh10]. Channel [OM11]. chaos [Cat12]. characteristic [Cle74g, Cle75d]. characters [Büh98a]. Chase [Cha85]. Chicago [Can90, Eve84, Roc87, Row05]. Chichester [Hun06, Pip04]. chief [C^{+}79a]. Christian [Mar99a, Hut98]. Church [Dea16, Sta15]. circuits [Cle74h, Cle75e]. circular [Cle69f]. Citizen [Ano99d]. Cl [Mar66]. claims [Wil72, Wil16]. Clarendon [Ano82]. class [Rei15]. classic [Fit16]. classica [De 78]. Classical [Har82b, Lan98a, Men08, Ozi94, SR05, Bes04, De 78, Heh10, Lan14, LR03, She04a, She04b]. Classification [Toy99, Cle71f]. Clausius [Bru58b, Cle5x, Dia94a, Dou16, Gar66, Gar70, Rey21]. Clay [Wes99]. Clerk [Ach13, Ano79a, Ano79b, Ano79c, Ano81a, Ano81b, Ano82, Ano83, Ano85, Ano31b, Ano74, Ano99c, Ano11b, Ano11a, Ano12, Ano13, Ano15, Ano16, Ano17b, Ano17a, Ano17c, Anoxxa, Anoxxb, Anoxxd, Anoxxc, Ber63, Bes04, Edi81, BP78, Bru71, Buc92, Büh98a, CG82, CG84, CG69, CG10, Cat16, Cha84, Cha21, Cle15, Cle79, Dar98a, Dar98b, Dar01a, Dar01b, Dib64, Dom80, Dom92, Dom97, Dom04, Dom18, Dou16, Duh02, Duh67, Duh15a, Eld00, Eve67b, Eve67a, Eve75, Eve84, Eve98, Eve99, Eve06, Fai31, Fis88, Fle31, Flo12, Flo13, FMW14, FSvL^{+}37, Ful86, Gar69, Gar79, Gar85b, Gee81, Gia70, Gla96, Gla01, Gol83, Gol13, Goo85a, Goo85b, Hal31, Har90, Har95a, Har98, Har02, Har14, Hen15, Hen86, Hun88a, Hun88b, Hun00a, Hun00b, Hun05a]. Clerk [Hun06, Hut98, Jam91, Jea31, Jon32, Jon49, Jon73, Jon80, Jon92, Kar14, Kno14, Kra13, Lam11, Lam03, Lan03, Lar26, LF08, Mac15, Mac85, Mah03, Mar66, Mat85a, Mat85b, Mau17, McC11, McC69, McN04, Mit17a, Mor91a, Mor91b, New55, Niv90a, Niv90b, Niv90c, Niv90d, Niv27, Niv52, Niv65, Niv03, Niv11, OR97a, OR97b, Ogi08a, Ogi08b, Pag38, Pip99, Pip04, Pla31a, QM15, Ran62, Ran63, RWT08, Rei08, Rei15, Ric14, Rod02, Row02, Seg92, Sei01, Sho14, Sie85a, Sie93a, Sie93b, Sim14, Smi82a, Smi82b, Smi85, Smi00a, Smi00b, Sta14, Sta12, Tai77, Tai80b, Tai11a, The86, The87, Tho31a, Tho31b, Tho11, Tol81, Tol82, Tro01, Tur53, Wal16, Wes99, Wil05, Yos99, Ano31a, AL30, Art16, Ati15, Chi64, Cra54, Dom63, Fer31, For99a, Har09, Hop56, Hop57a, JTR94]. Clerk [JTR95, Kle88, Lam31, Lar31, Lar36, Lar37, Lod27, Lod31, Lor23, Rid30, Tai73, Tai80a, Tai86, Tai95, Tai97, Tho93, Tho68, Tor00, Boo64, Jon64]. Clerk-Maxwell [Cle73e, JTR94, JTR95, Tai73, Tai80a, Tai86, Tai95, Tai97, Tho93, Tho68]. Clifford [Bri18, RdO07, RdO16]. Close [Whi10, Whi08]. cloth [Jam15]. Club [Jen70]. Co [Ano83, McC69]. coating [KM08]. coating/substrate [KM08]. cold [Sen21]. Collection [Spr06]. College [CJ63b, CJH64, Cle76a, Cle79, Hea29, Rod02]. Colliding [Tai97]. collision [Cle60b]. Collisions [Cle60g, Dia94a]. Collocation [Sch17]. collotype [Fit10, Fit16]. Color [Eva61, Cat14, Sto82]. colored [Gye17]. Colorimetry [Tur96]. Colour [Cle57a, Cle57h, Cle71b, Lon08, Tur83, Cle55b, Cle56e, Cle57b, Cle71c, Cle72e, Wil55, Bla83a, Bla83b, Sto82]. Colour-Blindness [Cle57a, Cle57h, Cle55b, Cle56e, Wil55]. colour-vision [Cle71c]. coloured [Wil55]. Colours [Cle61f, Cle11, Gar04, Cle55b, Cle56e, Cle57f, Cle57g, Cle60e, Cle60f, Cle61b, Cle61e]. Column [Cle73l]. Comet [Hug69]. commemoration [Tho31b, Tho11]. commemorative [Ano99c, Chi64, Dom63]. commencing [C^{+}79b, C^{+}79c]. Comment [Squ89]. Committee [C^{+}69, C^{+}64, C^{+}65, Cle66c, C^{+}68, C^{+}70a, C^{+}70b, C^{+}71a, C^{+}71b, C^{+}72, C^{+}73, Cle73q, C^{+}74, Cle74j, C^{+}75a, C^{+}75b, C^{+}76a, C^{+}76b, C^{+}77a, C^{+}77b, C^{+}77c, C^{+}78, C^{+}79a, C^{+}79b, C^{+}79c, C^{+}79d, C^{+}79e, C^{+}79]. Communication [App30, App68]. Comparing [KCP14]. comparison [Cle68b, Cle68c, Cle69b]. Complementary [SR05]. complete [Jon80]. Completeness [Sie75]. composition [Wil72, Wil16]. compound [Cle57f, Cle60f]. compressive [WG08]. Comput [Col98]. Computation [Bub01, Lai74, Ben03]. computer [Bro00a]. computing [LR03, She04a, She04b]. concept [O’H75]. conception [Ein31a]. Concepts [Čap76, Lon03]. Conceptual [All13, Har82a, Yan14]. concerning [Cle76d, Len19]. Condition [Cle72g, Wil79, Dia94b]. Conduction [Cle63]. Conductivity [Cle61d, Cle72i]. Conductors [Cle73p]. Conference [Sch96].

Confined [Cha21]. confins [Dar98c]. conformal [Wan08]. conjugate [Cle76m]. Conjuring [She18]. connected [Cle73o, Cle76f, Cle76l]. connections [VT89]. Consequences [Gal69, Gal79, Gal10]. considerations [Cle76d]. consistent [Hei70a]. constants [Gré19]. Constitution [Cle62b, Cle75f, Cle75g, Cle75h, Cle77b]. constructing [Ner84]. construction [Cle68a]. Constructive [Mul08]. Context [Nug16, Pip99, Sie95]. Contingencies [TS09]. contingency [Cle82b]. continuiteit [Cle74a]. continuity [Cle74a]. Continuum [Moy78]. Contracta [Cle69g]. Contribution [Hop56, EP81, Rey21]. Contributions [CG10, Wil68, Ano83, CG82, CG84]. contrived [HG12]. control [Ful86]. Conundrum [OM11]. Convergences [TS09]. Converter [Hos16]. Coordinates [Cle72g]. Cornell [Dar93a, Dar93b, Har93]. corrections [Cra54]. Correlation [Cle74b, Cle74c]. Correspondence [CG10, Cle90e, CS07, CS71, CS10, Lar10, Ano83, CG82, CG84, Cle76a, Dou16, Jam12, JTR94, JTR95, Lar07, Lar71, Cle76a]. corresponding [Cle72g]. cost [Fah96, Kra21]. Cotterill [Cha71]. Coulomb [Lan98a]. Coulson [Chi64]. countries [Rey21]. Course [Ano17b]. Covariant [Ive05]. Cover [Rau14b]. cracks [MGM08, WG08]. creation [Mar07]. Creativity [ADS83, Kim02, Eve83]. Critical [Coh94, Sim68, Sta09, Duh02, Duh67, Duh15a, Wil72, Wil16]. criticism [Mar07]. critique [Duh02, Duh67]. Crofton [Cha89]. Crossing [TS09]. Crystallography [RSB^{+}92]. crystals [Cle72d, Dea72]. Cultural [Bee96]. curmudgeon [Nah83]. Current [Att93a, Att93b, Buc93, Chr06, Hun92, KCP14, Mar93, Bro67, Lan14, Sel09, Sie91]. currents [Cle61g, Cle67d, Cle72i]. Curvature [Rap05]. Curves [Cle46c, Cle49, CF51]. Curvilinear [Cle72g]. Cybernetics [May71]. Cyclide [Cle68h]. cyclosis [Cle69c]. cylinder [Cle78h].

D [Bla83a, Bla83b, Spe65, Sto82, Tur83, CJ63b, Ive05]. ’Dafty [Ano06]. dales [Cle70e, Cle71d]. danger [Wil55]. Daniel [Att93a, Att93b, Buc93, Hun92, Mar93]. dash [Whe11]. Date [BP78]. day [Art13]. days [Gla31]. Deas [Cle72d]. Death [Ano79a]. decisive [Tho92]. découvert [D’A89]. Defence [Mul08]. degrees [GS06]. DeGroot [Lin71]. d’électricité [Cle84, Cle85c, Cle89]. Demon [Ain11, Ber08, Cha84, D’A12, Dea16, Dev02, Edw10, Ehr67, Goo85a, Goo85b, Gor83, Hos16, Lai74, Lef02, LR03, LMJ14, LC15, Mar99a, Mat85a, Mat85b, Mou08, NA02, OM11, RS11, Scu07, SGD^{+}02, SPW02, She04a, She04b, Sie85a, Tho79, Wei02, vB98, Ano99d, Fah96, Gol83, Kra21, Mad02, Sch14, Smi85, Sta15, Tai11b, Tho82, Tho89, All13, Ben03, Bro00b, Bub01, EN98, EN99, Mus99, Nor05, Rex17, Scu07, She99, She15, Col90, Dau70, Hei70b]. demons [Whi85]. demonstration [Edw10]. denudativen [Mar66]. Dependent [Den07]. Derivation [Dun08]. Derivations [Lin71]. Derivative [SR05]. Descartes [Cle46a, Whi10, Whi08]. descent [Cle54a]. Described [AVJ07, Pag38, Cla58, Cla59, Lar36, Lar37]. Describing [KS16]. Descriptio [Sim14]. Description [CF51, Cle56a, CJ63b, Pup25, CJH64]. descriptions [Lén29, Lén43]. deserves [Lam03]. design [Sta12]. detailed [Gor83]. Details [KS95]. detecting [Cle80d]. determinacy [Cat12]. determination [Hop57b, Hop58, Hop59a]. determine [Hug69]. determining [C^{+}77a, C^{+}79d, C^{+}79]. determinism [Cle82b]. Deutschland [Pla31a]. Development [Bru57a, Bru61, Bru62, Čap76, Gar66, Knu78, Nug16, Wat61, Ber74a, Ber74b, Bru57b, Bru58b, Bru58a, Ein31a, EP81, Har82a, Pol85]. Devonshire [Cle79c, Cle67]. Diagram [Cle10]. Diagrams [Cle68g, Cle70d, Cle72a, Cle77c, Cle64d, Cle68i, Cle69d, Cle72f, Cle76h]. Dialogue [Len19]. Diamagnetic [FM78]. dictionary [LB^{+}04]. did [Tur96]. Diderik [Cle74a]. didn’t [Nah83]. diese [Büh98a]. difference [Sen21]. differences [Col97, Col98]. different [Cle57g, Cle71c]. Differential [Cle51c, Sch17]. Difficulties [D’A00]. Diffraction [BWB19]. Diffusion [Cle61d, Cle73j, Cle76p, Cle77d, Cle60a]. digital [SBB^{+}08]. Dilemma [Squ89, Gré19]. Dimensional [D’A86, D’A01a, Lem17, KM08, Mit17a, VT89]. Dimensions [Cle72g, CG17]. Dirac [Bri18, BDP15, HB03, LU31, Rap05, RdO07, RdO16]. Direct [Sim66, Cle68b, Cle68c, Cle69b]. direction [Cle71e]. Disappearing [CPR11]. disc [Cle69f]. discontinuous [Cle73i, Cle76g]. discord [Sch14]. discourse [Cle73f]. discovered [D’A89]. discoverers [Wil72, Wil16]. Discovery [Ach87, Buc79a, D’A81, GCC81, Hun89, Sha73, OP87]. Discrete [Vis89]. Discs [Ogi08b, Ogi08a]. discussions [LdB12b, LdB12a]. disk [Cle78h]. Disperses [Mar99a, vB98]. Displacement [Att93a, Att93b, Buc93, Chr06, Hun92, KCP14, Knu78, Mar93, Bro67, Cle70f, Lan14, Sel09, Sie91]. disproving [Tho92]. dist [Sto82]. Distance [Cle72c, Cle72h, Cle75c]. Distribution [Cle79f, Den07, LL19, Tai97, Cle68k, Gye17, PS96, Pla94, Pla95, Tho91, Tho92]. Disvoery [Bri89]. divers [Rey21]. doctrine [Tho91, Tho92]. Documentation [Bri89, Hun89, OP87]. documented [Dar93c]. Does [Ain11, Cle82b]. doing [Buc95]. Domb [Boo64, Chi64, Jon64]. donnée [Rey21]. Don’t [Smi17]. Door [Cle74a, Mad02]. Dot [Whe11]. Dot-dash [Whe11]. Dot.com [Whe11]. Double [Cle74e, TC74, Cle72d, Cle74f, Dea72, Nap21]. d’ouvrages [Att93a, Att93b, Bla83a, Bla83b, Bre06a, Bre06b, Dar93a, Dar93b, Dar98a, Dar98b, Dar01a, Dar01b, Mat85a, Mat85b]. Down [Hos16]. Down-Converter [Hos16]. downwards [C^{+}69, C^{+}70b, C^{+}71b, C^{+}72, C^{+}73, Cle74j, C^{+}76a, C^{+}76b, C^{+}77b, C^{+}78, C^{+}79e]. drawing [Cle57d, Cle76h]. drawn [Cle56a]. Druck [vM82]. dry [C^{+}69, C^{+}70b, C^{+}71b, C^{+}72, C^{+}73, Cle74j, C^{+}76a, C^{+}76b, C^{+}77b, C^{+}78, C^{+}79e]. dual [Gol94]. Dualism [Men08]. Duhem [AB86]. Duke [Cle79c, Cle67]. durch [Cle76p]. during [Heh10]. Dutch [Cle74a]. Dynamic [Wat61]. Dynamical [Cle57c, Cle57i, Cle59a, Cle60a, Cle60b, Cle60g, Cle67a, Cle75f, Cle75g, Cle75h, Cle90b, Cle82, Cra54, QM15, Sim68, Tur56, Wis82, Cle60d, Cle64a, Cle65, Cle67b, Cle68j, Cle68k, C^{+}74, C^{+}75b, Cle27a, Mur07, Sim70, Tur55b].

Dynamics [Cle51d, Ogi08b, Moy77, Ogi08a]. Dzieciece [Nat32].

Early [Can90, Gla31, Ano91, Buc89, Dar93c, Hei79, Hei99, HG12, Jam91, Lam11, Sus68a, Sus68b, Sus69a, Sus69b]. Earth [Cle57c, Cle57i, Hug69, LS14]. EAT [Saa05]. Eaters [Nor05]. ecuaciones [Bun43b]. Ed [McC69, Buc92, Dom04, FSvL^{+}37, Hun05a, Jon92]. ed. [Boo64]. Eddington [Sta17]. Edinburgh [Cha84, Gee81, Smi85]. Edited [Chi64, Jon64, Cle79c, Cle67, Mac15, Sho14]. Editing [Fal15]. edition [McC69, Cle74b, Cle74c]. Editions [Spr06]. Editor [Cle70e]. Editorial [Cle77e]. Eds [She04a, She04b, Cat16, Har96, Hen15, Ric14]. education [Kim03, Zit02]. educator [Gol13]. Edward [Chi64]. EEs [Nah83]. Effect [Buc79a, Buc79b, WG08, RS11]. effectiveness [Bok15]. Eighth [C^{+}76a]. einigen [Cla58]. Einleitung [vM82]. Einstein [Bes04, Bri18, Sim06, Ari03, Ber74a, Ber74b, Dar00, EM84, FMM96, Hal80, Hou16, KS16, Ner84, RdO07, RdO16, Sen21, Sim06]. einzelnen [Cla58]. either [Nah83]. Elastic [Cle51a, Cle53, Cle60g, Cle60b]. elasticity [C^{+}79b, C^{+}79c]. elastodynamics [MGM08]. Electric [Buc79a, FSvL^{+}37, Knu78, Lar36, Pag38, Cle61g, Cle67d, CG68, Cle72i, Cle74h, Cle75e, D’A89, Duh15a, Lar37]. Electrical [App30, App68, CJ63a, Cle76o, Fal15, Hea92, Kra18, McC69, Wil68, Bat15, Bat55, CJ63b, CJH64, C^{+}64, CJ65a, CJ65b, C^{+}65, Cle66c, C^{+}68, C^{+}70a, C^{+}71a, Cle73k, Cle73q, C^{+}74, C^{+}75b, Cle76m, Cle78h, Cle79c, Cle0 , Cle21a, Cle67, Hop57b, Hop58, Hop59b, Mit17b, Pup25, Sch94, Duh02, Duh67]. Electricität [Cle83a]. Electricity [Ano11a, Bol93, Cha01, Cle71h, Cle73t, Cle81a, Cle81c, Cle83a, Cle84, Cle85c, Cle88a, Cle89, Cle92b, Cle04b, Cle53, Cle54, Cle55, Cle98a, Cle98b, Cle05, Cle13, Hei79, Hei99, Hei71, LM81, Ryn16, Sim68, Whi10, Whi53, Whi89, Wis79, Cle62c, Cle70c, Ell66, Hop59a, Kle10, Kle89, Pla98, Pla99, Sim70, Tho93, Tho68, Tur27, Tur72, Whi54, Whi87, Whi08, Wis77, Tai73, Ano82, Cle15]. Electrified [Cle73p]. électriques [D’A89, Duh02, Duh67]. electrodynamic [Dar93c, Her84]. Electrodynamical [Tai95]. Electrodynamical-Wave-Equations [Tai95]. Electrodynamics [Buc79a, Buc79b, D’A81, Dar00, Lan98a, Lan98b, Lan98c, Nug16, VT89, Heh10, Her84, Lan14, Ste05, Ste16, Var82, Wig90]. électrodynamique [Dar98c]. Electromagnetic [Ano11c, BWB19, Bri89, Buc93, Cha73a, Cle82, Cra54, D’A01a, Duh15d, Gia70, Hal08, Hea12, Hun88a, Hun88b, Hun89, Hun92, Hun15, Jam15, Mar93, Mor91a, Mor91b, Ner87, QM15, Sie75, Sie95, Sim66, Sim68, The87, Top70, Twe11, Wil97, Att93a, Att93b, Bin08, Buc85a, Cha86, Cle64a, Cle65, Cle68b, Cle68c, Cle69b, Cle70c, Cle27a, FM14, FMM96, Hen86, Hop59a, Lor23, Mit17b, Mor92, OP87, SE08, Sie85c, Sie91, Sim70, Wal85]. Electromagnetics [LF08, Sus68a]. Electromagnetism [Bar08, Hun84, Cha73b, Fle86, Rau14b]. Electron [D’A00, Lin71, O’H75]. electronics [Sus68a, Sus68b, Sus69a, Sus69b]. électrons [Dar98c]. electrostatic [Cle68b, Cle68c, Cle69b, Cle70c, Hop59a]. Electrostatics [Bro68, Duh15b, Duh15c, Duh15e, Cle73d, Fal17]. Èlektrichestve [Kle89]. Elektricität [Bol93, vM82]. Elektrizität [Cle15, Pla99, Pla98]. Elektrodynamik [Her84, Ste05]. elektrodynamischen [Her84]. elektromagnetischen [Kai87]. élémentaires [Cle91a]. elementarnoi [Kle88]. Elementary [Ano82, CJ63a, Cle81a, Cle88a, Cle05, Cle13, Tai77, CJ65a, CJ65b, Cle66a, Cle73k, Kle88, Mit17b]. Elements [Hei72, Cle73c]. elephants [Ano99d]. Elizabeth [Sch89, Sie85b, Har96]. elliptic [Cle69f]. Embrace [Wis79]. emission [Smi63]. Empire [SW89]. Empiricism [Mul08]. Empowered [Twe11]. Encounter [Bee96]. Encyclopedia [CF98]. endeavouring [C^{+}79a]. endnotes [Ber08]. Enduring [Lon16]. Energy [Cle79f, Duh15d, Har82a, Hos16, Moy77, SW89, Smi02, Edw10, Tho91, Tho92]. Engine [Dev02]. Engineering [LMJ14, Kim03]. England [Wat61]. English [Gal74, Gal70]. Enlightenment [Art16]. Enquiries [Har96, Sie96, CGBE95, GBE95]. Enquiry [Gal69, Gal79, Gal10]. enthusiasts [Rau14b]. Entropy [She04a, She04b, Fah96, LR03]. Entstehung [Kai87]. envelope [Cle67c]. environment [Lar31]. Epistemic [vDP17]. Equal [Tai97, Cle72g]. Equation [Bru61, Ger01]. Equations [AVJ07, Bri18, CF91, CPR11, Duh15d, Fle08, Ger01, GZ11, Ive05, KS16, LU31, Lin71, Rap05, Rau14b, Rau14a, Sch17, Sha73, SR05, Tai95, Art13, Bat15, Bat55, Bun43b, Cen89, Cle73o, Cle76l, Col97, Col98, Dar10, Dys90, Eng15, Heh10, Her84, HB03, Hou16, Kim03, KNP84, Mah15, Mar66, Mat87, MGM08, Mir99, Ozi94, RdO07, RdO16, Vis89, Vor08, Wal85, Yan95, Yan14]. Equilibrium [Cle51a, Cle53, Cle73l, Clexxa, Cle64b, Cle67c, Cle76j, Cle76k, Cle77h, Cle08a, Gye17, Lar08]. Equivalent [GZ11, C^{+}77a, C^{+}79d, C^{+}79]. eraser [LC15]. erasure [Kra21]. Erice [GCC81]. Errata [Ano74]. Erratum [Col98]. errors [Cat14]. Essay [Ano91, Cle59c, Fis88, Jam91, Cle73b]. Essays [ADS83, KS95, D’A01b]. establishment [Nah83, Sch94]. Estimate [LS14]. etc [Ano83, Wil51]. etc. [FSvL^{+}37]. Ether [Cle78d, Cle90a, Cle80d, Dar10]. étude [Duh02, Duh67]. EUR24.00 [Spr06]. Europe [Rod19]. evangelical [Sta12]. evening [Tho82]. events [Cle82b]. Everitt [Har96, Sie85b, Sch89]. every [Cle72g]. Everything [Hun06, Wil05, Mah03, Pip04]. Evidence [Ach13, Cle75f, Cle75g, Cle75h, Pla94, Pla95, Cat16, Har14, Kar14, Kra13, Sta14]. Evolution [Ein31b, Ein54b, Ein82b, Art13]. evolved [Whe11]. Exact [KS16]. exactly [Dia94b]. examination [Mit17b]. Except [KS95]. Excerpts [Cle15]. exhibiting [Cle57c, Cle57i, Cle60c, Cle61b]. Exorcising [OM11]. Exorcist [EN98, EN99]. Experiment [Can90, CG68, Mou08, AL30, Buc89, Cat12]. Experimental [Cle6x, Sim66, CJ63b, CJH64, Cle71a, Cle74k]. expérimentale [Cle74d, Cle74k]. experimentally [C^{+}77c]. Experimente [Ste05]. Experiments [Cle70c, Cle73j, Cle57b, C^{+}79b, C^{+}79c, Ste05, Ste16, Sus68b, Sus69b, Cle57a, Cle57h]. explains [Sen21]. Explanation [Tur56, Tur55b]. Explorative [Ste05]. Exploratory [Ste16, Ste05]. expositions [Dar93c]. expressed [Cle69f]. extended [Wig90]. extension [Cha89]. extracts [Ful86]. Extraordinary [Bea13, Gib19]. Eye [Cle57a, Cle57h, Cle76i].

F [Har96, Sch89, Sie85b]. F.R.S [Tro01]. F.R.S. [Ano81a, Cha89, CG68, Cle79c, Cle21a, Cle67, Gar79, Mac85, O’H75]. Faces [Bri18, Col90, RdO07, RdO16]. facsim [Ano83]. factorizability [PS96]. Familiar [Pag38, Lar37]. families [GS06]. Family [Tru14]. Famous [Ano17a, Lam03]. Faraday [Cle79d, Cle95, Cle12, Lam11, Ste05, Wil68, Ano65, Ber74a, Ber74b, Cle55a, Cle56c, Cle56d, Cle57d, Cle58a, Cle64c, Cle73s, Cle79e, Cle95, Cle12, Fle86, FM14, Jam12, Jam15, Kuw31, Mac64, Ner84, Ste05, Ste16, Twe11, Spe65]. February [Tho82]. Feldes [Mar66]. Feldtheorie [Kai87]. fellow [Nah83]. Ferdinand [Cle77j]. Fernand [Bre06a, Bre06b, Mye06, Spr06]. Feynman [Dys90]. Field [Ano11c, Buc79a, Cle82, D’A00, Gia70, Ive05, Jam15, Pei63, Ryn02, Sim68, SR05, Twe11, Wil97, Cle64a, Cle65, Cle27a, FM14, Gol94, Hen86, Mar66, Mit17b, Moy78, Ozi94, Sim70, The87, Hun88a, Hun88b, Mor91a, Mor91b, Cra54, QM15]. Fields [AVJ07, Arm04, Bee96, Ber74a, Ber74b]. Fifth [C^{+}73]. Figure [Cle72g]. Figures [Cle70d, Cle72a, Chi64, Cle56a, Cle64d, Cle69d, Cle72h, Cle72f, Jos07]. Films [Gar04]. Final [Cle73a, Cle73m, Cle68k]. finite [Col97, Col98, KM08]. Finsler [VT89]. First [C^{+}74, C^{+}77a, Duh15b, Wil62, Sus69b]. Fisher [Ryn16]. fisica [De 78]. físico [Bun43a, Bun43b]. Fixed [Cle57c, Cle57i]. fiziki [Pol85]. Fleeming [Mit17a]. Flood [Cat16, Hen15, Mac15, Ric14, Sho14]. flow [Wis77]. fluid [Cle70f, Cle74f, Cle74e]. fluids [Cle76f]. Focal [Cle73n]. Foci [CF51]. Foramen [Cle57g]. Forbes [Jam15, CF51]. Force [Cle58a, Smi02, Ber74a, Ber74b, Cle55a, Cle56c, Cle56d, Cle57d, Cle61a, Cle62a, Cle61c, Cle61h, Cle61g, Cle62c, Cle62d, Cle64c, Cle68b, Cle68c, Cle68k, Cle69b, Cle95, Cle12, Cle86, Eve06, Har82a, Lam11]. Forces [Cle70d, Cle72a, Cle73a, Cle73m, Cle74d, Cle74k, Cle64d, Cle68i, Cle69d, Cle72f, Cou63, Hei70b, Len19, Cle74c, Cle74b]. Forerunner [Tor00]. form [Cle56a, Cle57c, Cle57i, Cle60b, Heh10]. formed [Cle72d, Dea72]. forms [Cle57d]. formula [Mit17a]. Formulation [Chr06, Vis89, Wan08]. Fortune [Boh17]. Found [Har95b]. Foundation [Sim14]. Foundations [All13, Bar08, Bru67, Dor70]. Founders [ADS83, Fle86]. four [Rau14b]. Fourth [C^{+}72, C^{+}79]. frame [Var82]. Frames [Cle70d, Cle72a, Cle64b, Cle69d, Cle72f]. frameworks [Cha89]. Francis [Cle72d]. Frank [McC69]. Fraser [Fra57]. Fraternity [Jen70]. Frederick [Cle79a]. Free [Ano17b, She99, Cle82b]. Freedom [Jos14, Cle82b]. French [Cle74k, Cle84, Cle85c, Cle89, Cle91a, D’A89, Duh02, Duh67, Hal04, Poi99, Poi07, Rey21]. Frequency [LF08]. Fresnel [Rey21, Saa05]. Friction [Cle61d, Cle66e, Cle66b]. Friday [Tho82]. fridge [Sen21]. Friedrich [Hei72, vM82]. frontis [Chi64]. Function [Cle68g, Cle74g, Cle75d]. functionally [KM08]. Functions [Cle73p, Cle76m]. Fundamental [Cha21, Gré19, Her84]. Further [Hug69, CJH64].

G [Bri89, Chi64, Cle73c, FSvL^{+}37, Har96, Hun89, Sch89, Sie85b, vM82]. Gabriel [Lar07, Lar71, Lar10]. Galilean [HB03]. Galilei [Büh98b]. Galileo [Büh98b, Cro01]. Galvanometer [Cle19]. Gap [TS09]. Garber [Sch89, Sie85b]. Garnett [Ano83]. gas [Cle74a, Pla94, Pla95]. gas- [Cle74a]. Gase [Cle76p]. Gaseous [Cle73l, Cla58, Cla59]. Gases [Ber63, Bru57a, Bru61, Bru62, Bru71, Cle59a, Cle60a, Cle60b, Cle60g, Cle61d, Cle63, Cle66e, Cle67a, Cle73e, Cle73j, Cle76p, Cle79g, Cle90b, Den07, Gar70, GBE86, Por81, Bru57b, Bru58b, Bru58a, Cle60d, Cle66b, Cle67b, Cle68j, Cle68k, Cle78g, Cle78f, Cle07, Cle77a, Sch89]. gasförmiger [Cla58]. Gasmolecülen [Pla95]. Gasmolekülen [Pla94]. Gastheorie [vM82]. Gauge [Wan08, HB03, Yan14]. gegnerischen [Her84]. General [Cle58b, Cle76d, Sho14, Wan08]. generally [Hei88]. Generation [Hei72]. Genese [Kai87]. Genesis [Nug16]. Genius [For99a, Gal69, Gal79, Gal10, Pip99, Ano06, Gee81, Nah83]. Geographie [Mar66]. geography [Mar66]. Geometric [Cle72c]. geometrical [Cle72h, Cle74h, Cle74i, Cle75e]. George [Lar07, Lar71, Lar10, JTR94, JTR95, O’H75]. German [Bes04, Bol93, Büh98a, Büh98b, Cla58, Cle62a, Cle76p, Cle78k, Cle79j, Cle81b, Cle83a, Cle95, Cle12, Cle15, Cle86, Cle06, Dar93c, FSvL^{+}37, Hei88, Her84, Lén29, Lén43, Mar66, Pla94, Pla95, Pla98, Pla99, Pla03, Pla31a, Sim06, SB50, Ste05]. Germany [Ber08, Dar93c, Pla31a, Pla31b]. Gerstenberg [Kai87]. Gesamtausgabe [vM82]. Geschichte [Lén29, Lén43]. Geschwindigkeitsverteilungsgesetzes [Pla94]. Geschwindigkeitsvertheilungsgesetzes [Pla95]. giant [Bow63]. giants [Nah83]. Gibbs [Gar66]. give [Cle82b]. Glass [Har95b]. Glasses [Gar04]. Gleichungen [Mar66, Kai87]. Glenlair [Rid30]. global [Sch14]. Goal [Sie75]. god [Büh98a]. Going [Sta09]. golden [Sch14]. Goldman [Cha84, Sie85a, Smi85, Goo85a, Goo85b, Mat85a, Mat85b]. Good [Boh17]. Goodman [Wil79]. Gott [Büh98a]. Göttingen [Cle69c]. Governor [CL15]. Governors [Cle68e, Cle68f]. graded [KM08]. grand [Nah92]. graphical [Cle76h]. Grassmann [Tur96, Sch96]. Gravitation [Dew18]. Gravitational [SGD^{+}02]. Gravity [Cle73l, KNP84]. Graz [vM82]. Great [Cro01, Gar08, Lén33, LH33, Len70, Tho82, Büh98b, Pic08, Bes04, Lén29, Lén43]. greatest [Ano65, Ano11b]. Gresham [Flo13]. Grosse [Hei72, Bes04, Lén29, Lén43]. Group [Tai97]. Grove [CG68, Cle74c]. Grundgleichungen [Her84]. Guide [Fle08, Lem17, Rod19]. Guided [Wil97]. Günther [Sch96]. Guthrie [Cle73r, Kno11, Cle79a].

Half [Wil62]. Hall [Buc79a, Buc79b]. Hallyn [Mye06, Spr06, Bre06a, Bre06b]. Hamilton [Cle74g, Cle75d]. Handbook [Ano31b]. Handles [KS16]. Hanlon [Cle69g]. Hans [Mar99a]. happened [Hoo87]. hard [Dys99, Mac15]. hardback [Hun05a, Hun06, Pip99, Pip04]. hardcover [Can90, Eve84]. Harman [Buc92, Dom04, Hun05a, Jam91, Jon92, Pip99, Seg92, Yos99, Dar98a, Dar98b, Dar01a, Dar01b, Eve99, Hun00a, Hun00b, Sie93a, Sie93b, Smi00a, Smi00b]. Harmonic [Cle80a, Cle96]. Harris [Cha84, Smi85]. Harvey [She04a, She04b]. having [CF51]. Hawaii [CL15]. Hawking [Cro01, Pic08]. Heat [Cle61d, Cle63, Cle70g, Cle71g, Cle72j, Cle72k, Cle72l, Cle72m, Cle75i, Cle75j, Cle77l, Cle78k, Cle80e, Cle83b, Cle85b, Cle88c, Cle91b, Cle91c, Cle94, Cle97, Cle02, Cle04a, Cle08b, Cle21b, Cle70, Cle01, Har96, Sie96, Wat61, Cla59, C^{+}77a, C^{+}79d, C^{+}79, CGBE95, GBE95, Kle88, Kra21, Cle91a]. heavenly [Cle69e]. Heaviside [Buc85b, Nah83]. heavy [Cle54a]. heel [Mir99]. Heisenberg [Bes04, Bes04, Hei72]. Helmholtz [Bla83a, Bla83b, Sto82, Tur83, Bok15, Cle77j, Tur96]. helped [Rau14b]. Hendry [The87, Hun88a, Hun88b, Mor91a, Mor91b]. Henry [Cle77a, Cle79c, Cle0 , Cle21a, Cle67, McC69, Wil51, Wil72, Wil16, Fal15]. her [Saa05]. Herapath [Bru57a]. herausgegeben [vM82]. Hereditary [Gal69, Gal79, Gal10]. Hermann [Cle77j, Sch96]. Herneck [Hei72]. hero [Rin08]. Heroes [Gar85a, Ari03]. Hertz [Hun89, D’A89, D’A00, Hei71, OP87, Sus68a, Sus68b, Zit02, Bri89]. Hertzian [PV04, Poi99, Poi07]. Hertziennes [Poi99, Poi07]. Hestenes [Rap05]. Heterogeneous [Clexxa, Cle76j, Cle76k, Cle77h, Cle08a, Lar08]. heuristic [Cha86]. Heyden [Sto82]. hidden [Moy77]. High [Col97, Col98, Kra21]. Hildesheim [Kai87]. Hilger [Cha84, Sto82, The87, Tur83]. hills [Cle70e, Cle71d]. him [Nah83]. Historical [Bru71, Coh94, D’A01a, Duh02, Duh67, Rex17, Bun43a, Bun43b, Duh15a, Kim03]. histórico [Bun43a, Bun43b]. historique [Duh02, Duh67]. History [AB86, Ber63, CG17, Dar93a, Dar93b, Fit10, Fit16, Kim02, Nye02, SMO^{+}06, Sha73, Sim14, Whi10, Whi53, Whi89, Cat12, D’A01b, Fah99a, Fah99b, Fah01, Hea29, Heh10, Lén29, Lén33, LH33, Lén43, Len70, Sus68a, Sus68b, Sus69a, Sus69b, Whi54, Whi87, Whi08, Zit02]. holes [Bes04]. holography [SBB^{+}08]. Homogeneous [Cle73p]. Hon. [Cle0 ]. Honble. [Wil51, Wil72, Wil16]. Honor [KS95]. Honourable [McC69, GS06, Cle79c, Cle21a, Cle67]. honoured [Ano06]. Hopkins [Cle51b, Cra07]. hot [Sen21]. House [Mar99a, OR97b]. Howard [Ryn16]. Humanities [vDP17]. hundred [Mor95]. Hunt [Har93, Dar93a, Dar93b, Jor93]. Hutton [JTR94, JTR95]. Huxley [Dea16, Sta15]. Huygens [Wea16]. hypothesis [Pup25].

I. [Sie85a]. Ice [Cle72b]. Idea [Ein31b, Ein54b, Ein82b]. Ideas [Ein54a, Ein82a, Lar36, Pag38, D’A01b, Lar37, FSvL^{+}37]. II [Dar98a, Dar98b, Hun05a, vM82, Bru57b, Buc79b, Cle56d, Cle60a, Cle60b, Cle61g, Cle84b, Dia94b, EN99, Gol94, Har95a, Hop58, Hug69, JTR95, Kle89, Sus68b, Dom97]. III [Dom04, Hun05a, Bru58b, Cle58a, Cle60b, Cle62c, Har02, Hop59b, Sus69a]. illus [Jam15]. illustrate [Cle57e]. Illustration [Cat01]. Illustrations [Cle90b, Cle76h, Fit10, Fit16, Cle60a, Cle60b, Cle60g]. image [Cle68d, Sie85c]. Images [Wes99]. imagining [Ari03]. importance [Pla31a]. important [Wil51, Wil72, Wil16]. Improvement [Hop59b]. Inaugural [Cle56b, Cle56g, Cle90c, Cle79]. incease [C^{+}73]. including [Fah99a, Fah99b, Fah01, Wil51, Wil72, Wil16]. Incoming [CPR11]. incorporating [Kim03]. increase [C^{+}69, C^{+}70b, C^{+}71b, C^{+}72, Cle74j, C^{+}76a, C^{+}76b, C^{+}77b, C^{+}78, C^{+}79e]. independence [Gye17]. indeterminate [Cha71]. Index [Boh17, Cle69e, GS06, Jam15]. Induction [Cle76o, CG68, Cle72i]. Industrial [LF08]. Inequalities [Cle79g, Cle78g, Cle78f]. infinite [Cle72i]. Influence [Ati15, Ein31b, Ein54b, Ein82b, Cle69e, Ein31a, Pla31b]. Information [Dev02, LC15, Wei02, Edw10, Fah96, LR03, She04a, She04b]. Innovation [Mar93, Sie91, Att93a, Att93b, Buc93, Hun92]. Inquiry [Gib19, Wil72, Wil16]. Inspiring [Ati15]. Instability [Fed08]. Institute [She04a, She04b, McN04]. Institution [Tho82]. instrument [Cle56a, Cle57e, Cle60c, Cle61b, Cle75d]. Instruments [Cle58b, Cle76f, Cle66a]. Integral [Cle51c, MGM08]. integrals [Cle69f]. integration [Cle77g, Cle80c, NV19]. intended [Tho93, Tho68]. intense [SE08]. interaction [WG08]. Interatomic [Cou63]. interface [MGM08]. Interference [BWB19]. Internal [Cle66e, Cle61d, Cle66b]. Internet [Whe11]. interpretation [Tur55b]. Interrelations [AB86]. Interscience [Lin71]. Intertheoretic [Nug16]. Introduction [Cle73r]. Introductory [Cle6x, Cle71a]. invariable [Cle57c, Cle57i]. invariance [HB03]. Invariant [Chr06, Ive05]. invented [Kle10]. invention [Mit17a]. inverse [Fal17]. investigating [C^{+}69, C^{+}70b, C^{+}71b, C^{+}72, C^{+}73, Cle74j, C^{+}76a, C^{+}76b, C^{+}77b, C^{+}78, C^{+}79e]. invisible [Cat14]. involving [Cle73p]. ionosphere [App63]. Ions [Dar98c]. Irreversibility [Mou08, Col90]. Isaac [Bes04, Cat16, Har14, Kar14, Kra13, Sta14, Ach13, Cle76a]. ISBN [Ber08, Buc92, Can90, Cha84, Har93, Har96, Hen15, Hun89, Hun92, Hun05a, Hun06, Jam91, Mac15, Mar99a, Mar93, Pip99, Pip04, Roc87, Row05, Seg92, She04a, She04b, Smi85, Spr06, The87, Yos99]. isn [Lam03]. Italian [De 78]. Ithaca [Har93]. IV [Bru58a, Cle53, Cle57b, Cle62d, Sim70, Sus69b]. Ivan [Eve84, Gee81, Smi82a, Smi82b]. ix [Wil68].

J [Ber08, Bri89, Cle77k, Dar93a, Dar93b, Duh02, Duh67, Har93, Jor93, KS95, Ryn16, Top70]. J. [Ber84, Ber63, Cle69c, Duh02, Duh67, Duh15a, Hun89, Mar66, Roc87, Top70]. Jaffe [Hei72]. James [Ano82, Ano83, Ano11a, Bes04, Buc92, Büh98a, Cat16, Cha84, Dar98a, Dar98b, Dar01a, Dar01b, Dom92, Dom97, Dom04, Eve99, Fle31, Goo85a, Goo85b, Har09, Har14, Hun00a, Hun00b, Hun05a, Hun06, Jon92, Kar14, Kra13, Mat85a, Mat85b, Mau17, McC69, Pip99, Pip04, Seg92, Sie85a, Sie93a, Sie93b, Smi85, Smi00a, Smi00b, Sta14, Tai77, Wil05, Yos99, Ach13, Ano79a, Ano79b, Ano79c, Ano81a, Ano81b, Ano83, Ano85, Ano31b, Ano74, Ano99c, Ano11b, Ano12, Ano13, Ano15, Ano16, Ano17b, Ano17a, Ano17c, Anoxxa, Anoxxb, Anoxxd, Anoxxc, Edi81, BP78, Bru71, Büh98a, CG82, CG84, CG69, CG10, Cat16, Cha21, Cle15, Cle79, Dar10, Dib64, Dom80, Dom18, Dou16, Eld00, Eve67b, Eve67a, Eve75, Eve98, Eve06, Fai31, Fis88, Flo12, Flo13, FMW14]. James [Ful86, Gar69, Gar79, Gar85b, Gia70, Gla96, Gla01, Gol83, Gol13, Hal31, Har90, Har95a, Har98, Har02, Hen15, Hen86, Hut98, Tru14, Jea31, Jon32, Jon49, Jon73, Jon80, JTR94, JTR95, Kno14, Lam11, Lam03, Lan03, Lar26, LF08, Mac85, Mah03, Mar66, McC11, McN04, Mit17a, New55, Niv90a, Niv90b, Niv90c, Niv90d, Niv27, Niv52, Niv65, Niv03, Niv11, OR97a, OR97b, Ogi08a, Ogi08b, Pla31a, QM15, Ran62, Ran63, RWT08, Rei08, Rei15, Rod02, Row02, Sei01, Sie85a, Sim14, Sta12, Tai80b, Tai11a, The86, The87, Tho31a, Tho31b, Tho11, Tol81, Tol82, Tro01, Tur53, Wal16, Wes99, Wil05, Eve84, Gee81, Hun88a, Hun88b, Mac15, Mor91a, Mor91b, Ric14, Sho14, Smi82a, Smi82b]. Japanese [Hes86]. JCMB [Ano16]. Jed [Ner87, Can90]. Jenkin [Cha71, Mit17a]. Johannes [Cle74a]. John [Chi64, Hun88a, Hun88b, Hun06, Mor91a, Mor91b, Sim14, The87]. Johnstone [O’H75]. Join [McN04]. Joint [Sim14]. jointed [Cha89]. Joseph [FSvL^{+}37, Pag38]. Jungk [Hei72].

K.G. [Cle79c, Cle67]. Kagaku [Hes86]. Kaiser [vM82]. Kelland [Cle73r]. Kelvin [Tho10, CDKR16, Dou16, LS14, Mac64, SW89, Tho76, Tho05, Wis77, Spe65]. Kepler [Bre06a, Bre06b, Mye06, Spr06, Hal04]. Kind [Cle73a, Cle73m]. kinds [Cle60a, Cle69c]. Kinetic [Ach87, Ber63, Bru57a, Bru61, Bru62, Bru71, Cle73e, Cle73j, Cle77a, Gar66, Gar70, Gar08, Bru57b, Bru58b, Bru58a, Dor70, Tho92]. King [Ano17a, CJ63b, CJH64, Cle79, Hea29]. kinships [GS06]. Kirchhoff [Cle74h, Cle75e]. Klassiker [Bes04]. Klein [KS95]. know [Nah83]. Knowledge [Cha21]. Knox [Wea16]. Körper [Cla58]. Kraftlinien [Cle95, Cle62a, Cle12, Cle86].

L [Sto82, Cle64b, Cle70e]. laboratory [Gar31, Ano03, Cro74, Fit10, Fit16, Gla31, Kim02, Lod27]. Lady [Cle19]. land [C^{+}69, C^{+}70b, C^{+}71b, C^{+}72, C^{+}73, Cle74j, C^{+}76a, C^{+}76b, C^{+}77b, C^{+}78, C^{+}79e]. Landauer [Ben03, EN99, LC15, Nor05]. Länge [Cla58]. language [Ari03]. Laplacian [Rap05]. Largs [Tho10]. Larmor [FSvL^{+}37, Pag38]. lasers [Smi63]. Last [Ner87, Ano06, Buc85a, Heh10]. lata [Nat32]. late [Lar07, Lar71, Lar10]. Laue [RSB^{+}92]. Law [Pup25, Tai97, Cle68k, C^{+}77c, Fal17, FMM96, Pla94, Pla95, Den07, Lef02]. Laws [Cle58b, Gal69, Gal79, Gal10, Pic08]. layered [WG08]. Leadership [Kim02]. leading [Cro01]. Learn [Tur96]. Leben [Hei88]. Lebensbeschreibungen [Lén29, Lén43]. Lecons [Cle91a]. Lecture [Cle56b, Cle56g, Cle6x, Cle78i, Cle90c, Cle19, Flo13, Fra57, Tho82, Cle71a, Cle79, Dav60, Cle66e]. Lecturer [Lam31]. Lectures [Bol93, Cle84a, Cle84b, GCC81, Mac19, Chi64, Cle51b, Dom63, Sim06]. Leff [She04a, She04b]. Legacy [Lon16]. Lehrbuch [Cle83a]. length [Cla59]. Lengths [Cla58]. lessons [Cle91a]. Letter [Cle70e, Cle07, CG68]. Letters [Buc92, Dar98a, Dar98b, Dom04, Har90, Har95a, Har02, Hun05a, Jon92, Pag38, Sie93a, Sie93b, Dom92, Dom97, Jam91, JTR94, JTR95, Lar36, Lar37, Seg92]. Level [Sho14]. Lewis [Ano83]. LI [Cle61a]. Liboff [OM11]. Lichtes [vM82, Bol93]. Lie [Smi17]. Life [Ano83, CG82, CG84, CG10, Cat16, Cha84, Hei88, Hen15, Hun06, Tru14, Kno11, Ran62, Sho14, Sie85a, Wil05, CG69, Cro01, FMW14, Jon32, Lén29, Lén43, Mah03, McC11, Nat32, Pip04, Ran63, Ric14, Tho10, Tho76, Tho05, Toy99, Wil51, Wil72, Wil16]. Light [BWB19, Can90, D’A81, D’A86, D’A01a, Kra18, Mar93, Bol93, Bro67, Buc89, Cle57f, Cle57g, Cle62d, Cle68b, Cle68c, Cle69e, Cle69b, Cle72d, Cle74g, Dea72, Mah15, Rau14b, Sie91, Att93a, Att93b, Buc93, Hun92]. lightning [Cle77i]. Lignin [FSvL^{+}37]. Limitations [Cha73a]. Limits [Cha21, Cle77g, Cle80c, HB03]. linear [KNP84]. Lines [Cle58a, Cle73n, Cle55a, Cle56c, Cle56d, Cle57d, Cle61a, Cle62a, Cle61c, Cle61h, Cle61g, Cle62c, Cle62d, Cle64c, Cle69c, Cle95, Cle12, Cle86, Lam11]. linkage [Cle76h]. Liouville [Ain11]. liquid [Cle74a]. Liquides [Cle74d, Cle74k]. liquids [Cle74k]. Lise [Büh98b, Büh98b]. Listing [Cle69c]. Litchfield [Lit03, Lit10]. Literary [Jos07, Cle76a]. Literature [Wis82]. Lived [RSB^{+}92]. lives [Art16]. LL.D [Tro01]. loading [WG08]. localities [C^{+}69, C^{+}70b, C^{+}71b, C^{+}72, C^{+}73, Cle74j, C^{+}76a, C^{+}76b, C^{+}77b, C^{+}78, C^{+}79e]. Locality [PS96]. Localization [SE08]. Löcher [Bes04]. Logarithmorum [Sim14]. Logic [Tur56]. London [Ano83, Can90, Chi64, CJ63b, CJH64, Cle73r, Cle76a, Har93, Har96, Hun89, Jon64, McC69, Tro01, Cle79, Dom80, Hea29, QM15]. lonely [Mul04]. Long [Rau14b, Rau14a, Cle78h, NV19]. Looking [Har95b]. Looking-Glass [Har95b]. Lord [Sch26, SW89, Tho76, Tho05]. Lorentz [D’A00, Ive05, Ryn02, VT89, Yan95]. Lorenz [Kra18, McD18]. Loschmidt [Cle73j]. lotus [Nor05]. Ltd [Hun89, Tur83]. Ludvig [Kra18]. Ludwig [Cle77j, vM82]. luminiferous [Cle80d]. LX [Cle73f]. LXI [CJ65a]. lxvi [McC69]. LXVIII [Cle08a, Lar08]. LXX [CJ65b].

M [Att93a, Att93b, Buc92, Buc93, Dar98a, Dar98b, Dar01a, Dar01b, Dom04, Eve99, Hun92, Hun00a, Hun00b, Hun05a, Jam91, Jon92, Mar93, Pip99, Seg92, Sie85a, Sie93a, Sie93b, Smi00a, Smi00b, Yos99]. M. [Cha89, LdB12b, LdB12a]. M.A [Cle73b]. MacAdam [Sto82]. MacCullagh [Dar10]. MacDonald [Spe65]. Machine [Hei72]. machinery [Moy77]. machines [Bro00a]. Macmillan [Ano83]. Macroscopic [RS11, Yan95]. macroscopically [Wig90]. made [CJ63b, CJH64, Rey21]. magic [Cen89]. magnetic [Cle61h, Har08, Len19]. Magnetism [Cle71h, Cle73t, Cle81c, Cle83a, Cle84, Cle85c, Cle89, Cle92b, Cle04b, Cle53, Cle54, Cle55, Cle98a, Cle98b, Ryn16, Wis79, Cle62d, Kle89, Sim70, Tho93, Tho68, Tur27, Tur72, Wis77, Ano11a, Cle73d, Tai73, Cle15, LM81, Sim68, Tho68, Tho93]. magnetisme [Cle84, Cle85c, Cle89]. Magnetismus [Cle83a, Cle15]. Magnetizme [Kle89]. magneto [CG68]. magneto-electric [CG68]. magnets [Cle67d]. Mahon [Hun06, Jam15, Pip04]. maintenance [Cle67d]. make [Nah83]. Makers [Ano99a, Tur27, Tur72, Ano99b, Kle10, Mar99b]. Making [Mit17a, Cle68b, Cle68c, Cle69b]. Maksuelia [Kle88]. Maksvell [Pol85]. Man [Hun06, Wil05, Eve84, Mah03, Pip04]. Mance [Hop59b]. manuscripts [Cle79c, Cle67, Ful86]. Many [Bri18, RdO07, RdO16]. Marc [Mac15]. Marconi [Zit02]. Marian [Ber08]. marine [SBB^{+}08, Wil55]. Mark [Cat16, Hen15, Ric14, Sho14]. Martin [Cha84, Goo85a, Goo85b, Mat85a, Mat85b, KS95, Smi85]. Marx [Jos14]. masers [Smi63]. Mass [Fed08]. Master [Mul08, Cle76a]. Masters [War03, Row05]. mate [Ano06]. Material [Chr06, Cle79f]. materials [WG08]. math [Luc12]. Mathematical [Nye02, Bat15, Bat55, Cle71f, Cle73b, Fis88, Pla99, Row05, War03, Cle70b, Cle70a, Tur83]. Mathematically [GZ11]. Mathematician [Sch96]. mathematicians [Nah83]. Mathematics [Ati15, Row05, Twe11, Ari03, Ber84, Cat12, Cle74i, C^{+}79a, Cra07, FF91, Kim03, Sim14]. mathematischen [Pla99]. Matter [Cle73g, Cle76e, Cle78e, Cle88b, Cle92a, Cle20a, Cle20b, Cle20c, Cle25, Cle52, Cle96, Cle97, Cle02, Epp98, Har82a, Lar00, Cle79j, Cle81b, Cle06, Tai77]. Matthew [Dea16]. Max [Hei72, Hei88, Hei88, Hof08]. Maxwell [Ano79c, Ano81a, Ano82, Ano83, Ano31a, Ano74, Ano16, Att93a, Att93b, Ber08, Bes04, Bla83a, Bla83b, Bol93, Boo64, Bri18, Buc92, Buc93, Büh98a, Cat16, Chi64, Col98, Dar98a, Dar98b, Dea16, Dom97, Duh02, Duh67, Eve84, Flo13, For99a, FSvL^{+}37, Gar79, Hei72, Hen15, Her84, Hun88a, Hun88b, Hun92, Hun05a, Jam91, Jam15, Jon64, Jon92, Jos14, Lin71, Mac64, Mac85, Mar93, Mar66, Mau17, McC69, Mor91a, Mor91b, Ner87, Pag38, Pla31a, Ric14, Rid30, Roc87, Seg92, Sho14, Sie85a, Sie93a, Sie93b, Sim14, Smi82a, Smi82b, Spe65, Spr06, Sto82, The87, Tro01, Tur83, Wil68, Ach87, Ach13, AVJ07, Ain11, All13, Ano79a, Ano79b, Ano81b, Ano85, Ano31b, Ano65, Ano91, Ano99c, Ano99d, Ano08, Ano11b, Ano11a, Ano12, Ano13, Ano15, Ano17b]. Maxwell [Ano17a, Ano17c, Anoxxa, Anoxxb, Anoxxd, Anoxxc, AL30, AB86, Arm04, Art13, Ati15, Bat15, Bat55, Ben03, Ber84, Ber63, Bes04, BDP15, Edi81, Boh31, Boh17, Bok15, BP78, Bor67, Bro67, Bro68, Bro00b, Bru58a, BE69, Bru71, BEG83, Bub01, Buc85a, Buc85b, Büh98a, Bun43a, Bun43b, CG82, CG84, CG69, CG10, Cat01, Cat12, Cat14, Cat16, Cen89, Cha73a, Cha73b, Cha86, Cha01, Cha21, Cha71, Cha89, Cha85, CF91, Chr06, Coh52, Col97, Col90, CPR11, Cra54, D’A12, D’A81, D’A86, D’A89, D’A00, D’A01a, Dar93c, Dar10, Dau70, De 78, Dev02, Dew18, Dia94a, Dia94b, Dib64, Dom63, Dom80, Dom18, Dor70, Dou16, Duh02, Duh67, Duh15a, Duh15b, Duh15c, Duh15d, Duh15e, Dun08, Dys90, Dys99, EN98, EN99, Edw10, Ehr67, Ein31a, Ein31b, Ein54b]. Maxwell [Ein82b, EM84, Eld00, EP81, Eng15, Eva61, Eve67b, Eve67a, Eve67c, Eve75, Eve83, Eve84, Eve98, Eve06, Fah96, Fai31, Fal15, Fal17, Fed08, Fer31, Fis88, Fle08, Fle31, Fle86, Flo12, FMW14, FM14, FMM96, Fra57, Ful86, Gar66, Gar69, Gar70, GBE86, GBE95, Gar08, Gar85b, Gar31, Ger01, Gia70, GZ11, Gla96, Gla01, Gol94, Gol83, Gol13, Gor83, Gye17, Hal80, Hal31, Hal08, Hal04, Har88, Har90, Har95a, Har95b, Har96, Har98, Har02, Har09, Heh10, Hei70b, Hei70a, Hei71, Hen86, Her84, HB03, HG12, Hoo87, Hop56, Hop57a, Hop57b, Hop58, Hop59a, Hop59b, Hos16, Hou16, Hun06, Hun15, Hut98, Ive05, Tru14, Jam91, Jea31, Jon32, Jon49, Jon73, Jon80, JTR94, JTR95, KCP14, KS16, Kim03, KWK73, Kle88, Kno14, Knu78, KNP84, Kra21]. Maxwell [Kuw31, Lai74, Lam31, Lam11, Lam03, Lan03, Lan98a, Lan98b, Lan98c, Lan14, LU31, Lar26, Lar31, Lar36, Lar37, Lef02, LR03, LM81, LL19, Lod27, Lod31, Lon08, Lon16, Lor23, LF08, LMJ14, LC15, Mad02, Mah03, Mah15, Mar07, Mar66, Mat87, May71, McC11, McD18, McN04, Men08, Mir99, Mit17a, Mit17b, Mor92, Mou08, Moy77, Moy78, Mul04, Mul08, Mus99, Nah92, Nap21, Nat32, New55, NA02, Niv90a, Niv90b, Niv90c, Niv90d, Niv27, Niv52, Niv65, Niv03, Niv11, Nor05, OR97a, OR97b, Ogi08a, Ogi08b, OM11, Ozi94, Pei63, Pip99, Pip04, PS96, Pla94, Pla95, Pla98, Pla99, Pla03, Pla31a, Pla31b, Poi99, PV04, Poi07, Pol85, Por81, Por94, QM15, Ran62, Ran63, Rap05, RS11, Rau14b, Rau14a, RWT08, Rei08, Rei15, Rex17, Rey21, RSB^{+}92, Rod02]. Maxwell [RdO07, RdO16, Row02, Ryn02, Ryn16, Saa05, Sch14, Scu07, Sei01, Sel09, Sha73, SGD^{+}02, SPW02, She18, She99, She15, Sie75, Sie85c, Sie91, Sie95, Sim66, Sim68, Sim70, Smi17, SR05, Smi85, SB50, Squ89, Sta12, Sta15, Sta17, Tai73, Tai77, Tai80a, Tai80b, Tai86, Tai95, Tai97, Tai11a, Tai11b, TS09, The86, Tho79, Tho82, Tho89, Tho93, Tho31a, Tho31b, Tho68, Tho11, Tho91, Tho92, Tol81, Tol82, TS86, Top70, Tor00, Toy99, Tur53, Tur55a, Tur55b, Tur56, Tur96, Twe11, VT89, Vis89, Vor08, Wal16, Wal85, Wea16, Wei02, Wes99, Whi85, Wig90, Wil79, Wil05, Wis77, Wis82, Yag07, Yag14a, Yag14b, Yan95, Yan14, Yos99, Zit02, vB98, Bre06a, Bre06b, Cha84, Dar01a, Dar01b, Dom04, Eve99, Gee81, Goo85a, Goo85b, Har14]. Maxwell [Hun00a, Hun00b, Kar14, Kra13, Mac15, Mat85a, Mat85b, Mye06, Sie85a, Smi00a, Smi00b, Sta14, Dom92, Mar99a, Sch89, She04a, She04b, Sie85b, Sie96, Wil97, Wil05]. Maxwella [Nat32]. Maxwellian [Buc79a, Buc79b, Buc85b, Den07, Kra18, Myr11, Nug16, Var82]. Maxwellians [Bri89, Har93, Hun89, Hun84, Hun91, Hun94, Hun05b, OP87, Dar93a, Dar93b, Jor93]. maxwellienne [Dar98c]. Maxwells [Bol93, Cle15, vM82, Art16, Jos07]. Maxwell’sche [Pla03, Pla98, Pla99]. Maxwell’schen [Pla95, Kai87, Pla94, SB50]. McCartney [Cat16, Hen15, Mac15, Ric14, Sho14]. Mean [Cla58, Cle72c, Cla59, Cle72h]. meaning [Ner84]. means [Cle69f]. Measurement [Dev02, Hop57a, Hun15, Sto82, CJ63b, CJH64, Fal17, Gré19, Hop59b, Mit17a]. Measurements [CJ63a, CJ65a, CJ65b, Cle73k, Mit17b]. measuring [Cle56a]. Mechanical [LL19, Sie85c, Cha86, Cla59, Cle67d, C^{+}77a, C^{+}79d, C^{+}79, Mit17b]. Mechanics [All13, Har96, Sie96, Bru67, CGBE95, GBE95, Moy78, Myr11]. mechanische [Cla58]. Mechanism [Cha01]. media [Vor08]. medium [Cle54a]. Meeting [Sim14]. Meitner [Büh98b]. Meloid [Cle46b]. Member [Sei01]. Memoir [Lar07, Lar71, Lar10]. memories [Fle31]. Men [Jam15, Cra07, FM14, Gal74, Gal70, Kle10, Lén29, Lén33, LH33, Lén43, Len70]. Metal [Gar04, Pla03]. Metallic [Gar04]. Metalloptik [Pla03]. Metaphor [Cat01]. metaphysical [Cat12]. Metaphysics [Har82b, Cat12]. Method [Cat16, Har14, Kar14, Kra13, Sta14, Tur55a, Ach13, Bok15, Cle57d, Cle68b, Cle68c, Cle69b, Cle76h, Fal17, Hop59b, Jea31]. Methodology [Tur53, Cha73b, HG12]. Methods [Cat01, Sch17, EP81, Nah83]. Michael [Jam12]. Michaelson [AL30]. micro [GRGR08]. micro- [GRGR08]. Microphysics [Ner87, Roc87, Buc85a]. Microrotors [Ano99d]. microscopic [Yan95]. microwhiskers [GRGR08]. millennium [Ano99a, Ano99b, Mar99b, Sch14]. Minds [Bro00a, Pic08]. Mirifici [Sim14]. mittlere [Cla58]. Mixed [Row05]. mixture [Cle60e, Cle61b]. mixtures [Cle57f]. Mobile [She99]. mode [Cle80d]. model [Cha86]. Modeling [She15, HG12]. Models [Hes63, Hes66, Hes70, Hes86]. Modern [ADS83, Boh31, Chi64, Gib19, Jon64, Nye02, Pyl96, Boo64, Dom63, GS06, Gla96, Gla01, Gol13, Hei79, Hei99, Kle10, Whe11]. moderu [Hes86]. Modes [Hun15, Hei70a]. Modified [Gia70, Cle82c]. moléculaires [Cle74k, Cle74d]. Molecular [Att93a, Att93b, Buc93, Cle75f, Cle75g, Cle75h, Cle76f, Cle79h, Hei70b, Hun92, Mar93, Cla59, Cle61h, Cle61g, Cle62c, Cle62d, Cle74k, Cle79a, Dia94a, EP81, Sie91]. Molecularbewegung [Cla58]. Molecülen [Cla58]. Molecules [Cla58, Cle73a, Cle73m, Cle73h, GBE86, Sch89, Cat14, Cla59, Cle68k, Cle73f, CGBE95, GBE95, Pla94, Pla95, Har96, Sie96]. moments [Büh98b]. moral [Cat12]. Moseley [Hei72]. Motion [Cle57c, Cle57i, Cle59b, Cle59c, Cle73m, Cle73g, Cle76e, Cle78e, Cle79j, Cle81b, Cle88b, Cle92a, Cle20a, Cle20b, Cle20c, Cle25, Cle52, Cle96, Cle97, Cle02, Cle06, Bat15, Bat55, Cla59, Cle70f, Cle73o, Cle74f, Cle74e, Cle76i, Cle76l, Cle80d, Tai77]. Motions [Cle60g, Rap05, Cle60c, Cle69e]. moving [Cle60a, Hug69]. Mr [Cra07]. Mr. [CG68, Cle69g, C^{+}75a]. multiple [Cle77g, Cle80c]. multiverse [Bro00a]. Munchausen [She99]. Mutual [Wis79]. My [Hei72]. Mystery [Scu07, Bes04, Ber08]. Mystics [Ber08, Scu07].

N.Y [Jam15]. naked [Mar66]. Namesake [Ano08]. Nancy [Jam15]. nano [GRGR08]. nano-centipedes [GRGR08]. nanocomposites [GRGR08]. Napier [Sim14]. narrow [Cle74g, Cle78h]. Natural [Coh94, Dar01a, Dar01b, Dav60, Epp98, Eve99, Fra57, Har82b, Hun00a, Hun00b, Smi00a, Smi00b, Yos99, Cat12, Cle73c, CDKR16, Eve67a, Eve75, Har98, Pip99, Cle79b]. Naturalistic [Dea16, Sta15]. Nature [Cha01, Gia70, Hei71, Cle82a, Col90, Gal74, Gal70, Nap21]. Naturforscher [Hei72, Lén29, Lén43]. Naturforschung [Lén29, Lén43]. near [GS06]. nebst [Cla58]. nebulæ [Hug69]. necessity [Cat12, Cle82b]. Negative [Fed08]. Neohumanist [Sch96]. net [Chi64]. News [Kra21]. Newton [Bes04, Cat16, Har14, Kar14, Kra13, Sta14, Ach13, Büh98b, Dun08, Har88, Ozi94, TS09, Wea16, Jos14]. Newtonian [Rey21]. Newtonnienne [Rey21]. Newtons [Büh98b]. Nineteenth [Bla83a, Bla83b, Can90, Ner87, Sto82, Whi10, D’A01b, Har82a, Har85, LB^{+}04, Tur83, Whi08, Buc85a, Buc89, D’A81, Mac19, Wil62, Wil68]. nineteenth-century [Har82a, LB^{+}04]. Ninth [C^{+}77b]. No [Fal17, KS95]. nomenclature [C^{+}74, C^{+}75b]. Non [Kra18, Lin71, D’A89, Dor70, HB03, Vor08]. Non-Relativistic [Lin71, HB03]. non-speculative [Dor70]. non-stationary [Vor08]. Nonrelativistic [Var82]. nonstandard [Col97, Col98]. normal [Gye17]. Norman [FSvL^{+}37]. Note [D’A12, Tai95, Tai97, Cle68b, Cle68c, Cle69b, Tur55b, Tai86]. Notes [Ben03, Cle51c, Cle51b, Cle51d, Clexxa, Tho93, Tho68, Toy99, Dav60, Fra57]. Noteworthy [GS06]. Notions [Epp98]. November [Cle90c, OR97b]. novembre [LdB12b, LdB12a]. nu [Mit17b]. nuke [Ano99d]. null [Fal17, Gol94]. number [Hop59a]. Numbering [Hei72]. Numbers [Smi17]. numerical [NV19]. numerous [Chi64]. nurture [Gal74, Gal70].

O [Ber08]. ob [Kle89]. Obituary [Ano79c, Ano81a, Gar79, Tai80b]. Objection [McD18]. Objectivity [Wei02]. obrabotkie [Kle88]. Observation [RSB^{+}92]. observations [Hug69, Sim70]. Obtained [Cle59c, Yan95]. Occasional [CG10, Ano83, CG82, CG84]. occurrence [Cla59]. octobre [LdB12b, LdB12a]. O’Hara [Bri89, Hun89]. Ohm [C^{+}77c]. Oliver [Buc85b, Nah83]. ondes [D’A89]. One [Ger01, Gib19, Mar99b, Sta09, Cle60a, Cle68k, Hop59a, PV04]. One-Photon [Ger01]. Online [CF98]. only [Cle55a, Cle56c, Cle56d, Cle64a, Cle66b, Cle67b, Cle68b, Cle68c, Cle70b, Cle72a, Cle74k, Cle78g, Cle27a]. ont [Rey21]. Open [Bee96]. Operators [Rap05]. opinion [Cle82b]. Opinions [Ein54a, Ein82a]. opposing [Her84]. Optical [Can90, Cle58b, Bat15, Bat55, Buc89, Cle66a, Cle75d, Dar10]. Optics [BWB19, Knu78, Cle74i, Pla03]. Origin [Mou08, Tai11b]. original [Cle72g, Cle79c, Cle67]. Origins [For99a, Lar36, Lar37, May71, Por94, Pyl96, Tur96, Ste05, Ste16, Yan14, FSvL^{+}37, Pag38]. orthotropic [WG08]. Osborne [BE69]. oscillations [Poi99, PV04, Poi07]. Oskar [Kai87]. other [Cla59, Cle66b, Cle66e, Cle73p, Cle74i, Fit10, Fit16]. Outlook [Eld00]. Oval [CF51]. Ovals [Cle46a]. Oxford [Ano82, Hen15, Mac15, Wil68].

P [Cle73c, Dar98a, Dar98b, Dom04, Hun92, Hun00a, Hun00b, Jam91, Jon92, Sie93a, Sie93b, Smi00a, Smi00b, Sto82, Ano83]. P. [Buc92, Hun05a, Pip99, Yos99]. page [Cra54]. pain [Cle82c]. paper [Cle56a, Cle69c, Cle69g, Cle72d, QM15]. paperback [Can90, Row05]. Papers [Buc92, Cle73d, Dom04, Eve67c, Har90, Har95a, Har02, Hof08, Hun05a, Jon92, Kno11, Sch96, Sie93a, Sie93b, Cle0 , Cle21a, Cle27b, Cle51, Dar98a, Dar98b, Dom92, Dom97, Hea92, Jam91, Niv90a, Niv90b, Niv90c, Niv90d, Niv27, Niv52, Niv65, Niv03, Niv11, Seg92, Wil51, Wil72, Wil16]. para [Var82]. para-Maxwellian [Var82]. paradox [Cle77f, Cle80b, Den07]. Paradoxical [Cle90d, Cle78a]. Paraffin [Boh17]. Paragon [Pip04]. Paris [Spr06]. part [Sim70, Buc79a, Buc79b, Cle55a, Cle56d, Cle60a, Cle60b, Cle60g, Cle61h, Cle61g, Cle62c, Cle62d, EN98, EN99, JTR94, JTR95, PV04]. Partial [Sch17]. Particle [Bin08, Men08, Squ89]. Particles [Cha21, Cle60a, Wig90]. particular [Cle54a]. parts [Cle74i]. passage [Cle72d, Dea72]. Passes [vB98, Mar99a]. Paths [Cla58, Cla59]. Paul [Bla83a, Bla83b, Cha84, Smi85, Tur83]. pays [Rey21]. Peaucellier [Cle76h]. Peierls [Chi64]. Pencil [Cle73n]. People [Ano17a]. perceived [Cle57a, Cle57h]. perception [Cle57b]. Peregrinus [Hun89]. Perfectly [Cle60g, Cle60b]. performing [Yan95]. Pergamon [Wil68]. Periodic [CF91]. permanent [Cle67d]. Permittivity [Boh17]. Perpetuum [She99]. perplexed [Nah83]. Persistence [Sta17]. Personal [Har96, Sie96, CGBE95, GBE95]. persons [GS06]. Perspectives [Cat16, Coh94, Hen15, Ric14, Sho14, FMW14, Sel09]. Peter [Cle73r, Dar01a, Dar01b, Eve99, Hun89, Cat16, Har14, Kar14, Kno11, Kra13, Seg92, Sta14]. Phase [SGD^{+}02]. Phenomena [Cle57c, Cle57i, Cha86, Cle61h]. Philadelphia [Sto82]. Philip [Cle73r]. Philoponos [Tor00]. philosopher [Eve67a, Eve75, Fis88]. philosophical [Sel09]. Philosophy [AB86, Cle90d, CF98, Dav60, Epp98, Eve99, Fra57, Har82b, Hun00a, Hun00b, Pyl96, Smi00a, Smi00b, Yos99, Cat12, Dar01a, Dar01b, Har98, Pip99, Cle78a, Cle79b, Cle73c]. Photograph [Eva61]. Photon [Ger01]. Phys [Col98]. Physical [AVJ07, Cle70b, Cle70a, Cle73p, Cle74b, Cle74c, Cle76n, Cle85a, Ein31b, Ein54b, Ein82b, Nye02, Toy99, Tur55a, Wil62, Bok15, Cat12, Cle61a, Cle62a, Cle61c, Cle61h, Cle61g, Cle62c, Cle62d, Cle71f, Cle82b, Cle84a, Cle84b, Cle86, Ein31a, Gré19, Lan14, Nap21, Bun43a, Bun43b]. Physicist [Ano15, Rod02, Eve67a, Eve75, Fis88, Jon80]. physicists [Bes04, Cro01, Har85, Mac19, Gar85a]. Physics [Bea13, Boh31, Cle6x, Cle79a, Cle79h, Har82b, Har88, Hop56, McC69, Pip04, Por81, Por94, She04a, She04b, Mor95, Buc95, Büh98b, Cle71a, Cle74i, Cle76f, C^{+}79a, D’A01b, De 78, EP81, Eve06, FF91, Fis88, FM14, Gla96, Gla01, Har82a, Har85, Hei79, Hei99, Kuw31, Lon03, Pla31a, Pla31b, Pol85, Rey21, Tai77, War03, Jam15, Mau17, Row05, Cle73b, Hei72]. Physik [Büh98b, Pla31a]. physikalische [Cle62a, Cle86]. Physiker [Bes04]. physique [Rey21, Rey21]. PIA09857 [Ano08]. pin [Cha89]. pin-jointed [Cha89]. Pioneers [App30, App68]. pl [Ano83]. Planck [Hei88, FMM96, Hei88, Hof08]. plane [Cle56a, Cle72h, Cle72i]. planetary [NV19]. plankton [SBB^{+}08]. Plasma [Mau17]. plasmas [SE08]. Plateau [Cle74d]. plates [Chi64, Wil68]. platometer [Cle56a]. Plurality [CF51]. poetic [McC11]. Poetry [Coh52, Lan03, Row02]. Poinsôt [Cle57e]. Point [Cle57c, Cle57i, Cle69f]. Points [Cle79f, Cle71c]. polarised [Cle62d, Cle72d, Dea72]. Polish [Nat32]. Polyuronides [FSvL^{+}37]. Popularized [Wil05]. Popularizing [Gee81]. Portraits [SGD^{+}02, Fit10, Fit16]. positive [Cle71e]. possession [Cle79c, Cle67]. possible [Cle80d]. Potential [Bor67, Rap05, Cle69f]. Potentials [McD18]. Pourquoi [D’A89]. power [Kle10]. powerful [Nah83]. Pp [Buc92, Can90, Chi64, Har93, Har96, Hen15, Hun89, Hun05a, Hun06, Jam91, Jon64, Mar93, McC69, Pip99, Pip04, Roc87, Row05, Smi85, Spr06, The87, Tur83, Wil68, Yos99, Ber08, Eve84, Hun92, Jam15, Lin71, Mac15, Mar99a, Seg92, She04a, She04b, Sto82, Gee81]. Practical [Cle79h, Smi63, Cle79a]. Practice [Sta09, Buc95]. Pre [Lan98a, Lan98b, Lan98c]. Pre-Maxwell [Lan98a, Lan98b, Lan98c]. Precise [CG17]. precision [Gré19]. Predicting [GRGR08]. preferred [Var82]. Prehistory [Sus69a]. Preliminary [Cle73q]. present [Art13, Wil55]. presentation [QM15]. Press [Ano82, Buc92, Can90, Chi64, Dom04, Eve84, Har93, Hen15, Hun05a, Jam91, Jon64, Mac15, Mar93, Pip99, Roc87, Row05, Seg92, Wil68, Yos99]. Presses [Har96]. Preußischen [Sim06]. Price [Chi64]. Pricha [Hun89, Bri89]. Primary [Cle61f, Cle11, Cle61e]. Principia [Har88]. principle [Ben03, Nor05]. Principles [BWB19, Cle73b, PV04]. Prize [Mau17, Cle59c]. probability [Gye17]. Problem [Har82b, Lan98a, SB50, Wil79, Cle73i, Cle74h, Cle75e, Cle76g]. Problems [Cle54b, LF08, Wil62, Cle76m]. Proceedings [Sim06]. process [Cle60a]. processes [Kle88]. procure [C^{+}79a]. producing [Cle69a]. proefschrift [Cle74a]. Proell [RS11]. Professor [CF51, Fle31, Fra57, Tho93, Tho68, Tro01]. Professors [Cle73c]. progrès [Rey21]. progress [C^{+}79a, Cle82b, Lén33, LH33, Len70, Rey21]. Prometheus [Jam15]. Proof [Ive05, LL19, Cle73o, Cle76l, Dys90]. proofs [Cra54]. Propagation [BWB19]. properties [GRGR08]. proposals [Fah99a, Fah99b, Fah01, Hop58]. proposes [CL15]. Proposition [Hut98, Dia94b]. prort [Ano83]. protection [Cle77i]. protesters [Ano99d]. proves [Dia94b]. Prussian [Sim06]. Psychophysik [Cle82d]. publication [Mar07, Sim14]. publicly [GS06]. Published [Kno11]. Publishing [Cha84, Smi85]. Pure [Cle69a, D’A00]. Pure-Field [D’A00]. purpose [C^{+}69, C^{+}70b, C^{+}71b, C^{+}72, C^{+}73, Cle74j, C^{+}75a, C^{+}76a, C^{+}76b, C^{+}77a, C^{+}77b, C^{+}78, C^{+}79a, C^{+}79d, C^{+}79e, C^{+}79]. Put [Ain11]. Puzzle [Bro00b]. pyrometer [Cle73q, C^{+}75a]. Pythagorean [Ber08, Scu07].

Quadratic [Cle73p]. quanta [LdB12b, LdB12a]. Quantenrätsel [Bes04]. quantities [Cle71f]. Quantum [Ber08, Bes04, BDP15, Ger01, Lan98c, Rap05, Scu07, Bro00a, LR03, She04a, She04b]. Quarter [Ner87, Buc85a]. quatercentenary [And07]. Quaternions [Cle73r, Cle73r]. que [Rey21]. quest [Bro00a]. Question [Cle69f, Cle71e]. Quetelet [Por94]. Quotations [Cle84c, EM84].

R [Chi64, CG68, Lin71, Wil68, Sch17]. Radiation [Bri89, Hun89, FMM96, OP87, Smi63]. Radio [App63, Bow63]. Radiometer [BE69]. radiotelegraphy [Sus69a, Sus69b]. railway [Wil55]. Rainbow [TC74]. Randall [Chi64]. Random [Mar99a]. Rankine [Moy77]. rapports [LdB12b, LdB12a]. Rarefied [Cle79g, Den07, Cle78g, Cle78f]. rate [C^{+}69, C^{+}70b, C^{+}71b, C^{+}72, C^{+}73, Cle74j, C^{+}76a, C^{+}76b, C^{+}77b, C^{+}78, C^{+}79e]. Ratio [Cha85, Cle70c, Mit17b]. ratio-of-unit [Mit17b]. Ray [RSB^{+}92]. Rayleigh [Sch94, Sch26]. Raymond [Cat16, Mac15, Ric14, Sho14, Hen15]. rayonnement [LdB12b, LdB12a]. razvitie [Pol85]. re [Hop58, Mit17b]. re-determination [Hop58]. re-examination [Mit17b]. readership [Sho14]. real [Cle68d, Cle82a]. real-image [Cle68d]. Realist [Gia70, Saa05]. Reality [Ein31b, Ein54b, Ein82b, Ein31a, Lan14, Sie85c]. reasoning [Lon03]. received [Ano82]. Reciprocal [Cle68g, Cle70d, Cle72a, Cle64d, Cle69d, Cle72f]. Recollections [Tho36, Tho37, Tho75]. Reconsidering [Saa05]. Record [And07]. recorded [GS06]. Rede [Cle78i]. reduction [Cat14]. reference [Cle57f, Var82]. Reflecting [Cle19]. Reflections [Luc12, Tho36, Tho37, Yag07, Yag14a, Yag14b, Tho75]. reform [Cra07]. Refracted [Cle73n]. refracting [Cle72d, Dea72]. refraction [Cle69e, Cle74f, Cle74e]. Refractive [Boh17]. Refusal [McN04]. regarding [Tho91, Tho92]. regions [Cat12, Cle69c]. Reihe [Kai87]. relating [Ful86]. Relation [Cle73j, Cle55b, Cle56e, Cle68g, Cle74i]. Relations [CJ63a, Cle60f, CJ65a, CJ65b, Cle73k, Mit17b]. Relationship [KWK73]. relationships [Her84]. Relative [Boh17]. Relativistic [Lin71, HB03]. Relativitätstheorie [Bes04]. Relativity [Bes04, Dun08, Ell66, Hal08, VT89, Wan08]. relevant [Luc12]. religion [The86]. Religious [Eld00, Sta17]. Remarks [CF51, Cle57a, Cle57h, Cle69g, Cle71f, Cla59, Cle5x]. Remote [Hos16]. Report [C^{+}64, C^{+}65, Cle66c, C^{+}68, C^{+}69, C^{+}70a, C^{+}71a, C^{+}71b, C^{+}72, C^{+}73, Cle73q, C^{+}75a, C^{+}77c, C^{+}79a, C^{+}79b, C^{+}70b, C^{+}74, Cle74j, C^{+}75b, C^{+}76a, C^{+}76b, C^{+}77a, C^{+}77b, C^{+}78, C^{+}79c, C^{+}79d, C^{+}79e, C^{+}79]. Reports [Cle79i, C^{+}79a]. representation [Hei70b, Hei70a]. Representing [Twe11]. Reprint [McC69, Fit16, Cle73d]. required [Fal17]. Research [LF08, Lén29, Lén43]. Researches [Fal15, McC69, Wil55, Cle79c, Cle21a, Cle67, Tho93, Tho68]. resistance [CJ63b, CJH64, C^{+}64, C^{+}65, Cle66c, C^{+}68, C^{+}70a, C^{+}71a, Hop57b, Hop58, Hop59b]. resisting [Cle54a]. Resolution [Squ89]. Results [Cle61d]. Retarded [McD18]. retina [Cle71c]. return [Nor05]. réunion [LdB12b, LdB12a]. Rev [Cle73b]. reveal [Col90]. Reversals [Wes99]. reversible [Ben03]. Review [Ano83, Ber08, Boo64, Bri18, Bri89, Bru71, Buc92, Buc93, Can90, Cha84, Chi64, Cle73c, Cle73b, Cle73d, Cle74b, Cle74a, Cle74c, Cle74d, Cle76a, Cle77a, Cle78a, Cle78b, Cle78c, Cle79a, Cle79b, Dea16, Dom97, Dom04, Eve84, Eve99, Goo85a, Goo85b, Har93, Har96, Har14, Hei72, Hun88a, Hun88b, Hun89, Hun92, Hun00a, Hun00b, Hun06, Jam15, Jon64, Jon92, Jor93, Jos07, Jos14, Kar14, Kra13, Lin71, Mac15, Mar99a, Mar93, McC69, Mor91a, Mor91b, Mye06, Ner87, Pag38, Rex17, Ric14, Roc87, Ryn16, Sch89, Seg92, She04a, She04b, Sho14, Sie85b, Sie85a, Sie93a, Sie93b, Sie96, Smi82a, Smi82b, Smi85, Smi00a, Smi00b, Spe65, Spr06, Sta14, Tai73, Tai77, The87, Tur83, Wil68, Wil97, Wil05, Yos99, Cle73r, Dom92, Fis88, Pip99]. review [Pip04, Row05]. Reviews [Cat16, Hen15, Hun05a, Mir99, Sto82, Ano91, Ano04, Jam91, FSvL^{+}37]. revisiting [Sel09]. revolution [Dar93c, Kuw31]. Revolutionized [Jam15, FM14]. Rex [She04a, She04b]. Reynolds [BE69]. Rezensionen [Kai87, vM82]. rhetorical [Hal04]. rhétoriques [Bre06a, Bre06b, Hal04, Mye06, Spr06]. Richard [Lit03, Lit10]. Rigid [Clexxb]. Rings [Cle62b, Fed08, Gar08, Tis17, BEG83, Chi31, Cle59b, Cle59c, Cle60c, Mur07, NV19, Sie85b]. Rise [Can90, Buc89, Row05, War03]. Road [All13, Rau14b, Rau14a]. Robert [Ber08, Hei72]. Role [D’A01a, Cha86, HB03, Sim70]. Rolling [Cle49]. Roman [vM82]. roofs [Cle68i]. Rotation [FM78, Cle57e, Cle71e]. route [Dar10]. Routledge [CF98]. Royal [QM15, Tho82]. rules [Cle74h, Cle75e]. Russian [Kle89, Pol85].

S [Lin71]. safety [Gré19]. Saturn [Sie85b, BEG83, Chi31, Cle59b, Cle59c, Cle60c, Cle62b, Fed08, Gar08, Mur07, Tis17]. Saunders [Wea16]. Savior [Anoxxc]. say [Dia94b]. Scalar [Rap05]. schen [Her84]. Scholar [Ano15, Sch96]. schrieb [Büh98a]. Schriften [Hei88]. Schrödinger [Hou16]. Schwarzen [Bes04]. Science [Ano99d, Anoxxc, AB86, ADS83, Bee96, CG10, Chi64, Cle70b, Cle70a, Cle76n, Dar93a, Dar93b, Dea16, Gib19, Hei88, Hes63, Hes66, Lon08, Nye02, Pyl96, Tur27, Ber84, Bre06a, Bre06b, CG82, CG84, Cle79i, Cle82b, Cle84a, Cle84b, Dom63, Gal74, GS06, Gal70, Gar85a, Hal04, Hes70, Hes86, Lam03, Lén29, Lén33, LH33, Lén43, Len70, Mye06, Pic08, Pup25, Rin08, Spr06, Sta12, Sta15, Tur72, Zit02, Wil68, Ano83, Dea16, Jon64, Boo64]. Sciences [Coh94, Nye02, Sim06, Toy99, Tur83, Wil62, vDP17, Cle85a]. scientiarum [Kai87]. Scientific [Ach87, Ano81b, ADS83, Buc92, Buc95, Cat01, Cat16, Cha21, Cle73s, Cle77e, Cle77j, CS07, CS71, CS10, Dar98a, Dar98b, Dom04, Eve67c, Gia70, GCC81, Har90, Har95a, Har02, Har14, Hun05a, Tru14, Jon92, Jos14, Kar14, Kno11, Kra13, Lar10, Niv90a, Niv90b, Niv90c, Niv90d, Niv27, Niv52, Niv65, Rod19, Sie93a, Sie93b, Sta14, Sta17, Tai80a, Ach13, Ano03, Cat12, Cle76a, Cle76d, Cle21a, Cle27b, Cle51, Dom92, Dom97, Eve83, Jam91, Lar07, Lar31, Lar71, Lén29, Lén33, LH33, Lén43, Len70, Ner84, Niv03, Niv11, Seg92, Sel09, Wil51, Wil72, Wil16]. Scientist [Ano15, Sch96]. Scientists [Cro35, Cro40, Jen70, H^{+}69, LB^{+}04]. Scope [Sho14]. Scotland [Gee81]. Scottish [Art16, Rei08, Rod02]. Screen [Cle69a]. Scully [Ber08]. Second [C^{+}70b, C^{+}75b, C^{+}79c, Duh15c, Den07, Lef02]. Second-Law [Den07]. Sections [Cle70b, Cle70a]. secular [C^{+}79b, C^{+}79c]. seen [Cle72b]. seiner [Pla31a]. Seite [Pla99]. Seiten [Kai87]. Selected [Lan03, Lar10]. Selection [CG10, Ano83, CG82, CG84, C^{+}74, C^{+}75b, Hei88]. selections [Cle76a]. Self [Gol94]. Self-dual [Gol94]. Senior [For99b]. sense [Mit17a]. sensibility [Cle57g]. sensible [Cle78h]. separate [Cla58, Cla59]. September [C^{+}65, Cle66c]. Sequel [Cle78a, Tho93, Tho68]. Series [Ano82, Dar93a, Dar93b]. Sesquicentennial [QM15, Sch96, Ber84]. Seuil [Spr06]. seules [Cle74d, Cle74k]. Seventh [C^{+}76b]. Sexl [vM82]. shall [Cle72g]. Shaped [Gib19]. sheet [Cle72i]. Sherman [Sto82, Tur83, Bla83a, Bla83b]. Shermann [Bla83a, Bla83b]. shift [Saa05]. Short [RSB^{+}92]. Short-Lived [RSB^{+}92]. side [Pla99]. Siegel [Hun92, Mar93, Att93a, Att93b, Buc93]. Siemens [Cle73q, C^{+}75a]. signals [Wil55]. Significado [Bun43a, Bun43b]. Significance [Tor00, Bun43a, Bun43b]. Simpson [Jos07, Wil97]. since [Pei63]. Sings [Clexxb]. Sir [Chi64, Cle73c, Cle73d, FSvL^{+}37, Lar10, Lar07, Lar71]. Sitzungsberichte [Sim06]. Six [Chi64, Duh15d, Dom63]. Sixth [Cle74b, Cle74j, Cle74c]. Sketch [CG10, Ano83, CG82, CG84]. Slamming [Mad02]. Small [Gar08]. Smith [Chi64]. soap [Cle74d]. soap-bubbles [Cle74d]. Social [Coh94, Por81, Por94]. Societies [Cle80d]. Society [Gib19, QM15, Sim14]. Software [Mir99]. Sohn [vM82]. solar [Cle80d]. Solid [D’A12, SPW02]. Solid-State [D’A12, SPW02]. Solids [Cle51a, Cle53]. solution [Cle73i, Cle74h, Cle75e, Cle76g, Cle76m, Col97, Col98, Hou16]. Solutions [Cle54b, Clexxa, CPR11, KS16]. Solvay [LdB12b, LdB12a]. Some [Ano03, Coh94, Fle31, H^{+}69, Sim70, Cla59, Cle57c, Cle57i, Fah99a, Fah99b, Fah01, Hug69, Tho91, Tur55b]. someone [Dia94b]. Sorting [Tho79, Tho82, Tho89]. soumis [Cle74d, Cle74k]. Sound [Cle79h, Cle79a]. Sources [Wil62]. sous [LdB12b, LdB12a]. Space [Čap76, Cle68g, Epp98, Ryn02, SGD^{+}02, Wea16, VT89, Cle69c]. Spanish [Bun43a, Bun43b, Cle96]. Spannungen [SB50]. Special [CS07, CS71, CS10, Dun08, Hal08, Cle79i, VT89]. Species [RSB^{+}92]. spectra [Cle72d, Dea72, Hug69]. Spectrum [Cle69a, Cle60e, Cle60f, Cle61b]. speculative [Dor70]. Speed [Den07, LL19]. Speed-Dependent [Den07]. Spheres [Cle60g, Tai97]. spherical [Cle67c]. Spin [Arm04, Wan08]. Spin- [Arm04]. Spinor [LU31]. Spirit [Gib19, Jos14]. Spline [Sch17]. spotlights [Heh10]. Springer [Sto82]. Springs [ADS83]. square [Fal17]. Stability [Cle59b, Cle59c, Ful86]. Stacks [Ano11a]. Standard [Ive05]. standards [C^{+}64, C^{+}65, Cle66c, C^{+}68, C^{+}70a, C^{+}71a, Sch94]. Standpoint [Gia70]. Stanley [Dea16]. stars [Hug69]. State [Bru61, Cle73a, Cle73m, D’A12, SPW02, Cle74a]. Static [Boh17]. statical [Cle62c]. statically [Cha71]. statically-indeterminate [Cha71]. Statics [Cle51d, Cle74k, Cle76h]. stationary [Vor08]. Statique [Cle74k, Cle74d]. Statistical [All13, Har96, Myr11, Por81, Por94, Sie96, Bru67, CGBE95, EP81, GBE95, Hei70b]. Statistics [Por94, Cat14, Dia94a, FMM96]. steam [Kle10]. Stephen [Har96, Sch89, Sie85b, vM82]. stereograms [Cle68a]. stereoscope [Cle68d]. Sternstunden [Büh98b]. stiffness [Cle64b]. still [Luc12]. Stimulated [Smi63]. Stokes [Lar07, Lar71, Lar10]. Stoney [O’H75]. stories [Buc95]. Story [Goo85a, Goo85b, Rau14b, Gol83, Gye17, Mat85a, Mat85b, Smi85]. Strategies [Cat16, Har14, Kar14, Kra13, Sta14, Ach13]. Strength [Clexxa, Mat87]. strengths [Cle74k]. Stress [Cle68g]. Stresses [Cle79g, Cle78g, Cle78f]. Structure [Čap76, CF91]. Structures [Bre06a, Bre06b, Spr06, Cha71, Hal04, Mye06]. Student [Fle08, Lem17]. Studie [Kai87]. Studies [Bru71, Har85, SBB^{+}08]. Study [AB86, Bri89, Gia70, Hun89, Sim68, SW89, Dia94b, Duh02, Duh67, Duh15a, Hei79, Hei99, Kim03, Mor92, OP87, Wil97]. subaqueous [Fah99a, Fah99b, Fah01]. Subject [Cle73a, Cle73m, Cle74k]. subjected [Cle73l]. Subjects [Gar08]. Substance [Har82b]. Substances [Cle76p, Clexxa, Cle76j, Cle76k, Cle77h, Cle08a, Lar08]. Substanz [Cle79j, Cle81b, Cle06]. Substanzen [Cle76p]. substrate [KM08]. suggestions [Cle57c, Cle57i]. Summary [Cle56g]. summation [Cle77g, Cle80c]. Sun [Har08, Hug69]. Supersymmetric [Rap05]. supplement [Wil55]. Supplementing [Kno11]. Suppressing [Ano99d]. Surface [Cle72b, Hop57a, Gol94]. surfaces [Cle56f, Cle66d, Cle68a, Cle69c]. Survey [Por81]. swap [CL15]. symmetrical [Cle75d]. symmetries [Wan08]. symposium [Ber84]. System [Cle57c, Cle57i, Cle73a, Cle73m, Cle73p, Cle79f, Cle73o, Cle76l, Cle80d, KM08, Wil55]. Systems [Rap05, Cle68k]. Szilard [Dev02, EN98, EN99].

t [D’A89, Lam03]. Tait [Cle73r, Cle79b, Cle73c, Cle78b, Cle78c, Dou16, Kno11, Moy77]. taken [Fra57]. Taking [Kra21]. Taming [Mus99]. Teaching [KCP14]. technology [Ber84, Zit02]. Teil [vM82]. telecommunications [Whe11]. telegraph [Whe11]. telegraphs [Fah99a, Fah99b, Fah01]. telegraphy [Fah99a, Fah99b, Fah01, Lod31, PV04]. telephone [Cle78j, Cle78i]. telescope [CL15]. telescopes [Bow63]. Temperature [Cle73l, Cle79g, C^{+}69, C^{+}70b, C^{+}71b, C^{+}72, C^{+}73, Cle74j, C^{+}76a, C^{+}76b, C^{+}77b, Cle78g, Cle78f, C^{+}78, C^{+}79e]. ten [Mac19, VT89, Nat32]. ten-dimensional [VT89]. tend [Cle82b]. tendency [Gye17]. Tension [Hop57a]. Tenth [C^{+}78]. tenue [LdB12b, LdB12a]. Teoría [Bun43a]. Teoriia [Kle88]. teploty [Kle88]. term [NV19]. Test [Sim66, Tho91, Tho92]. test-case [Tho92]. testing [C^{+}75a, C^{+}77c, VT89]. Text [Sie95, Chi64]. Thank [Smi17]. Their [Čap76, Cle61d, Cle68g, Gal74, Gal70]. Theistic [Dea16, Sta15]. them [Kim03, Nah83, Pic08]. Theme [Sto82]. theology [Cat12]. Theorem [Cle79f, Tai86, Dia94b, Ain11]. Theoretical [Boh31, Lon03, Cle57d, Cle74k, D’A01b, Mar66, Pla31a, Pla31b]. theoretische [Pla31a]. theoretischen [Mar66]. Theorie [Bol93, Pla98, Pla99, vM82, Poi99, Poi07, LdB12b, LdB12a, Cle77k, Cle78k, Pla03]. Theories [Cle62b, Cle73p, D’A00, Duh02, Duh67, Mul08, SR05, Whi10, Whi53, Whi89, Buc95, Duh15a, Ner84, Whi54, Whi87, Whi08, Duh02, Duh67]. théorique [Cle74d, Cle74k]. Theory [Ach87, Ber63, Bla83a, Bla83b, Bol93, BWB19, Bru57a, Bru61, Bru62, Bru71, Buc79b, Buc93, Can90, Cha73a, Cle49, Cle59a, Cle60a, Cle60b, Cle60g, Cle61d, Cle61f, Cle67a, Cle70g, Cle71g, Cle72j, Cle72k, Cle72l, Cle72m, Cle73e, Cle73j, Cle73p, Cle75i, Cle75j, Cle76o, Cle77a, Cle77l, Cle80e, Cle83b, Cle85b, Cle88c, Cle90b, Cle91b, Cle91c, Cle94, Cle97, Cle02, Cle04a, Cle08b, Cle21b, Cle70, Cle82, Cle01, Cle11, Cra54, D’A01a, Fed08, Gar66, Gar70, Gar08, Gia70, Hal08, Hun88a, Hun88b, Hun92, Hun15, Kra18, Mar93, Mor91a, Mor91b, Ner87, QM15, Ryn02, Sie75, Sie95, Sim66, Sim68, The87, Top70, Twe11, Wat61, Wis82, Att93a, Att93b, Bes04, Bro67, Bru57b, Bru58b, Bru58a, Buc85a]. theory [Buc89, Cat12, Cha71, Cha89, Cla59, Cle55b, Cle56e, Cle57e, Cle57f, Cle60d, Cle60f, Cle61e, Cle61h, Cle61g, Cle62c, Cle62d, Cle64a, Cle65, Cle66a, Cle67b, Cle67d, Cle68b, Cle68c, Cle68i, Cle68j, Cle68k, Cle69b, Cle75d, Cle77f, Cle80b, Cle07, Cle27a, Dar93c, Dor70, Dys99, FMM96, Hea12, Hen86, Lor23, Mit17b, Mor92, Moy78, Ozi94, Pei63, Pla98, Pla99, Pla03, Poi99, PV04, Poi07, Rau14b, Row05, Saa05, Sie85c, Sie91, Sim70, Tur83, War03, Wig90, Yan14, Cle78k, Kle88, Bun43a, Lin71, Sto82]. There [Har95b, Cle82a]. therefrom [Hug69]. Thermodyamics [Lef02]. Thermodynamic [She18]. Thermodynamics [Bub01, Clexxa, Gar66, Gar69, Smi02, Cle91a, CDKR16, Myr11, Cle78b, Cle78c]. thermodynamique [Cle91a]. thermométrie [Cle91a]. thermometry [Cle91a]. these [Büh98a, Hug69]. thesis [Cle74a]. thickness [Cle78h, KM08]. Third [C^{+}71b, C^{+}79d, Duh15e]. Thomas [Wil97, Jos07]. Thomson [Cle73c, Tho10, Cle73d, Cle79b, Cle19, Lar36, Lar37, Mit17a, Moy77, Moy78, Top70, Pag38]. those [CF51]. Thought [Jos07, Mou08]. thousand [Mar99b]. Three [Cle72g, Cle11, Fle86, KM08, OM11, Cle61e]. Three-Channel [OM11]. Three-dimensional [KM08]. Time [Čap76, Mar99a, SR05, vB98]. Times [Wea16, Cro01]. Todhunter [Cle76a]. together [Cla59]. Tolstoy [Eve84, Gee81, Sie85a, Smi82a, Smi82b]. Tom [Kle89]. too [Saa05]. Top [Cle57c, Cle57i]. Topological [Bar08, Epp98]. Topology [Epp98]. Total [SR05]. Traité [Cle84, Cle85c, Cle89]. Traktat [Kle89]. Transactions [Cle69c]. Transcript [Flo12]. Transformation [Cle72g, Cle56f, Cle66d, Cle76m]. Transformations [Ive05]. Transition [D’A00]. travaux [Bre06a, Bre06b, Spr06]. Traveling [Rod02]. Travelling [Rod19]. treasury [FF91]. ’Treatise [Tho68, Ano82, Ano11a, Cle71h, Cle73t, Cle77a, Cle81a, Cle81c, Cle88a, Cle92b, Cle04b, Cle53, Cle54, Cle55, Cle98a, Cle98b, Cle05, Cle13, Ryn16, Sim70, Cle83a, Cle84, Cle85c, Cle89, Kle89, LM81, Sim68, Tho93, Yag07, Yag14a, Yag14b]. tribute [Ano65]. Tricker [Wil68]. Trifocal [Cle46c]. Trinity [Cle76a]. Truth [KS95]. turns [Edw10]. Twelfth [C^{+}79e]. twentieth [D’A01b]. twistors [KNP84]. Two [Bes04, Col90, GZ11, Jam15, Kno11, SR05, Cle60a, Cle68k, Cle72h, FM14, JTR94, JTR95, PV04]. type [VT89]. Types [SR05].

U [Hun92, vM82]. Ueber [Cla58, Cle62a, Cle76p, Cle95, Cle12, Pla95]. Ukrainian [Kle88]. Uncertainty [Bes04]. Undead [She18]. undergraduate [Kim03]. underground [C^{+}69, C^{+}70b, C^{+}71b, C^{+}72, C^{+}73, Cle74j, C^{+}76a, C^{+}76b, C^{+}77b, C^{+}78, C^{+}79e]. underlying [Wal85]. understand [Dys99]. understandable [Hei88]. Understanding [Cat01].

Underwater [SBB^{+}08]. unequal [Cle57g]. Unification [Buc79b, Ryn02, Mor92, Nah92, Sta12]. uniform [Cle69f, Cle72i]. unit [Cle70c, Hop57b, Hop58, Hop59a, Mit17b]. Units [CG17, C^{+}74, C^{+}75b, Hop59a]. universe [Sen21, Var82, Cle78a]. University [Buc92, Can90, Chi64, Dom04, Eve84, Har93, Har96, Hen15, Hun05a, Jam91, Mac15, Mar93, Pip99, Roc87, Row05, Seg92, Yos99, And07, Cle59c]. Unlocking [Cen89]. Unmasking [NA02]. Unpublished [Cle5x, Cle0 ]. unreasonable [Bok15]. Unresolved [SGD^{+}02]. Unschärfe [Bes04]. Unseen [Cle78a]. unsung [Rin08]. upon [Cla59, C^{+}79b, C^{+}79c, Ive05]. Ursprünge [Ste05]. use [Cle67d]. uses [Lam11]. using [Col97, Col98, Wan08, Zit02].

V [Chi64, Kle88, Bru61, Cle60g, Ful86]. value [Cle70c]. Variable [Clexxa]. variations [Cle73i, Cle76g, Sto82]. various [C^{+}69, C^{+}70b, C^{+}71b, C^{+}72, C^{+}73, Cle74j, C^{+}76a, C^{+}76b, C^{+}77b, C^{+}78, C^{+}79e, Rey21]. VCH [Ber08]. Vector [Bor67]. vekov [Pol85]. Velocity [D’A81, D’A86, Tai97, Cle68k, Pla94, Pla95]. Vena [Cle69g]. Verlag [Kai87]. Verlagsanstalt [vM82]. vers [Dar98c]. version [HG12]. Versions [GZ11]. versus [Gré19]. VI [Ano03, Bru62]. Victoria [McN04]. Victorian [Jen70, Lam11]. View [Gia70, Ber74a, Ber74b, Lon03, Myr11]. Vieweg [vM82]. viii [Gee81, Cle65]. Virtues [Sta17, vDP17]. Viscosity [Cle66e, Cle66b, Bru62]. viscous [Cle74f, Cle74e]. Visible [Boh17, Cat14]. Vision [Bla83a, Bla83b, Cle71b, Sto82, Cle71c, Cle72e, Tur83, Jos14]. Visionary [Sch96]. Visit [OR97b]. vistas [Ber84]. vloeistofiocstand [Cle74a]. Vol [Dom97, Dom04, Dar98a, Dar98b, Seg92]. Vol. [Har90, Har95a, Har02]. vols. [Cle76a]. Voltages [SB50]. Volume [Buc92, Cle21a, Hun05a, Jam91, Jon92, Sie93a, Sie93b, Sta09, Cle69c, Kle89, Tho31b, Tho11, Hun05a, Nye02]. Volumes [Kno11]. Vorlesungen [vM82, Bol93]. Vortices [Att93a, Att93b, Buc93, Hun92, Mar93, Cle61h, Cle61g, Cle62c, Cle62d, Sie91]. Vorträge [Sim06].

W [Bri89, Cha89, Cle73c, Cle73d, Har96, Sch89, Sie85b]. W. [CG68, Lar36]. Waals [Cle74a]. Walter [vM82]. War [Büh98a, Mul04]. Wärme [Cle77k, Cle78k]. Wärmetheorie [Cla58]. Warmth [Mar99a, vB98]. Warwick [Row05]. Was [LS14, Fal17, Büh98a]. water [C^{+}69, C^{+}70b, C^{+}71b, C^{+}72, C^{+}73, Cle74j, C^{+}76a, C^{+}76b, C^{+}77b, C^{+}78, C^{+}79e, Wil72, Wil16]. Waterston [Bru57b]. Watson [Cle77a]. Watt [JTR94, JTR95]. Wave [Bri89, Can90, Hun89, Men08, Squ89, Tai95, Bat15, Bat55, Buc89, Hou16, OP87]. wave-motion [Bat15, Bat55]. Wave/Particle [Squ89]. waves [Bin08, D’A89, Poi99, Poi07, SE08]. Weber [D’A81]. Wege [Cla58]. Weighting [Den07]. Weinheim [Ber08]. welche [Cla58]. werden [Cla58]. Werner [Bes04, Bes04]. Wet [Wes99]. Weyl [Rap05, BDP15]. Whatever [Hoo87]. whether [Hug69]. Whewell [Cle76a, Cle76a, Cle90e]. which [Cle59c, Cle73i, Cle76g]. Whitaker [Cat16, Hen15, Mac15, Ric14, Sho14]. Who [Hun06, Wil05, Büh98a, Kle10, Mah03, Pip04]. whose [GS06]. Wiesbaden [vM82]. Wiley [Ber08, Hun06, Pip04]. Wiley-VCH [Ber08]. Will [She99, Cle82b, Dia94b]. William [Ano83, Cle77a, Pag38, Cle76a, Lar37, Mit17a, Tho10]. Willibald [Hun89]. wire [Fah99a, Fah99b, Fah01]. Wireless [SMO^{+}06, Fah99a, Fah99b, Fah01, Lod31, PV04]. wires [C^{+}79b, C^{+}79c]. Wissenschaft [Hei88]. Wissenschaften [Sim06]. Wissenschaftsgeschichte [Kai87]. wissenschaftshistorische [Kai87]. without [Cle57d, Cle67d]. women [Cle84a, Cle84b]. woodenheaded [Nah83]. Work [Kno11, LF08, Ran62, Tai80a, Cle67d, FMW14, Hop57b, Hop58, Hop59b, Jon32, Ran63, Hen15, Cat16, Sho14, Ric14]. Working [Wes99]. world [Ari03, Ber74a, Ber74b, CDKR16, FF91, Rin08]. Wormholes [KS16]. Worthies [Cle73s, Cle77e, Cle77j, Ano81b]. Wranglers [For99b, Har85]. Wrath [EN98, EN99]. Writings [CG10, Cle90e, Ano83, CG82, CG84, Cle76a, Hei88]. written [Cle79c, Cle67]. Wrong [LS14]. wrote [Büh98a].

X [Hun06, Kai87, The87, Chi64, Hen15, Mar93, Cla59, Jen70, Pip04, RSB^{+}92]. X-Ray [RSB^{+}92]. xiii [Har93, Tur83, Cle66e]. xiv [Hun89, Pip99, Row05, Yos99, Cle62d, EN98, EN99]. xix [The87, Pol85]. XLII [CG68]. XLIV [Cle61g]. XLV [Cle64d]. xv [Roc87, Cle68j]. XVIII [Ano81b, Cle57h]. xx [Hun06, Pip04, Pol85]. XXII [Cle68k]. xxiv [Can90]. XXV [Cle61h]. xxvii [Hun05a, Jam91]. xxviii [Buc92, Hun05a, Seg92]. xxx [Hun05a]. XXXIV [Cle57i].

Year [Ano11a, Cle59c, Ano11b]. Years [Ano11a, RWT08, Mor95, Eng15, Heh10, Nat32]. yellow [Cle57f]. York [Hun92, Lin71, Mar99a, Sto82]. Young [Bla83a, Bla83b, Sto82, Tur83].

Z [Can90, Ner87, Roc87]. Zeichen [Büh98a]. Zombies [She18]. Zum [SB50]. zur [Kai87, Kai87]. zurückgelegt [Cla58]. Zwei [Bes04]. zwischen [Her84].

References

Ariew:1986:DMC

[AB86] Roger Ariew and Peter Barker. Duhem on Maxwell: A case-study in the interrelations of history of science and philosophy of science. PSA: Proceedings of the Biennial Meeting of the Philosophy of Science Association, 1986:145–156, 1986. ISSN 0270-8647 (print), 2327-9486

        (electronic). URL http://www.jstor.org/stable/193116.

Achinstein:1987:SDM

[Ach87] Peter Achinstein. Scientific discovery and Maxwell’s kinetic theory. Philosophy of Science, 54(3):409–434, September 1987. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http:/

/www.jstor.org/stable/187582.

Achinstein:2013:EMS

[Ach13] Peter Achinstein. Evidence and method: scientific strategies of Isaac Newton and James Clerk Maxwell. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-992185-7 (hardcover), 0-19-933289-4 (e-book). xv + 177 pp. LCCN Q174.8 .A238 2013.

Aris:1983:SSC

[ADS83] Rutherford Aris, H. Ted (Howard Ted) Davis, and Roger H. Stuewer, editors. Springs of Scientific Creativity: Essays on Founders of Modern Science. University of Minnesota Press, Minneapolis, MN, USA, 1983. ISBN 0-8166-1087-8. viii + 342 pp. LCCN QC7 .S77 1983.

Ainsworth:2011:WCD

[Ain11] Peter M. Ainsworth. What chains does Liouville’s Theorem put on Maxwell’s demon? Philosophy of Science, 78(1):149–164, January 2011. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/10.1086/658113.

Appleyard:1930:CMM

[AL30] Rollo Appleyard and Philipp Lénard. Clerk Maxwell and the Michaelson experiment. Nature, 125(3154):566–567, April 12, 1930. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL https://www.nature.com/nature/journal/v125/
n3154/pdf/125566a0.pdf.

Allori:2013:RMD

[All13] Valia Allori. The road to Maxwell’s Demon: Conceptual foundations of statistical mechanics. International Studies in the Philosophy of Science, 27(4):453–456, 2013. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/02698595.2013.868179.

Anderson:1907:RCQ

[And07] P. J. (Peter John) Anderson. Record of the celebration of the quatercentenary of the University of Aberdeen. University of Aberdeen, Aberdeen, Scotland, 1907. xi + 656 + 2 pp. LCCN LF978 1906.

Appendix G contains an address by Royal Society President Lord Rayleigh with the statement: “within living memory the University has numbered among its professors the world-renowned path-finder James Clerk Maxwell”.

Anonymous:1879:DJC

[Ano79a] Anonymous. Death of James Clerk Maxwell, at Cambridge. The Scotsman, ??(??):4c–d–??, November 7, 1879.

Anonymous:1879:JCM

[Ano79b] Anonymous. James Clerk Maxwell. Proceedings of the American Academy of Arts and Sciences, 15:391–394, 1879. CODEN PAAAAV.

ISSN 0065-6836. URL http://www.jstor.org/stable/25138589.

Anonymous:1879:OJC

[Ano79c] Anonymous. Obituary: James Clerk Maxwell (1831–1879). The Times [London, UK], ??(??):6c–??, November 7, 1879. ISSN 0140-0460, 0956-1382.

Anonymous:1881:OJC

[Ano81a] Anonymous. [obituary:] James Clerk Maxwell, F.R.S. Proceedings of the Royal Society of London, 33(??):i–xvi, January 1, 1881. CODEN PRSLAZ. ISSN 0370-1662 (print), 2053-9126 (electronic).

Anonymous:1881:SWX

[Ano81b] Anonymous. Scientific worthies. XVIII. — James Clerk Maxwell. Nature, 24(626):601, October 27, 1881. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.

com/nature/journal/v24/n626/pdf/024601a0.pdf.

Anonymous:1882:BRL

[Ano82] Anonymous. Books received: … An Elementary Treatise on Electricity, by James Clerk Maxwell, Clarendon Press Series, Oxford, 1881. Science, os-3(80):14, January 14, 1882. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://science.

sciencemag.org/content/os-3/80/14.5. See [Cle81a].

Anonymous:1883:BRJ

[Ano83] Anonymous. Book review: James Clerk Maxwell: The Life of James Clerk Maxwell: with a selection from his correspondence and occasional writings, and a sketch of his contributions to science. By Lewis Campbell and William Garnett. London. Macmillan & Co., 1882. 16 + 662 p., 3 prort., 4 pl., facsim., etc. 8s. Science, ns-1(13):360–365, May 4, 1883. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.jstor.org/stable/10.2307/1759814.

Anonymous:1885:JCM

[Ano85] Anonymous. James Clerk Maxwell. Science, ns-5(115):317–318, April 17, 1885. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-

9203 (electronic). URL http://www.jstor.org/stable/10.2307/

1760607. Book review of [CG84].

Anonymous:1903:SSC

[Ano03] Anonymous. Some scientific centres. VI. — The Cavendish Laboratory. Nature, 69(1780):128–132, December 10, 1903. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://

www.nature.com/nature/journal/v69/n1780/pdf/069128a0.pdf.

Anonymous:1931:CMC

[Ano31a] Anonymous. The Clerk Maxwell Centenary celebrations. Nature, 128 (3232):604–607, October 10, 1931. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.

com/nature/journal/v128/n3232/pdf/128604a0.pdf.

Anonymous:1931:HJC

[Ano31b] Anonymous. Handbook for the James Clerk Maxwell centenary catalogue. ????, ????, 1931. ???? pp.

Anonymous:1965:MGT

[Ano65] Anonymous. Maxwell’s greatest tribute to Faraday. IEEE Spectrum, 2(10):5–8, October 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:EJC

[Ano74] Anonymous. Errata: James Clerk Maxwell at Aberdeen, 1856–1860. Notes and Records of the Royal Society of London, 28(2):344, April 1974. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531339. See [Jon73].

Anonymous:1991:ERM

[Ano91] Anonymous. Essay reviews: Maxwell’s early career. Notes and Records of the Royal Society of London, 45(2):266–270, July 1991. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:1999:MMa

[Ano99a] Anonymous. 1000 Makers of the millennium. Dorling Kindersley, London, UK, 1999. ISBN 0-7894-4709-6, 0-7513-5664-6. 256 pp. LCCN CT107 .A16 1999a.

Anonymous:1999:MMb

[Ano99b] Anonymous. 1000 makers of the millennium. DK Pub., New York, NY, USA, 1999. ISBN 0-7894-4709-6, 0-7513-5664-6. 256 pp. LCCN CT107 .A16 1999.

Anonymous:1999:JCM

[Ano99c] Anonymous. James Clerk Maxwell commemorative booklet. James Clerk Maxwell Foundation, Edinburgh, Scotland, 1999. ???? pp. Produced on the occasion of the Fourth International Congress on Industrial and Applied Mathematics coming to Edinburgh in July 1999.

Anonymous:1999:SCM

[Ano99d] Anonymous. Science and the citizen: Microrotors and Maxwell’s demon. Suppressing anti-nuke protesters. Ants against elephants. Scientific American, 280(2):24–??, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:2004:BR

[Ano04] Anonymous. Book reviews. Notes and Records of the Royal Society of London, 58(1):101–106, January 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:2006:DGH

[Ano06] Anonymous. ’Dafty’ genius honoured at last by his alma mate. The Scotsman Web site., November 8, 2006. URL http://

www.scotsman.com/news/education/dafty-genius-honoured-at-
last-by-his-alma-mater-1-724315.

Anonymous:2008:PMN

[Ano08] Anonymous. PIA09857: Maxwell’s namesake. NASA JPL Web

site, March 13, 2008. URL https://photojournal.jpl.nasa.gov/
catalog/PIA09857.

Anonymous:2011:YSY

[Ano11a] Anonymous. 150 years in the stacks: Year 13 — 1873: A Treatise on Electricity and Magnetism by James Clerk Maxwell. MIT Library Web site., January 19, 2011. URL http://libraries.mit.

edu/150books/2011/01/19/1873/.

Anonymous:2011:JCM

[Ano11b] Anonymous. 1861: James Clerk Maxwell’s greatest year. Web site.,

April 18, 2011. URL https://www.kcl.ac.uk/newsevents/news/
newsrecords/2011/04Apr/JamesClerkMaxwell.aspx.

Anonymous:2011:EF

[Ano11c] Anonymous. The electromagnetic field. Web site., 2011. URL http:

//www-history.mcs.st-and.ac.uk/Projects/Johnson/Chapters/
Ch4_4.html;   https://web.archive.org/web/20110827131533/

.

Anonymous:2012:JCM

[Ano12] Anonymous. James Clerk Maxwell biography. Web video., May 29,

2012. URL https://www.youtube.com/watch?v=LjY1x5CDvD4.

Anonymous:2013:JCM

[Ano13] Anonymous. James Clerk Maxwell (1831–1879). National Library of Scotland Scottish Science Hall of Fame Web site., 2013. URL

http://digital.nls.uk/scientists/biographies/james-clerk-
maxwell/index.html.

Anonymous:2015:JCM

[Ano15] Anonymous. James Clerk Maxwell: Scholar, physicist, scientist (1831–1879). Web site., 2015. URL https://www.biography.com/

people/james-c-maxwell-9403463.

Anonymous:2016:JCM

[Ano16] Anonymous. James Clerk Maxwell Building (JCMB). University of Edinburgh Web site., April 13, 2016. URL http://www.ph.ed.ac.

uk/about/locations/jcmb.

Anonymous:2017:FKP

[Ano17a] Anonymous. Famous King’s people: James Clerk Maxwell. King’s

College Web sit, 2017. URL https://www.kcl.ac.uk/aboutkings/
history/famouspeople/jamesclerkmaxwell.aspx.

Anonymous:2017:FCJ

[Ano17b] Anonymous. Free course: James Clerk Maxwell. Open University

OpenLearn Web site, 2017. URL http://www.open.edu/openlearn/
science-maths-technology/science/physics-and-astronomy/james-
clerk-maxwell/content-section-0.

Anonymous:2017:JCM

[Ano17c] Anonymous. James Clerk Maxwell. BBC program Web site.,

2017. URL http://www.bbc.co.uk/history/people/james_clerk_
maxwell.

Anonymous:20xx:ABJ

[Anoxxa] Anonymous. Archives biographies: James Clerk Maxwell. The Institution of Engineering and Technology Web site., 20xx. URL http:

//www.theiet.org/resources/library/archives/biographies/
maxwell.cfm.

Anonymous:20xx:JCMa

[Anoxxb] Anonymous. James Clerk Maxwell. Web site., 20xx. URL https://

www.famousscientists.org/james-clerk-maxwell/.

Anonymous:20xx:JCMc

[Anoxxc] Anonymous. James Clerk Maxwell — his science and his savior. Audio recording of 4m25s speech about JCM., 20xx. URL https://

archive.org/details/JamesClerkMaxwell-HisScienceAndHisSavior.

Anonymous:20xx:JCMb

[Anoxxd] Anonymous. James Clerk Maxwell: Biography. Web site, 20xx. URL

http://ethw.org/James_Clerk_Maxwell.

Appleyard:1930:PEC

[App30] Rollo Appleyard. Pioneers of Electrical Communication. MacMillan Publishing Company, New York, NY, USA, 1930. ix + 347 pp. LCCN QC514 .A6.

Appleton:1963:RI

[App63] E. V. Appleton. Radio and the ionosphere. In Clerk Maxwell and modern science; six commemorative lectures [Dom63], page ?? LCCN QC71 .D63.

Appleyard:1968:PEC

[App68] Rollo Appleyard. Pioneers of Electrical Communication. Essay index reprint series. Books for Libraries Press, Freeport, NY, USA, 1968. ix + 347 pp. LCCN QC514 .A6 1968.

Arianrhod:2003:EHI

[Ari03] Robyn Arianrhod. Einstein’s heroes: imagining the world through the language of mathematics. University of Queensland Press, St Lucia, Queensland, Australia, 2003. ISBN 0-7022-3408-7. viii + 323 pp. LCCN QC19.6 .A75 2003. US$24.00.

Armour:2004:SMF

[Arm04] Rollin S. Armour, Jr. Spin-1/2 Maxwell fields. Foundations of Physics, 34(5):815–842, May 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/B%3AFOOP.0000022188.90097.10.

Arthur:2013:EME

[Art13] J. W. Arthur. The evolution of Maxwell’s equations from 1862 to the present day. IEEE Antennas Propagat. Mag., 55(??):61–81, June 2013.

Arthur:2016:BLC

[Art16] John W. Arthur. Brilliant lives: the Clerk Maxwells and the Scottish Enlightenment. John Donald, an imprint of Birlinn Ltd, Edinburgh, Scotland, 2016. ISBN 1-906566-97-6 (paperback). xv + 358 + 20 pp. LCCN B1402.E55 A78 2016.

Atiyah:2015:IBC

[Ati15] Sir Michael Atiyah. Inspiring brilliance: Clerk Maxwell’s influence on mathematics. Web video, November 11, 2015. URL https://ieeetv.

ieee.org/ieeetv-specials/inspiring-brilliance-clerk-maxwells-
influence-on-mathematics?rf=events72—.

Atten:1993:AOB

[Att93a] Michel Atten. Analyses d’ouvrages: Innovation in Maxwell’s electromagnetic theory. Molecular Vortices, Displacement Current, and Light par Daniel M. Siegel. Revue d’Histoire des Sciences, 46(4):545– 547, October 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633313.

Atten:1993:AOI

[Att93b] Michel Atten. Analyses d’ouvrages: Innovation in Maxwell’s electromagnetic theory. Molecular Vortices, Displacement Current, and Light par Daniel M. Siegel. Revue d’Histoire des Sciences, 46(4):545– 547, October 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633313.

Ahmetaj:2007:PFD

[AVJ07] Skender Ahmetaj, Ahmet Veseli, and Gani Jashari. Physical fields described by Maxwell’s equations. AIP Conference Proceedings, 899 (??):706–??, 2007. CODEN APCPCS. ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465.

Barrett:2008:TFE

[Bar08] T. W. (Terence William) Barrett. Topological Foundations of Electromagnetism, volume 26 of World Scientific series in contemporary chemical physics. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2008. ISBN 981-277-996-5. ix + 185 pp. LCCN QC760 .B282 2008. URL http://www.

zentralblatt-math.org/zmath/en/search/?an=1171.78002.

Bateman:1915:MAE

[Bat15] Harry Bateman. The mathematical analysis of electrical and optical wave-motion on the basis of Maxwell’s equations. Cambridge University Press, Cambridge, UK, 1915. vi + 1 + 159 + 1 pp. LCCN QC661 .B28.

Bateman:1955:MAE

[Bat55] Harry Bateman. The mathematical analysis of electrical and optical wave-motion on the basis of Maxwell’s equations. Dover Publications, Inc., New York, NY, USA, 1955. 159 pp. LCCN QC661 .B28 1955.

Bisio:2015:WDM

[BDP15] Alessandro Bisio, Giacomo Mauro D’Ariano, and Paolo Perinotti. Weyl, Dirac and Maxwell quantum cellular automata. Foundations of Physics, 45(10):1203–1221, October 2015. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-015-9927-0.

Brush:1969:MOR

[BE69] S. G. Brush and C. W. F. Everitt. Maxwell, Osborne Reynolds, and the radiometer. Historical Studies in the Physical Sciences, 1(??): 105–125, ???? 1969. CODEN HSPSAS. ISSN 0073-2672. URL http:

//www.jstor.org/stable/27757296. See [Cle79g].

Bearden:2013:EP

[Bea13] Tom Bearden. Extraordinary physics. Web site., 2013.

Beer:1996:OFS

[Bee96] Gillian Beer. Open Fields: Science in Cultural Encounter. Clarendon Press, New York, NY, USA, 1996. ISBN 0-19-818369-0 (hardcover), 0-19-818635-5 (paperback). viii + 341 pp. LCCN PR468.S34 B44

1996. URL http://www.loc.gov/catdir/enhancements/fy0639/
95043234-d.html.

Brush:1983:MSR

[BEG83] Stephen G. Brush, C. W. F. (C. W. Francis) Everitt, and Elizabeth Garber, editors. Maxwell on Saturn’s rings. MIT Press, Cambridge, MA, USA, 1983. ISBN 0-262-13190-0. xii + 199 pp. LCCN QB405 .M38 1983.

Bennett:2003:NLP

[Ben03] Charles H. Bennett. Notes on Landauer’s principle, reversible computation, and Maxwell’s Demon. Studies in History and Philosophy of Modern Physics, 34(3):501–510, September 2003. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S135521980300039X.

Bernstein:1963:JCM

[Ber63] Henry T. Bernstein. J. Clerk Maxwell on the history of the kinetic theory of gases, 1871. Isis, 54(2):206–216, June 1963. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228538.

Berkson:1974:FFDa

[Ber74a] William Berkson. Fields of force; the development of a world view from Faraday to Einstein. Routledge and K. Paul, London, UK, 1974. ISBN 0-7100-7626-6. xiii + 370 pp. LCCN QC173.7 .B47 1974b.

Berkson:1974:FFDb

[Ber74b] William Berkson. Fields of force; the development of a world view from Faraday to Einstein. Wiley, New York, NY, USA, 1974. ISBN 0-470-07029-3. xiii + 370 pp. LCCN QC173.7 .B47 1974.

Berger:1984:JCM

[Ber84] Melvyn S. (Melvyn Stuart) Berger, editor. J. C. Maxwell, the sesquicentennial symposium: new vistas in mathematics, science, and technology. North-Holland Publishing Co., Amsterdam, The Netherlands, 1984. ISBN 0-444-86707-4 (Elsevier). LCCN QC669 .J17 1984.

Beretta:2008:DQP

[Ber08] Gian Paolo Beretta. Book review: The Demon and the Quantum: From the Pythagorean Mystics to Maxwell’s Demon and Quantum Mystery, Robert J. Scully, with endnotes by Marian O. Scully, Wiley-VCH, Weinheim, Germany, 2007. $45.00 (271 pp.). ISBN 978-3-527-40688-3. Physics Today, 61(11):57, November 2008. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Beste:2004:GPQ

[Bes04] Dieter Beste, editor. Grosse Physiker: Quantenrätsel — Werner Heisenberg und die Unschärfe; Relativitätstheorie — Albert Einstein und die Schwarzen Löcher; Zwei Klassiker —- Isaac Newton und James Clerk Maxwell. (German) [Great physicists: Quantum mystery — Werner Heisenberg and Uncertainty; Relativity theory — Albert Einstein and black holes; Two classical authors — Isaac Newton and James Clerk Maxwell]. Spektrum der Wissenschaft Dossier. Spektrum-der-Wiss.-Verl.-Ges., Heidelberg, Germany, 2004. ISBN 3-936278-75-X. 82 pp. LCCN ????

Bingham:2008:PAE

[Bin08] R. Bingham. Particle acceleration by electromagnetic waves. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 366(1871):1749–1756, May 2008. CO-DEN PTRMAD, PTMSFB. ISSN 1364-503X (print), 1471-2962 (electronic).

Blay:1983:AOB

[Bla83a] Michel Blay. Analyses d’ouvrages: Colour Vision in the Nineteenth Century. The Young–Helmholtz–Maxwell Theory, Bristol par Paul D. Shermann; Paul D. Sherman. Revue d’Histoire des Sciences, 36(2):199–201, avril 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632388.

Blay:1983:AOC

[Bla83b] Michel Blay. Analyses d’ouvrages: Colour Vision in the Nineteenth Century. The Young–Helmholtz–Maxwell Theory, Bristol par Paul D. Shermann; Paul D. Sherman. Revue d’Histoire des Sciences, 36(2):199–201, avril 1983. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23632388.

Bohr:1931:MMT

[Boh31] Niels Bohr. Maxwell and modern theoretical physics. Nature, 128 (3234):691–692, October 24, 1931. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v128/n3234/pdf/128691a0.pdf.

Bohren:2017:MGF

[Boh17] Craig F. Bohren. Maxwell’s good fortune: the visible refractive index and static relative permittivity of paraffin. American Journal of Physics, 85(3):165, March 2017. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Bokulich:2015:MHU

[Bok15] Alisa Bokulich. Maxwell, Helmholtz, and the unreasonable effectiveness of the method of physical analogy. Studies in History and Philosophy of Science Part A, 50(??):28–37, April 2015. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0039368114000909.

Boltzmann:1891:VMT

[Bol93] Ludwig Boltzmann. Vorlesungen über Maxwells Theorie der Elektricität und des Lichtes. (German) [Lectures on Maxwell ’s Theory of Electricity and light]. ????, Leipzig, Germany, 1891–1893. ???? pp.

Booker:1964:BRC

[Boo64] H. G. Booker. Book review: Clerk Maxwell and modern science, C. Domb, ed. IEEE Spectrum, 1(6):204, June 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bork:1967:MVP

[Bor67] Alfred M. Bork. Maxwell and the vector potential. Isis, 58(2):210–222, Summer 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228226.

Bowen:1963:RAG

[Bow63] E. G. Bowen. Radio astronomy and giant telescopes. In Clerk Maxwell and modern science; six commemorative lectures [Dom63], page ?? LCCN QC71 .D63.

Bolton:1978:DBJ

[BP78] H. C. Bolton and W. C. Price, F.R.S. The date of birth of James Clerk Maxwell. Notes and Records of the Royal Society of London, 32(2):213–214, March 1978. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531729.

Brenner:2006:AOB

[Bre06a] Anastasios Brenner. Analyses d’ouvrages: Les Structures rhétoriques de la science : De Kepler à Maxwell, ^{≪}Des travaux^{≫}par Fernand Hallyn. Revue d’Histoire des Sciences, 59(2):355–357, juillet–décembre 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634331.

Brenner:2006:AOS

[Bre06b] Anastasios Brenner. Analyses d’ouvrages: Les Structures rhétoriques de la science : De Kepler à Maxwell, ^{≪}Des travaux^{≫}par Fernand Hallyn. Revue d’Histoire des Sciences, 59(2):355–357, juillet–décembre 2006. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23634331.

Brittain:1989:BRH

[Bri89] James E. Brittain. Book review: Hertz and the Maxwellians: A Study and Documentation of the Disvoery of Electromagnetic Wave Radiation, 1873–1894 by J. G. O’Hara; W. Pricha. Isis, 80(1):187–188, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234427.

Bridges:2018:BRM

[Bri18] Thomas J. Bridges. Book review: The Many Faces of Maxwell, Dirac and Einstein Equations: A Clifford Bundle Approach. Contemporary Physics, 59(3):323, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bromberg:1967:MDC

[Bro67] Joan Bromberg. Maxwell’s displacement current and his theory of light. Archive for History of Exact Sciences, 4(3):218–234, January 1967. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=4&issue=3&spage=218.

Bromberg:1968:ME

[Bro68] Joan Bromberg. Maxwell’s electrostatics. American Journal of Physics, 36(2):142–151, February 1968. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Brown:2000:MMM

[Bro00a] J. R. (Julian Russell) Brown. Minds, machines, and the multiverse: the quest for the quantum computer. Simon and Schuster, New York, NY, USA, 2000. ISBN 0-684-81481-1, 0-7432-4263-7. 396 pp. LCCN QA76.889 .B76 2000. URL http://catdir.loc.gov/

catdir/bios/simon054/99056638.html;   http://catdir.loc.
gov/catdir/description/simon032/99056638.html;   http://
catdir.loc.gov/catdir/enhancements/fy0705/99056638-s.html;
 http://catdir.loc.gov/catdir/enhancements/fy0707/99056638-
t.html.

Brown:2000:PMD

[Bro00b] J. R. (Julian Russell) Brown. The puzzle of Maxwell’s Demon. In Minds, machines, and the multiverse: the quest for the quantum computer [Bro00a], page ?? ISBN 0-684-81481-1, 0-7432-4263-7. LCCN QA76.889 .B76 2000. URL http://catdir.loc.gov/

catdir/bios/simon054/99056638.html;   http://catdir.loc.
gov/catdir/description/simon032/99056638.html;   http://
catdir.loc.gov/catdir/enhancements/fy0705/99056638-s.html;
 http://catdir.loc.gov/catdir/enhancements/fy0707/99056638-
t.html.

Brush:1957:DKTa

[Bru57a] S. G. Brush, A.B., D.Phil. The development of the kinetic theory of gases. I. Herapath. Annals of Science, 13(3):188–198, 1957. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brush:1957:DKTb

[Bru57b] S. G. Brush, A.B.D.Phil. The development of the kinetic theory of gases II. Waterston. Annals of Science, 13(4):273–282, 1957. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brush:1958:DKTb

[Bru58a] S. G. Brush, A.B., D.Phil. The development of the kinetic theory of gases IV. Maxwell. Annals of Science, 14(4):243–255, 1958. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brush:1958:DKTa

[Bru58b] S. G. Brush, A.B.D.Phil. The development of the kinetic theory of gases III. Clausius. Annals of Science, 14(3):185–196, 1958. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brush:1961:DKT

[Bru61] Stephen G. Brush. Development of the kinetic theory of gases. V. The equation of state. American Journal of Physics, 29(9):593–605, September 1961. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Brush:1962:DKT

[Bru62] Stephen G. Brush. Development of the kinetic theory of gases. VI. Viscosity. American Journal of Physics, 30(4):269–281, April 1962. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). Comments on [Cle60g].

Brush:1967:FSM

[Bru67] Stephen G. Brush. Foundations of statistical mechanics 1845– 1915. Archive for History of Exact Sciences, 4(3):145–183, January 1967. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=4&issue=3&spage=145.

Brush:1971:JCM

[Bru71] Stephen G. Brush. James Clerk Maxwell and the kinetic theory of gases: A review based on recent historical studies. American Journal of Physics, 39(6):631–640, June 1971. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://aapt.scitation.

org/doi/abs/10.1119/1.1986248.

Bub:2001:MDT

[Bub01] Jeffrey Bub. Maxwell’s Demon and the thermodynamics of computation. Studies in History and Philosophy of Modern Physics, 32(4):569– 579, December 2001. CODEN ???? ISSN 1355-2198 (print), 1879-

2502 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219801000235.

Buchwald:1979:HEMa

[Buc79a] J. Z. Buchwald. The Hall effect and Maxwellian electrodynamics in the 1880’s. Part I: the discovery of a new electric field. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 23(1):51–99, September 1979. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Buchwald:1979:HEMb

[Buc79b] J. Z. Buchwald. The Hall effect and Maxwellian electrodynamics in the 1880’s. Part II: the unification of theory, 1881–1893. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 23(2):118–162, December 1979. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Buchwald:1985:MMA

[Buc85a] Jed Z. Buchwald. From Maxwell to microphysics: aspects of electromagnetic theory in the last quarter of the Nineteenth Century. University of Chicago Press, Chicago, IL, USA, 1985. ISBN 0-226-07882-5 (hardcover), 0-226-07883-3 (paperback). xiii + 339 pp. LCCN QC670 .B83 1985. US$55.00.

Buchwald:1985:OHM

[Buc85b] Jed Z. Buchwald. Oliver Heaviside, Maxwell’s apostle and Maxwellian apostate. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 28(3):288–330, October 1985. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Buchwald:1989:RWT

[Buc89] Jed Z. Buchwald. The rise of the wave theory of light: optical theory and experiment in the early Nineteenth Century. University of Chicago Press, Chicago, IL, USA, 1989. ISBN 0-226-07886-8 (paperback), 0-226-07884-1 (hardcover). xxiv + 474 pp.

LCCN QC403 .B83 1989. URL http://www.loc.gov/catdir/
description/uchi052/88018647.html;   http://www.loc.gov/
catdir/enhancements/fy0608/88018647-t.html;   http://www.
loc.gov/catdir/enhancements/fy0609/88018647-b.html.

Buchwald:1992:BRP

[Buc92] J. Z. Buchwald. Book review: P. M. Harman (ed.). The Scientific Letters and Papers of James Clerk Maxwell. Volume I: 1846–1862. Cambridge: Cambridge University Press, 1990. Pp. xxviii + 748. ISBN 0-521-25625-9. £135.00, $195. British Journal for the History of Science, 25(3):369–371, September 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.jstor.

org/stable/4027269.

Buchwald:1993:BRI

[Buc93] Jed Z. Buchwald. Book review: Innovation in Maxwell’s Electromagnetic Theory: Molecular Vortices, Displacement Current, and Light by Daniel M. Siegel. Isis, 84(2):395–396, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/236285.

Buchwald:1995:SPT

[Buc95] Jed Z. Buchwald, editor. Scientific practice: theories and stories of doing physics. University of Chicago Press, Chicago, IL, USA, 1995. ISBN 0-226-07889-2 (hardcover), 0-226-07890-6 (paperback). xiii + 398 pp. LCCN QC33 .S33 1995. URL http://www.

loc.gov/catdir/description/uchi052/94049715.html;   http:
//www.loc.gov/catdir/toc/uchi051/94049715.html.

Buhrke:1998:LWG

[Büh98a] Thomas Bührke. “War es ein Gott, der diese Zeichen schrieb?” James Clerk Maxwell (1831–1879). (German) [“Was it a god who wrote these characters?” James Clerk Maxwell (1831–1879)]. In Newtons Apfel: Sternstunden der Physik; von Galilei bis Lise Meitner. (German) [Newton’s apple: great moments of physics; from Galileo to Lise Meitner] [Büh98b], pages 60–83. ISBN 3-406-44402-4, 3-406-42002-8 (paperback). LCCN ????

Buhrke:1998:NAS

[Büh98b] Thomas Bührke. Newtons Apfel: Sternstunden der Physik; von Galilei bis Lise Meitner. (German) [Newton’s apple: great moments of physics; from Galileo to Lise Meitner], volume 1202 of Beck’sche Reihe. Verlag C. H. Beck, München, Germany, 1998. ISBN 3-406-44402-4, 3-406-42002-8 (paperback). 258 pp. LCCN ????

Bunge:1943:SFHa

[Bun43a] Mario Bunge. Significado físico e histórico de la Teoría de Maxwell. (Spanish) [Physical and historical significance of the Maxwell’s theory]. Tall. Gráficos Suárez y Dini., Buenos Aires, Argentina, 1943. 16 pp.

Bunge:1943:SFHb

[Bun43b] Mario Augusto Bunge. Significado físico e histórico de las ecuaciones de Maxwell. (Spanish) [Physical and historical significance of the Maxwell equations]. Universidad Obrera Argentina, Buenos Aires, Argentina, 1943. ???? pp.

Born:2019:POE

[BWB19] Max Born, Emil Wolf, and A. B. (Avadh Behari) Bhatia. Principles of Optics: Electromagnetic Theory of Propagation, Interference, and Diffraction of Light. Cambridge University Press, Cambridge, UK, seventh (expanded) anniversary edition, 60th anniversary edition, 2019. ISBN 1-108-47743-7 (hardcover), 1-108-76991-8 (e-pub). 990 (est.) pp. LCCN QC355.2 .B67

2019. URL https://assets.cambridge.org/97811084/77437/toc/
9781108477437_toc.pdf;   https://www.cambridge.org/core/
books/principles-of-optics/.

ClerkMaxwell:1864:RCA

[C^{+}64] James Clerk Maxwell et al. Report of the Committee appointed by the British Association on standards of electrical resistance. Report of the British Association for the Advancement of Science, ??(??): 111–176, ???? 1864.

ClerkMaxwell:1865:RCS

[C^{+}65] James Clerk Maxwell et al. Report of the Committee on standards of electrical resistance, Bath, September 1864. Report of the British Association for the Advancement of Science, ??(??):345–367, ????

1865. URL https://hdl.handle.net/2027/njp.32101076796364?
urlappend=%3Bseq=431.

ClerkMaxwell:1868:RCS

[C^{+}68] James Clerk Maxwell et al. Report of the Committee on standards of electrical resistance. Report of the British Association for the Advancement of Science, ??(??):474–522, ???? 1868.

ClerkMaxwell:1869:RCP

[C^{+}69] James Clerk Maxwell et al. Report of the committee for the purpose of investigating the rate of increase of underground temperature downwards in various localities, of dry land and under water. Report of the British Association for the Advancement of Science, ?? (??):510–514, ???? 1869. Read at Norwich, August 1868.

ClerkMaxwell:1870:RCS

[C^{+}70a] James Clerk Maxwell et al. Report of the Committee on standards of electrical resistance. Report of the British Association for the Advancement of Science, ??(??):434–438, ???? 1870.

ClerkMaxwell:1870:SRC

[C^{+}70b] James Clerk Maxwell et al. Second report of the Committee for the purpose of investigating the rate of increase of underground temperature downwards in various localities, of dry land and under water. Report of the British Association for the Advancement of Science, ?? (??):176–189, ???? 1870.

ClerkMaxwell:1871:RCS

[C^{+}71a] James Clerk Maxwell et al. Report of the Committee on standards of electrical resistance. Report of the British Association for the Advancement of Science, ??(??):14–15, ???? 1871.

ClerkMaxwell:1871:TRC

[C^{+}71b] James Clerk Maxwell et al. Third report of the Committee for the purpose of investigating the rate of increase of underground temperature downwards in various localities, of dry land and under water. Report of the British Association for the Advancement of Science, ?? (??):29–41, ???? 1871.

ClerkMaxwell:1872:FRC

[C^{+}72] James Clerk Maxwell et al. Fourth report of the Committee for the purpose of investigating the rate of increase of underground temperature downwards in various localities, of dry land and under water. Report of the British Association for the Advancement of Science, ?? (??):14–25, ???? 1872.

ClerkMaxwell:1873:FRC

[C^{+}73] James Clerk Maxwell et al. Fifth report of the Committee for the purpose of investigating the rate of incease of underground temperature downwards in various localities, of dry land and under water. Report of the British Association for the Advancement of Science, ?? (??):128–134, ???? 1873.

ClerkMaxwell:1874:FRC

[C^{+}74] James Clerk Maxwell et al. First report of the Committee for the selection and nomenclature of dynamical and electrical units. Report of the British Association for the Advancement of Science, ??(??): 222–225, ???? 1874.

ClerkMaxwell:1875:RCA

[C^{+}75a] James Clerk Maxwell et al. Report of the Committee appointed for the purpose of testing the new pyrometer of Mr. Siemens. Report of the British Association for the Advancement of Science, ??(??): 242–249, ???? 1875.

ClerkMaxwell:1875:SRC

[C^{+}75b] James Clerk Maxwell et al. Second report of the Committee for the selection and nomenclature of dynamical and electrical units. Report of the British Association for the Advancement of Science, ??(??):255, ???? 1875.

ClerkMaxwell:1876:ERC

[C^{+}76a] James Clerk Maxwell et al. Eighth report of the Committee for the purpose of investigating the rate of increase of underground temperature downwards in various localities, of dry land and under water.

Report of the British Association for the Advancement of Science, ?? (??):156–162, ???? 1876.

ClerkMaxwell:1876:SRC

[C^{+}76b] James Clerk Maxwell et al. Seventh report of the Committee for the purpose of investigating the rate of increase of underground temperature downwards in various localities, of dry land and under water. Report of the British Association for the Advancement of Science, ?? (??):14–18, ???? 1876.

ClerkMaxwell:1877:FRC

[C^{+}77a] James Clerk Maxwell et al. First report of the Committee appointed for the purpose of determining the mechanical equivalent of heat. Report of the British Association for the Advancement of Science, ?? (??):275, ???? 1877.

ClerkMaxwell:1877:NRC

[C^{+}77b] James Clerk Maxwell et al. Ninth report of the Committee for the purpose of investigating the rate of increase of underground temperature downwards in various localities, of dry land and under water. Report of the British Association for the Advancement of Science, ?? (??):204–211, ???? 1877.

ClerkMaxwell:1877:RCT

[C^{+}77c] James Clerk Maxwell et al. Report of the Committee for testing experimentally Ohm’s law. Report of the British Association for the Advancement of Science, ??(??):36–63, ???? 1877.

ClerkMaxwell:1878:TRC

[C^{+}78] James Clerk Maxwell et al. Tenth report of the Committee for the purpose of investigating the rate of increase of underground temperature downwards in various localities, of dry land and under water. Report of the British Association for the Advancement of Science, ?? (??):194–199, ???? 1878.

ClerkMaxwell:1879:RCA

[C^{+}79a] James Clerk Maxwell et al. Report of the Committee appointed for the purpose of endeavouring to procure reports on the progress of the chief branches of mathematics and physics. Report of the British Association for the Advancement of Science, ??(??):37–40, ???? 1879.

ClerkMaxwell:1879:RCC

[C^{+}79b] James Clerk Maxwell et al. Report of the Committee for commencing secular experiments upon the elasticity of wires. Report of the British Association for the Advancement of Science, ??(??):103–104, ???? 1879.

ClerkMaxwell:1879:SRC

[C^{+}79c] James Clerk Maxwell et al. Second report of the Committee for commencing secular experiments upon the elasticity of wires. Report of the British Association for the Advancement of Science, ??(??):33–36, ???? 1879.

ClerkMaxwell:1879:TRCa

[C^{+}79d] James Clerk Maxwell et al. Third report of the Committee appointed for the purpose of determining the mechanical equivalent of heat. Report of the British Association for the Advancement of Science, ?? (??):102, ???? 1879.

ClerkMaxwell:1879:TRCb

[C^{+}79e] James Clerk Maxwell et al. Twelfth report of the Committee for the purpose of investigating the rate of increase of underground temperature downwards in various localities, of dry land and under water. Report of the British Association for the Advancement of Science, ?? (??):40–46, ???? 1879.

ClerkMaxwell:1979:FRC

[C^{+}79] James Clerk Maxwell et al. Fourth report of the Committee appointed for the purpose of determining the mechanical equivalent of heat. Report of the British Association for the Advancement of Science, ?? (??):36, ???? 1979.

Cantor:1990:BRJ

[Can90] Geoffrey Cantor. Book review: Jed Z. Buchwald. The Rise of the Wave Theory of Light: Optical Theory and Experiment in the Early Nineteenth Century. Chicago and London: University of Chicago Press, 1989. Pp. xxiv + 474. ISBN 0-226-07884-1 (hardcover) and 0-226-07886-8 (paperback). £59.95, $86.25 (hardcover) £19.95, $28.75 (paperback). British Journal for the History of Science, 23(3):365–367, September 1990. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.jstor.org/

stable/4026786.

Capek:1976:CST

[Čap76] Milič Čapek. The Concepts of Space and Time: Their Structure and Their Development, volume 22 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1976. ISBN 90-277-0375-2, 94-010-1727-1 (e-book). ISSN 0068-0346. lvii + 570 pp. LCCN BD632. URL

https://link.springer.com/book/10.1007/978-94-010-1727-5.

Cat:2001:UMM

[Cat01] Jordi Cat. On understanding: Maxwell on the methods of illustration and scientific metaphor. Studies in History and Philosophy of Modern Physics, 32(3):395–441, September 2001. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219801000181.

Cat:2012:RPM

[Cat12] Jordi Cat. Into the ‘regions of physical and metaphysical chaos’: Maxwell’s scientific metaphysics and natural philosophy of action (agency, determinacy and necessity from theology, moral philosophy and history to mathematics, theory and experiment). Studies in History and Philosophy of Science Part A, 43(1):91–104, March 2012. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0039368111001002.

Cat:2014:MCS

[Cat14] Jordi Cat. Maxwell’s color statistics: From reduction of visible errors to reduction to invisible molecules. Studies in History and Philosophy of Science Part A, 48(??):60–75, December 2014. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0039368114000582.

Cat:2016:BRP

[Cat16] Jordi Cat. Book reviews: Peter Achinstein. Evidence and Method: Scientific Strategies of Isaac Newton and James Clerk Maxwell; Raymond Flood; Mark McCartney; Andrew Whitaker, eds. James Clerk Maxwell: Perspectives on His Life and Work. Isis, 107(3):642–644, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Collins:2016:KTN

[CDKR16] M. W. (Michael W.) Collins, R. C. (Richard C.) Dougal, C. S. (Carola S.) König, and I. S. (Ivan S.) Ruddock, editors. Kelvin, thermodynamics and the natural world, volume 10 of WIT series on design and nature. WIT Press, Southampton, UK, 2016. ISBN 1-84564-149-3 (hardcover), 1-84564-387-9 (e-book). ISSN 1755-8336. xxvi + 454 pp. LCCN QC16.K3 K36 2016.

Cendes:1989:UMM

[Cen89] Z. J. Cendes. Unlocking the magic of Maxwell’s equations. IEEE Spectrum, 26(4):29–33, April 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

ClerkMaxwell:1851:DOC

[CF51] James Clerk Maxwell and George Forbes. 1. On the description of oval curves and those having a plurality of foci, with remarks by Professor Forbes. Proceedings of the Royal Society of Edinburgh, 2(??):89–91, April 1851. CODEN PRSEAE. ISSN 0080-4541 (print), 2053-5902 (electronic). Paper presented on 6 April 1846 by his tutor, Philip Kelland: JCM was viewed as too young to do so, being then 14 years and 10 months old.

Chen:1991:MEP

[CF91] Xinfu Chen and Avner Friedman. Maxwell’s equations in a periodic structure. Transactions of the American Mathematical Society, 323(2):465–507, February 1991. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (electronic). URL http://www.ams.

org/journals/tran/1991-323-02/S0002-9947-1991-1010883-1;
 http://www.jstor.org/stable/2001542.

Craig:1998:REP

[CF98] Edward Craig and Luciano Floridi, editors. Routledge Encyclopedia of Philosophy Online. Routledge, London, UK, version 2.0 edition, 1998. ISBN 0-415-16916-X (CD-ROM), 0-415-19608-6 (user guide), 0-415-16917-8 (10 vol set + CD-ROM). LCCN B51. URL http://

www.rep.routledge.com/.

ClerkMaxwell:1868:EME

[CG68] James Clerk Maxwell and W. R. Grove. XLII. On Mr. Grove’s “experiment in magneto-electric induction”. In a letter to W. R. Grove, F.R.S. Philosophical Magazine (Series 4 (1851–1875)), 35(238):360– 363, ???? 1868. ISSN 0031-8086. URL http://www.tandfonline.

com/doi/abs/10.1080/14786446808639993. Letter to W. R. Grove,

F.R.S., 8 Palace Gardens Terrace, W. March 27, 1868.

Campbell:1882:LJC

[CG82] Lewis Campbell and William Garnett. The Life of James Clerk Maxwell: with a selection from his correspondence and occasional writings, and a sketch of his contributions to science. MacMillan Publishing Company, New York, NY, USA, 1882. xvi + 662 pp. LCCN QC16.M4 C2 1882. URL https://

archive.org/details/lifejamesclerkm01garngoog;   https://
www.biodiversitylibrary.org/item/71775.

Campbell:1884:LJC

[CG84] Lewis Campbell and William Garnett. The Life of James Clerk Maxwell: with a selection from his correspondence and occasional writings, and a sketch of his contributions to science. MacMillan Publishing Company, New York, NY, USA, new abridged and revised edition, 1884. xv + 421 pp. LCCN QC16.M4 C2 1884. With selections from his correspondence and occasional writings.

Campbell:1969:LJC

[CG69] Lewis Campbell and William Garnett. The life of James Clerk Maxwell, volume 85 of The Sources of Science. Johnson Reprint Corp., New York, NY, USA, 1969. xxv + xvi + 662 pp. LCCN QC16.M4 C2 1969. With a new preface and appendix with letters and by Robert H. Kargon.

Campbell:2010:LJC

[CG10] Lewis Campbell and William Garnett. The Life of James Clerk Maxwell: With a Selection from his Correspondence and Occasional Writings and a Sketch of his Contributions to Science. Cambridge Library Collection — Physical Sciences. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-511-70905-6 (e-book), 1-108-01370-8 (paperback). 4 + xvi + 662 pp. LCCN ????

Cooper:2017:PDH

[CG17] Malcolm (Malcolm J.) Cooper and Jim Grozier. Precise Dimensions: a History of Units from 1791–2018. IOP expanding physics, [IOP release 4]. Institute of Physics Publishers, Bristol, UK, 2017. ISBN 0-7503-1485-0 (hardcover), 0-7503-1486-9 (mobi), 0-7503-1487-7 (ebook). ISSN 2053-2563. 181 (est.) pp. LCCN QC86 .C667 2017eb.

URL http://iopscience.iop.org/book/978-0-7503-1487-9.

ClerkMaxwell:1995:MHS

[CGBE95] James Clerk Maxwell, Elizabeth Garber, Stephen G. Brush, and C. W. F. (C. W. Francis) Everitt. Maxwell on heat and statistical mechanics: on “avoiding all personal enquiries” of molecules. Lehigh University Press, Bethlehem, PA, USA, 1995. ISBN 0-934223-34-3. 550 pp. LCCN QC310.2 .M39 1995.

Chase:1885:CMR

[Cha85] Pliny Earle Chase. The Chase–Maxwell ratio. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 22(120):375–376, 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://www.jstor.

org/stable/982997.

Charlton:1971:MJC

[Cha71] T. M. Charlton. Maxwell, Jenkin and Cotterill and the theory of statically-indeterminate structures. Notes and Records of the Royal Society of London, 26(2):233–246, December 1971. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/531166.

Chalmers:1973:LME

[Cha73a] A. F. Chalmers. The limitations of Maxwell’s electromagnetic theory. Isis, 64(4):469–483, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229644.

Chalmers:1973:MMH

[Cha73b] A. F. Chalmers. Maxwell’s methodology and his application of it to electromagnetism. Studies in History and Philosophy of Science Part A, 4(2):107–164, August 1973. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368173900010.

Chambers:1984:BRD

[Cha84] R. G. Chambers. Book review: The Demon in the Aether: The Life of James Clerk Maxwell, Martin Goldman, 1983 Edinburgh: Paul Harris Publishing in association with Adam Hilger, ISBN 0-86228-026-5. Physics Bulletin, 35(5):201, ???? 1984. CODEN PHSBB4.

ISSN 0031-9112. URL http://iopscience.iop.org/article/10.
1088/0031-9112/35/5/025.

Chalmers:1986:HRM

[Cha86] Alan F. Chalmers. The heuristic role of Maxwell’s mechanical model of electromagnetic phenomena. Studies in History and Philosophy of Science Part A, 17(4):415–427, December 1986. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368186900026.

Charlton:1989:EMT

[Cha89] T. M. Charlton. An extension of Maxwell’s theory of pin-jointed frameworks by M. W. Crofton, F.R.S. Notes and Records of the Royal Society of London, 43(1):53–56, January 1989. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/531717.

Chalmers:2001:MMN

[Cha01] Alan Chalmers. Maxwell, mechanism, and the nature of electricity. Physics in Perspective (PIP), 3(4):425–438, November 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http:

//link.springer.com/article/10.1007/PL00000539.

Charalampous:2021:CAJ

[Cha21] Charis Charalampous. The confined atom: James Clerk Maxwell on the fundamental particles and the limits of scientific knowledge. Perspectives on Science, 29(2):189–214, March–April 2021. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL

https://muse.jhu.edu/article/787908.

Chisholm:1931:SR

[Chi31] Ian Stewart Chisholm. Saturn’s rings. The Edinburgh Academy Chronicle, 38 (supplement)(??):4–??, July 1931.

Chilton:1964:BRC

[Chi64] D. Chilton. Book review: Clerk Maxwell and Modern Science. Six commemorative lectures by Sir Edward V. Appleton, E. G. Bowen, C. A. Coulson, R. E. Peierls, Sir John Randall, and R. A. Smith. Edited by C. Domb. Pp. x + 118, frontis., 5 plates and numerous text figures. University of London, The Athlone Press. 1963. Price 25s. net. British Journal for the History of Science, 2(1):76, June 1964. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/stable/4025088.

Christov:2006:MIF

[Chr06] Christo I. Christov. On the material invariant formulation of Maxwell’s displacement current. Foundations of Physics, 36(11):1701– 1717, November 2006. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/
10.1007/s10701-006-9075-7.

ClerkMaxwell:1863:ACE

[CJ63a] James Clerk Maxwell and Fleeming Jenkin. Appendix C: On the elementary relations between electrical measurements. In Report of 1863 to the British Association, page ?? ????, ????, 1863.

ClerkMaxwell:1863:ADD

[CJ63b] James Clerk Maxwell and Fleeming Jenkin. Appendix D: Description of an experimental measurement of electrical resistance, made at King’s College [London]. In Report of 1863 to the British Association, pages 163–176. ????, ????, 1863. Parts I, III, and IV by Maxwell. Part II by Jenkin.

ClerkMaxwell:1865:LERa

[CJ65a] James Clerk Maxwell and Fleeming Jenkin. LXI. On the elementary relations between electrical measurements. Philosophical Magazine (Series 4 (1851–1875)), 29(198):436–460, ???? 1865. ISSN

0031-8086. URL http://www.tandfonline.com/doi/abs/10.1080/
14786446508643901.

ClerkMaxwell:1865:LERb

[CJ65b] James Clerk Maxwell and Fleeming Jenkin. LXX. On the elementary relations between electrical measurements. Philosophical Magazine (Series 4 (1851–1875)), 29(199):507–525, ???? 1865. ISSN

0031-8086. URL http://www.tandfonline.com/doi/abs/10.1080/
14786446508643913.

ClerkMaxwell:1864:ADF

[CJH64] James Clerk Maxwell, Fleeming Jenkin, and Charles Hockin. Appendix A: Description of a further experimental measurement of electrical resistance made at King’s College [London]. Report of the British Association for the Advancement of Science, ??(??):

350–351, ???? 1864. URL https://hdl.handle.net/2027/njp.
32101076796364?urlappend=%3Bseq=438.

Cho:2015:HGP

[CL15] Adrian Cho and Ilima Loomis. Hawaii’s Governor proposes telescope swap. Science news Web site., June 2, 2015. URL

http://www.sciencemag.org/news/2015/06/hawaii-s-governor-

proposes-telescope-swap. Discussion of the religious controversy over the 13 telescopes on Hawaii’s Mauna Kea volcano, one of which is the James Clerk Maxwell Telescope, and possibly slated for removal in an appeasement effort.

Clausius:1858:MLW

[Cla58] Rudolf Julius Emmanuel Clausius. Ueber die mittlere Länge der Wege, welche bei der Molecularbewegung gasförmiger Körper von den einzelnen Molecülen zurückgelegt werden; nebst einigen anderen Bemerkungen über die mechanische Wärmetheorie. (German) [On the mean lengths of paths described by separate molecules of gaseous bodies]. Annalen der Physik (Berlin), 181(10):239–258, ???? 1858. CODEN ANPYA2. ISSN 1521-3889. URL http://onlinelibrary.

wiley.com/doi/10.1002/andp.18581811005/full. See also English

translation in [Cla59].

Clausius:1859:XML

[Cla59] Rudolf Julius Emmanuel Clausius. X. On the mean length of the paths described by the separate molecules of gaseous bodies on the occurrence of molecular motion: together with some other remarks upon the mechanical theory of heat. Philosophical Magazine (Series 4 (1851–1875)), 17(112):81–91, February 1859. CODEN PHMAA4.

ISSN 0031-8086. URL http://www.tandfonline.com/doi/abs/10.

1080/14786445908642626. English translation of [Cla58].

ClerkMaxwell:1846:DO

[Cle46a] James Clerk Maxwell. Descartes’ ovals. Unpublished notes, cited in [Tai80a]., 1846.

ClerkMaxwell:1846:MA

[Cle46b] James Clerk Maxwell. The meloid and apioid. Unpublished notes, cited in [Tai80a]., 1846.

ClerkMaxwell:1846:TC

[Cle46c] James Clerk Maxwell. Trifocal curves. Unpublished notes, cited in [Tai80a]., 1846.

ClerkMaxwell:1849:TRC

[Cle49] James Clerk Maxwell. On the theory of rolling curves. Transactions of the Royal Society of Edinburgh, 16(5):519–540, ???? 1849. CODEN TRSEAO. ISSN 0080-4568 (print), 2053-5945 (electronic). URL

https://hdl.handle.net/2027/mdp.39015035450207?urlappend=
%3Bseq=577.

ClerkMaxwell:1851:EES

[Cle51a] James Clerk Maxwell. 1. On the equilibrium of elastic solids. Proceedings of the Royal Society of Edinburgh, 2:294–296, 1851. CODEN PRSEAE. ISSN 0080-4541 (print), 2053-5902 (electronic).

ClerkMaxwell:1851:NHB

[Cle51b] James Clerk Maxwell. Notes from Hopkins booklets and lectures. Six notebooks. University Library, Cambridge, Add. MSS 7655, v, m/5– 10., 1851.

ClerkMaxwell:1851:NDI

[Cle51c] James Clerk Maxwell. Notes on differential & integral calculus. University Library, Cambridge, Add. MSS 7655, v, m/7, f. 9., 1851.

ClerkMaxwell:1851:NSD

[Cle51d] James Clerk Maxwell. Notes on statics dynamics. University Library, Cambridge, Add. MSS 7655, v, m/7, f. 9., 1851.

ClerkMaxwell:1853:IEE

[Cle53] James Clerk Maxwell. IV. On the equilibrium of elastic solids. Transactions of the Royal Society of Edinburgh, 20 (part 1):87– 120, 1853. CODEN TRSEAO. ISSN 0080-4568 (print), 2053-5945

(electronic). URL https://hdl.handle.net/2027/hvd.hwhq7y?
urlappend=%3Bseq=113.

ClerkMaxwell:1854:PCD

[Cle54a] James Clerk Maxwell. On a particular case of the descent of a heavy body in a resisting medium. Cambridge and Dublin Mathematical Journal, 9(??):145–148, May 1854. URL http://gdz.sub.uni-

goettingen.de/en/dms/loader/img/?PID=PPN600493962_0009%7CLOG_

0025. See [Tai80a, page 319, column 2] for an description of this work applied to a falling slip of paper.

ClerkMaxwell:1854:SP

[Cle54b] James Clerk Maxwell. Solutions of problems. Cambridge and Dublin Mathematical Journal, 8(??):188–189, February 1854. Contains invention of the fisheye lens.

ClerkMaxwell:1855:FLF

[Cle55a] James Clerk Maxwell. On Faraday’s lines of force [Part I abstract only]. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 1 (part xi)(??):160–162, December 10, 1855. CODEN PCPSA4. ISSN 0008-1981. URL https://hdl.handle.net/

2027/njp.32101018907830?urlappend=%3Bseq=170. Full paper in

[Cle58a].

ClerkMaxwell:1855:TCR

[Cle55b] James Clerk Maxwell. On the theory of colours in relation to colourblindness. Letter to Dr G. Wilson, reprinted in [Wil55]., January 4, 1855.

ClerkMaxwell:1856:DNF

[Cle56a] James Clerk Maxwell. Description of a new form of the platometer, an instrument for measuring the areas of plane figures drawn on paper. Transactions of the Royal Scottish Society of Arts [1839–1925], 4(??): 420–429, ???? 1856.

ClerkMaxwell:1856:ILA

[Cle56b] James Clerk Maxwell. Inaugural lecture at Aberdeen. Manuscript unpublished in JCM’s lifetime, but held in the Cambridge University Library, and reprinted in [Jon73, pages 69–81] and [Har90, pages 419– 431]. A summary appeared in [Cle56g], November 3, 1856.

ClerkMaxwell:1856:FLFa

[Cle56c] James Clerk Maxwell. On Faraday’s lines of force [abstract only]. Philosophical Magazine (Series 4 (1851–1875)), 9(61):404–405, ????

1856. ISSN 0031-8086. URL http://www.tandfonline.com/doi/
abs/10.1080/14786445508641892. Full paper in [Cle58a].

ClerkMaxwell:1856:FLFb

[Cle56d] James Clerk Maxwell. On Faraday’s lines of force [Part II abstract only]. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 1 (part xii)(??):163–166, February 11, 1856. CODEN PCPSA4. ISSN 0008-1981. URL https://hdl.handle.net/

2027/njp.32101018907830?urlappend=%3Bseq=173. Full paper in

[Cle58a].

ClerkMaxwell:1856:TCR

[Cle56e] James Clerk Maxwell. On the theory of colours in relation to colourblindness. Transactions of the Royal Scottish Society of Arts [1839– 1925], 4(??):394–400, ???? 1856. First printed in [Wil55].

ClerkMaxwell:1856:TSB

[Cle56f] James Clerk Maxwell. On the transformation of surfaces by bending. Transactions of the Cambridge Philosophical Society, 9(??):445–470,

???? 1856. URL http://biodiversitylibrary.org/page/2437845.

ClerkMaxwell:1856:SIL

[Cle56g] James Clerk Maxwell. [Summary of inaugural lecture at Aberdeen]. Aberdeen Journal, ??(??):??, November 5, 1856.

ClerkMaxwell:1857:ECP

[Cle57a] James Clerk Maxwell. 1. Experiments on colour as perceived by the eye, with remarks on colour-blindness. Proceedings of the Royal Society of Edinburgh, 3:299–301, 1857. CODEN PRSEAE. ISSN 0080-4541 (print), 2053-5902 (electronic). URL http://lhldigital.

lindahall.org/cdm/ref/collection/color/id/29023.

ClerkMaxwell:1857:IAE

[Cle57b] James Clerk Maxwell. IV: Account of experiments on the perception of colour. Philosophical Magazine (Series 4 (1851–1875)), 14(90):40–47, ???? 1857. ISSN 0031-8086. URL http://www.

tandfonline.com/doi/abs/10.1080/14786445708642349. This pa-

per demonstrates that “every sensation of colour in a perfect human eye is distinguished by three, and only three, elementary quantities, so that in mathematical language, the quality of a colour may be expressed as a function of three independent variables.”.

ClerkMaxwell:1857:DTE

[Cle57c] James Clerk Maxwell. On a dynamical top, for exhibiting the phenomena of the motion of a system of invariable form about a fixed point; with some suggestions as to the Earth’s motion. Proceedings of the Royal Society of Edinburgh, 3(??):503–504, January 1857. CO-DEN PRSEAE. ISSN 0080-4541 (print), 2053-5902 (electronic).

ClerkMaxwell:1857:MDT

[Cle57d] James Clerk Maxwell. On a method of drawing the theoretical forms of Faraday’s lines of force without calculation. Report of the British Association for the Advancement of Science, ??(??):12–??, ???? 1857.

ClerkMaxwell:1857:IIP

[Cle57e] James Clerk Maxwell. On an instrument to illustrate Poinsôt’s theory of rotation. Report of the British Association for the Advancement of Science, ??(??):27–28, ???? 1857.

ClerkMaxwell:1857:TCC

[Cle57f] James Clerk Maxwell. ‘on the theory of compound colours with reference to mixtures of blue and yellow light. Report of the British Association for the Advancement of Science, ??(??):12–13, ???? 1857.

ClerkMaxwell:1857:USF

[Cle57g] James Clerk Maxwell. On the unequal sensibility of the Foramen Centrale to light of different colours. Report of the British Association for the Advancement of Science, ??(??):12–??, ???? 1857.

ClerkMaxwell:1857:XEC

[Cle57h] James Clerk Maxwell. XVIII. Experiments on colour, as perceived by the eye, with remarks on colour-blindness. Transactions of the Royal Society of Edinburgh, 21(2):275–298, January 1857. CODEN TRSEAO. ISSN 0080-4568 (print), 2053-5945 (electronic). URL

https://hdl.handle.net/2027/mdp.39015005660991?urlappend=
%3Bseq=331.

ClerkMaxwell:1857:XDT

[Cle57i] James Clerk Maxwell. XXXIV: On a dynamical top, for exhibiting the phenomena of the motion of a system of invariable form about a fixed point; with some suggestions as to the Earth’s motion. Transactions of the Royal Society of Edinburgh, 21(4):559– 570, ???? 1857. CODEN TRSEAO. ISSN 0080-4568 (print),

2053-5945 (electronic). URL https://hdl.handle.net/2027/mdp.
39015005660991?urlappend=%3Bseq=639.

ClerkMaxwell:1858:IFL

[Cle58a] James Clerk Maxwell. III. On Faraday’s lines of force. Transactions of the Cambridge Philosophical Society, 10 (part 1)(??):27–

83, ???? 1858. URL http://biodiversitylibrary.org/page/
2424191;   http://www.scribd.com/doc/39568221/maxwell-on-

faraday-s-lines-of-force. See extensive commentary about the philosophy and theology in this paper [Lam11].

ClerkMaxwell:1858:GLO

[Cle58b] James Clerk Maxwell. On the general laws of optical instruments. The Quarterly Journal of Pure and Applied Mathematics [1857–1927], 2

(??):233–246, ???? 1858. URL https://hdl.handle.net/2027/uc1.
%24b417524?urlappend=%3Bseq=251.

ClerkMaxwell:1859:IDT

[Cle59a] James Clerk Maxwell. On the dynamical theory of gases. In ????, editor, [Proceedings of the August 1859 meeting of the British Association for the Advancement of Science: Transactions of Section A], pages 9–?? ????, ????, 1859.

ClerkMaxwell:1859:SMSa

[Cle59b] James Clerk Maxwell. On the stability of the motion of Saturn’s rings. Monthly Notices of the Royal Astronomical Society, 19(??):297–384, ???? 1859. CODEN MNRAA4. ISSN 0035-8711 (print), 1365-2966 (electronic). Adams Prize Essay, 1856, for Cambridge University.

ClerkMaxwell:1859:SMSb

[Cle59c] James Clerk Maxwell. On the Stability of the Motion of Saturn’s rings: an Essay which Obtained the Adams Prize for the Year 1856, in the University of Cambridge. MacMillan Publishing Company, New York, NY, USA, 1859. viii + 71 pp. URL https://archive.org/

details/onstabilityofmot00maxw.

ClerkMaxwell:185x:UAC

[Cle5x] James Clerk Maxwell. [unpublished answer to Clausius’s remarks]. Cited in [Gar70, footnote 52, page 314]., 185x.

ClerkMaxwell:1860:IIDa

[Cle60a] James Clerk Maxwell. II. Illustrations of the dynamical theory of gases. Part II. on the process of diffusion of two or more kinds of moving particles among one another. Philosophical Magazine (Series 4 (1851–1875)), 20(130):21–33, July 1860. ISSN 0031-8086. URL http:

//www.tandfonline.com/doi/abs/10.1080/14786446008642902.

ClerkMaxwell:1860:IIDb

[Cle60b] James Clerk Maxwell. II. Illustrations of the dynamical theory of gases. Part III. on the collision of perfectly elastic bodies of any form. Philosophical Magazine (Series 4 (1851–1875)), 20(130):33–37, July 1860. ISSN 0031-8086.

ClerkMaxwell:1860:IEM

[Cle60c] James Clerk Maxwell. On an instrument for exhibiting the motions of Saturn’s rings. Report of the British Association for the Advancement of Science, ??(??):62–??, ???? 1860.

ClerkMaxwell:1860:DTG

[Cle60d] James Clerk Maxwell. On the dynamical theory of gases. Report of the British Association for the Advancement of Science, ??(??):9–??, ???? 1860.

ClerkMaxwell:1860:MCS

[Cle60e] James Clerk Maxwell. On the mixture of the colours of the spectrum. Report of the British Association for the Advancement of Science, ?? (??):15–??, ???? 1860.

ClerkMaxwell:1860:TCC

[Cle60f] James Clerk Maxwell. On the theory of compound colours, and the relations of the colours of the spectrum. Philosophical transactions of the Royal Society of London, 150(??):57–84, ???? 1860. CODEN PTRSAV. ISSN 0370-2316 (print), 2053-9207 (electronic).

ClerkMaxwell:1860:VID

[Cle60g] James Clerk Maxwell. V. Illustrations of the dynamical theory of gases. Part I. On the motions and collisions of perfectly elastic spheres. Philosophical Magazine (Series 4 (1851–1875)), 19(124):19–32, Jan-

uary 1860. ISSN 0031-8086. URL http://www.tandfonline.com/
doi/abs/10.1080/14786446008642818. See comments [Bru62].

ClerkMaxwell:1861:LPL

[Cle61a] James Clerk Maxwell. LI. On physical lines of force. Philosophical Magazine (Series 4 (1851–1875)), 21(141):338–348, ???? 1861. ISSN

0031-8086. URL http://www.tandfonline.com/doi/abs/10.1080/
14786446108643056.

ClerkMaxwell:1861:IEM

[Cle61b] James Clerk Maxwell. On an instrument for exhibiting any mixture of the colours of the spectrum. Report of the British Association for the Advancement of Science, ??(??):16–??, ???? 1861.

ClerkMaxwell:1861:PLF

[Cle61c] James Clerk Maxwell. On physical lines of force. The London, Edinburgh, and Dublin Philosophical Magazine and Journal of Science: series 4, XXI(??):161–175, ???? 1861.

ClerkMaxwell:1861:RBT

[Cle61d] James Clerk Maxwell. On the results of Bernoulli’s theory of gases as applied to their internal friction, their diffusion, and their conductivity for heat. Report of the British Association for the Advancement of Science, pages 15–16, 1861.

ClerkMaxwell:1861:TTP

[Cle61e] James Clerk Maxwell. On the theory of three primary colours. Notices of the Proceedings of the Royal Institution of Great Britain [1851– 1928], 11(??):370–374, ???? 1861. Lecture at the Royal Institution of Great Britain, May 17, 1861.

ClerkMaxwell:1861:TPC

[Cle61f] James Clerk Maxwell. The theory of the primary colours. British Journal of Photography, ??(??):??, August 9, 1861. CODEN BR-

JFAM. ISSN 0007-1196. URL http://notesonphotographs.org/
index.php?title=%22The_Theory_of_the_Primary_Colours.%22_
The_British_Journal_of_Photography,_August_9,_1861.

ClerkMaxwell:1861:XPLb

[Cle61g] James Clerk Maxwell. XLIV. On physical lines of force. Part II. The theory of molecular vortices applied to electric currents. Philosophical Magazine (Series 4 (1851–1875)), 21(140):281–291, ???? 1861. ISSN

0031-8086. URL http://www.tandfonline.com/doi/abs/10.1080/
14786446108643056.

ClerkMaxwell:1861:XPLa

[Cle61h] James Clerk Maxwell. XXV. On physical lines of force. Part I. The theory of molecular vortices applied to magnetic phenomena. Philosophical Magazine (Series 4 (1851–1875)), 21(139):161–175, 1861.

ISSN 0031-8086. URL http://www.tandfonline.com/doi/abs/10.
1080/14786446108643033.

ClerkMaxwell:1861:PKG

[Cle62a] James Clerk Maxwell. Ueber physikalische Kraftlinien. (German) [On physical lines of force]. ????, Leipzig, Germany, 1861–1862. LCCN ???? Translation to German by Ludwig Boltzmann.

ClerkMaxwell:1862:TCS

[Cle62b] James Clerk Maxwell. 2. On theories of the constitution of Saturn’s rings. Proceedings of the Royal Society of Edinburgh, 4(??):99–101, ???? 1862. CODEN PRSEAE. ISSN 0080-4541 (print), 2053-5902 (electronic).

ClerkMaxwell:1862:IPL

[Cle62c] James Clerk Maxwell. III. On physical lines of force. Part III. The theory of molecular vortices applied to statical electricity. Philosophical Magazine (Series 4 (1851–1875)), 23(151):12–24, January 1862.

ISSN 0031-8086. URL http://www.tandfonline.com/doi/abs/10.
1080/14786446208643207.

ClerkMaxwell:1862:XPL

[Cle62d] James Clerk Maxwell. XIV. On physical lines of force. Part IV. The theory of molecular vortices applied to the action of magnetism on polarised light. Philosophical Magazine (Series 4 (1851–1875)), 23 (152):85–95, January 1862. ISSN 0031-8086. URL http://www.

tandfonline.com/doi/abs/10.1080/14786446208643219.

ClerkMaxwell:1863:CHG

[Cle63] James Clerk Maxwell. On the conduction of heat in gases. Unpublished work, reprinted in [Gar66, pages 279–297], and in [Niv11]., 1863.

ClerkMaxwell:1864:DTE

[Cle64a] James Clerk Maxwell. A dynamical theory of the electromagnetic field [abstract only]. Proceedings of the Royal Society of London, 13 (??):531, ???? 1864. CODEN PRSLAZ. ISSN 0370-1662 (print), 2053-9126 (electronic).

ClerkMaxwell:1864:CES

[Cle64b] James Clerk Maxwell. L. On the calculation of the equilibrium and stiffness of frames. Philosophical Magazine (Series 4 (1851–1875)), 27(182):294–299, ???? 1864. ISSN 0031-8086. URL http://www.

tandfonline.com/doi/abs/10.1080/14786446408643668.

ClerkMaxwell:1864:FLF

[Cle64c] James Clerk Maxwell. On Faraday’s lines of force. Transactions of the Cambridge Philosophical Society, X(??):27–83, ???? 1864.

ClerkMaxwell:1864:XRF

[Cle64d] James Clerk Maxwell. XLV. On reciprocal figures and diagrams of forces. Philosophical Magazine (Series 4 (1851–1875)), 27(182):250– 261, ???? 1864. ISSN 0031-8086. URL http://www.tandfonline.

com/doi/abs/10.1080/14786446408643663.

ClerkMaxwell:1865:VDT

[Cle65] James Clerk Maxwell. VIII. A dynamical theory of the electromagnetic field. Philosophical transactions of the Royal Society of London, 155:459–512, January 1, 1865. CODEN PTRSAV. ISSN 0370-2316 (print), 2053-9207 (electronic). URL http://upload.

wikimedia.org/wikipedia/commons/1/19/A_Dynamical_Theory_
of_the_Electromagnetic_Field.pdf;   http://www.jstor.org/

stable/108892. See [Eng15] for a 150-year retrospective on this paper that introduced ‘Maxwell’s equations’.

ClerkMaxwell:1866:ETO

[Cle66a] James Clerk Maxwell. On the elementary theory of optical instruments. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 1 (part xii)(??):175, ???? 1866. CO-DEN PCPSA4. ISSN 0008-1981.

ClerkMaxwell:1866:VIF

[Cle66b] James Clerk Maxwell. On the viscosity or internal friction of air and other gases [abstract only]. Proceedings of the Royal Society of London, 15(??):14–17, ???? 1866. CODEN PRSLAZ. ISSN 0370-1662 (print), 2053-9126 (electronic). Full paper in [Cle66e].

ClerkMaxwell:1866:RCS

[Cle66c] James Clerk Maxwell. Report of the Committee on standards of electrical resistance, Birmingham, September 1865. Report of the British Association for the Advancement of Science, ??(??):308–313, ???? 1866.

ClerkMaxwell:1866:TSB

[Cle66d] James Clerk Maxwell. The transformation of surfaces by bending. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 1 (part x [127–146])(??):134–136, ???? 1866. CO-DEN PCPSA4. ISSN 0008-1981.

ClerkMaxwell:1866:XBL

[Cle66e] James Clerk Maxwell. XIII. the Bakerian Lecture. On the viscosity or internal friction of air and other gases. Philosophical transactions of the Royal Society of London, 156(??):249–268, ???? 1866. CO-DEN PTRSAV. ISSN 0370-2316 (print), 2053-9207 (electronic). The Bakerian Lecture for 1866, read 8 February 1866.

ClerkMaxwell:1867:DTGb

[Cle67a] James Clerk Maxwell. On the dynamical theory of gases. Philosophical transactions of the Royal Society of London, 157(??):49–88, ???? 1867. CODEN PTRSAV. ISSN 0370-2316 (print), 2053-9207 (electronic).

ClerkMaxwell:1867:DTGa

[Cle67b] James Clerk Maxwell. On the dynamical theory of gases [abstract only]. Proceedings of the Royal Society of London, 15(??):167–171, ?? 1867. CODEN PRSLAZ. ISSN 0370-1662 (print), 2053-9126 (electronic). Full paper in [Cle67a].

ClerkMaxwell:1867:ESE

[Cle67c] James Clerk Maxwell. On the equilibrium of a spherical envelope. The Quarterly Journal of Pure and Applied Mathematics [1857–1927], 8 (??):325–333, ???? 1867.

ClerkMaxwell:1867:TME

[Cle67d] James Clerk Maxwell. On the theory of the maintenance of electric currents by mechanical work without the use of permanent magnets. Proceedings of the Royal Society of London, 15(??):397–402, ???? 1867. CODEN PRSLAZ. ISSN 0370-1662 (print), 2053-9126 (electronic).

ClerkMaxwell:1868:CSS

[Cle68a] James Clerk Maxwell. The construction of stereograms of surfaces. Proceedings of the London Mathematical Society. First Series, s1-2: 57–58, January 23, 1868. CODEN PLMTAL. ISSN 0024-6115 (print),

1460-244X (electronic). URL http://onlinelibrary.wiley.com/
doi/10.1112/plms/s1-2.1.50/full.

ClerkMaxwell:1868:MMDa

[Cle68b] James Clerk Maxwell. On a method of making a direct comparison of electrostatic and electromagnetic force; with a note on the electromagnetic theory of light [abstract only]. Philosophical Magazine (Series 4 (1851–1875)), 36(243):316–317, ???? 1868. ISSN

0031-8086. URL http://www.tandfonline.com/doi/abs/10.1080/

14786446808640067. Full paper in [Cle69b].

ClerkMaxwell:1868:MMDb

[Cle68c] James Clerk Maxwell. ‘on a method of making a direct comparison of electrostatic and electromagnetic force; with a note on the electromagnetic theory of light [abstract only]. Proceedings of the Royal Society of London, 16(??):449–450, ???? 1868. CODEN PRSLAZ. ISSN 0370-1662 (print), 2053-9126 (electronic). Full paper in [Cle69b].

ClerkMaxwell:1868:RIS

[Cle68d] James Clerk Maxwell. On a real-image stereoscope. Report of the British Association for the Advancement of Science, ??(??):11–??, ???? 1868.

ClerkMaxwell:1868:Ga

[Cle68e] James Clerk Maxwell. On governors. Proceedings of the Royal Society of London, 16:270–283, 1868. CODEN PRSLAZ. ISSN 0370-1662 (print), 2053-9126 (electronic).

ClerkMaxwell:1868:Gb

[Cle68f] James Clerk Maxwell. On governors. Philosophical Magazine (Series 4 (1851–1875)), 35(238):385–398, ???? 1868. ISSN

0031-8086. URL http://www.tandfonline.com/doi/abs/10.1080/
14786446808639996.

ClerkMaxwell:1868:RDS

[Cle68g] James Clerk Maxwell. On reciprocal diagrams in space, and their relation to Airy’s function of stress. Proceedings of the London Mathematical Society. First Series, s1-2:58–61, January 23, 1868. CODEN PLM-TAL. ISSN 0024-6115 (print), 1460-244X (electronic). URL http://

onlinelibrary.wiley.com/doi/10.1112/plms/s1-2.1.58/full.

ClerkMaxwell:1868:C

[Cle68h] James Clerk Maxwell. On the cyclide. The Quarterly Journal of Pure and Applied Mathematics [1857–1927], 9:111–126,

1868.  URL https://hdl.handle.net/2027/uc1.%24b417531?
urlappend=%3Bseq=127.

ClerkMaxwell:1868:TDF

[Cle68i] James Clerk Maxwell. On the theory of diagrams of forces as applied to roofs and bridges. Report of the British Association for the Advancement of Science, ??(??):156–??, ???? 1868.

ClerkMaxwell:1868:XDTa

[Cle68j] James Clerk Maxwell. XV. On the dynamical theory of gases. Philosophical Magazine (Series 4 (1851–1875)), 35(235):129–145, ????

1868. ISSN 0031-8086. URL http://www.tandfonline.com/doi/
abs/10.1080/14786446808639951. See also continuation [Cle68k].

ClerkMaxwell:1868:XDTb

[Cle68k] James Clerk Maxwell. XXII. On the dynamical theory of gases. On the final distribution of velocity among the molecules of two systems acting on one another according to any law of force. Philosophical Magazine (Series 4 (1851–1875)), 35(236):185–217, ???? 1868. ISSN

0031-8086. URL http://www.tandfonline.com/doi/abs/10.1080/

14786446808639963. Continuation of [Cle68j].

ClerkMaxwell:1869:BAP

[Cle69a] James Clerk Maxwell. 4. On the best arrangement for producing a pure spectrum on a screen. Proceedings of the Royal Society of Edinburgh, 6(??):238–242, ???? 1869. CODEN PRSEAE. ISSN 0080-4541 (print), 2053-5902 (electronic).

ClerkMaxwell:1869:MMD

[Cle69b] James Clerk Maxwell. On a method of making a direct comparison of electrostatic with electromagnetic force; with a note on the electromagnetic theory of light. Philosophical transactions of the Royal Society of London, 158(??):643–657, ???? 1869. CODEN PTRSAV. ISSN 0370-2316 (print), 2053-9207 (electronic).

ClerkMaxwell:1869:JBL

[Cle69c] James Clerk Maxwell. On J. B. Listing’s paper, “On the kinds of cyclosis in lines, surfaces, and regions of space”, in the 10th volume of the Göttingen Transactions. Proceedings of the London Mathematical Society. First Series, s1-2:165–166, February 11, 1869. CODEN PLM-TAL. ISSN 0024-6115 (print), 1460-244X (electronic). URL http://

onlinelibrary.wiley.com/doi/10.1112/plms/s1-2.1.161/full.

ClerkMaxwell:1869:RFF

[Cle69d] James Clerk Maxwell. On reciprocal figures, frames, and diagrams of forces. Transactions of the Royal Society of Edinburgh, 26:1–40, 1869. CODEN TRSEAO. ISSN 0080-4568 (print), 2053-5945 (electronic).

ClerkMaxwell:1869:IMH

[Cle69e] James Clerk Maxwell. On the influence of the motions of the heavenly bodies on the index of refraction of light. Philosophical transactions of the Royal Society of London, 158(??):532–535, ???? 1869. CODEN PTRSAV. ISSN 0370-2316 (print), 2053-9207 (electronic). Embedded in [Hug69].

ClerkMaxwell:1869:QCP

[Cle69f] James Clerk Maxwell. Question — “can the potential of a uniform circular disc at any point be expressed by means of elliptic integrals?”. Proceedings of the London Mathematical Society. First Series, s1-3(??):8, November 11, 1869. CODEN PLMTAL. ISSN 0024-6115 (print), 1460-244X (electronic). URL http://onlinelibrary.

wiley.com/doi/10.1112/plms/s1-3.1.6/full.

ClerkMaxwell:1869:RMH

[Cle69g] James Clerk Maxwell. Remarks on Mr. Hanlon’s paper on “The Vena Contracta”. Proceedings of the London Mathematical Society. First Series, s1-3:6–9, November 11, 1869. CODEN PLMTAL. ISSN 0024-6115 (print), 1460-244X (electronic). URL http://onlinelibrary.

wiley.com/doi/10.1112/plms/s1-3.1.6/full.

ClerkMaxwell:186x:ILE

[Cle6x] James Clerk Maxwell. Introductory lecture on experimental physics. Unpublished lecture., 186x.

ClerkMaxwell:1870:AMPb

[Cle70a] James Clerk Maxwell. Address to the Mathematical and Physical Science Sections of the British Association. Report of the British Association for the Advancement of Science, ??(??):1–9, ???? 1870.

ClerkMaxwell:1870:AMPa

[Cle70b] James Clerk Maxwell. Address to the Mathematical and Physical Science Sections of the British Association [abstract only]. Nature, 2 (47):419–422, September 22, 1870. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.

com/nature/journal/v2/n47/pdf/002419a0.pdf.

ClerkMaxwell:1870:EVR

[Cle70c] James Clerk Maxwell. Experiments on the value of v, the ratio of the electromagnetic to the electrostatic unit of electricity. Report of the British Association for the Advancement of Science, ??(??):436–438, ???? 1870.

ClerkMaxwell:1870:RFF

[Cle70d] James Clerk Maxwell. I. on reciprocal figures, frames, and diagrams of forces. Transactions of the Royal Society of Edinburgh, 26(1):1–40, January 1870. CODEN TRSEAO. ISSN 0080-4568 (print), 2053-5945

(electronic). URL http://digital.nls.uk/scientists/archive/

74629052. This paper was awarded the Keith Prize of the Royal Society of Edinburgh.

ClerkMaxwell:1870:LEH

[Cle70e] James Clerk Maxwell. Letter to the Editor: L. On hills and dales. Philosophical Magazine (Series 4 (1851–1875)), 40(269):421– 427, ???? 1870. ISSN 0031-8086. URL http://www.tandfonline.

com/doi/abs/10.1080/14786447008640422. See also [Cle71d].

ClerkMaxwell:1870:DCF

[Cle70f] James Clerk Maxwell. On the displacement in a case of fluid motion. Proceedings of the London Mathematical Society. First Series, s1-3: 82–87, March 10, 1870. CODEN PLMTAL. ISSN 0024-6115 (print),

1460-244X (electronic). URL http://onlinelibrary.wiley.com/
doi/10.1112/plms/s1-3.1.82/full.

ClerkMaxwell:1870:TH

[Cle70g] James Clerk Maxwell. Theory of Heat. Longmans, Green, and Co., London, UK, 1870. xii + 312 pp. LCCN ???? URL https://hdl.

handle.net/2027/nyp.33433057781498;   https://hdl.handle.
net/2027/nyp.33433069099806.

ClerkMaxwell:1871:ILE

[Cle71a] James Clerk Maxwell. Introductory lecture on experimental physics. MacMillan Publishing Company, New York, NY, USA, 1871. ???? pp.

ClerkMaxwell:1871:CV

[Cle71b] James Clerk Maxwell. On colour vision. Nature, 4(79):13–16, May 4, 1871. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL https://www.nature.com/nature/journal/v4/n79/
pdf/004013b0.pdf.

ClerkMaxwell:1871:CVD

[Cle71c] James Clerk Maxwell. On colour-vision at different points of the retina. Report of the British Association for the Advancement of Science, ??(??):40–41, ???? 1871.

ClerkMaxwell:1871:HD

[Cle71d] James Clerk Maxwell. On hills and dales. Report of the British Association for the Advancement of Science, ??(??):17–18, ???? 1871.

ClerkMaxwell:1871:QPD

[Cle71e] James Clerk Maxwell. Question as to positive direction of rotation. Proceedings of the London Mathematical Society. First Series, s1-3(??):279–280, May 11, 1871. CODEN PLMTAL. ISSN 0024-6115 (print), 1460-244X (electronic). URL http://onlinelibrary.

wiley.com/doi/10.1112/plms/s1-3.1.269/full.

ClerkMaxwell:1871:RMC

[Cle71f] James Clerk Maxwell. Remarks on the mathematical classification of physical quantities. Proceedings of the London Mathematical Society. First Series, s1-3:224–232, March 9, 1871. CO-DEN PLMTAL. ISSN 0024-6115 (print), 1460-244X (electronic).

URL http://onlinelibrary.wiley.com/doi/10.1112/plms/s1-
3.1.224/abstract.

ClerkMaxwell:1871:TH

[Cle71g] James Clerk Maxwell. Theory of Heat. Longmans, Green, and Co., London, UK, 1871. xii + 312 pp. LCCN ???? URL https://hdl.

handle.net/2027/nyp.33433057781498;   https://hdl.handle.
net/2027/nyp.33433069099806.

ClerkMaxwell:1871:TEM

[Cle71h] James Clerk Maxwell. A Treatise on Electricity and Magnetism. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1871.

ClerkMaxwell:1872:RFFa

[Cle72a] James Clerk Maxwell. 1. On reciprocal figures, frames, and diagrams of forces [abstract only]. Proceedings of the Royal Society of Edinburgh, 7:53–56, 1872. CODEN PRSEAE. ISSN 0080-4541 (print), 2053-5902 (electronic). Printed in full in [Cle72f].

ClerkMaxwell:1872:BSS

[Cle72b] James Clerk Maxwell. 5. On a bow seen on the surface of ice. Proceedings of the Royal Society of Edinburgh, 7:69, 1872. CODEN PRSEAE. ISSN 0080-4541 (print), 2053-5902 (electronic).

ClerkMaxwell:1872:GMDa

[Cle72c] James Clerk Maxwell. 6. On geometric mean distance. Proceedings of the Royal Society of Edinburgh, 7:613, 1872. CODEN PRSEAE. ISSN 0080-4541 (print), 2053-5902 (electronic).

ClerkMaxwell:1872:APF

[Cle72d] James Clerk Maxwell. Addition to a paper by Francis Deas on spectra formed by the passage of polarised light through double refracting crystals. Transactions of the Royal Society of Edinburgh, 26(1): 185–188, ???? 1872. CODEN TRSEAO. ISSN 0080-4568 (print),

2053-5945 (electronic). URL https://hdl.handle.net/2027/umn.
31951d00649878j?urlappend=%3Bseq=217. See [Dea72].

ClerkMaxwell:1872:CV

[Cle72e] James Clerk Maxwell. On colour vision. Proceedings of the Royal Institution of Great Britain, 6(??):260–271, ???? 1872. CODEN PIG-BAI. ISSN 0035-8959.

ClerkMaxwell:1872:RFFb

[Cle72f] James Clerk Maxwell. On reciprocal figures, frames, and diagrams of forces. Transactions of the Royal Society of Edinburgh, 26(1):1–40 + plates I–III, ???? 1872. CODEN TRSEAO. ISSN 0080-4568 (print),

2053-5945 (electronic). URL https://hdl.handle.net/2027/umn.
31951d00649878j?urlappend=%3Bseq=31.

ClerkMaxwell:1872:CTF

[Cle72g] James Clerk Maxwell. On the condition that, in the transformation of any figure by curvilinear coordinates in three dimensions, every angle in the new figure shall be equal to the corresponding angle in the original figure. Proceedings of the London Mathematical Society. First Series, s1-4:117–119, May 9, 1872. CODEN PLMTAL. ISSN 0024-6115 (print), 1460-244X (electronic). URL http://onlinelibrary.

wiley.com/doi/10.1112/plms/s1-4.1.117/full.

ClerkMaxwell:1872:GMDb

[Cle72h] James Clerk Maxwell. On the geometrical mean distance of two figures in a plane. Transactions of the Royal Society of Edinburgh, 26:729– 733, 1872. CODEN TRSEAO. ISSN 0080-4568 (print), 2053-5945 (electronic).

ClerkMaxwell:1872:IEC

[Cle72i] James Clerk Maxwell. On the induction of electric currents in an infinite plane sheet of uniform conductivity. Proceedings of the Royal Society of London, 20(??):160–168, ???? 1872. CODEN PRSLAZ. ISSN 0370-1662 (print), 2053-9126 (electronic).

ClerkMaxwell:1872:THa

[Cle72j] James Clerk Maxwell. Theory of Heat. D. Appleton and Co., New York, NY, USA, 1872. xii + 312 pp. LCCN ????

ClerkMaxwell:1872:THb

[Cle72k] James Clerk Maxwell. Theory of Heat. Longmans, Green, and Co., London, UK, second edition, 1872. xii + 312 pp. LCCN

???? URL https://hdl.handle.net/2027/hvd.32044009528449;
 https://hdl.handle.net/2027/mdp.39015005777670;   https:
//hdl.handle.net/2027/uc1.c039633434.

ClerkMaxwell:1872:THc

[Cle72l] James Clerk Maxwell. Theory of Heat. Longmans, Green, and Co., London, UK, third edition, 1872. xii + 313 pp. LCCN

????  URL https://babel.hathitrust.org/cgi/pt?id=mdp.
39015064437489%3Bview=1up%3Bseq=7;   https://hdl.handle.
net/2027/mdp.39015064437489.

ClerkMaxwell:1872:THd

[Cle72m] James Clerk Maxwell. Theory of Heat. Longmans, Green, and Co., London, UK, fourth edition, 1872. xii + 333 pp. LCCN ???? URL

https://hdl.handle.net/2027/umn.319510005598689.

ClerkMaxwell:1873:FSSa

[Cle73a] James Clerk Maxwell. 0n the final state of a system of molecules subject to forces of any kind. In Report of 1873 to the British Association, page ???? ????, ????, 1873.

ClerkMaxwell:1873:BREb

[Cle73b] James Clerk Maxwell. Book review: An essay on the mathematical principles of physics. By the Rev. James Challis, M.A., &c. Nature, 8(197):279–280, August 7, 1873. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.

com/nature/journal/v8/n197/pdf/008279a0.pdf.

ClerkMaxwell:1873:BREa

[Cle73c] James Clerk Maxwell. Book review: Elements of natural philosophy. By Professors Sir W. Thomson and P. G. Tait. Nature, 7 (178):399–400, March 27, 1873. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v7/n178/pdf/007399a0.pdf.

ClerkMaxwell:1873:BRR

[Cle73d] James Clerk Maxwell. Book review: Reprint of Papers on electrostatics and magnetism. By Sir W. Thomson. Nature, 7(169):218–221, January 23, 1873. CODEN NATUAS. ISSN 0028-0836 (print), 1476-

4687 (electronic). URL https://www.nature.com/nature/journal/
v7/n169/pdf/007218a0.pdf.

ClerkMaxwell:1873:CMK

[Cle73e] James Clerk Maxwell. Clerk-Maxwell’s kinetic theory of gases. Nature, 8(187):85, May 29, 1873. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.

com/nature/journal/v8/n187/pdf/008085a0.pdf.

ClerkMaxwell:1873:LDM

[Cle73f] James Clerk Maxwell. LX. A discourse on molecules. Philosophical Magazine (Series 4 (1851–1875)), 46(308):453–469, ???? 1873. ISSN

0031-8086. URL http://www.tandfonline.com/doi/abs/10.1080/
14786447308640975.

ClerkMaxwell:1873:MM

[Cle73g] James Clerk Maxwell. Matter and Motion. Society for Promoting Christian Knowledge, London, UK, 1873. ???? pp. LCCN ????

ClerkMaxwell:1873:M

[Cle73h] James Clerk Maxwell. Molecules. Nature, 8(204):437–441, September 25, 1873. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL https://www.nature.com/nature/journal/v8/
n204/pdf/008437a0.pdf.

ClerkMaxwell:1873:PCV

[Cle73i] James Clerk Maxwell. On a problem in the calculus of variations, in which the solution is discontinuous. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 2:294–295, 1873. CODEN PCPSA4. ISSN 0008-1981.

ClerkMaxwell:1873:LED

[Cle73j] James Clerk Maxwell. On Loschmidt’s experiments on diffusion in relation to the kinetic theory of gases. Nature, 8(198):298–300, August 14, 1873. CODEN NATUAS. ISSN 0028-0836 (print), 1476-

4687 (electronic). URL http://www.nature.com/nature/journal/
v8/n198/pdf/008298a0.pdf.

ClerkMaxwell:1873:ERB

[Cle73k] James Clerk Maxwell. On the elementary relations between electrical measurements. In Fleeming Jenkin, editor, Reports of the Committee on Electrical Standards, pages 59–96. [British Association??], London, UK, 1873. LCCN ????

ClerkMaxwell:1873:ETG

[Cle73l] James Clerk Maxwell. On the equilibrium of temperature of a gaseous column subjected to gravity. Nature, 8(208):527–528, October 23, 1873. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v8/n208/
pdf/008527b0.pdf.

ClerkMaxwell:1873:FSSb

[Cle73m] James Clerk Maxwell. On the final state of a system of molecules in motion subject to forces of any kind. Nature, 8(208):537–538, October 23, 1873. CODEN NATUAS. ISSN 0028-0836 (print), 1476-

4687 (electronic). URL http://www.nature.com/nature/journal/
v8/n208/pdf/008537a0.pdf.

ClerkMaxwell:1873:FLR

[Cle73n] James Clerk Maxwell. On the focal lines of a refracted pencil. Proceedings of the London Mathematical Society. First Series, s1-4:337– 341, April 10, 1873. CODEN PLMTAL. ISSN 0024-6115 (print),

1460-244X (electronic). URL http://onlinelibrary.wiley.com/
doi/10.1112/plms/s1-4.1.337/full.

ClerkMaxwell:1873:PEM

[Cle73o] James Clerk Maxwell. On the proof of the equations of motion of a connected system. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 2:292–294, 1873. CODEN PCPSA4. ISSN 0008-1981.

ClerkMaxwell:1873:TSE

[Cle73p] James Clerk Maxwell. On the theory of a system of electrified conductors, and other physical theories involving homogeneous quadratic functions. Proceedings of the London Mathematical Society. First Series, s1-4:334–336, April 10, 1873. CODEN PLMTAL. ISSN 0024-6115 (print), 1460-244X (electronic). URL http://onlinelibrary.

wiley.com/doi/10.1112/plms/s1-4.1.334/full.

ClerkMaxwell:1873:PRC

[Cle73q] James Clerk Maxwell. Preliminary report of the Committee on Siemens’s electrical pyrometer. Report of the British Association for the Advancement of Science, ??(??):134–135, ?? 1873. Read at Brighton, August 1872.

ClerkMaxwell:1873:QRP

[Cle73r] James Clerk Maxwell. Quaternions: [review of Philip Kelland and Peter Guthrie Tait’s Introduction to Quaternions, London (1873)]. Nature, 9(217):137–138, December 1873.

ClerkMaxwell:1873:SWF

[Cle73s] James Clerk Maxwell. Scientific worthies: [Faraday]. Nature, 8 (203):397–399, September 11, 1873. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.

com/nature/journal/v8/n203/pdf/008397a0.pdf.

ClerkMaxwell:1873:TEM

[Cle73t] James Clerk Maxwell. A Treatise on Electricity and Magnetism. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1873. xxxi + 1 + 464 + I–XIII (vol. 1), xxiii + 1 + 456 + XIV–XX + 32

(vol. 2) pp. URL https://hdl.handle.net/2027/uva.x002226063.

Preface dated: 1 February 1873.

ClerkMaxwell:1874:BRCb

[Cle74a] James Clerk Maxwell. Book review: Over de continuiteit van den gas- en vloeistofiocstand: Academisch proefschrift. Door Johannes Diderik van der Waals. (Dutch) [On the continuity of the gas and liquid state: Academic thesis]. Nature, 10(259):477–480, October 15, 1874. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL https://www.nature.com/nature/journal/v10/
n259/pdf/010477a0.pdf.

ClerkMaxwell:1874:BRCa

[Cle74b] James Clerk Maxwell. Book review: The Correlation of Physical Forces. Sixth edition. Nature, 10(251):302–304, August 20, 1874. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v10/n251/
pdf/010302a0.pdf.

ClerkMaxwell:1874:BRG

[Cle74c] James Clerk Maxwell. Book review: Grove’s Correlation of Physical Forces, sixth edition. Nature, 10(251):302–304, August 20, 1874. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL https://www.nature.com/nature/journal/v10/
n251/pdf/010302a0.pdf.

ClerkMaxwell:1874:BRP

[Cle74d] James Clerk Maxwell. Book review: Plateau on soap-bubbles: Statique expérimentale et théorique des Liquides soumis aux seules Forces moléculaires. Nature, 10(242):119–212, June 11, 1874. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https:/

/www.nature.com/nature/journal/v10/n242/pdf/010119a0.pdf.

ClerkMaxwell:1874:DRVb

[Cle74e] James Clerk Maxwell. Double refraction in a viscous fluid in motion. Annales de Chimie et de Physique, ??(??):??, ???? 1874. CODEN ACPHAA. ISSN 0365-1444 (print), 2437-3621 (electronic). Reprint of [Cle74f].

ClerkMaxwell:1874:DRVa

[Cle74f] James Clerk Maxwell. On double refraction in a viscous fluid in motion. Proceedings of the Royal Society of London, 22(??):46–47, ???? 1874. CODEN PRSLAZ. ISSN 0370-1662 (print), 2053-9126 (electronic).

ClerkMaxwell:1874:HCF

[Cle74g] James Clerk Maxwell. On Hamilton’s characteristic function for a narrow beam of light. Proceedings of the London Mathematical Society. First Series, s1-6:182–190, January 8, 1874. CODEN PLM-TAL. ISSN 0024-6115 (print), 1460-244X (electronic). URL http://

onlinelibrary.wiley.com/doi/10.1112/plms/s1-6.1.182/full.

ClerkMaxwell:1874:AKR

[Cle74h] James Clerk Maxwell. On the application of Kirchhoff’s rules for electric circuits to the solution of a geometrical problem. Nature, 10(255):411, September 17, 1874. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.

com/nature/journal/v10/n255/pdf/010410a0.pdf.

ClerkMaxwell:1874:RGO

[Cle74i] James Clerk Maxwell. On the relation of geometrical optics to other parts of mathematics and physics. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 2:338–340, 1874. CODEN PCPSA4. ISSN 0008-1981.

ClerkMaxwell:1874:SRC

[Cle74j] James Clerk Maxwell. Sixth report of the Committee for the purpose of investigating the rate of increase of underground temperature downwards in various localities, of dry land and under water. Report of the British Association for the Advancement of Science, ?? (??):252–256, ???? 1874.

ClerkMaxwell:1874:SET

[Cle74k] James Clerk Maxwell. Statique expérimentale et théorique des liquides soumis aux seules forces moléculaires. (French) [Experimental and theoretical statics of liquids subject only to molecular strengths]. Nature, 10(242):119–121, June 18, 1874. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v10/n242/pdf/010119a0.pdf.

ClerkMaxwell:1875:Aa

[Cle75a] James Clerk Maxwell. Atom. In Encyclopaedia Britannica, volume 3, pages 36–49. ????, ????, ninth edition, 1875. LCCN ????

ClerkMaxwell:1875:Ab

[Cle75b] James Clerk Maxwell. Attraction. In Encyclopaedia Britannica, volume 3, pages 63–65. ????, ????, ninth edition, 1875. LCCN ????

ClerkMaxwell:1875:AD

[Cle75c] James Clerk Maxwell. On action at a distance. Proceedings of the Royal Institution of Great Britain, 7(??):44–54, ???? 1875. CODEN PIGBAI. ISSN 0035-8959.

ClerkMaxwell:1875:AHC

[Cle75d] James Clerk Maxwell. On the application of Hamilton’s characteristic function to the theory of an optical instrument symmetrical about its axis. Proceedings of the London Mathematical Society. First Series, s1-6:117–122, April 8, 1875. CODEN PLMTAL. ISSN 0024-6115 (print), 1460-244X (electronic). URL http://onlinelibrary.

wiley.com/doi/10.1112/plms/s1-6.1.117/full.

ClerkMaxwell:1875:AKR

[Cle75e] James Clerk Maxwell. On the application of Kirchhoff’s rules for electric circuits to the solution of a geometrical problem. Report of the British Association for the Advancement of Science, ??(??):18–19, ???? 1875.

ClerkMaxwell:1875:DEMa

[Cle75f] James Clerk Maxwell. On the dynamical evidence of the molecular constitution of bodies. Nature, 11(279):357–359, March 4, 1875. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL https://www.nature.com/nature/journal/v11/
n279/pdf/011357a0.pdf.

ClerkMaxwell:1875:DEMb

[Cle75g] James Clerk Maxwell. On the dynamical evidence of the molecular constitution of bodies. Nature, 11(280):374–377, March 11, 1875. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL https://www.nature.com/nature/journal/v11/
n280/pdf/011374a0.pdf.

ClerkMaxwell:1875:DEMc

[Cle75h] James Clerk Maxwell. On the dynamical evidence of the molecular constitution of bodies. Journal of the Chemical Society, 13(??):493– 508, ???? 1875. CODEN JCSOA9. ISSN 0368-1769 (print), 2050-5574 (electronic).

ClerkMaxwell:1875:THa

[Cle75i] James Clerk Maxwell. Theory of Heat. Longmans, Green, and Co., London, UK, fourth edition, 1875. xii + 333 pp. LCCN ???? URL

https://hdl.handle.net/2027/nyp.33433090920426.

ClerkMaxwell:1875:THb

[Cle75j] James Clerk Maxwell. Theory of Heat. D. Appleton and Co., New York, NY, USA, fourth edition, 1875. xii + 333 pp. LCCN ????

ClerkMaxwell:1876:BRW

[Cle76a] James Clerk Maxwell. Book review: Whewell’s writings and correspondence: Isaac Todhunter. William Whewell, master of Trinity College, Cambridge. An account of his writings with selections from his literary and scientific correspondence. London, 1876. 2 vols. Nature, 14(349):206–208, July 6, 1876. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.

com/nature/journal/v14/n349/pdf/014206a0.pdf.

ClerkMaxwell:1876:BA

[Cle76b] James Clerk Maxwell. The British Association. Nature, 14(361): 479, September 28, 1876. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.com/

nature/journal/v14/n361/pdf/014476a0.pdf.

ClerkMaxwell:1876:CA

[Cle76c] James Clerk Maxwell. Capillary action. In Encyclopaedia Britannica, volume 5, pages 56–71. ????, ????, ninth edition, 1876. LCCN ????

ClerkMaxwell:1876:GCC

[Cle76d] James Clerk Maxwell. General considerations concerning scientific apparatus. In ????, editor, South Kensington Museum. Handbook to the special loan collection of scientific apparatus, pages 1–21. ????, London, UK, 1876.

ClerkMaxwell:1876:MM

[Cle76e] James Clerk Maxwell. Matter and Motion. Manuals of elementary science. Society for Promoting Christian Knowledge, London, UK, 1876. viii + 128 pp.

ClerkMaxwell:1876:MPI

[Cle76f] James Clerk Maxwell. Molecular physics — instruments connected with fluids. In ????, editor, South Kensington Museum. Handbook to the special loan collection of scientific apparatus, pages 87–92. ????, London, UK, 1876.

ClerkMaxwell:1876:PCV

[Cle76g] James Clerk Maxwell. On a problem in the calculus of variations in which the solution is discontinuous. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 2 (part xiv) (??):294–295, ???? 1876. CODEN PCPSA4. ISSN 0008-1981.

ClerkMaxwell:1876:BMD

[Cle76h] James Clerk Maxwell. On Bow’s method of drawing diagrams in graphical statics with illustrations from Peaucellier’s linkage. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 2 (part xvii):407–414, 1876. CODEN PCPSA4. ISSN 0008-1981.

ClerkMaxwell:1876:CME

[Cle76i] James Clerk Maxwell. On the centre of motion of the eye. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 2 (part xv)(??):365–366, ???? 1876. CODEN PCPSA4. ISSN 0008-1981.

ClerkMaxwell:1876:EHSa

[Cle76j] James Clerk Maxwell. On the equilibrium of heterogeneous substances. In ????, editor, [Proceedings of the South Kensington Conferences in Connexion with the Loan Collection of Scientific Apparatus], pages 144–150. ????, ????, 1876. Address [on Professor Willard Gibbs’ Thermodynamic formulation for coexistent phases] delivered on 24 May 1876.

ClerkMaxwell:1876:EHSb

[Cle76k] James Clerk Maxwell. On the equilibrium of heterogeneous substances. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 2:427–430, 1876. CODEN PCPSA4. ISSN 0008-1981. Address delivered on 24th May 1876 at the South Kensington Museum, Conferences held in connection with the Special Loan Collection of Scientific Apparatus.

ClerkMaxwell:1876:PEM

[Cle76l] James Clerk Maxwell. On the proof of the equations of motion of a connected system. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 2 (part xiv)(??):292–294, ???? 1876. CODEN PCPSA4. ISSN 0008-1981.

ClerkMaxwell:1876:SEP

[Cle76m] James Clerk Maxwell. On the solution of electrical problems by the transformation of conjugate functions. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 2 (part xiii)(??):242–243, ???? 1876. CODEN PCPSA4. ISSN 0008-1981.

ClerkMaxwell:1876:RAP

[Cle76n] James Clerk Maxwell. Recent advances in physical science. Nature, 13(337):461–463, April 13, 1876. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.

com/nature/journal/v13/n337/pdf/013461b0.pdf.

ClerkMaxwell:1876:TEI

[Cle76o] James Clerk Maxwell. Theory of electrical induction. Nature, 14 (341):27–28, May 11, 1876. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.com/

nature/journal/v14/n341/pdf/014027b0.pdf.

ClerkMaxwell:1876:DGD

[Cle76p] James Clerk Maxwell. Ueber die Diffusion der Gase durch absorbierende Substanzen. (German) [On the diffusion of gases by absorbing substances]. Nature, 14(341):24–25, May 11, 1876. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL https://www.nature.com/nature/journal/v14/n341/pdf/
014024a0.pdf.

ClerkMaxwell:1877:BRT

[Cle77a] James Clerk Maxwell. Book review: A Treatise on the Kinetic Theory of Gases, by William Henry Watson. Nature, 16(404):242–246, July 26, 1877. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL https://www.nature.com/nature/journal/v16/
n404/pdf/016242a0.pdf.

ClerkMaxwell:1877:CB

[Cle77b] James Clerk Maxwell. Constitution of bodies. In Encyclopaedia Britannica, volume 6, pages 310–313. ????, ????, ninth edition, 1877. LCCN ????

ClerkMaxwell:1877:Da

[Cle77c] James Clerk Maxwell. Diagrams. In Encyclopaedia Britannica, volume 7, pages 149–153. ????, ????, ninth edition, 1877. LCCN ????

ClerkMaxwell:1877:Db

[Cle77d] James Clerk Maxwell. Diffusion. In Encyclopaedia Britannica, volume 7, pages 214–221. ????, ????, ninth edition, 1877. LCCN ????

ClerkMaxwell:1877:ESW

[Cle77e] James Clerk Maxwell. Editorial: Scientific worthies. Nature, 15 (384):389–391, March 8, 1877. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.

com/nature/journal/v15/n384/pdf/015389a0.pdf.

ClerkMaxwell:1877:PTA

[Cle77f] James Clerk Maxwell. On a paradox in the theory of attraction. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 3:34–39, 1877. CODEN PCPSA4. ISSN 0008-1981.

ClerkMaxwell:1877:AMI

[Cle77g] James Clerk Maxwell. On approximate multiple integration between limits of summation. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 3:39–47, 1877. CODEN PCPSA4. ISSN 0008-1981.

ClerkMaxwell:1877:EHS

[Cle77h] James Clerk Maxwell. On the equilibrium of heterogeneous substances. In South Kensington Museum. Conferences held in connection with the Special Loan Collection of Scientific Apparatus. 1876. (Volume 1.), pages 145–150. Chapman and Hall [for the Lords of the Committee of Council on Education], Piccadilly, London, UK, 1877.

ClerkMaxwell:1877:PBL

[Cle77i] James Clerk Maxwell. On the protection of buildings from lightning. Report of the British Association for the Advancement of Science, ?? (??):43–45, ???? 1877.

ClerkMaxwell:1877:SWH

[Cle77j] James Clerk Maxwell. Scientific worthies: Hermann Ludwig Ferdinand Helmholtz. Nature, 15(384):389–391, March 3, 1877. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL https://www.nature.com/nature/journal/v15/n384/pdf/
015389a0.pdf.

ClerkMaxwell:1877:TWJ

[Cle77k] James Clerk Maxwell. Theorie der Wärme von J. C. Maxwell. Maruschke & Berendt, Breslau, Germany, 1877. xii + 324 pp. LCCN ???? German translation by F. Auerbach of [Cle75i].

ClerkMaxwell:1877:TH

[Cle77l] James Clerk Maxwell. Theory of Heat. ????, ????, fifth edition, 1877. xii + 333 pp. LCCN ????

ClerkMaxwell:1878:BRP

[Cle78a] James Clerk Maxwell. Book review: Paradoxical Philosophy, A Sequel to the “Unseen Universe”. Nature, 19(477):141–143, December 19, 1878. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL https://www.nature.com/nature/journal/v19/
n477/pdf/019141a0.pdf.

ClerkMaxwell:1878:BRTa

[Cle78b] James Clerk Maxwell. Book review: Tait’s Thermodynamics. Nature, 17(431):257–259, January 31, 1878. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.

com/nature/journal/v17/n431/pdf/017257a0.pdf.

ClerkMaxwell:1878:BRTb

[Cle78c] James Clerk Maxwell. Book review: Tait’s Thermodynamics. Nature, 17(432):278–280, February 7, 1878. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.

com/nature/journal/v17/n432/pdf/017278a0.pdf.

ClerkMaxwell:1878:E

[Cle78d] James Clerk Maxwell. Ether. In Encyclopaedia Britannica, volume 8, pages 568–572. ????, ????, ninth edition, 1878. LCCN ????

ClerkMaxwell:1878:MM

[Cle78e] James Clerk Maxwell. Matter and Motion. D. Van Nostrand, New York, NY, USA, 1878. viii + 9–224 pp. URL https:

//hdl.handle.net/2027/hvd.32044019356450;   https://hdl.
handle.net/2027/mdp.39015031226510;   https://hdl.handle.
net/2027/mdp.39015049745477.

ClerkMaxwell:1878:SRGb

[Cle78f] James Clerk Maxwell. On stresses in rarefied gases arising from inequalities of temperature. Proceedings of the Royal Society of London, 27:304–308, 1878. CODEN PRSLAZ. ISSN 0370-1662 (print), 2053-9126 (electronic).

ClerkMaxwell:1878:SRGa

[Cle78g] James Clerk Maxwell. On stresses in rarefied gases arising from inequalities of temperature [abstract only]. Nature, 18(445):54–55, September 9, 1878. CODEN NATUAS. ISSN 0028-0836 (print), 1476-

4687 (electronic). URL https://www.nature.com/nature/journal/
v18/n445/pdf/018054a0.pdf. Full paper in [Cle79g].

ClerkMaxwell:1878:ECL

[Cle78h] James Clerk Maxwell. On the electrical capacity of a long narrow cylinder, and of a disk of sensible thickness. Proceedings of the London Mathematical Society. First Series, s1-9:94–102, March 14, 1878. CODEN PLMTAL. ISSN 0024-6115 (print), 1460-

244X (electronic). URL http://onlinelibrary.wiley.com/doi/
10.1112/plms/s1-9.1.94/full.

ClerkMaxwell:1878:Tb

[Cle78i] James Clerk Maxwell. The Rede lecture: The telephone. Nature, 18 (449):159–163, June 6, 1878. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.com/

nature/journal/v18/n449/pdf/018159a0.pdf. Sir Robert Rede

lecture.

ClerkMaxwell:1878:Ta

[Cle78j] James Clerk Maxwell. The telephone. The Electrician: the Oldest Weekly Illustrated Journal of Electrical Engineering, Industry, Science and Finance, 1(??):16–17, 29–??, June 1, 1878. CODEN ELE-TAU. ISSN 0367-0805. Sir Robert Rede lecture.

ClerkMaxwell:1878:TWG

[Cle78k] James Clerk Maxwell. Theorie der Wärme. (German) [Theory of Heat]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1878. German translation by F. Neesen of [Cle72l].

ClerkMaxwell:1879:BRG

[Cle79a] James Clerk Maxwell. Book review: (Guthrie’s physics) Practical physics, molecular physics, and sound. By Frederick Guthrie. Nature, 19(484):311–312, February 6, 1879. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.

com/nature/journal/v19/n484/pdf/019311a0.pdf.

ClerkMaxwell:1879:BRT

[Cle79b] James Clerk Maxwell. Book review: Thomson and Tait’s Natural Philosophy. Nature, 20(505):213–216, July 3, 1879. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.

nature.com/nature/journal/v20/n505/pdf/020213a0.pdf.

ClerkMaxwell:1879:ERH

[Cle79c] James Clerk Maxwell. The electrical researches of the Honourable Henry Cavendish, F.R.S.: written between 1771 and 1781, edited from the original manuscripts in the possession of the Duke of Devonshire, K.G. Cambridge University Press, Cambridge, UK, 1879. lxvi +

454 pp. LCCN ???? URL http://catalog.hathitrust.org/api/
volumes/oclc/320238869.html.

ClerkMaxwell:1879:Fa

[Cle79d] James Clerk Maxwell. Faraday. In Encyclopaedia Britannica, volume 9, pages 29–31. ????, ????, ninth edition, 1879. LCCN ????

ClerkMaxwell:1879:Fb

[Cle79e] James Clerk Maxwell. Faraday. In Encyclopaedia Britannica, volume 10, pages 173–175. ????, ????, ninth edition, 1879. LCCN ????

ClerkMaxwell:1879:BTA

[Cle79f] James Clerk Maxwell. On Boltzmann’s theorem on the average distribution of energy in a system of material points. Transactions of the Cambridge Philosophical Society, 12(??):549–570, ???? 1879.

ClerkMaxwell:1879:SRG

[Cle79g] James Clerk Maxwell. On the stresses in rarefied gases arising from inequalities in temperature. Philosophical transactions of the Royal Society of London, 170 (part i)(??):231–256, ???? 1879. CODEN PTRSAV. ISSN 0370-2316 (print), 2053-9207 (electronic). Appendix (pages 249–256) added May 1879. See [BE69] for the history of the explanation in this article of how a radiometer works.

ClerkMaxwell:1879:PPM

[Cle79h] James Clerk Maxwell. Practical physics, molecular physics, and sound. Nature, 19(484):311–312, February 6, 1879. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https:/

/www.nature.com/nature/journal/v19/n484/pdf/019311a0.pdf.

ClerkMaxwell:1879:RSB

[Cle79i] James Clerk Maxwell. Reports on special branches of science. Report of the British Association for the Advancement of Science, pages 37– 38, 1879. In Report of the Committee appointed for the purpose of endeavouring to procure reports on the progress of the chief branches of mathematics and physics.

ClerkMaxwell:1879:SBG

[Cle79j] James Clerk Maxwell. Substanz und Bewegung. (German) [Matter and Motion]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1879. xiv + 144 pp. LCCN QC73 M53 1879. URL http://pi.lib.

uchicago.edu/1001/dig/pres/2015-0839.

ClerkMaxwell:1880:HA

[Cle80a] James Clerk Maxwell. Harmonic analysis. In Encyclopaedia Britannica, volume 11, pages 481–482. ????, ????, ninth edition, 1880. LCCN ????

ClerkMaxwell:1880:PTA

[Cle80b] James Clerk Maxwell. On a paradox in the theory of attraction. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 3 (part i)(??):34–39, ???? 1880. CODEN PCPSA4. ISSN 0008-1981.

ClerkMaxwell:1880:AMI

[Cle80c] James Clerk Maxwell. On approximate multiple integration between limits by summation. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 3 (part i)(??):39–49, 1880. CODEN PCPSA4. ISSN 0008-1981.

ClerkMaxwell:1880:SAP

[Cle80d] James Clerk Maxwell. Societies and academies: On a possible mode of detecting a motion of the solar system through the luminiferous ether. Nature, 21(535):314–315, January 29, 1880. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.

nature.com/nature/journal/v21/n535/pdf/021314a0.pdf. Letter

to Mr D. P. Todd, Director of the Nautical Almanac Office, Washington, U.S. Communicated by Prof. Stoke, Sec. R.S. Letter is dated: Cavendish Laboratory, Cambridge, 19th March, 1879.

ClerkMaxwell:1880:TH

[Cle80e] James Clerk Maxwell. Theory of Heat. Longmans, Green, and Co., London, UK, sixth edition, 1880. xii + 333 pp. LCCN ???? URL

https://hdl.handle.net/2027/uc1.c047124312.

ClerkMaxwell:1881:ETE

[Cle81a] James Clerk Maxwell. An Elementary Treatise on Electricity. Clarendon Press, New York, NY, USA, 1881. xvi + 208 pp. LCCN ????

URL https://hdl.handle.net/2027/hvd.32044080810559. Edited

by William Garnett.

ClerkMaxwell:1881:SBG

[Cle81b] James Clerk Maxwell. Substanz und Bewegung. (German) [Matter and Motion]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1881. ???? pp. LCCN QC73 M53 1881. URL http://pi.lib.

uchicago.edu/1001/dig/pres/2015-0839.

ClerkMaxwell:1881:TEM

[Cle81c] James Clerk Maxwell. A Treatise on Electricity and Magnetism. Clarendon Press, New York, NY, USA, second edition, 1881. xiv + 493 (vol. 1), xiv + 471 (vol. 2) pp. URL https://

hdl.handle.net/2027/hvd.yl11wt;   https://hdl.handle.net/

2027/hvd.yl11wu. Edited by William Davidson Niven.

ClerkMaxwell:1882:TRA

[Cle82a] James Clerk Maxwell. Are there real analogies in nature? In The Life of James Clerk Maxwell: with a selection from his correspondence and occasional writings, and a sketch of his contributions to science [CG82], pages 235–244. LCCN QC16.M4 C2 1882. URL https://

archive.org/details/lifejamesclerkm01garngoog;   https://
www.biodiversitylibrary.org/item/71775.

ClerkMaxwell:1882:DPP

[Cle82b] James Clerk Maxwell. Does the progress of physical science tend to give any advantage to the opinion of necessity (or determinism) over that of the contingency of events and the freedom of the will?’ (science and free will.). In The Life of James Clerk Maxwell: with a selection from his correspondence and occasional writings, and a sketch of his contributions to science [CG82], pages 434–444. LCCN QC16.M4 C2 1882. URL https:/

/archive.org/details/lifejamesclerkm01garngoog;   https:/
/www.biodiversitylibrary.org/item/71775.

ClerkMaxwell:1882:MAP

[Cle82c] James Clerk Maxwell. On modified aspects of pain. In The Life of James Clerk Maxwell: with a selection from his correspondence and occasional writings, and a sketch of his contributions to science [CG82], pages 444–452. LCCN QC16.M4 C2 1882. URL https://

archive.org/details/lifejamesclerkm01garngoog;   https://
www.biodiversitylibrary.org/item/71775.

ClerkMaxwell:1882:P

[Cle82d] James Clerk Maxwell. Psychophysik. In The Life of James Clerk Maxwell: with a selection from his correspondence and occasional writings, and a sketch of his contributions to science [CG82], pages 452–463. LCCN QC16.M4 C2 1882. URL https:/

/archive.org/details/lifejamesclerkm01garngoog;   https:/
/www.biodiversitylibrary.org/item/71775.

ClerkMaxwell:1883:LEM

[Cle83a] James Clerk Maxwell. Lehrbuch der Electricität und des Magnetismus. (German) [Treatise on Electricity and Magnetism]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1883. xx + 555 (vol. 1), xiv + 637 (vol. 2)

pp. URL https://hdl.handle.net/2027/hvd.32044079972188;
 https://hdl.handle.net/2027/hvd.32044091959031. Two vol-

umes. Translation to German by Dr. B. [or P.??] Weinstein.

ClerkMaxwell:1883:TH

[Cle83b] James Clerk Maxwell. Theory of Heat. ????, ????, seventh edition, 1883. ???? pp. LCCN ????

ClerkMaxwell:1884:TEM

[Cle84] James Clerk Maxwell. Traité d’électricité et de magnetisme. (French) [Treatise on Electricity and Magnetism]. ????, ????, 1884. ???? pp. LCCN ???? French translation by Gustave Richard, with preface by William Garnett on the Professor Maxwell’s works on electricity.

ClerkMaxwell:1885:PS

[Cle85a] James Clerk Maxwell. Physical sciences. In Encyclopaedia Britannica, volume ??, page ?? ????, ????, ninth edition, 1885. LCCN ????

ClerkMaxwell:1885:TH

[Cle85b] James Clerk Maxwell. Theory of Heat. ????, ????, eighth edition, 1885. ???? pp. LCCN ????

ClerkMaxwell:1885:TEM

[Cle85c] James Clerk Maxwell. Traité d’électricité et de magnetisme. (French) [Treatise on Electricity and Magnetism], volume 1. Gauthier-Villars, Paris, France, 1885. xx + 578 pp. LCCN ???? URL https://hdl.

handle.net/2027/mdp.39015030994100;   https://hdl.handle.

net/2027/mdp.39015030994118. First of two volumes. French translation by G. Séligmann-Lui of the second edition. With notes and clarifications by MM. Cornu, Potier, adn Sarrau, Professors at l’École Polytechnique.

ClerkMaxwell:1888:ETE

[Cle88a] James Clerk Maxwell. An Elementary Treatise on Electricity. ????, London, UK, second edition, 1888. ???? pp. LCCN ???? Edited by William Garnett.

ClerkMaxwell:1888:MM

[Cle88b] James Clerk Maxwell. Matter and Motion. Cambridge University Press, Cambridge, UK, 1888. ISBN 0-511-70932-3. ???? pp.

ClerkMaxwell:1888:TH

[Cle88c] James Clerk Maxwell. Theory of Heat. Longmans, Green, and Co., London, UK, ninth edition, 1888. xii + 333 pp. LCCN ???? URL

https://hdl.handle.net/2027/mdp.39015002183815.

ClerkMaxwell:1889:TEM

[Cle89] James Clerk Maxwell. Traité d’électricité et de magnetisme. (French) [Treatise on Electricity and Magnetism], volume 2. Gauthier-Villars, Paris, France, 1889. 651 pp. LCCN ???? URL https://hdl.

handle.net/2027/mdp.39015030994100;   https://hdl.handle.
net/2027/mdp.39015030994118. Second of two volumes. French

translation by G. Séligmann-Lui of the second edition. With notes and clarifications by MM. Cornu, Potier, adn Sarrau, Professors at l’École Polytechnique.

ClerkMaxwell:1890:E

[Cle90a] James Clerk Maxwell. Ether. In Niven [Niv90a], pages 2:763–2:775.

URL http://www.archive.org/download/scientificpapers01maxwuoft/
scientificpapers01maxwuoft.pdf;   http://www.biodiversitylibrary.
org/item/59553;   https://hdl.handle.net/2027/umn.31951p011043735.

ClerkMaxwell:1890:IDT

[Cle90b] James Clerk Maxwell. Illustrations of the dynamical theory of gases. In Niven [Niv90a], pages 1:377–1:409. URL http://www.archive.

org/download/scientificpapers01maxwuoft/scientificpapers01maxwuoft.
pdf;   http://www.biodiversitylibrary.org/item/59553;   https:
//hdl.handle.net/2027/umn.31951p011043735.

ClerkMaxwell:1890:ILA

[Cle90c] James Clerk Maxwell. Inaugural lecture, Aberdeen, 3 November 1856. In Niven [Niv90a], pages 1:430–?? URL http://www.archive.org/

download/scientificpapers01maxwuoft/scientificpapers01maxwuoft.
pdf;   http://www.biodiversitylibrary.org/item/59553;   https:
//hdl.handle.net/2027/umn.31951p011043735. Cited elsewhere as

being in this volume, but not found there. Page 430 is in the middle of [Cle60f].

ClerkMaxwell:1890:PP

[Cle90d] James Clerk Maxwell. Paradoxical philosophy. In Niven [Niv90a],

pages 2:756–2:762. URL http://www.archive.org/download/
scientificpapers01maxwuoft/scientificpapers01maxwuoft.pdf;
 http://www.biodiversitylibrary.org/item/59553;   https:/
/hdl.handle.net/2027/umn.31951p011043735.

ClerkMaxwell:1890:WWC

[Cle90e] James Clerk Maxwell. Whewell’s writings and correspondence. In Niven [Niv90a], pages 2:528–2:532. URL http://www.archive.org/

download/scientificpapers01maxwuoft/scientificpapers01maxwuoft.
pdf;   http://www.biodiversitylibrary.org/item/59553;   https:
//hdl.handle.net/2027/umn.31951p011043735.

ClerkMaxwell:1891:CLE

[Cle91a] James Clerk Maxwell. La chaleur. Leçons élémentaires sur la thermométrie, la calorimétrie, la thermodynamique. (French) [Heat. Basic lessons on thermometry, calorimetry, thermodynamics]. ????, Paris, France, 1891. 432 pp. Traduction d’après la 8^{e} édition anglaise par G. Mouret, précédée d’une préface par A. Potier.

ClerkMaxwell:1891:TH

[Cle91b] James Clerk Maxwell. Theory of Heat. ????, ????, tenth edition, 1891. xiv + 343 pp. LCCN ????

ClerkMaxwell:1891:THT

[Cle91c] James Clerk Maxwell. Theory of Heat. Longmans, Green and Co., London, UK, tenth edition, 1891. With corrections and additions by Lord Rayleigh.

ClerkMaxwell:1892:MM

[Cle92a] James Clerk Maxwell. Matter and Motion. D. Van Nostrand, New York, NY, USA, 1892. viii + 9–224 pp.

ClerkMaxwell:1892:TEM

[Cle92b] James Clerk Maxwell. A Treatise on Electricity and Magnetism. Clarendon Press, New York, NY, USA, third edition, 1892. xxxii + 557 (vol. 1), xxiv + 515 (vol. 1) pp. URL https://hdl.

handle.net/2027/hvd.32044015466394;   https://hdl.handle.

net/2027/hvd.hs1nk7. Two volumes. Edited by J. J. Thomson.

ClerkMaxwell:1894:TH

[Cle94] James Clerk Maxwell. Theory of Heat. Longmans, Green, and Co., London, UK, new edition, 1894. xiv + 348 pp. LCCN ????

URL https://hdl.handle.net/2027/mdp.39015011543942. Based

on the Tenth Edition (1891). With corrections and additions by Lord Rayleigh, M.A., D.C.L., LL.D., Secretary of the Royal Society, Professor of Natural Philosophy in the Royal Institution, and late Professor of Experimental Physics in the University of Cambridge.

ClerkMaxwell:1895:FKG

[Cle95] James Clerk Maxwell. Ueber Faraday’s Kraftlinien (1855). (German) [On Faraday’s lines of force], volume 69 of Ostwald’s Klassiker der exacten Wissenschaften. Engelmann, Leipzig, Germany, 1895. 130

pp. URL https://hdl.handle.net/2027/hvd.32044021161674;
https://hdl.handle.net/2027/osu.32435011243508. Edited by

Ludwig Boltzmann.

ClerkMaxwell:1896:AAS

[Cle96] James Clerk Maxwell. Análisis armonico. (Spanish) [Harmonic analysis]. Archivo de Mat., 1:92–97, 1896.

ClerkMaxwell:1897:TH

[Cle97] James Clerk Maxwell. Theory of Heat. Longmans, Green, and Co., London, UK, 1897. xiv + 348 pp. LCCN ???? URL https:

//hdl.handle.net/2027/chi.11552642. Based on the Tenth Edi-

tion (1891). With corrections and additions by Lord Rayleigh, M.A., D.C.L., LL.D., Secretary of the Royal Society, Professor of Natural Philosophy in the Royal Institution, and late Professor of Experimental Physics in the University of Cambridge.

ClerkMaxwell:18xx:NEH

[Clexxa] James Clerk Maxwell. Notes on the equilibrium of heterogeneous substances: “on thermodynamics of solutions of variable strength”. Reprinted in [Gar69]., 18xx.

ClerkMaxwell:18xx:RBS

[Clexxb] James Clerk Maxwell. Rigid body sings. Web site., 18xx. URL http:

//ww3.haverford.edu/physics-astro/songs/rigid.htm.

ClerkMaxwell:1902:TH

[Cle02] James Clerk Maxwell. Theory of Heat. Longmans, Green, and Co., London, UK, new edition, 1902. xiv + 348 pp. LCCN ???? URL

https://hdl.handle.net/2027/uc1.b3141125. Based on the Tenth

Edition (1891). With corrections and additions by Lord Rayleigh, M.A., D.C.L., LL.D., Secretary of the Royal Society, Professor of Natural Philosophy in the Royal Institution, and late Professor of Experimental Physics in the University of Cambridge.

ClerkMaxwell:1904:TH

[Cle04a] James Clerk Maxwell. Theory of Heat. Longmans, Green, and Co., London, UK, new edition, 1904. xiv + 348 pp. LCCN ????

URL https://hdl.handle.net/2027/nyp.33433069098410. Based

on the Tenth Edition (1891). With corrections and additions by Lord Rayleigh, M.A., D.C.L., LL.D., Secretary of the Royal Society, Professor of Natural Philosophy in the Royal Institution, and late Professor of Experimental Physics in the University of Cambridge.

ClerkMaxwell:1904:TEM

[Cle04b] James Clerk Maxwell. A Treatise on Electricity and Magnetism. Clarendon Press, New York, NY, USA, third edition, 1904. xxxii + 553 (vol. 1), xxiv + 525 (vol. 2) pp. LCCN ???? URL https://hdl.

handle.net/2027/hvd.32044018703090;   https://hdl.handle.
net/2027/hvd.32044018703108.

ClerkMaxwell:1907:LTG

[Cle07] James Clerk Maxwell. [letter on a theory of gases]. In Larmor [Lar07], pages 2:8–2:11. LCCN QC16.S8 A2. This letter of May 1859 was inspired by a February 1859 English translation [Cla59] of an 1858 paper in German [Cla58], and for a few years, Clerk Maxwell and Clausius competed in improving their kinetic theories of gases.

ClerkMaxwell:1908:LEH

[Cle08a] James Clerk Maxwell. LXVIII. On the equilibrium of heterogeneous substances (1876). Philosophical Magazine (Series 6 (1901 - 1925)), 16(95):818–824, ???? 1908. CODEN PHMAA4. ISSN 1941-5982 (print), 1941-5990 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1080/14786441108636559.

ClerkMaxwell:1908:TH

[Cle08b] James Clerk Maxwell. Theory of Heat. ????, ????, ???? edition, 1908. ???? pp. LCCN ????

ClerkMaxwell:1910:D

[Cle10] James Clerk Maxwell. Diagram. In Encyclopaedia Britannica, volume 9, pages 146–149. ????, ????, eleventh edition, 1910. LCCN ????

ClerkMaxwell:1911:CA

[Cle11] James Clerk Maxwell. Capillary action. In Encyclopaedia Britannica, volume 5, pages 256–275. ????, ????, eleventh edition, 1911. LCCN ????

ClerkMaxwell:1912:FKG

[Cle12] James Clerk Maxwell. Ueber Faraday’s Kraftlinien. (German) [On Faraday’s lines of force], volume 69 of Ostwald’s Klassiker der exacten Wissenschaften. Engelmann, Leipzig, Germany, second edition, 1912. 130 pp. Second edition editing by Ludwig Boltzmann. German translation of [Cle58a].

ClerkMaxwell:1915:AJC

[Cle15] James Clerk Maxwell. Auszüge aus James Clerk Maxwells Elektrizität und magnetismus. (German) [Excerpts from James Clerk Maxwell’s Electricity and Magnetism]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1915. xxxi + 1 + 182 pp. LCCN ???? URL

https://hdl.handle.net/2027/mdp.39015067270093. Translation

to German by Hilde Barkhausen. Edited by Fritz Emde.

ClerkMaxwell:1919:LLT

[Cle19] James Clerk Maxwell. Lecture to a lady on Thomson’s reflecting galvanometer. In Lectures on ten British physicists of the Nineteenth Century [Mac19], page 61. LCCN QC15 .M3. URL https://

archive.org/stream/lecturesontenbri00macfrich/lecturesontenbri00macfrich_
djvu.txt.

ClerkMaxwell:1920:MMa

[Cle20a] James Clerk Maxwell. Matter and Motion. Society for Promoting Christian Knowledge, London, UK, 1920. xv + 163 pp.

LCCN QA846 M38 1920. URL https://hdl.handle.net/2027/
mdp.39015064574646.

ClerkMaxwell:1920:MMb

[Cle20b] James Clerk Maxwell. Matter and Motion. MacMillan Publishing Company, New York, NY, USA, 1920. xv + 163 pp. LCCN QA846 M38 1920. Reprinted with notes and appendices by Sir Joseph Larmor. (Society for Promoting Christian Knowledge).

ClerkMaxwell:1920:MMc

[Cle20c] James Clerk Maxwell. Matter and Motion. Dover Publications, Inc., New York, NY, USA, 1920. 163 pp. LCCN QA846 M38 1920.

ClerkMaxwell:1921:SPH

[Cle21a] James Clerk Maxwell, editor. The scientific papers of the Honourable Henry Cavendish, F.R.S.: Volume I: The electrical researches. Cambridge University Press, Cambridge, UK, 1921. xxvii + 452 pp. LCCN Q113 .C38 v.1. Revised by Sir Joseph Larmor, F.R.S., M.P.. First published in 1879.

ClerkMaxwell:1921:TH

[Cle21b] James Clerk Maxwell. Theory of Heat. Longmans, Green, and Co., London, UK, 1921. xiv + 348 pp. LCCN ???? URL https:

//hdl.handle.net/2027/uc1.b5300289. Based on the Tenth Edi-

tion (1891). With corrections and additions by Lord Rayleigh, M.A., D.C.L., LL.D., Secretary of the Royal Society, Professor of Natural Philosophy in the Royal Institution, and late Professor of Experimental Physics in the University of Cambridge.

ClerkMaxwell:1925:MM

[Cle25] James Clerk Maxwell. Matter and Motion. Sheldon Press, London, UK, 1925. xv + 163 pp. LCCN QA846.M5 1925. Reprinted with notes and appendices by Sir Joseph Larmor.

ClerkMaxwell:1927:DTE

[Cle27a] James Clerk Maxwell. A dynamical theory of the electromagnetic field [abstract only]. Nature, 119(2986):125–127, January 22, 1927. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL https://www.nature.com/nature/journal/v119/
n2986/pdf/119125a0.pdf.

ClerkMaxwell:1927:SP

[Cle27b] James Clerk Maxwell. The scientific papers. J. Hermann, Paris, France, 1927. xxxii + 607 (vol. I), viii + 806 (vol. II) pp. Edited by W. D. Niven. Vol. I, Vol. II. Photographic reprint by arrangement with the Cambridge University Press.

ClerkMaxwell:1951:SP

[Cle51] James Clerk Maxwell. The scientific papers. Dover Publications, Inc., New York, NY, USA, 1951. xxix + 806 pp.

ClerkMaxwell:1952:MM

[Cle52] James Clerk Maxwell. Matter and Motion. Dover Publications, Inc., New York, NY, USA, 1952. xiii + 163 + 1 pp. LCCN QA846

.M38. URL https://hdl.handle.net/2027/umn.31951000528326w.

Reprinted with notes and appendices by Sir Joseph Larmor.

ClerkMaxwell:1953:TEM

[Cle53] James Clerk Maxwell. A Treatise on Electricity and Magnetism. Academic Reprints, Stanford, CA, USA, third edition, 1953. xxxiv + 533 (vol. 1), xxiv + 515 (vol. 2) pp. URL https://hdl.

handle.net/2027/inu.30000117692750;   https://hdl.handle.
net/2027/inu.32000004186518.

ClerkMaxwell:1954:TEM

[Cle54] James Clerk Maxwell. A Treatise on Electricity and Magnetism. Dover Publications, Inc., New York, NY, USA, third edition, 1954. xxxii + 506 + 13 (vol. I), xxiv + 500 + 7 (vol. II) pp. URL https://

hdl.handle.net/2027/mdp.39015012340405. Unabridged third edi-

tion, with two original volumes bound as one.

ClerkMaxwell:1955:TEM

[Cle55] James Clerk Maxwell. A Treatise on Electricity and Magnetism. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, third edition, 1955. ???? pp. LCCN ???? URL https://hdl.handle.

net/2027/uc1.b4045360;   https://hdl.handle.net/2027/uc1.
b4045361.

ClerkMaxwell:1967:ERH

[Cle67] James Clerk Maxwell, editor. The electrical researches of the Honourable Henry Cavendish, F.R.S.: written between 1771 and 1781, edited from the original manuscripts in the possession of the Duke of Devonshire, K.G. Frank Cass, London, UK, 1967. 66 + 454 pp. LCCN ???? Reprint of [Cle79c].

ClerkMaxwell:1970:TH

[Cle70] James Clerk Maxwell. Theory of Heat. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, third edition, 1970. ISBN 0-8371-4097-8. xii + 313 pp. LCCN QC255 .M4 1970. URL https:/

/hdl.handle.net/2027/ien.35556027873199.

ClerkMaxwell:1979:JCM

[Cle79] James Clerk Maxwell. James Clerk Maxwell’s inaugural lecture at King’s College London. American Journal of Physics, 47(11):928– 933, November 1979. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

ClerkMaxwell:1982:DTE

[Cle82] James Clerk Maxwell. A Dynamical Theory of the Electromagnetic Field. Scottish Academic Press, Edinburgh, UK, 1982. ISBN 0-7073-0324-9 (paperback). xiii + 103 pp. LCCN QC670 .M38

1982. URL https://hdl.handle.net/2027/mdp.39015017148431;
 https://hdl.handle.net/2027/mdp.39076002001753. With an

appreciation by Albert Einstein. Edited and introduced by Thomas F. Torrance.

ClerkMaxwell:1984:LWPa

[Cle84a] James Clerk Maxwell. Lectures to women on physical science I. Physics Bulletin, 35(5):179, May 1984. CODEN PHSBB4. ISSN 0031-

9112. URL http://stacks.iop.org/0031-9112/35/i=5/a=005.

ClerkMaxwell:1984:LWPb

[Cle84b] James Clerk Maxwell. Lectures to women on physical science II. Physics Bulletin, 35(6):219, 1984. CODEN PHSBB4. ISSN 0031-

9112. URL http://stacks.iop.org/0031-9112/35/i=6/a=003.

ClerkMaxwell:1984:QM

[Cle84c] James Clerk Maxwell. Quotations from Maxwell. In Berger [Ber84], pages 1–9. ISBN 0-444-86707-4 (Elsevier). LCCN QC669 .J17 1984.

ClerkMaxwell:1986:PKG

[Cle86] James Clerk Maxwell. Über physikalische Kraftlinien. (German) [On physical lines of force]. Wissenschaftliche Buchgesellschaft, Darmstadt, West Germany, 1986. ISBN 3-534-07019-4. viii + 146 pp. URL

https://hdl.handle.net/2027/uc1.%24b564335. Translated from

the English by Ludwig Boltzmann.

ClerkMaxwell:1996:MM

[Cle96] James Clerk Maxwell. Matter and motion. In Pyle [Pyl96], chapter 7, page ?? ISBN 0-415-13267-3 (set). LCCN QA9 .C63 1996. URL

https://hdl.handle.net/2027/mdp.39015037757609. 12 volumes.

ClerkMaxwell:1997:MM

[Cle97] James Clerk Maxwell. Matter and Motion. ElecBook, ????, 1997. ISBN 1-84327-116-8. ???? pp.

ClerkMaxwell:1998:TEMa

[Cle98a] James Clerk Maxwell. A Treatise on Electricity and Magnetism, volume 1 of Oxford Classic Texts in the Physical Sciences. Clarendon Press, New York, NY, USA, 1998. ISBN 0-19-850373-3 (paperback). xxxii + 521 pp. With prefaces by W. D. Niven and J. J. Thomson, Reprint of the unabridged third (1891) edition.

ClerkMaxwell:1998:TEMb

[Cle98b] James Clerk Maxwell. A Treatise on Electricity and Magnetism, volume 2 of Oxford Classic Texts in the Physical Sciences. Clarendon Press, New York, NY, USA, 1998. ISBN 0-19-850374-1 (paperback). xxiv + 507 pp. Reprint of the unabridged 3rd edition of 1891.

ClerkMaxwell:2001:TH

[Cle01] James Clerk Maxwell. Theory of Heat. Dover Publications, Inc., New York, NY, USA, 2001. ISBN 0-486-41735-2 (paperback), 0-486-17406-9 (e-book). xx + 364 pp. LCCN QC255 .M4 2001.

ClerkMaxwell:2002:MM

[Cle02] James Clerk Maxwell. Matter and Motion. Great Minds Series. Prometheus Books, Amherst, NY, USA, 2002. ISBN 1-57392-989-1 (paperback). xvi + 163 pp. LCCN QA846 .M39 2002. With notes and appendices by Joseph Larmor. Reprint of the 1876 original [Cle76e].

ClerkMaxwell:2005:ETE

[Cle05] James Clerk Maxwell, editor. An Elementary Treatise on Electricity. Dover Publications, Inc., New York, NY, USA, 2005. ISBN 0-486-43884-8 (paperback). xxii + 234 pp. LCCN QC518

.M44 2005. URL http://www.loc.gov/catdir/enhancements/
fy0711/2004058284-d.html;   https://hdl.handle.net/2027/

pst.000058527900. Edited by William Garnett. Introduction and notes by Peter Pesic.

ClerkMaxwell:2006:SBG

[Cle06] James Clerk Maxwell. Substanz und Bewegung. (German) [Matter and Motion]. Edition classic. VDM, Müller, Saarbrücken, Germany, 2006. ISBN 3-86550-868-5, 3-86550-869-3. xiv + 144 pp. LCCN ????

ClerkMaxwell:2011:TTP

[Cle11] James Clerk Maxwell. On the theory of three primary colours. In Niven [Niv11], pages 445–450. ISBN 1-108-01225-6 (vol. 1), 1-108-01538-7 (vol. 2), 1-108-01227-2 (set). URL http://

notesonphotographs.org/index.php?title=Maxwell,_J._Clerk.
_%22On_the_Theory_of_Three_Primary_Colours.%22. Two-volume

set. Reprint of the 1890 hardcover edition.

ClerkMaxwell:2013:ETE

[Cle13] James Clerk Maxwell. An Elementary Treatise on Electricity. Dover Books on Physics. Dover Publications, Inc., New York, NY, USA, second edition, 2013. ISBN 0-486-43884-8, 0-486-17463-8 (e-book). 367 pp. LCCN QC518 .M44 2015. URL http://ebookcentral.

proquest.com/lib/ucm/detail.action?docID=1919764;   http:
//public.eblib.com/choice/publicfullrecord.aspx?p=1919764.

ClerkMaxwell:1880:UEP

[Cle0 ] James Clerk Maxwell. On the unpublished electrical papers of the Hon. Henry Cavendish. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 3 (part iii)(??):86–89, ???? 1880-. CODEN PCPSA4. ISSN 0008-1981. This may be the last paper presented by JCM during his lifetime.

Cohen:1952:MP

[Coh52] I. Bernard Cohen. Maxwell’s poetry. Scientific American, 186 (3):62–63, March 1952. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v186/n3/pdf/scientificamerican0352-
62.pdf.

Cohen:1994:NSS

[Coh94] I. Bernard Cohen, editor. The Natural Sciences and the Social Sciences: Some Critical and Historical Perspectives, volume 150 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1994. ISBN 0-7923-2223-1, 94-017-3391-0 (e-book). ISSN 0068-0346. xxxvi + 404 pp. LCCN Q174 .B67 vol. 150. URL https://link.

springer.com/book/10.1007/978-94-017-3391-5.

Collier:1990:TFM

[Col90] John D. Collier. Two faces of Maxwell’s demon reveal the nature of irreversibility. Studies in History and Philosophy of Science Part A, 21(2):257–268, June 1990. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368190900254.

Cole:1997:HAS

[Col97] James B. Cole. High accuracy solution of Maxwell’s equations using nonstandard finite differences. Computers in Physics, 11(3):287– ??, May 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/

1.168620. See erratum [Col98].

Cole:1998:EHA

[Col98] J. B. Cole. Erratum: “High accuracy solution of Maxwell’s equations using nonstandard finite differences” [Comput. Phys. 11, 287 (1997)]. Computers in Physics, 12(2):181–??, March 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168634. See [Col97].

Coulson:1963:IF

[Cou63] C. A. Coulson. Interatomic forces. In Clerk Maxwell and modern science; six commemorative lectures [Dom63], page ?? LCCN QC71 .D63.

Colombini:2011:IDS

[CPR11] Ferruccio Colombini, Vesselin Petkov, and Jeffrey Rauch. Incoming and disappearing solutions for Maxwell’s equations. Proceedings of the American Mathematical Society, 139(6):2163–2173, 2011. CO-DEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic). URL

http://www.jstor.org/stable/41291775.

Cranefield:1954:CMC

[Cra54] Paul F. Cranefield, Ph.D. Clerk Maxwell’s corrections to the page proofs of A Dynamical Theory of the Electromagnetic Field. Annals of Science, 10(4):359–362, 1954. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Craik:2007:MHM

[Cra07] Alex D. D. Craik. Mr Hopkins’ men: Cambridge reform and British mathematics in the 19th century. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 1-84628-790-1 (hardcover), 1-84628-791-X (e-book). xiv + 405 + 24 pp.

LCCN QA14.G7 C73 2007. URL http://www.loc.gov/catdir/
enhancements/fy0825/2006940325-b.html;   http://www.loc.
gov/catdir/enhancements/fy0825/2006940325-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0825/2006940325-t.html.

Crowther:1935:BSC

[Cro35] James Gerald Crowther. British Scientists of the 19th century. Penguin, London, UK, 1935. LCCN ????

Crowther:1940:BSC

[Cro40] James Gerald Crowther. British Scientists of the 19th century. Penguin, London, UK, 1940. LCCN ????

Crowther:1974:CL

[Cro74] James Gerald Crowther. The Cavendish Laboratory 1874–1974. Science History Publications, New York, NY, USA, 1974. ISBN 0-88202-029-3. xvi + 464 pp. LCCN QC51.G72 C353 1974.

Cropper:2001:GPL

[Cro01] William H. Cropper. Great physicists: the life and times of leading physicists from Galileo to Hawking. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2001. ISBN 0-19-513748-5. xii + 500

pp. LCCN QC15 .C76 2001. URL http://www.loc.gov/catdir/
enhancements/fy0637/2001021611-d.html;   http://www.loc.
gov/catdir/enhancements/fy0724/2001021611-b.html.

ClerkMaxwell:1907:SSC

[CS07] James Clerk Maxwell and George Gabriel Stokes. [special scientific correspondence]. In Larmor [Lar07], page ?? LCCN QC16.S8 A2. Two volumes.

ClerkMaxwell:1971:SSC

[CS71] James Clerk Maxwell and George Gabriel Stokes. [special scientific correspondence]. In Larmor [Lar71], page ?? LCCN QC16.S8 A2 1971.

ClerkMaxwell:2010:SSC

[CS10] James Clerk Maxwell and George Gabriel Stokes. [special scientific correspondence]. In Larmor [Lar10], pages 1–45. ISBN 0-511-70712-6 (e-book), 1-108-00892-5 (paperback). LCCN ????

DAgostino:1981:WMD

[D’A81] Salvo D’Agostino. Weber and Maxwell on the discovery of the velocity of light in Nineteenth Century electrodynamics. In Grmek et al. [GCC81], pages 281–293. ISBN 90-277-1123-2, 94-010-1284-9 (ebook). ISSN 0068-0346. LCCN ???? URL https://link.springer.

com/chapter/10.1007/978-94-010-1284-3_14.

DAgostino:1986:MDA

[D’A86] S. D’Agostino. Maxwell’s dimensional approach to the velocity of light. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 29(3):178–204, September 1986. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

DAgostino:1989:PHN

[D’A89] S. D’Agostino. Pourquoi Hertz et non pas Maxwell, a-t-il découvert les ondes électriques?. (French) [Why Hertz, and not Maxwell, discovered electric waves]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 32(1):66–76, April 1989. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

DAgostino:2000:DTM

[D’A00] Salvo D’Agostino. On the difficulties of the transition from Maxwell’s and Hertz’s pure-field theories to Lorentz’s electron. Physics in Perspective (PIP), 2(4):398–410, December 2000. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s000160050052.

DAgostino:2001:HRD

[D’A01a] Salvo D’Agostino. A historical role for dimensional analysis in Maxwell’s electromagnetic theory of light. In A history of the ideas of theoretical physics: essays on the nineteenth and twentieth century physics [D’A01b], chapter 3, pages 45–75. ISBN 1-4020-0244-0, 94-010-9034-3 (e-book). ISSN 0068-0346. LCCN QC5.53;

QC6. URL https://link.springer.com/chapter/10.1007/978-
94-010-9034-6_3.

DAgostino:2001:HIT

[D’A01b] Salvo D’Agostino. A history of the ideas of theoretical physics: essays on the nineteenth and twentieth century physics, volume 213 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 1-4020-0244-0, 94-010-9034-3 (e-book). ISSN 0068-0346. xviii + 381 pp. LCCN QC5.53; QC6.

DAbramo:2012:NSS

[D’A12] Germano D’Abramo. A note on solid-state Maxwell demon. Foundations of Physics, 42(3):369–376, March 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-011-9607-7.

Darrigol:1993:AOB

[Dar93a] Olivier Darrigol. Analyses d’ouvrages: The Maxwellians^{≪}Cornell History of Science Series ^{≫}par Bruce J. Hunt. Revue d’Histoire des Sciences, 46(2–3):305, avril–septembre 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.

org/stable/23633832.

Darrigol:1993:AOM

[Dar93b] Olivier Darrigol. Analyses d’ouvrages: The Maxwellians^{≪}Cornell History of Science Series ^{≫}par Bruce J. Hunt. Revue d’Histoire des Sciences, 46(2–3):305, avril–septembre 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.

org/stable/23633832.

Darrigol:1993:ERG

[Dar93c] Olivier Darrigol. The electrodynamic revolution in Germany as documented by early German expositions of “Maxwell’s theory”. Archive for History of Exact Sciences, 45(3):189–280, September 1993. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.jstor.org/stable/41133949;   http://www.
springerlink.com/openurl.asp?genre=article&issn=0003-9519&
volume=45&issue=3&spage=189.

Darrigol:1998:AOB

[Dar98a] Olivier Darrigol. Analyses d’ouvrages: The Scientific Letters and papers of James Clerk Maxwell, vol. II : 1862–1873 par P. M. Harman. Revue d’Histoire des Sciences, 51(1):157, January 1998. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633921.

Darrigol:1998:AOS

[Dar98b] Olivier Darrigol. Analyses d’ouvrages: The Scientific Letters and papers of James Clerk Maxwell, vol. II: 1862–1873 par P. M. Harman. Revue d’Histoire des Sciences, 51(1):157, January 1998. CO-DEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23633921.

Darrigol:1998:CEM

[Dar98c] Olivier Darrigol. Aux confins de l’électrodynamique maxwellienne : Ions et électrons vers 1897. Revue d’Histoire des Sciences, 51(1):5–34, January 1998. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23633905.

Darrigol:2000:EAE

[Dar00] Olivier Darrigol. Electrodynamics from Ampère to Einstein. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-850594-9. xix + 532 pp. LCCN 0.2dar a0165 a4100.

Darrigol:2001:AOB

[Dar01a] Olivier Darrigol. Analyses d’ouvrages: The Natural philosophy of James Clerk Maxwell par Peter M. Harman. Revue d’Histoire des Sciences, 54(1):124–125, January 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.

org/stable/23633621.

Darrigol:2001:AON

[Dar01b] Olivier Darrigol. Analyses d’ouvrages: The Natural philosophy of James Clerk Maxwell par Peter M. Harman. Revue d’Histoire des Sciences, 54(1):124–125, January 2001. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.

org/stable/23633621.

Darrigol:2010:JME

[Dar10] Olivier Darrigol. James MacCullagh’s ether: An optical route to Maxwell’s equations? European Physical Journal H, 35(2):133–172, November 2010. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/10.
1140/epjh/e2010-00009-3.

Daub:1970:MD

[Dau70] Edward E. Daub. Maxwell’s demon. Studies in History and Philosophy of Science Part A, 1(3):213–227, November 1970. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0039368170900105.

Davidson:1860:NPL

[Dav60] A. Davidson. Natural philosophy lecture notes. Cambridge University Library Ms Add 8791., 1860. Written for a course given by James Clerk Maxwell at King’s College, London.

DeMarzo:1978:MFC

[De 78] Carlo De Marzo. Maxwell e la fisica classica. (Italian) [Maxwell and classical physics], volume 488 of Universale Laterza. Laterza, Roma, Italy, 1978. 225 pp. LCCN QC16.M4 D4. L3900.

Deas:1872:SFP

[Dea72] Francis Deas. On spectra formed by the passage of polarised light through double refracting crystals. Transactions of the Royal Society of Edinburgh, 26(1):177–185, ???? 1872. CODEN TRSEAO. ISSN 0080-4568 (print), 2053-5945 (electronic). URL https://hdl.

handle.net/2027/umn.31951d00649878j?urlappend=%3Bseq=209.

See addition [Cle72d].

Dear:2016:BRM

[Dea16] Peter Dear. Book review: Matthew Stanley. Huxley’s Church and Maxwell’s Demon: From Theistic Science to Naturalistic Science. Isis, 107(2):429–431, June 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Denur:2007:SDW

[Den07] Jack Denur. Speed-dependent weighting of the Maxwellian distribution in rarefied gases: A Second-Law Paradox? Foundations of Physics, 37(12):1685–1706, December 2007. CODEN FNDPA4.

ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-007-9160-6.

Devereux:2002:SEM

[Dev02] Michael Devereux. Szilard’s Engine: Measurement, information, and Maxwell’s demon. In Daniel P. Sheehan, editor, Quantum Limits to the Second Law: First International Conference on Quantum Limits to the Second Law. San Diego, California (USA), 29–31 July 2002. AIP Conference Proceedings, volume 643 of American Institute of Physics Conference Series, pages 279–284. American Institute of Physics, Woodbury, NY, USA, November 2002. ISBN 0-7354-0098-9.

URL http://adsabs.harvard.edu/abs/2002AIPC..643..279D.

Dewar:2018:MG

[Dew18] Neil Dewar. Maxwell gravitation. Philosophy of Science, 85(2):249– 270, April 2018. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Dias:1994:CMS

[Dia94a] Penha Maria Cardoso Dias. Clausius and Maxwell: The statistics of molecular collisions (1857–1862). Annals of Science, 51(3):249– 261, 1994. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dias:1994:WSS

[Dia94b] Penha Maria Cardoso Dias. “Will someone say exactly what the Htheorem proves?” A study of Burbury’s Condition A and Maxwell’s Proposition II. Archive for History of Exact Sciences, 46(4):341–366, December 1994. CODEN AHESAN. ISSN 0003-9519 (print), 1432-

0657 (electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0003-9519&volume=46&issue=4&spage=
341.

Dibner:1964:JCM

[Dib64] B. Dibner. James Clerk Maxwell. IEEE Spectrum, 1(12):50–56, December 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Domb:1963:CMM

[Dom63] Cyril Domb. Clerk Maxwell and modern science; six commemorative lectures. University of London, Athlone Press, London, UK, 1963. vii + 118 pp. LCCN QC71 .D63.

Domb:1980:JCM

[Dom80] C. Domb, F.R.S. James Clerk Maxwell in London: 1860–1865. Notes and Records of the Royal Society of London, 35(1):67–103 + 1, July 1980. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531602.

Domb:1992:BRS

[Dom92] Cyril Domb. Book review: The scientific letters and papers of James Clerk Maxwell. Journal of Statistical Physics, 67(3–4):837– 838, May 1992. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.1007/
BF01049732.

Domb:1997:BRS

[Dom97] C. Domb. Book review: The scientific letters and papers of James Clerk Maxwell, Vol. II. Journal of Statistical Physics, 88(5–6):1419– 1422, September 1997. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/article/
10.1007/BF02732444.

Domb:2004:BRS

[Dom04] Cyril Domb. Book review: The Scientific Letters and Papers of James Clerk Maxwell. P. M. Harman, ed. Vol. III, Cambridge University Press, Cambridge. Journal of Statistical Physics, 116(5–6):1703–1706, September 2004. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/10.
1023/B%3AJOSS.0000041912.31649.5a.

Domb:2018:JCM

[Dom18] Cyril Domb. James Clerk Maxwell. Encyclopedia Britannica arti-

cle, April 18, 2018. URL https://www.britannica.com/biography/
James-Clerk-Maxwell.

Dorling:1970:MAA

[Dor70] Jon Dorling. Maxwell’s attempts to arrive at non-speculative foundations for the kinetic theory. Studies in History and Philosophy of Science Part A, 1(3):229–248, November 1970. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368170900117.

Dougal:2016:KMC

[Dou16] R. C. Dougal. Kelvin, Maxwell, Clausius and Tait: the correspondence of James Clerk Maxwell. In Collins et al. [CDKR16], pages 135–151. ISBN 1-84564-149-3 (hardcover), 1-84564-387-9 (ebook). ISSN 1755-8336. LCCN QC16.K3 K36 2016. URL https:

//www.witpress.com/Secure/elibrary/papers/9781845641498/
9781845641498005FU1.pdf.

Duhem:1902:TEJ

[Duh02] Pierre Maurice Marie Duhem. Les théories électriques de J. Clerk Maxwell: étude historique et critique. (French) [Electrical Theories of J. Clerk Maxwell: Historical and critical study]. A. Hermann, Paris, France, 1902. 3 + 228 pp. LCCN QC518 .D87. URL http://

gallica.bnf.fr/ark:/12148/bpt6k55434315.

Duhem:1967:TEJ

[Duh67] Pierre Maurice Marie Duhem. Les théories électriques de J. Clerk Maxwell: étude historique et critique. (French) [Electrical Theories of J. Clerk Maxwell: Historical and critical study]. Readex Microprint, New York, NY, USA, 1967. ???? pp. LCCN ????

Duhem:2015:ETJ

[Duh15a] Pierre Maurice Marie Duhem. The electric theories of J. Clerk Maxwell: a historical and critical study, volume 314 of Boston Studies in the Philosophy and History of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. ISBN 3-319-18514-4 (hardcover), 3-319-18515-2 (e-book). ISSN 0068-0346. xv + 185 pp. LCCN QC518 .D84 2015. URL https://link.

springer.com/book/10.1007/978-3-319-18515-6. English transla-

tion by Alan Aversa.

Duhem:2015:FEM

[Duh15b] Pierre Maurice Marie Duhem. The first electrostatics of Maxwell. In The electric theories of J. Clerk Maxwell: a historical and critical study [Duh15a], chapter 3, pages 35–42. ISBN 3-319-18514-4 (hardcover), 3-319-18515-2 (e-book). ISSN 0068-0346. LCCN QC518 .D84

2015. URL https://link.springer.com/chapter/10.1007/978-3-

319-18515-6_3. English translation by Alan Aversa.

Duhem:2015:SEMa

[Duh15c] Pierre Maurice Marie Duhem. The second electrostatics of Maxwell. In The electric theories of J. Clerk Maxwell: a historical and critical study [Duh15a], chapter 4, pages 43–63. ISBN 3-319-18514-4 (hardcover), 3-319-18515-2 (e-book). ISSN 0068-0346. LCCN QC518 .D84

2015. URL https://link.springer.com/chapter/10.1007/978-3-

319-18515-6_4. English translation by Alan Aversa.

Duhem:2015:SEMb

[Duh15d] Pierre Maurice Marie Duhem. The six equations of Maxwell and electromagnetic energy. In The electric theories of J. Clerk Maxwell: a historical and critical study [Duh15a], chapter 7, pages 101–140. ISBN 3-319-18514-4 (hardcover), 3-319-18515-2 (e-book). ISSN 0068-0346. LCCN QC518 .D84 2015. URL https://link.springer.

com/chapter/10.1007/978-3-319-18515-6_7. English translation

by Alan Aversa.

Duhem:2015:TEM

[Duh15e] Pierre Maurice Marie Duhem. The third electrostatics of Maxwell. In The electric theories of J. Clerk Maxwell: a historical and critical study [Duh15a], chapter 5, pages 65–78. ISBN 3-319-18514-4 (hardcover), 3-319-18515-2 (e-book). ISSN 0068-0346. LCCN QC518 .D84

2015. URL https://link.springer.com/chapter/10.1007/978-3-

319-18515-6_5. English translation by Alan Aversa.

Dunstan:2008:DSR

[Dun08] D. J. Dunstan. Derivation of special relativity from Maxwell and Newton. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 366(1871):1861–1865, May 2008. CODEN PTRMAD, PTMSFB. ISSN 1364-503X (print), 1471-2962

(electronic). URL http://www.jstor.org/stable/25190793.

Dyson:1990:FPM

[Dys90] Freeman J. Dyson. Feynman’s proof of the Maxwell equations. American Journal of Physics, 58(3):209–211, 1990. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://link.aip.

org/link/?AJP/58/209/1.

Dyson:1999:WMT

[Dys99] Freeman J. Dyson. Why is Maxwell’s theory so hard to understand? In James Clerk Maxwell Commemorative Booket for The Fourth International Congress on Industrial and Applied Mathematics, page ?? James Clerk Maxwell Foundation, Edinburgh, UK, 1999. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?arnumber=4458240;
http://www.clerkmaxwellfoundation.org/DysonFreemanArticle.
pdf;   http://www.damtp.cam.ac.uk/user/tong/em/dyson.pdf.

Board:1981:ABJ

[Edi81] Editorial Board. On the 150th anniversary of the birth of James Clerk Maxwell. Soviet Physics. Uspekhi, 24(11):875, November 1981. CODEN SOPUAP. ISSN 0038-5670 (print), 2169-5296 (elec-

tronic). URL http://stacks.iop.org/0038-5670/24/i=11/a=E01;
 http://ufn.ru/en/articles/1981/11/a/.

Edwards:2010:MDD

[Edw10] Lin Edwards. Maxwell’s demon demonstration turns information into energy. Web site., November 15, 2010. URL https://phys.org/

news/2010-11-maxwell-demon-energy.html.

Ehrenberg:1967:MD

[Ehr67] W. Ehrenberg. Maxwell’s demon. Scientific American, 217(5): 103–110, November 1967. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v217/n5/pdf/scientificamerican1167-
103.pdf.

Einstein:1931:MID

[Ein31a] Albert Einstein. Maxwell’s influence on the development of the conception of physical reality. In Thomson [Tho31b], page ?? LCCN QC16.M4 J3. Essays by Sir J. J. Thomson, Max Planck, Albert Einstein, Joseph Larmor, James Jeans, William Garnett, Ambrose Fleming, Oliver Lodge, R. T. Glazebrook, and Horace Lamb.

Einstein:1931:MIE

[Ein31b] Albert Einstein. Maxwell’s influence on the evolution of the idea of physical reality. In J. J. (Joseph John) Thomson, editor, James Clerk Maxwell: A Commemoration Volume, 1831–1931, page ?? Cambridge University Press, Cambridge, UK, 1931. LCCN QC16.M4 J3. English translation in [Ein82a, pp. 266–270].

Einstein:1954:IO

[Ein54a] Albert Einstein. Ideas and Opinions. Crown Publishers, New York, NY, USA, 1954. vi + 377 pp. LCCN AC35 .E526 1954.

Einstein:1954:MIE

[Ein54b] Albert Einstein. Maxwell’s influence on the evolution of the idea of physical reality. In Ideas and Opinions [Ein54a], pages 266–270. LCCN AC35 .E526 1954.

Einstein:1982:IO

[Ein82a] Albert Einstein. Ideas and Opinions. Three Rivers Press, New York, NY, USA, 1982. ISBN 0-517-88440-2. 377 pp. LCCN AC35 .E526 1982x.

Einstein:1982:MIE

[Ein82b] Albert Einstein. Maxwell’s influence on the evolution of the idea of physical reality. In Ideas and Opinions [Ein82a], pages 266–270. ISBN 0-517-88440-2. LCCN AC35 .E526 1982x.

Elder:2000:ROJ

[Eld00] Kingsley Elder. The religious outlook of James Clerk Maxwell. IOP History of Physics Group Newsletter, 13:41–47, Spring 2000. ISSN

1756-168X. URL http://www.iop.org/activity/groups/subject/
hp/newsletter/archive/file_66506.pdf.

Elliott:1966:RE

[Ell66] R. S. Elliott. Relativity and electricity. IEEE Spectrum, 3(3):140–152, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Einstein:1984:EQA

[EM84] Albert Einstein and James Clerk Maxwell. Einstein’s quotations about Maxwell. In Berger [Ber84], pages 11–14. ISBN 0-444-86707-4 (Elsevier). LCCN QC669 .J17 1984.

Earman:1998:EXW

[EN98] John Earman and John D. Norton. Exorcist XIV: The wrath of Maxwell’s Demon. Part I. from Maxwell to Szilard. Studies in History and Philosophy of Modern Physics, 29(4):435–471, December 1998. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1355219898000239.

Earman:1999:EXW

[EN99] John Earman and John D. Norton. Exorcist XIV: The wrath of Maxwell’s Demon. Part II from Szilard to Landauer and beyond. Studies in History and Philosophy of Modern Physics, 30(1):1–40, March 1999. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1355219898000264.

Engheta:2015:YME

[Eng15] N. Engheta. 150 years of Maxwell’s equations. Science, 349(6244): 136–137, July 2015. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). Retrospective on [Cle65].

Elyashevich:1981:MCD

[EP81] M. A. El’yashevich and T. S. Prot’ko. Maxwell’s contribution to the development of molecular physics and statistical methods. Soviet Physics. Uspekhi, 24(11):876–903, November 1981. CODEN SOP-

UAP. ISSN 0038-5670. URL http://ufn.ru/en/articles/1981/
11/b/.

Epple:1998:TMS

[Epp98] Moritz Epple. Topology, matter, and space, I: Topological notions in 19th-century natural philosophy. Archive for History of Exact Sciences, 52(4):297–392, February 1998. CODEN AHE-SAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=0003-
9519&volume=52&issue=4&spage=297.

Evans:1961:MCP

[Eva61] Ralph M. Evans. Maxwell’s color photograph. Scientific American, 205(5):118–128, November 1961. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v205/n5/pdf/scientificamerican1161-
118.pdf.

Everitt:1967:JCMb

[Eve67a] C. W. F. (C. W. Francis) Everitt. James Clerk Maxwell: physicist and natural philosopher. C. Scribner’s Sons, New York, NY, USA, 1967. ISBN 0-684-14253-8. 205 + 18 pp. LCCN QC16.M4 E8 1975.

Everitt:1967:JCMa

[Eve67b] C. W. Francis Everitt. [James Clerk Maxwell]. In Encyclopedia of philosophy, page ?? Macmillan Library Reference USA, New York, NY, USA, 1967.

Everitt:1967:MSP

[Eve67c] C. W. Francis Everitt. Maxwell’s scientific papers. Applied Optics, 6(4):639–649, April 1, 1967. CODEN APOPAI. ISSN 0003-

6935. URL https://www.osapublishing.org/ao/abstract.cfm?
uri=ao-6-4-639.

Everitt:1975:JCM

[Eve75] C. W. F. (C. W. Francis) Everitt. James Clerk Maxwell: physicist and natural philosopher. DSB editions. Charles Scribner’s Sons, New York, NY, USA, 1975. ISBN 0-684-14254-6, 0-684-14253-8 (paperback). 205

pp. LCCN QC16.M4 E8 1975. URL http://catalog.hathitrust.
org/api/volumes/oclc/2074768.html.

Everitt:1983:MSC

[Eve83] C. W. Francis Everitt. Maxwell’s scientific creativity. In Aris et al. [ADS83], pages 71–141. ISBN 0-8166-1087-8. LCCN QC7 .S77 1983.

Everitt:1984:BRM

[Eve84] C. W. Francis Everitt. Book review: Maxwell, the man: James Clerk Maxwell: a biography. Ivan Tolstoy. University of Chicago Press, 1983. 184 pp. $6.95, $17.00, hardcover. The Physics Teacher, 22(4):264–266, April 1984. CODEN PHTEAH. ISSN 0031-921X (print), 1943-4928 (electronic).

Everitt:1998:JCM

[Eve98] C. W. Francis Everitt. [James Clerk Maxwell]. In Craig and Floridi [CF98], page ?? ISBN 0-415-16916-X (CD-ROM), 0-415-19608-6 (user guide), 0-415-16917-8 (10 vol set + CD-ROM). LCCN B51. URL

http://www.rep.routledge.com/.

Everitt:1999:BRN

[Eve99] C. W. Francis Everitt. Book review: The Natural Philosophy of James Clerk Maxwell, by Peter M. Harman. Physics Today, 52(8):61, August 1999. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Everitt:2006:JCM

[Eve06] C. W. Francis Everitt. James Clerk Maxwell: a force for physics. Physics World, 19(12):32, December 2006. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://stacks.iop.

org/2058-7058/19/i=12/a=36.

Fahie:1899:HWTa

[Fah99a] J. J. (John Joseph) Fahie. A history of wireless telegraphy, 1838– 1899: including some bare-wire proposals for subaqueous telegraphs. Blackwood, Edinburgh, UK, 1899. xvii + 325 pp. LCCN TK5711 .F2.

Fahie:1899:HWTb

[Fah99b] J. J. (John Joseph) Fahie. A history of wireless telegraphy, 1838–1899: including some bare-wire proposals for subaqueous telegraphs. Dodd, Mead and Co., New York, NY, USA, 1899. xvii + 325 pp. LCCN TK5711 .F2.

Fahie:1901:HWT

[Fah01] J. J. (John Joseph) Fahie. A history of wireless telegraphy, 1838– 1899: including some bare-wire proposals for subaqueous telegraphs. Blackwood, Edinburgh, UK, 1901. LCCN TK5711 .F2.

Fahn:1996:MDE

[Fah96] Paul N. Fahn. Maxwell’s demon and the entropy cost of information. Foundations of Physics, 26(1):71–93, January 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:

//link.springer.com/article/10.1007/BF02058888.

Fairgrieve:1931:JCM

[Fai31] Mungo McCallum Fairgrieve. James Clerk Maxwell. The Edinburgh Academy Chronicle, 38(??):81–83, 105–109, June/July 1931.

Falconer:2015:ECM

[Fal15] I. Falconer. Editing Cavendish: Maxwell and the electrical researches of Henry Cavendish. ArXiv e-prints, April 2015. URL http://

adsabs.harvard.edu/abs/2015arXiv150407437F.

Falconer:2017:NAM

[Fal17] Isobel Falconer. No actual measurement … was required: Maxwell and Cavendish’s null method for the inverse square law of electrostatics. Studies in History and Philosophy of Science Part A, 65–66(??): 74–86, October/December 2017. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368117301188.

Fedele:2008:MTS

[Fed08] Renato Fedele. From Maxwell’s theory of Saturn’s rings to the negative mass instability. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 366(1871): 1717–1733, May 2008. CODEN PTRMAD, PTMSFB. ISSN 1364-503X (print), 1471-2962 (electronic). URL http://www.jstor.org/

stable/25190780.

Ferguson:1931:CMC

[Fer31] Allan Ferguson. The Clerk Maxwell centenary celebrations. Nature, 128(3232):604–608, October 10, 1931. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.

com/nature/journal/v128/n3232/pdf/128604a0.pdf.

Ferris:1991:WTP

[FF91] Timothy Ferris and Clifton Fadiman, editors. The world treasury of physics, astronomy, and mathematics. Little, Brown and Co., Boston, MA, USA, 1991. ISBN 0-316-28129-8. xv + 859 pp. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Fisch:1988:ERP

[Fis88] Menachem Fisch. Essay review: A physicists’ philosopher —James Clerk Maxwell on mathematical physics. Journal of Statistical Physics, 51(1–2):309–319, April 1988. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01015334.

Fitzpatrick:1910:HCL

[Fit10] Thomas Cecil Fitzpatrick. A History of the Cavendish Laboratory, 1871–1910: with 3 portraits in collotype and 8 other illustrations. Longmans, Green and Co., London, UK, 1910. x + 1 + 342 pp. LCCN QC51 .C3. A collection of three essays covering the period of the work of Maxwell, Lord Rayleigh, and J. J. Thomson, as directors of the laboratory.

Fitzpatrick:2016:HCL

[Fit16] Thomas Cecil Fitzpatrick. A History of the Cavendish Laboratory, 1871–1910: with 3 portraits in collotype and 8 other illustrations (classic reprint). Forgotten Books, London, UK, 2016. ISBN 1-

332-58925-1. LCCN ???? URL https://forgottenbooks.com/en/
books/AHistoryoftheCavendishLaboratory18711910_10050996.

Fleming:1931:SMP

[Fle31] Ambrose Fleming. Some memories [of Professor James Clerk Maxwell]. In Thomson [Tho31b], page ?? LCCN QC16.M4 J3. Essays by Sir J. J. Thomson, Max Planck, Albert Einstein, Joseph Larmor, James Jeans, William Garnett, Ambrose Fleming, Oliver Lodge, R. T. Glazebrook, and Horace Lamb.

Fletcher:1986:AFM

[Fle86] Colin R. Fletcher. Ampère, Faraday, and Maxwell: Three founders of electromagnetism. Historia Mathematica, 13(3):296, August 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086086901011.

Fleisch:2008:SGM

[Fle08] Daniel A. Fleisch. A Student’s Guide to Maxwell’s Equations. Cambridge University Press, Cambridge, UK, 2008. ISBN 0-521-87761-X (hardcover), 0-521-70147-3 (paperback). ix + 134 pp.

LCCN QC670 .F56 2008. URL http://www.loc.gov/catdir/
enhancements/fy0803/2007037901-b.html;   http://www.loc.
gov/catdir/enhancements/fy0803/2007037901-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0803/2007037901-t.html.

Flood:2012:JCM

[Flo12] Raymond Flood. James Clerk Maxwell transcript. Gresham College Lecture, October 31, 2012. URL https://www.gresham.

ac.uk/lecture/transcript/download/james-clerk-maxwell/;
https://www.youtube.com/watch?v=v40OcJ7rfSE.

Flood:2013:JCM

[Flo13] Raymond Flood. James Clerk Maxwell (BSHM–Gresham Lecture 2012). BSHM Bulletin: Journal of the British Society for the History of Mathematics, 28(2):84–96, 2013. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/17498430.2013.755073.

Forbes:1878:DR

[FM78] George Forbes and J. Clerk Maxwell. 3. On diamagnetic rotation. Proceedings of the Royal Society of Edinburgh, 9:85–92, 1878. CODEN PRSEAE. ISSN 0080-4541 (print), 2053-5902 (electronic).

Forbes:2014:FME

[FM14] Nancy Forbes and Basil Mahon. Faraday, Maxwell, and the electromagnetic field: how two men revolutionized physics. Prometheus Books, Amherst, NY, USA, 2014. ISBN 1-61614-942-6 (hardcover), 1-61614-943-4 (e-book). 320 pp. LCCN QC16.F2 F67 2014.

Franca:1996:MET

[FMM96] H. M. França, A. Maia, Jr., and C. P. Malta. Maxwell electromagnetic theory, Planck’s radiation law, and Bose–Einstein statistics. Foundations of Physics, 26(8):1055–1068, August 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF02061403.

Flood:2014:JCM

[FMW14] Raymond Flood, Mark McCartney, and Andrew Whitaker, editors. James Clerk Maxwell: perspectives on his life and work. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-966437-4. x + 364 + 4 pp. LCCN QC16.M4 J36 2014. URL

http://www.loc.gov/catdir/enhancements/fy1410/2013942195-
b.html;   http://www.loc.gov/catdir/enhancements/fy1410/
2013942195-d.html;   http://www.loc.gov/catdir/enhancements/
fy1410/2013942195-t.html.

Forfar:1999:OCM

[For99a] David O. Forfar. Origins of the Clerk (Maxwell) genius. In James Clerk Maxwell Commemorative Booket for The Fourth International Congress on Industrial and Applied Mathematics, page ?? James Clerk Maxwell Foundation, Edinburgh, UK, 1999. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?arnumber=4458240;
http://www.clerkmaxwellfoundation.org/Maxwell_-_Origins_
of_Genius.pdf;   http://www.damtp.cam.ac.uk/user/tong/em/
dyson.pdf.

Forfar:1999:WBS

[For99b] David O. Forfar. What became of the Senior Wranglers? In James Clerk Maxwell Commemorative Booket for The Fourth International Congress on Industrial and Applied Mathematics, page ?? James Clerk Maxwell Foundation, Edinburgh, UK, 1999. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?arnumber=4458240;
http://www.clerkmaxwellfoundation.org/WranglersWhatBecame2008_
1_24.pdf;   http://www.damtp.cam.ac.uk/user/tong/em/dyson.
pdf.

Fraser:1857:LNP

[Fra57] Angus Fraser. Lecture notes of Professor Maxwell taken by Angus Fraser on natural philosophy. Aberdeen University Special Libraries and Archives Ms. AMCS/3/17., 1857.

Freudenberg:1937:BNG

[FSvL^{+}37] Karl Freudenberg, Arnold Sommerfeld, Max von Laue, A. v. Muralt, and E. Ries. Besprechungen: Norman, A. G., The Biochemistry of Cellulose, the Polyuronides, Lignin, etc.. Sir Joseph Larmor (ed.), Origins of Clerk Maxwell’s Electric Ideas. (German) [Reviews]. Die Naturwissenschaften, 25(47):766–768, November 19, 1937. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL

http://link.springer.com/article/10.1007/BF01789839.

Fuller:1986:JCM

[Ful86] A. T. Fuller. James Clerk Maxwell’s Cambridge manuscripts: extracts relating to control and stability. V. International Journal of Control, 43(3):805–818, 1986. CODEN IJCOAZ. ISSN 0020-7179 (print), 1366-5820 (electronic).

Galton:1869:HGE

[Gal69] Sir Francis Galton. Hereditary Genius: an Enquiry into its Laws and Consequences. MacMillan Publishing Company, New York, NY, USA, 1869. viii + 390 pp. LCCN ???? URL http://galton.org/books/

hereditary-genius/text/pdf/galton-1869-genius-v3.pdf.

Galton:1874:EMS

[Gal74] Francis Galton. English men of science: their nature and nurture. The Social history of science. MacMillan Publishing Company, New York, NY, USA, 1874. ???? pp. LCCN Q149.G7 G35 1874.

Galton:1970:EMS

[Gal70] Francis Galton. English men of science: their nature and nurture, volume 3 of The Social history of science. Cass, London, second [first reprinted] edition, 1970. ISBN 0-7146-1622-2. xi + iii + xiv + 270 pp. LCCN Q149.G7 G35 1970. With a new introduction by Ruth Schwartz Cowan.

Galton:1979:HGE

[Gal79] Sir Francis Galton. Hereditary Genius: an Enquiry into its Laws and Consequences. Friedmann, London, UK, 1979. ISBN 0-904014-46-0. viii + 390 pp. LCCN ???? Introduced by H. J. Eysenck.

Galton:2010:HGE

[Gal10] Sir Francis Galton. Hereditary Genius: an Enquiry into its Laws and Consequences. Nord Press, ????, second edition, 2010. ISBN 1-4455-8123-X (paperback). 408 pp. LCCN ????

Garnett:1879:OJC

[Gar79] William Garnett. [obituary:] James Clerk Maxwell, F.R.S. Nature, 21(524):43–46, November 13, 1879. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.

com/nature/journal/v21/n524/pdf/021043a0.pdf.

Garnett:1885:HSP

[Gar85a] William Garnett. Heroes of science. Physicists. Society for Promoting Christian Knowledge, London, UK, 1885. vii + 1 + 339 + 1 pp. LCCN QC15 .G3 1885.

Garnett:1885:JCM

[Gar85b] William Garnett. James Clerk Maxwell. In Heroes of science. Physicists [Gar85a], pages 278–308. LCCN QC15 .G3 1885.

Garnett:1904:CMG

[Gar04] J. C. Maxwell Garnett. Colours in metal glasses and in metallic films. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 203(359–371):385–420, January 1904. CODEN PTRMAD, PTMSFB. ISSN 1364-503X (print), 1471-

2962 (electronic). URL https://en.wikipedia.org/wiki/James_
Clerk_Maxwell_Garnett;   https://royalsocietypublishing.
org/doi/abs/10.1098/rsta.1904.0024.

Garnett:1931:ML

[Gar31] William Garnett. Maxwell’s laboratory. In Thomson [Tho31b], page ?? LCCN QC16.M4 J3. Essays by Sir J. J. Thomson, Max Planck, Albert Einstein, Joseph Larmor, James Jeans, William Garnett, Ambrose Fleming, Oliver Lodge, R. T. Glazebrook, and Horace Lamb.

Garber:1966:MCG

[Gar66] Elizabeth Anne Wolfe Garber. Maxwell, Clausius and Gibbs: Aspects of the Development of Kinetic Theory and Thermodynamics. Ph.D. thesis, Case Institute of Technology, Cleveland, OH, USA, November

1966. ii + 320 pp. URL http://search.proquest.com/pqdtglobal/
docview/302345662.

Garber:1969:JCM

[Gar69] Elizabeth W. Garber. James Clerk Maxwell and thermodynamics. American Journal of Physics, 37(2):146–155, February 1969. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL

http://aapt.scitation.org/doi/abs/10.1119/1.1975430. Con-

tains reprint of [Clexxa] in Appendix on pages 154–155.

Garber:1970:CMK

[Gar70] Elizabeth Wolfe Garber. Clausius and Maxwell’s kinetic theory of gases. Historical Studies in the Physical Sciences, 2(??):299–319, ????

1970. CODEN HSPSAS. ISSN 0073-2672. URL http://www.jstor.

org/stable/27757309.

Garber:2008:SGS

[Gar08] Elizabeth Garber. Subjects great and small: Maxwell on Saturn’s rings and kinetic theory. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 366(1871): 1697–1705, May 2008. CODEN PTRMAD, PTMSFB. ISSN 1364-503X (print), 1471-2962 (electronic). URL http://www.jstor.org/

stable/25190778.

Garber:1986:MMG

[GBE86] Elizabeth Garber, Stephen G. Brush, and C. W. F. (C. W. Francis) Everitt, editors. Maxwell on Molecules and Gases. MIT Press, Cambridge, MA, USA, 1986. ISBN 0-262-07094-4 (hardcover). xvii + 565

pp. LCCN QC16.M4 A4 1986. URL http://mitpress.mit.edu/
books/maxwell-molecules-and-gases.

Garber:1995:MHS

[GBE95] Elizabeth Garber, Stephen G. Brush, and C. W. F. Everitt, editors. Maxwell on heat and statistical mechanics: on “avoiding all personal enquiries” of molecules. Lehigh University Press, Bethlehem, PA, USA, 1995. ISBN 0-934223-34-3. 550 pp. LCCN QC310.2.

Grmek:1981:SDE

[GCC81] Mirko Dražen Grmek, Robert S. Cohen, and Guido Cimino, editors. On Scientific Discovery: The Erice Lectures 1977, volume 34 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1981. ISBN 90-277-1123-2, 94-010-1284-9 (e-book). ISSN 0068-0346.

LCCN ???? URL https://link.springer.com/book/10.1007/
978-94-010-1284-3.

Gee:1981:PSG

[Gee81] Brian Gee. Popularizing Scotland’s genius: James Clerk Maxwell. Ivan Tolstoy. 1981 Edinburgh: Canongate. viii + 184 pp. Physics Bulletin, 32(12):401, 1981. CODEN PHSBB4. ISSN 0031-9112. URL

http://stacks.iop.org/0031-9112/32/i=12/a=039.

Gersten:2001:MEO

[Ger01] Alexander Gersten. Maxwell equations — the one-photon quantum equation. Foundations of Physics, 31(8):1211–1231, August 2001. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/
A%3A1017551920941.

Gianelli:1970:SEF

[Gia70] Arthur Francis Gianelli. Study of the Electromagnetic Field Theory of James Clerk Maxwell from the Standpoint of a Modified Realist View of the Nature of Scientific Theory. Ph.D. thesis, Department of Philosophy, St. John’s University, St. John’s, NL, Canada, 1970. URL

http://search.proquest.com/pqdtglobal/docview/302529768.

Gibson:2019:SIH

[Gib19] Susannah Gibson. The Spirit of Inquiry: How One Extraordinary Society Shaped Modern Science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2019. ISBN 0-19-883337-7 (hardcover). xxi + 377 pp. LCCN Q41.C194 G537 2019.

Glazebrook:1896:JCM

[Gla96] Richard T. Glazebrook, F.R.S. James Clerk Maxwell and modern physics. The Century Science Series. MacMillan Publishing Company, New York, NY, USA, 1896. i–viii + 9– 224 + 16 pp. LCCN QC16.M4 G5 1896. URL https:

//archive.org/details/in.ernet.dli.2015.37733;   https://
archive.org/details/in.ernet.dli.2015.59485.

Glazebrook:1901:JCM

[Gla01] Richard Glazebrook. James Clerk Maxwell and modern physics. The Century Science Series. Cassell and Company, Limited, London, UK, 1901. vi + 1 + 9–224 pp. LCCN QC16.M4 G5.

Glazebrook:1931:EDC

[Gla31] R. T. Glazebrook. Early days at the Cavendish Laboratory. In Thomson [Tho31b], page ?? LCCN QC16.M4 J3. Essays by Sir J. J. Thomson, Max Planck, Albert Einstein, Joseph Larmor, James Jeans, William Garnett, Ambrose Fleming, Oliver Lodge, R. T. Glazebrook, and Horace Lamb.

Goldman:1983:DAS

[Gol83] Martin Goldman. The demon in the aether: the story of James Clerk Maxwell. P. Harris, Edinburgh, Scotland, 1983. ISBN 0-86228-026-5. 224 + 12 pp. LCCN QC16.M4 G65 1983. US$15.00.

Goldberg:1994:SDM

[Gol94] Joshua N. Goldberg. Self-dual Maxwell field on a null surface. II. Foundations of Physics, 24(4):467–476, April 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:

//link.springer.com/article/10.1007/BF02058058.

Golin:2013:JCM

[Gol13] Genrikh Golin. James Clerk Maxwell, a modern educator. Physics Today, 66(6):8–10, June 2013. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Gooding:1985:BRB

[Goo85a] David Gooding. Book review: The Demon in the Aether: The Story of James Clerk Maxwell by Martin Goldman. Isis, 76(2):281, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/231813.

Gooding:1985:BRD

[Goo85b] David Gooding. Book review: The Demon in the Aether: The Story of James Clerk Maxwell by Martin Goldman. Isis, 76(2):281, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/231813.

Gordon:1983:MDD

[Gor83] L. G. M. Gordon. Maxwell’s demon and detailed balancing. Foundations of Physics, 13(10):989–997, October 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00729519.

Gregis:2019:AAM

[Gré19] Fabien Grégis. Assessing accuracy in measurement: the dilemma of safety versus precision in the adjustment of the fundamental physical constants. Studies in History and Philosophy of Science Part A, 74(??):42–55, April 2019. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL https://www.sciencedirect.

com/science/article/pii/S0039368117302972.

Guz:2008:PPM

[GRGR08] I. A. Guz, A. A. Rodger, A. N. Guz, and J. J. Rushchitsky. Predicting the properties of micro- and nanocomposites: from the microwhiskers to the bristled nano-centipedes. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 366(1871):1827–1833, May 2008. CODEN PTRMAD, PTMSFB. ISSN 1364-503X (print), 1471-2962 (electronic).

Galton:1906:NFM

[GS06] Francis Galton and Edgar Schuster. Noteworthy families (modern science): an index to kinships in near degrees between persons whose achievements are honourable, and have been publicly recorded, volume 1 of London. University. Eugenics record office. Publications. J. Murray, London, UK, 1906. xlii + 96 pp. LCCN HQ753 .G2.

Gyenis:2017:MND

[Gye17] Balázs Gyenis. Maxwell and the normal distribution: a colored story of probability, independence, and tendency toward equilibrium. Studies in History and Philosophy of Modern Physics, 57(??):53– 65, February 2017. CODEN ???? ISSN 1355-2198 (print), 1879-

2502 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219817300102.

Gill:2011:TME

[GZ11] Tepper L. Gill and Woodford W. Zachary. Two mathematically equivalent versions of Maxwell’s equations. Foundations of Physics, 41(1): 99–128, January 2011. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/
10.1007/s10701-009-9331-8.

Harre:1969:SCB

[H^{+}69] Rom Harré et al., editors. Some 19th century British scientists, volume 5 of Commonwealth and international library. Liberal studies division. Science and society. Pergamon, New York, NY, USA, 1969.

viii + 259 pp. LCCN ???? URL http://catalogue.bnf.fr/ark:/
12148/cb35188849c.

Hall:1931:JCM

[Hal31] Edwin H. Hall. James Clerk Maxwell. The Scientific Monthly, 33(6): 500–511, 1931. CODEN SCMOAA. ISSN 0096-3771 (print), 2327-

7513 (electronic). URL http://www.jstor.org/stable/15224.

Halevi:1980:EM

[Hal80] P. Halevi. Einstein and Maxwell. Physics Today, 33(2):15, February 1980. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Hallyn:2004:SRS

[Hal04] Fernand Hallyn. Les structures rhétoriques de la science: de Kepler à Maxwell. (French). [The rhetorical structures of science: from Kepler to Maxwell]. Des travaux. Seuil, Paris, France, 2004. ISBN 2-02-063249-7. 321 pp. LCCN Q174.8 .H35 2004.

Hall:2008:MET

[Hal08] Graham Hall. Maxwell’s electromagnetic theory and special relativity. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 366(1871):1849–1860, May 2008. CODEN PTRMAD, PTMSFB. ISSN 1364-503X (print), 1471-2962

(electronic). URL http://www.jstor.org/stable/25190792.

Harman:1982:EFM

[Har82a] P. M. (Peter Michael) Harman. Energy, force, and matter: the conceptual development of nineteenth-century physics. Cambridge history of science. Cambridge University Press, Cambridge, UK, 1982. ISBN 0-521-24600-8, 0-521-28812-6 (paperback). ix + 182 pp. LCCN QC7

.H257. URL http://www.loc.gov/catdir/description/cam032/
81017029.html;   http://www.loc.gov/catdir/enhancements/
fy0632/81017029-t.html.

Harman:1982:MNP

[Har82b] P. M. (Peter Michael) Harman. Metaphysics and Natural Philosophy: the Problem of Substance in Classical Physics. Harvester Press, Brighton, Sussex, UK, 1982. ISBN 0-389-20305-X (hardcover). xvi + 168 pp. LCCN QC7 .H26 1982.

Harman:1985:WPS

[Har85] P. M. (Peter Michael) Harman, editor. Wranglers and physicists: studies on Cambridge physics in the nineteenth century. Manchester University Press, Manchester, UK, 1985. ISBN 0-7190-1756-4. LCCN QC9.G7 W73 1985.

Harman:1988:NMP

[Har88] P. M. Harman. Newton to Maxwell: The ‘Principia’ and British physics. Notes and Records of the Royal Society of London, 42(1):75– 96, January 1988. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/531370.

Newton’s ıt Principia and its legacy.

Harman:1990:SLP

[Har90] P. M. (Peter Michael) Harman, editor. The Scientific Letters and Papers of James Clerk Maxwell. Vol. I: 1846–1862, volume I. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-25625-9 (hardcover). xxviii + 748 pp. LCCN QC21.2; QC670 .M385 1990.

URL http://www.gbv.de/dms/bowker/toc/9780521256254.pdf;
 http://www.loc.gov/catdir/description/cam023/89000452.html;
 http://www.loc.gov/catdir/enhancements/fy0732/89000452-
b.html;   http://www.loc.gov/catdir/toc/cam021/89000452.
html;   http://www.zentralblatt-math.org/zmath/en/search/
?an=0744.01034.

Harman:1993:BRB

[Har93] P. M. Harman. Book review: Bruce J. Hunt, The Maxwellians. Ithaca and London: Cornell University Press, 1991. Pp. xiii + 266. ISBN 0-8014-2641-3. $38.45. British Journal for the History of Science, 26(1):117–118, March 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.jstor.org/

stable/4027380.

Harman:1995:SLP

[Har95a] P. M. Harman, editor. The Scientific Letters and Papers of James Clerk Maxwell. Vol. II: 1862–1873, volume II. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-25626-7 (hardcover). xxx + 999 pp.

Harman:1995:TLG

[Har95b] Peter M. Harman. Through the looking-glass, and what Maxwell found there. In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], chapter 4, pages 79–93. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346. LCCN ???? URL https:/

/link.springer.com/chapter/10.1007/978-94-011-0217-9_4.

Harman:1996:BRE

[Har96] P. M. Harman. Book review: Elizabeth Carber, Stephen G. Brush and C. W. F. Everitt (eds.), Maxwell on Heat and Statistical Mechanics: On ‘Avoiding All Personal Enquiries of Molecules’. London: Associated University Presses, 1995, Pp. 550. ISBN 0-934223-34-3. £45.00. British Journal for the History of Science, 29(1):107–109, March 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X

(electronic). URL http://www.jstor.org/stable/4027529.

Harman:1998:NPJ

[Har98] P. M. (Peter Michael) Harman. The natural philosophy of James Clerk Maxwell. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-56102-7, 0-521-00585-X (paperback). xiv + 232 pp. LCCN QC16.M4 H37 1998. URL http://www.

loc.gov/catdir/description/cam028/97035703.html;   http:

//www.loc.gov/catdir/toc/cam021/97035703.html;   http://
www.zentralblattmath.org/zmath/en/search/?an=1005.01006.

Harman:2002:SLP

[Har02] P. M. Harman, editor. The Scientific Letters and Papers of James Clerk Maxwell. Vol. III: 1874–1879, volume III. Cambridge University Press, Cambridge, UK, 2002. ISBN 0-521-25627-5. xxvii + 932 pp. LCCN QC21.2; QC670 .M385 1990. URL http://www.

zentralblatt-math.org/zmath/en/search/?an=0744.01034.

Harrison:2008:MS

[Har08] Richard A. Harrison. The magnetic Sun. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 366(1871):1735–1748, May 2008. CODEN PTRMAD, PTMSFB. ISSN 1364-503X (print), 1471-2962 (electronic).

Harman:2009:MJC

[Har09] P. M. Harman. Maxwell, James Clerk (1831–1879). In Oxford Dictionary of National Biography, page ?? Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. URL https:

//www.oxforddnb.com/view/10.1093/ref:odnb/9780198614128.
001.0001/odnb-9780198614128-e-5624.

Harper:2014:BRP

[Har14] William Harper. Book review: Peter Achinstein, Evidence and Method: Scientific Strategies of Isaac Newton and James Clerk Maxwell. Philosophy of Science, 81(4):684–687, October 2014. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/678556.

Holland:2003:NRL

[HB03] Peter Holland and Harvey R. Brown. The non-relativistic limits of the Maxwell and Dirac equations: the role of Galilean and gauge invariance. Studies in History and Philosophy of Modern Physics, 34 (2):161–187, June 2003. CODEN ???? ISSN 1355-2198 (print), 1879-

2502 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219803000054.

Heaviside:1892:EP

[Hea92] Oliver Heaviside. Electrical papers. MacMillan Publishing Company, New York, NY, USA, 1892. ???? pp. LCCN QC503 .H53.

Heaviside:1894:ET

[Hea12] Oliver Heaviside. Electromagnetic theory. “The Electrician” series. “The Electrician” Printing and Publishing Company, Limited, London, UK, 1894–1912. ???? pp. LCCN QC661 .H44.

Hearnshaw:1929:CHK

[Hea29] F. J. C. (Fossey John Cobb) Hearnshaw. The centenary history of King’s College, London, 1828–1928. G. G. Harrap and Company, Ltd., London, UK, 1929. 524 + 1 pp. LCCN LF435 .H4.

Hehl:2010:CFM

[Heh10] Friedrich W. Hehl. On the changing form of Maxwell’s equations during the last 150 years — spotlights on the history of classical electrodynamics. 24 lecture slides., February 22, 2010. URL http://www.

thp.uni-koeln.de/gravitation/mitarbeiter/MaxwellUCL2.pdf.

Heimann:1970:MMC

[Hei70a] P. M. Heimann. Maxwell and the modes of consistent representation. Archive for History of Exact Sciences, 6(3):171–213, January 1970. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=6&issue=3&spage=171.

Heimann:1970:MFS

[Hei70b] P. M. Heimann. Molecular forces, statistical representation and Maxwell’s demon. Studies in History and Philosophy of Science Part A, 1(3):189–211, November 1970. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368170900099.

Heimann:1971:MHN

[Hei71] P. M. Heimann. Maxwell, Hertz, and the nature of electricity. Isis, 62(2):149–157, Summer 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229238.

Heilbron:1972:BRP

[Hei72] John L. Heilbron. Book review: Physics in My Generation by Max Born; Bahnbrecher des Atomzeitalters. Grosse Naturforscher von Maxwell bis Heisenberg by Friedrich Herneck; Moseley and the Numbering of the Elements by Bernard Jaffe; The Big Machine by Robert Jungk. Isis, 63(1):111–112, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/229207;   http://www.jstor.org/stable/pdfplus/
229207.pdf.

Heilbron:1979:ECS

[Hei79] J. L. Heilbron. Electricity in the 17th and 18th centuries: a study of early modern physics. University of California Press, Berkeley, CA, USA, 1979. ISBN 0-520-03478-3. xiv + 606 pp. LCCN QC507 .H48.

Heilbron:1988:MPL

[Hei88] J. L. Heilbron. Max Planck: Ein Leben für die Wissenschaft 1858– 1947. Mit einer Auswahl der allgemeinverständlichen Schriften von Max Planck. (German) [Max Planck: A Life in Science 1858–1947. With a selection of generally understandable writings of Max Planck]. S. Hirzel Verlag, Stuttgart, Germany, 1988. ISBN 3-7776-0392-9. 474 pp. LCCN QC16.P5 H455 1988.

Heilbron:1999:ECS

[Hei99] J. L. Heilbron. Electricity in the 17th and 18th centuries: a study of early modern physics. Dover books on physics. Dover Publications, Inc., New York, NY, USA, 1999. ISBN 0-486-40688-1 (paperback). xxx + 606 pp. LCCN QC507 .H48 1999.

Hendry:1986:JCM

[Hen86] John Hendry. James Clerk Maxwell and the theory of the electromagnetic field. Adam Hilger Ltd., Bristol, UK, 1986. ISBN 0-85274-563-X. xix + 305 pp. LCCN QC665.E4 H46 1986. URL http://www.loc.

gov/catdir/enhancements/fy0745/86196851-d.html.

Hendrickson:2015:BRR

[Hen15] Kenneth E. Hendrickson. Book reviews: Raymond Flood, Mark Mc-Cartney and Andrew Whitaker (eds.), James Clerk Maxwell: Perspectives on His Life and Work. Oxford: Oxford University Press, 2014, Pp. x + 364. ISBN 978-0-19-966437-5. £39.99. British Journal for the History of Science, 48(3):520–521, September 2015. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hertz:1884:BZM

[Her84] Heinrich Hertz. Über die Beziehungen zwischen den Maxwell schen elektrodynamischen Grundgleichungen und den Grundgleichungen der gegnerischen Elektrodynamik. (German) [On the relationships between Maxwell’s basic electrodynamic equations and the fundamental equations of opposing electrodynamics]. Annalen der Physik und Chemie, 23(??):84 103, ???? 1884. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Hesse:1963:MAS

[Hes63] Mary B. Hesse. Models and Analogies in Science, volume 14 of Newman history and philosophy of science series. Sheed and Ward, London, UK, 1963. 145 pp. LCCN ????

Hesse:1966:MAS

[Hes66] Mary B. Hesse. Models and Analogies in Science. University of Notre Dame Press, Notre Dame, IN, USA, 1966. 184 pp. LCCN Q175 .H43.

Hesse:1970:MAS

[Hes70] Mary B. Hesse. Models and analogies in science. University of Notre Dame Press, Notre Dame, IN, USA, 1970. 184 pp. LCCN ????

Hesse:1986:KMA

[Hes86] Mary B. Hesse. Kagaku moderu anarojī. (Japanese) [Models and analogies in science]. Baifukan, Tokyo, Japan, 1986. ISBN 4-563-02036-2. 200 pp. LCCN ???? Translation to Japanese by Kiyoshi Takada.

Hon:2012:MCA

[HG12] Giora Hon and Bernard R. Goldstein. Maxwell’s contrived analogy: an early version of the methodology of modeling. Studies in History and Philosophy of Modern Physics, 43(4):236–257, November 2012. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1355219812000524.

Hoffmann:2008:MPA

[Hof08] Dieter Hoffmann, editor. Max Planck: Annalen Papers. Wiley-VCH, Weinheim, Germany, 2008. ISBN 3-527-40819-3 (hardcover). xv + 857

pp. LCCN QC484 .M39 2008. URL http://d-nb.info/987214098/

04. All of the references in this book to volumes of Beiblätter zu den Annalen der Physik are too large by 10: change XXII and XXIII to XII and XIII, respectively.

Hood:1987:WHM

[Hoo87] C. Gregory Hood. Whatever happened to Maxwell? American Journal of Physics, 55(7):616–618, July 1987. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Hopley:1956:CMC

[Hop56] Ian B. Hopley. Clerk Maxwell’s Contribution to Physics. Ph.D. dissertation, University of London, London, UK, 1956. URL http://

search.proquest.com/pqdtglobal/docview/1797583302. Two vol-

umes.

Hopley:1957:CMA

[Hop57a] I. B. Hopley, M.Sc., Ph.D. Clerk Maxwell’s apparatus for the measurement of surface tension. Annals of Science, 13(3):180–187, 1957. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hopley:1957:MWE

[Hop57b] I. B. Hopley, M.Sc., Ph.D. Maxwell’s work on electrical resistance I. The determination of the absolute unit of resistance. Annals of Science, 13(4):265–272, 1957. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hopley:1958:MWE

[Hop58] I. B. Hopley, M.Sc., Ph.D. Maxwell’s work on electrical resistance. II. Proposals for the re-determination of the B.A. unit of 1863. Annals of Science, 14(3):197–210, 1958. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hopley:1959:MDN

[Hop59a] I. B. Hopley, M.Sc., Ph.D. Maxwell’s determination of the number of electrostatic units in one electromagnetic unit of electricity. Annals of Science, 15(2):91–108, 1959. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hopley:1959:MWE

[Hop59b] I. B. Hopley, M.Sc., Ph.D. Maxwell’s work on electrical resistance III. Improvement on Mance’s method for the measurement of battery resistance. Annals of Science, 15(1):51–55, 1959. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hossenfelder:2016:RMD

[Hos16] S. Hossenfelder. The remote Maxwell demon as energy downconverter. Foundations of Physics, 46(4):505–516, April 2016. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/s10701-015-
9981-7.

Houston:2016:BWS

[Hou16] Louis M. Houston. A de Broglie wave solution to the Maxwell– Schrödinger–Einstein equations. Theor. Math. Appl., 6(3):1–13, 2016. ISSN 1792-9687 (print), 1792-9709 (electronic).

Huggins:1869:FOS

[Hug69] William Huggins. Further observations on the spectra of some of the stars and nebulæ, with an attempt to determine therefrom whether these bodies are moving towards or from the Earth, also observations on the spectra of the Sun and of Comet II. Philosophical transactions of the Royal Society of London, 158(??):529–564, ???? 1869. CODEN PTRSAV. ISSN 0370-2316 (print), 2053-9207 (electronic). Contains [Cle69e].

Hunt:1984:ME

[Hun84] Bruce James Hunt. The Maxwellians (Electromagnetism). Ph.D. thesis, The Johns Hopkins University, Baltimore, MD, USA, 1984. URL

http://search.proquest.com/pqdtglobal/docview/303320139.

Hunt:1988:BRB

[Hun88a] Bruce J. Hunt. Book review: James Clerk Maxwell and the Theory of the Electromagnetic Field by John Hendry. Isis, 79(4):735–736, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234808.

Hunt:1988:BRJ

[Hun88b] Bruce J. Hunt. Book review: James Clerk Maxwell and the Theory of the Electromagnetic Field by John Hendry. Isis, 79(4):735–736, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234808.

Hunt:1989:BRJ

[Hun89] Bruce J. Hunt. Book review: J. G. O’Hara & Willibald Pricha. Hertz and the Maxwellians: A Study and Documentation of the Discovery of Electromagnetic Wave Radiation, 1873–1894. London: Peter Peregrinus Ltd., 1987. Pp. xiv + 154. ISBN 0-86341-101-0. £24.00. British Journal for the History of Science, 22(3):392–393, September 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/stable/4026909.

Hunt:1991:M

[Hun91] Bruce J. Hunt. The Maxwellians. Cornell University Press, Ithaca, NY, USA, 1991. ISBN 0-8014-2641-3 (hardcover), 0-8014-8234-8 (paperback). x + 266 pp. LCCN QC670.H939.

Hunt:1992:BRI

[Hun92] Bruce J. Hunt. Book review: Innovation in Maxwell’s Electromagnetic Theory: Molecular Vortices Displacement Current, and Light, by Daniel M. Siegel, Cambridge U. P., New York, 1991. 225 pp. $49.50. ISBN 0-521-35365-3. Physics Today, 45(8):68, August 1992. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Hunt:1994:M

[Hun94] Bruce J. Hunt. The Maxwellians. Cornell history of science series. Cornell University Press, Ithaca, NY, USA, 1994. ISBN 0-8014-8234-8, 0-8014-2641-3. x + 266 pp. LCCN ????

Hunt:2000:BRB

[Hun00a] Bruce J. Hunt. Book review: The Natural Philosophy of James Clerk Maxwell by P. M. Harman. Isis, 91(2):370–371, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/236982.

Hunt:2000:BRN

[Hun00b] Bruce J. Hunt. Book review: The Natural Philosophy of James Clerk Maxwell by P. M. Harman. Isis, 91(2):370–371, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/236982.

Hunt:2005:BRP

[Hun05a] Bruce J. Hunt. Book reviews: P. M. Harman (ed.), The Scientific Letters and Papers of James Clerk Maxwell. Volume I: 1846–1862. Cambridge: Cambridge University Press, 1990. Pp. xxviii + 748. ISBN 0-521-25625-9. $195.00 (hardback). Volume II: 1862–1873. Cambridge: Cambridge University Press, 1995. Pp. xxx + 999. ISBN 0-521-25626-7. $285.00 (hardback). Volume III: 1873–1879. Cambridge: Cambridge University Press, 2002. Pp. xxvii + 932. ISBN 0-521-25627-5. £210.00, $315.00 (hardback). British Journal for the History of Science, 38(4):485–487, December 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.jstor.

org/stable/4028469.

Hunt:2005:M

[Hun05b] Bruce J. Hunt. The Maxwellians. Cornell University Press, Ithaca, NY, USA, 2005. ISBN 0-8014-8234-8 (paperback). x + 266 pp. LCCN QC670 .H86 2005.

Hunt:2006:BRB

[Hun06] Bruce J. Hunt. Book review: Basil Mahon, The Man Who Changed Everything: The Life of James Clerk Maxwell. Chichester: John Wiley, 2003. Pp. xx + 226. ISBN 0-470-86088-X. £18.99 (hardback). British Journal for the History of Science, 39(2):303, June 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

http://www.jstor.org/stable/4028627.

Hunt:2015:MMM

[Hun15] Bruce J. Hunt. Maxwell, measurement, and the modes of electromagnetic theory. Historical Studies in the Natural Sciences, 45(2):303– 339, April 2015. CODEN ???? ISSN 1939-1811 (print), 1939-182X

(electronic). URL http://www.jstor.org/stable/10.1525/hsns.
2015.45.2.303.

Hutchinson:1998:JCM

[Hut98] Ian Hutchinson. James Clerk Maxwell and the Christian proposition. Web document., January 1998. URL http://silas.psfc.mit.edu/

maxwell/.

Ivezic:2005:PME

[Ive05] Tomislav Ivezić. The proof that Maxwell equations with the 3D E and B are not covariant upon the Lorentz transformations but upon the standard transformations: The new Lorentz invariant field equations. Foundations of Physics, 35(9):1585–1615, September 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:

//link.springer.com/article/10.1007/s10701-005-6484-y.

James:1991:ERM

[Jam91] Frank A. J. L. James. Essay reviews: Maxwell’s early career [P. M. Harman, The scientific letters and papers of James Clerk Maxwell. Volume 1. 1846–1862. Cambridge University Press, 1990. Pp. xxvii + 748, £125.00. ISBN 0-521-25625-9]. Notes and Records of the Royal Society of London, 45(2):266–270, July 1991. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/531705.

James:1991:CMF

[Jam12] Frank A. J. L. James, editor. The correspondence of Michael Faraday. [History of technology series]. Institution of Electrical Engineers, London, UK, 1991–2012. ISBN 0-86341-248-3 (vol. 1), 0-86341-249-1 (vol. 2), 0-86341-250-5 (vol. 3), 0-86341-251-3 (vol. 4). ???? pp. LCCN QC16.F2 A4 1991.

James:2015:BRN

[Jam15] Frank A. J. L. James. Book review: Nancy Forbes and Basil Mahon, Faraday, Maxwell, and the Electromagnetic Field: How Two Men Revolutionized Physics, 320 pp., illus., bibl., index. Amherst, N.Y.: Prometheus Books, 2014. $25.95 (cloth). Isis, 106(2):462–463, June 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/682792.

Jeans:1931:JCM

[Jea31] James Jeans. James Clerk Maxwell’s method. In Thomson [Tho31b], page ?? LCCN QC16.M4 J3. Essays by Sir J. J. Thomson, Max Planck, Albert Einstein, Joseph Larmor, James Jeans, William Garnett, Ambrose Fleming, Oliver Lodge, R. T. Glazebrook, and Horace Lamb.

Jensen:1970:XCF

[Jen70] J. Vernon Jensen. The X Club: Fraternity of Victorian scientists. British Journal for the History of Science, 5(1):63–72, June 1970. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/stable/4025353.

Jones:1932:LWJ

[Jon32] Edward Taylor Jones. The life and work of James Clerk Maxwell 1831–1879. Proceedings of the Royal Philosophical Society of Glasgow, 60(??):54–77, ???? 1932. CODEN PHIJBC. ISSN 0031-8078.

Jones:1949:JCM

[Jon49] R. V. Jones. James Clerk Maxwell. Aberdeen University Review, 33 (??):3–19, ???? 1949. ISSN 0001-320X.

Jones:1964:BRC

[Jon64] R. V. Jones. Book review: Clerk Maxwell and Modern Science. Edited by C. Domb. London: Athlone Press. 1963. Pp. 118. Physics Bulletin, 15(2):41, 1964. CODEN PHSBB4. ISSN 0031-9112. URL http://

stacks.iop.org/0031-9112/15/i=2/a=007.

Jones:1973:JCM

[Jon73] R. V. Jones, F.R.S. James Clerk Maxwell at Aberdeen, 1856–1860. Notes and Records of the Royal Society of London, 28(1):57–81, June 1973. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531113. See erratum

[Ano74].

Jones:1980:CPJ

[Jon80] R. V. Jones. The complete physicist: James Clerk Maxwell 1831– 1879. In Yearbook of the Royal Society of Edinburgh, page ???? Royal Society of Edinburgh, Edinburgh, Scotland, 1980.

Jones:1992:BRS

[Jon92] D. S. Jones. Book review: The Scientific Letters and Papers of James Clerk Maxwell Volume 1: 1846–1862. (P. M. Harman ed.). SIAM Review, 34(3):511–513, September 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1993:BRM

[Jor93] D. W. Jordan. Book review: The Maxwellians by Bruce J. Hunt. Isis, 84(3):595–596, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235697.

Josefowicz:2007:BRT

[Jos07] Diane Greco Josefowicz. Book review: Thomas K. Simpson: Figures of Thought: A Literary Appreciation of Maxwell’s. Isis, 98(2):417, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/521486.

Josefowicz:2014:BRN

[Jos14] Diane Greco Josefowicz. Book review: Newton, Maxwell, Marx: Spirit, Freedom, and the Scientific Vision. Isis, 105(2):422–423, June 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/677990.

Jones:1994:CBJ

[JTR94] Jean Jones, Hugh S. Torrens, and Eric Robinson. The correspondence between James Hutton (1726–1797) and James Watt (1736– 1819) with two letters from Hutton to George Clerk-Maxwell (1715– 1784): Part I. Annals of Science, 51(6):637–653, 1994. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Jones:1995:CBJ

[JTR95] Jean Jones, Hugh S. Torrens, and Eric Robinson. The correspondence between James Hutton (1726–1797) and James Watt (1736– 1819) with two letters from Hutton to George Clerk-Maxwell (1715– 1784). Part II. Annals of Science, 52(4):357–382, 1995. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kaiser:1987:ROB

[Kai87] Walter Kaiser. Rezensionen: Oskar Blumtritt: Zur Genese der elektromagnetischen Feldtheorie: Eine wissenschaftshistorische Studie zur Entstehung der Maxwellschen Gleichungen. (arbor scientiarum, Beiträge zur Wissenschaftsgeschichte, Reihe A: Abhandlungen, Bd 10). Hildesheim: Gerstenberg Verlag 1986. X und 413 Seiten. Berichte zur Wissenschaftsgeschichte, 10(4):240–242, ???? 1987. CODEN BE-WID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Karam:2014:BRP

[Kar14] Ricardo Karam. Book review: Peter Achinstein: Evidence and Method: Scientific Strategies of Isaac Newton and James Clerk Maxwell. Science & Education (Springer), 23(10):2137–2148, October 2014. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic).

Karam:2014:CTA

[KCP14] Ricardo Karam, Debora Coimbra, and Maurício Pietrocola. Comparing teaching approaches about Maxwell’s displacement current. Science & Education (Springer), 23(8):1637–1661, August 2014. CO-DEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic).

Kim:2002:LCH

[Kim02] Dong-Won Kim. Leadership and Creativity: A History of the Cavendish Laboratory, 1871–1919, volume 5 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 2002. ISBN 90-481-5956-3 (hardcover), 94-017-2055-X (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). xxiv + 225 pp. LCCN QC51.G72 K45 2002.

Kim:2003:MEH

[Kim03] Young Sam Kim. Maxwell’s equations and a historical study of incorporating them into undergraduate mathematics and engineering education. Ph.D. thesis, Columbia University, New York, NY, USA,

2003. URL http://search.proquest.com/pqdtglobal/docview/
305336978.

KlerkMaksuellia:1888:TTV

[Kle88] Dzheĭms Klerk Maksuellia. Teoriia teploty v elementarnoi obrabotkie Klerk Maksuelia. (Ukrainian) [Theory of heat in elementary processes Clerk Maxwell]. Tip. I. N. Kushnereva i ko., Kiev, Ukraine, 1888. vi

+ 292 pp. URL https://hdl.handle.net/2027/uc1.%24b564085.

Translation to Ukrainian by A. L. Korol’kova.

KlerkMaksvell:1989:TOE

[Kle89] Dzheĭms Klerk Maksvell. Traktat ob Èlektrichestve i Magnetizme. Tom I, II. (Russian) [Treatise on electricity and magnetism, volume 1 and 2]. Klassiki Nauki [Classics of Science]. Nauka, Moscow, Russia, 1989. ISBN 5-02-000042-6. 416 (vol. I), 437 (vol. II) pp. URL https:

//hdl.handle.net/2027/uc1.b3764366. Translated from the third

English edition by B. M. Bolotovskiĭ, I. L. Burshteĭn, M. A. Miller and E. V. Suvorov, Translation edited and with an afterword by M. L. Levin, Miller and Suvorov, with a preface by Bolotovskiĭ, Burshteĭn, Miller, Suvorov and Levin.

Klein:2010:PMS

[Kle10] Maury Klein. The power makers: steam, electricity, and the men who invented modern America. Bloomsbury Press, New York, NY, USA, 2010. ISBN 1-59691-834-9 (e-book). 638 pp.

LCCN T20 .K585 2010. URL http://excerpts.cdn.overdrive.
com/FormatType-410/2183-1/304/09D/B9/PowerMakers.epub;
http://rbdigital.oneclickdigital.com.

Kashtalyan:2008:TDA

[KM08] M. Kashtalyan and M. Menshykova. Three-dimensional analysis of a functionally graded coating/substrate system of finite thickness. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 366(1871):1821–1826, May 2008. CODEN PTRMAD, PTMSFB. ISSN 1364-503X (print), 1471-2962 (electronic).

Knott:1911:LSW

[Kno11] Cargill Gilston Knott, editor. Life and Scientific Work of Peter Guthrie Tait, Supplementing the Two Volumes of Scientific Papers Published in 1898 and 1900. Cambridge University Press, Cambridge, UK, 1911. x + 379 pp.

Knott:1914:JCM

[Kno14] Cargill Gilston Knott. James Clerk Maxwell. In Famous Edinburgh students, pages 175–179. ????, Edinburgh, Scotland, 1914. LCCN ????

Kozameh:1984:MEL

[KNP84] Carlos N. Kozameh, Ezra T. Newman, and John R. Porter. Maxwell’s equations, linear gravity, and twistors. Foundations of Physics, 14 (11):1061–1081, November 1984. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01882489.

Knudsen:1978:EDD

[Knu78] Ole Knudsen. Electric displacement and the development of optics after Maxwell. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 22(1):53–60, September 1978. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kragh:2013:BRP

[Kra13] Helge Kragh. Book review: Peter Achinstein: Evidence and Method: Scientific Strategies of Isaac Newton and James Clerk Maxwell. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 44(2):405–408, December 2013. CODEN JG-PSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http:/

/link.springer.com/article/10.1007/s10838-013-9238-0.

Kragh:2018:LLH

[Kra18] Helge Kragh. Ludvig Lorenz and his non–Maxwellian electrical theory of light. Physics in Perspective (PIP), 20(3):221–253, September 2018. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Krakovsky:2021:NTH

[Kra21] Marina Krakovsky. News: Taking the heat: Maxwell’s demon and the high cost of erasure. Communications of the ACM, 64(6):18–20, June 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317

(electronic). URL https://dl.acm.org/doi/10.1145/3460214.

Kox:1995:NTE

[KS95] A. J. Kox and Daniel M. Siegel. No Truth Except in the Details: Essays in Honor of Martin J. Klein, volume 167 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1995. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346. xxv + 382

+ 13 + 1 pp. LCCN ???? URL https://link.springer.com/book/
10.1007/978-94-011-0217-9.

Khlestkov:2016:ESE

[KS16] Yu. A. Khlestkov and L. A. Sukhanova. Exact solutions to the Einstein–Maxwell equations describing wormholes and handles. Foundations of Physics, 46(6):668–688, June 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-016-9998-6.

Kuwaki:1931:RPF

[Kuw31] Ayao Kuwaki. The revolution in physics by Faraday and Maxwell. Proc. Phys.-Math. Soc. Japan, 5(??):??, ???? 1931.

Klein:1973:MBR

[KWK73] Martin J. Klein, Hugh C. Wolfe, and J. Kestin. The Maxwell– Boltzmann relationship. AIP Conference Proceedings, 11(1):297–??, 1973. CODEN APCPCS. ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465.

Laing:1974:MDC

[Lai74] Richard Laing. Maxwell’s demon and computation. Philosophy of Science, 41(2):171–178, June 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186866.

Lamb:1931:CML

[Lam31] Horace Lamb. Clerk Maxwell as lecturer. In Thomson [Tho31b], page ?? LCCN QC16.M4 J3. Essays by Sir J. J. Thomson, Max Planck, Albert Einstein, Joseph Larmor, James Jeans, William Garnett, Ambrose Fleming, Oliver Lodge, R. T. Glazebrook, and Horace Lamb.

Lambton:2003:ASW

[Lam03] Christopher Lambton. Appliance of science. Why isn t James Clerk Maxwell as famous as he deserves to be? The Scotsman, ??(??):9a–9b, October 4, 2003.

Lambert:2011:UAJ

[Lam11] Kevin Lambert. The uses of analogy: James Clerk Maxwell’s ‘On Faraday’s lines of force’ and early Victorian analogical argument.

British Journal for the History of Science, 44(1):61–88, 2011. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). See [Cle58a].

Land:1998:CCP

[Lan98a] M. C. Land. The classical Coulomb problem in pre-Maxwell electrodynamics. Foundations of Physics, 28(9):1489–1497, September 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/
A%3A1018865413499.

Land:1998:PME

[Lan98b] M. C. Land. Pre-Maxwell electrodynamics. Foundations of Physics, 28(9):1479–1487, September 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1023/A%3A1018813429428.

Land:1998:PMQ

[Lan98c] M. C. Land. Pre-Maxwell quantum electrodynamics. Foundations of Physics, 28(9):1499–1506, September 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1018817530337.

Lancashire:2003:SPJ

[Lan03] Ian Lancashire. Selected poetry of James Clerk Maxwell (1831–1879).

Web site, 2003. URL https://tspace.library.utoronto.ca/html/
1807/4350/poet400.html.

Landini:2014:APR

[Lan14] M. Landini. About the physical reality of ‘Maxwell’s displacement current’ in classical electrodynamics. Progress in Electromagnetics Research (PIER), 144:329–343, 2014. CODEN PELREX. ISSN 1070-4698 (print), 1559-8985 (electronic). URL http://www.jpier.org/

PIER/pier.php?paper=13111501.

Larmor:1900:AM

[Lar00] Joseph Larmor. Aether and matter. Cambridge University Press, Cambridge, UK, 1900. ???? pp. LCCN QC671 .L32.

Larmor:1907:MSC

[Lar07] Joseph Larmor, editor. Memoir and scientific correspondence of the late Sir George Gabriel Stokes, Bart. Cambridge University Press, Cambridge, UK, 1907. ???? pp. LCCN QC16.S8 A2. Two volumes.

Larmor:1908:LEH

[Lar08] Sir Joseph Larmor. LXVIII. On the equilibrium of heterogeneous substances (1876). Philosophical Magazine (Series 6 (1901 - 1925)), 16(95):818–824, ???? 1908. CODEN PHMAA4. ISSN 1941-5982 (print), 1941-5990 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1080/14786441108636559. This is a reprint of [Cle76j].

Larmor:1926:JCM

[Lar26] Sir Joseph Larmor. James Clerk Maxwell. Nature, 118(2981 (supplement)):44–47, December 18, 1926. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.

com/nature/journal/v118/n2981supp/pdf/118044a0.pdf.

Larmor:1931:SEC

[Lar31] Joseph Larmor. The scientific environment of Clerk Maxwell. In Thomson [Tho31b], page ?? LCCN QC16.M4 J3. Essays by Sir J. J. Thomson, Max Planck, Albert Einstein, Joseph Larmor, James Jeans, William Garnett, Ambrose Fleming, Oliver Lodge, R. T. Glazebrook, and Horace Lamb.

Larmor:1936:OCM

[Lar36] Joseph Larmor. The origins of Clerk Maxwell’s electric ideas, as described in letters to W. Thomson. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 32(??):695–750, ???? 1936. CODEN PCPSA4. ISSN 0008-1981.

Larmor:1937:OCM

[Lar37] Sir Joseph Larmor, editor. Origins of Clerk Maxwell’s electric ideas, as described in familiar letters to William Thomson. Cambridge University Press, Cambridge, UK, 1937. iii + 56 pp. LCCN QC3 .M52.

URL https://hdl.handle.net/2027/wu.89055051577.

Larmor:1971:MSC

[Lar71] Joseph Larmor, editor. Memoir and scientific correspondence of the late Sir George Gabriel Stokes, Bart., volume 139 of The Sources of Science. Johnson Reprint Corp., New York, NY, USA, 1971. ???? pp. LCCN QC16.S8 A2 1971.

Larmor:2010:MSC

[Lar10] Joseph Larmor, editor. Memoir and Scientific Correspondence of the Late Sir George Gabriel Stokes, Bart.: Selected and Arranged by Joseph Larmor, volume 2 of Cambridge Library Collection - Physical Sciences. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-511-70712-6 (e-book), 1-108-00892-5 (paperback). 520 pp. LCCN ????

Lightman:2004:DNC

[LB^{+}04] Bernard (Bernard V.) Lightman, John (John Hedley) Brooke, et al., editors. The dictionary of nineteenth-century British scientists. Thoemmes Continuum, Bristol, UK, 2004. ISBN 1-85506-999-7 (set). xxxix + 2256 (4 volumes) pp. LCCN Q141.

Lutz:2015:IMD

[LC15] Eric Lutz and Sergio Ciliberto. Information: From Maxwell’s demon to Landauer’s eraser. Physics Today, 68(9):30–35, September 2015. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Langevin:1912:TDR

[LdB12a] Paul Langevin and Louis de Broglie, editors. La théorie du rayonnement et les quanta: rapports et discussions de la réunion tenue à Bruxelles du 30 octobre au 3 novembre 1911 / sous les auspices de m. E. Solvay. Gauthier-Villars, Paris, France, 1912. URL https://

archive.org/details/lathoriedurayo00inst.

Langevin:1912:GWH

[LdB12b] Paul Langevin and Louis de Broglie, editors. La théorie du rayonnement et les quanta: rapports et discussions de la réunion tenue à Bruxelles du 30 octobre au 3 novembre 1911/sous les auspices de m. E. Solvay. Gauthier-Villars, Paris, France, 1912. URL https://

archive.org/details/lathoriedurayo00inst.

Leff:2002:MDS

[Lef02] Harvey S. Leff. Maxwell’s demon and the Second Law [of thermodyamics]. AIP Conference Proceedings, 643(1):408–??, 2002. CO-DEN APCPCS. ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465.

Lemons:2017:SGD

[Lem17] Don S. (Don Stephen) Lemons. A Student’s Guide to Dimensional Analysis. Cambridge University Press, Cambridge, UK, 2017. ISBN 1-107-16115-0 (hardcover), 1-316-61381-X (paperback), 1-316-67616-1 (e-book). xii + 102 pp. LCCN TA347.D5 L46 2017.

Lenard:1929:GNG

[Lén29] Philipp Lénard. Grosse Naturforscher: eine Geschichte der Naturforschung in Lebensbeschreibungen. (German) [Great men of science: a history of scientific research in life descriptions]. J. F. Lehmann, München, Germany, 1929. 324 + 16 pp. LCCN Q141 .L6.

Lenard:1933:GMSa

[Lén33] Philipp Lénard. Great men of science: a history of scientific progress. G. Bell and Sons Ltd., London, 1933. xix + 389 pp. LCCN Q141 .L62.

Lenard:1943:GNG

[Lén43] Philipp Lénard. Grosse Naturforscher: eine Geschichte der Naturforschung in Lebensbeschreibungen. (German) [Great men of science: a history of scientific research in life descriptions]. J. F. Lehmann, München, sixth edition, 1943. 348 pp. LCCN Q141 .L6 1943.

Lenard:1970:GMS

[Len70] Philipp Lenard. Great men of science: a history of scientific progress. Essay index reprint series. Books for Libraries Press, Freeport, NY, USA, 1970. ISBN 0-8369-1614-X. xx + 389 pp. LCCN Q141 .L62 1970.

Lentze:2019:DCM

[Len19] Georg Lentze. Dialogue concerning magnetic forces. Studies in History and Philosophy of Modern Physics, 68(??):158–162, November 2019. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1355219819300528.

Lowther:2008:ARW

[LF08] D. A. Lowther and E. M. Freeman. The application of the research work of James Clerk Maxwell in electromagnetics to industrial frequency problems. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 366(1871): 1807–1820, May 2008. CODEN PTRMAD, PTMSFB. ISSN 1364-503X (print), 1471-2962 (electronic). URL http://www.jstor.org/

stable/25190787.

Lenard:1933:GMSb

[LH33] Philipp Lenard and H. (Henry) Hatfield. Great men of science: a history of scientific progress. MacMillan Publishing Company, New York, NY, USA, 1933. xx + 1 + 389 pp. LCCN Q141 .L62 1933a. URL

http://books.google.com/books?id=ntk5AAAAMAAJ;   http://
catalog.hathitrust.org/api/volumes/oclc/1394475.html.

Lindsay:1971:BRM

[Lin71] R. Bruce Lindsay. Book review: The Maxwell Equations: Non-Relativistic and Relativistic Derivations from Electron Theory, by S. R. DeGroot, 179 pp., Interscience, New York, 1969, $8.95. Physics Today, 24(7):47, July 1971. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Litchfield:1903:RBL

[Lit03] Henrietta Litchfield. Richard Buckley Litchfield. [privately printed], London, UK, 1903. ???? pp. LCCN ????

Litchfield:1910:RBL

[Lit10] Henrietta Litchfield. Richard Buckley Litchfield. [privately printed], London, UK, 1910. ???? pp. LCCN ????

Lin:2019:MPM

[LL19] Tsung-Wu Lin and Hejie Lin. Mechanical proof of the Maxwell speed distribution. International Journal of Statistics and Probability, 8(2): 90–??, ???? 2019. CODEN ???? ISSN 1927-7032 (print), 1927-7040

(electronic). URL http://www.ccsenet.org/journal/index.php/
ijsp/article/view/0/38323.

Levin:1981:MTE

[LM81] M. L. Levin and M. A. Miller. Maxwell’s Treatise on Electricity and Magnetism. Soviet Physics. Uspekhi, 24(11):904–913, November 1981. CODEN SOPUAP. ISSN 0038-5670. URL http://ufn.ru/en/

articles/1981/11/c/.

Lu:2014:EMD

[LMJ14] Zhiyue Lu, Dibyendu Mandal, and Christopher Jarzynski. Engineering Maxwell’s demon. Physics Today, 67(8):60–61, August 2014. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). See comments [She15].

Lodge:1927:CMC

[Lod27] Oliver Lodge. Clerk Maxwell and the Cavendish Laboratory. Nature, 119(2984):46, January 8, 1927. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.

com/nature/journal/v119/n2984/pdf/119046a0.pdf.

Lodge:1931:CMW

[Lod31] Oliver Lodge. Clerk Maxwell and wireless telegraphy. In Thomson [Tho31b], page ?? LCCN QC16.M4 J3. Essays by Sir J. J. Thomson, Max Planck, Albert Einstein, Joseph Larmor, James Jeans, William Garnett, Ambrose Fleming, Oliver Lodge, R. T. Glazebrook, and Horace Lamb.

Longair:2003:TCP

[Lon03] Malcolm S. Longair. Theoretical concepts in physics: an alternative view of theoretical reasoning in physics. Cambridge University Press, Cambridge, UK, second edition, 2003. ISBN 0-521-52878-X (paperback), 0-521-82126-6. xvii + 569 pp. LCCN QC20 .L64 2003. URL

http://www.loc.gov/catdir/description/cam031/2002073612.
html;   http://www.loc.gov/catdir/samples/cam041/2002073612.
html;   http://www.loc.gov/catdir/toc/cam031/2002073612.
html.

Longair:2008:MSC

[Lon08] Malcolm S. Longair. Maxwell and the science of colour. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 366(1871):1685–1696, May 2008. CODEN PTR-MAD, PTMSFB. ISSN 1364-503X (print), 1471-2962 (electronic).

URL http://www.jstor.org/stable/25190777.

Longair:2016:MEL

[Lon16] Malcolm Longair. Maxwell’s Enduring Legacy. Cambridge University Press, Cambridge, UK, 2016. ISBN 1-107-08369-9. xxi + 664

pp. LCCN ???? URL http://www.cambridge.org/us/academic/
subjects/physics/general-and-classical-physics/maxwells-
enduring-legacy-scientific-history-cavendish-laboratory.

Lorentz:1923:CME

[Lor23] Hendrik A. Lorentz. Clerk Maxwell’s electromagnetic theory. Cambridge University Press, Cambridge, UK, 1923. 35 pp. Rede lecture for 1923.

Leff:2003:MDE

[LR03] Harvey S. Leff and Andrew F. Rex, editors. Maxwell’s Demon 2: entropy, classical and quantum information, computing. Institute of Physics Publishers, Bristol, UK, 2003. ISBN 0-7503-0759-5. xvi + 485 pp. LCCN QA318.M35 M38 2003. URL http://www.loc.

gov/catdir/enhancements/fy0806/2007299804-d.html;   http:
//www.loc.gov/catdir/toc/fy0802/2007299804.html.

Lovatt:2014:WWK

[LS14] Ian Lovatt and M. Qasim Syed. Why was Kelvin’s estimate of the Earth’s age wrong? The Physics Teacher, 52(5):286–287, May 2014. CODEN PHTEAH. ISSN 0031-921X (print), 1943-4928 (electronic).

Laporte:1931:ASA

[LU31] Otto Laporte and George E. Uhlenbeck. Application of spinor analysis to the Maxwell and Dirac equations. Physical Review, 37(11):1380– 1397, June 1931. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Lucky:2012:MSR

[Luc12] Robert Lucky. Is math still relevant? [reflections]. IEEE Spectrum, 49 (3):23, March 2012. CODEN IEESAM. ISSN 0018-9235 (print), 1939-

9340 (electronic). URL https://ieeexplore.ieee.org/document/
6156859.

Macaulay:1885:JCM

[Mac85] James Macaulay. James Clerk Maxwell, F.R.S. In Short biographies for all the people, by various writers, page ?? ????, London, UK, 1885. LCCN ????

Macfarlane:1919:LTB

[Mac19] Alexander Macfarlane. Lectures on ten British physicists of the Nineteenth Century, volume 20 of Mathematical monographs. Wiley, New York, NY, USA, 1919. 144 pp. LCCN QC15 .M3. URL https://

archive.org/stream/lecturesontenbri00macfrich/lecturesontenbri00macfrich_
djvu.txt.

MacDonald:1964:FMK

[Mac64] D. K. C. MacDonald. Faraday, Maxwell, and Kelvin, volume 28 of Science study series. Heinemann, London, UK, 1964. 143 pp. LCCN ????

MacGregor:2015:BRJ

[Mac15] Peter MacGregor. Book review: James Clerk Maxwell, edited by [Raymond Flood,] Marc McCartney and Andrew Whitaker, pp 364, £39.99 (hard), ISBN 978-0-19-966437-5, Oxford University Press (2014). The Mathematical Gazette, 99(546):564, November 2015. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Maddox:2002:MDS

[Mad02] Sir John Maddox. Maxwell’s demon: Slamming the door. Nature, 417(6892):903, June 27, 2002. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Mahon:2003:MWC

[Mah03] Basil Mahon. The man who changed everything: the life of James Clerk Maxwell. Wiley, New York, NY, USA, 2003. ISBN 0-470-86088-X, 0-470-86171-1 (paperback). xx + 226 + 16 pp. LCCN QC16.M4

M34 2003. URL http://www.loc.gov/catdir/bios/wiley047/
2004426225.html;   http://www.loc.gov/catdir/description/
wiley042/2004426225.html;   http://www.loc.gov/catdir/toc/
wiley041/2004426225.html.

Mahon:2015:HME

[Mah15] Basil Mahon. How Maxwell’s equations came to light. Nature Photonics, 9:2–4, January 2015. CODEN NPAHBY. ISSN 1749-4885 (print), 1749-4893 (electronic). URL http://adsabs.harvard.edu/

abs/2015NaPho...9....2M.

Martens:1966:JCM

[Mar66] Robert Martens. J. Cl. Maxwell’s Gleichungen des denudativen Feldes: Über Anfänge der theoretischen Geographie um 1870. (German) [James Clerk Maxwell’s equations of the naked field: Beginning of theoretical geography around 1870]. Sudhoffs Archiv: Zeitschrift für Wissenschaftsgeschichte, 50(3):286–296, ???? 1966. CODEN

SUARAH. ISSN 0039-4564. URL http://www.jstor.org/stable/
20775518.

Marsden:1993:BRD

[Mar93] Ben Marsden. Book review: Daniel M. Siegel, Innovation in Maxwell’s Electromagnetic Theory: Molecular Vortices, Displacement Current, and Light. Cambridge: Cambridge University Press, 1991. Pp. x + 225. ISBN 0-521-35365-3. £30.00, $49.50. British Journal for the History of Science, 26(1):116–117, March 1993. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4027379.

March:1999:BRM

[Mar99a] Robert H. March. Book review: Maxwell’s Demon: Why Warmth Disperses and Time Passes, by Hans Christian von Baeyer, Random House, New York, 1998. 207 pp. $25.00 ISBN 0-679-43342-2. Physics Today, 52(3):84, March 1999. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Martin:1999:MMO

[Mar99b] Linda Martin, editor. 1000 makers of the millennium: One thousand makers of the millennium. DK Pub., New York, NY, USA, 1999. ISBN 0-7894-4709-6. 256 pp. LCCN CT107 .A16 1999.

Marston:2007:MCA

[Mar07] Philip L. Marston. Maxwell and creation: Acceptance, criticism, and his anonymous publication. American Journal of Physics, 75(8):731– 740, August 2007. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Mathieu:1985:AOB

[Mat85a] J.-P. Mathieu. Analyses d’ouvrages: The Demon in the Aether. The story of James Clerk Maxwell par Martin Goldman. Revue d’Histoire des Sciences, 38(2):173, avril–juin 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.

org/stable/23632521.

Mathieu:1985:AOD

[Mat85b] J.-P. Mathieu. Analyses d’ouvrages: The Demon in the Aether. The story of James Clerk Maxwell par Martin Goldman. Revue d’Histoire des Sciences, 38(2):173, avril–juin 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.

org/stable/23632521.

Matthews:1987:SME

[Mat87] N. F. J. Matthews. On the strength of Maxwell’s equations. Journal of Mathematical Physics, 28(4):810–814, April 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i4/p810_s1.

Mauel:2017:AJC

[Mau17] Michael E. Mauel. Announcement: The 2016 James Clerk Maxwell Prize for Plasma Physics. Physics of Plasmas, 24(5):055401, May 2017. CODEN PHPAEN. ISSN 1070-664X (print), 1089-7674 (electronic), 1527-2419.

Mayr:1971:MOC

[May71] Otto Mayr. Maxwell and the origins of cybernetics. Isis, 62(4):424– 444, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/229816.

McCormmach:1969:BRA

[McC69] Russell McCormmach. Book review: Astronomy and Physics: The Electrical Researches of the Honourable Henry Cavendish. Ed. by James Clerk Maxwell. Reprint of 1879 edition. London: Frank Cass & Co. Pp. lxvi + 454. 1967. £6 6s. British Journal for the History of Science, 4(4):408–409, December 1969. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.jstor.

org/stable/4025240.

McCartney:2011:PLJ

[McC11] Mark McCartney. The poetic life of James Clerk Maxwell. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 26(1):29–43, 2011. CODEN ???? ISSN 1749-8430 (print), 1749-

8341 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17498430903516138.

McDonald:2018:MOL

[McD18] Kirk T. McDonald. Maxwell’s objection to Lorenz’s retarded potentials. Report, Joseph Henry Laboratories, Princeton University, Princeton, NJ 08544, USA, April 9, 2018. URL http://

www.physics.princeton.edu/~mcdonald/examples/maxwell.pdf.

Original version dates 26 October 2009.

McNat:2004:JCM

[McN04] Jerrold L. McNat. James Clerk Maxwell’s refusal to join the Victoria Institute. Perspectives on Science and Christian Faith, 56(3):204–215, September 2004. ISSN 0892-2675. URL http://www.asa3.org/ASA/

PSCF/2004/PSCF9-04McNatt.pdf.

Mendonca:2008:MCW

[Men08] J. T. Mendonça. Maxwell and the classical wave particle dualism. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 366(1871):1771–1780, May 2008. CODEN PTRMAD, PTMSFB. ISSN 1364-503X (print), 1471-2962

(electronic). URL http://www.jstor.org/stable/25190784.

Menshykov:2008:BIE

[MGM08] O. V. Menshykov, I. A. Guz, and V. A. Menshykov. Boundary integral equations in elastodynamics of interface cracks. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 366(1871):1835–1839, May 2008. CODEN PTR-MAD, PTMSFB. ISSN 1364-503X (print), 1471-2962 (electronic).

Mirotznik:1999:BME

[Mir99] M. S. Mirotznik. Bringing Maxwell’s equations to heel [software reviews]. IEEE Spectrum, 36(8):82–84, August 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mitchell:2017:MSA

[Mit17a] Daniel Jon Mitchell. Making sense of absolute measurement: James Clerk Maxwell, William Thomson, Fleeming Jenkin, and the invention of the dimensional formula. Studies in History and Philosophy of Modern Physics, 58(??):63–79, May 2017. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1355219815300393.

Mitchell:2017:WNR

[Mit17b] Daniel Jon Mitchell. What’s nu? A re-examination of Maxwell’s ‘ratio-of-units’ argument, from the mechanical theory of the electromagnetic field to ‘On the elementary relations between electrical measurements’. Studies in History and Philosophy of Science Part A, 65–66(??):87–98, October/December 2017. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0039368117301607.

Morrison:1991:BRB

[Mor91a] Margaret Morrison. Book review: James Clerk Maxwell and the Theory of the Electromagnetic Field by John Hendry. Philosophy of Science, 58(3):505–507, September 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187950.

Morrison:1991:BRJ

[Mor91b] Margaret Morrison. Book review: James Clerk Maxwell and the Theory of the Electromagnetic Field by John Hendry. Philosophy of Science, 58(3):505–507, September 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187950.

Morrison:1992:STU

[Mor92] Margaret Morrison. A study in theory unification: The case of Maxwell’s electromagnetic theory. Studies in History and Philosophy of Science Part A, 23(1):103–145, March 1992. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368192900285.

Anonymous:1995:HYM

[Mor95] Dennis Moralee. A hundred years and more of Cambridge physics. Cambridge University Physics Society, Cambridge, UK, third edition, 1995. ISBN 0-9507343-1-4. 47 pp. LCCN ????

Moue:2008:TEM

[Mou08] Aspasia S. Moue. The thought experiment of Maxwell’s demon and the origin of irreversibility. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 39(1):69–84, September 2008. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/40390690.

Moyer:1977:EDH

[Moy77] Donald Franklin Moyer. Energy, dynamics, hidden machinery: Rankine, Thomson and Tait, Maxwell. Studies in History and Philosophy of Science Part A, 8(3):251–268, ???? 1977. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368177900085.

Moyer:1978:CMF

[Moy78] Donald Franklin Moyer. Continuum mechanics and field theory: Thomson and Maxwell. Studies in History and Philosophy of Science Part A, 9(1):35–50, March 1978. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0039368178900201.

Muller:2004:MLW

[Mul04] F. A. Muller. Maxwell’s lonely war. Studies in History and Philosophy of Modern Physics, 35(1):109–119, March 2004. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S135521980300090X.

Muller:2008:DCE

[Mul08] F. A. Muller. In defence of constructive empiricism: Maxwell’s master argument and aberrant theories. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 39(1):131–156, September 2008. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/40390694.

Murray:2007:SDR

[Mur07] Carl D. Murray. Saturn’s dynamical rings. Physics Today, 60(8): 74–75, August 2007. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Musser:1999:TMD

[Mus99] George Musser. Taming Maxwell’s Demon. Scientific American, 280(2):24, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n2/pdf/scientificamerican0299-
24.pdf.

Myers:2006:BRF

[Mye06] Greg Myers. Book review: Fernand Hallyn: Les structures rhétoriques de la science: De Kepler à Maxwell. Isis, 97(2):339–340, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/507353.

Myrvold:2011:SMT

[Myr11] Wayne C. Myrvold. Statistical mechanics and thermodynamics: a Maxwellian view. Studies in History and Philosophy of Modern Physics, 42(4):237–243, November 2011. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219811000487.

Nieuwenhuizen:2002:UMD

[NA02] Theo M. Nieuwenhuizen and Armen E. Allahverdyan. Unmasking Maxwell’s demon. AIP Conference Proceedings, 643(1):436–??, 2002. CODEN APCPCS. ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465.

Nahin:1983:AGO

[Nah83] Paul J. Nahin. Among the giants: Oliver Heaviside: genius and curmudgeon: he called establishment mathematicians ‘woodenheaded’ when his powerful new methods perplexed them. His fellow EEs didn’t know what to make of him either. IEEE Spectrum, 20(7):63–69, July 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Nahin:1992:MGU

[Nah92] P. J. Nahin. Maxwell’s grand unification. IEEE Spectrum, 29(3):45, March 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Nappo:2021:DNM

[Nap21] Francesco Nappo. The double nature of Maxwell’s physical analogies. Studies in History and Philosophy of Science Part A, 89(??):212–225, October 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-

2510 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368121001199.

Natanson:1932:DLM

[Nat32] Władysław Natanson. Dzieciece lata Maxwella. (polish) [Ten years in Maxwell’s life]. Przeglad Współczesny, 126(??):??, October 1932.

Nersessian:1984:FEC

[Ner84] Nancy J. Nersessian. Faraday to Einstein: constructing meaning in scientific theories. Science and philosophy. M. Nijhoff, Dordrecht, The Netherlands, 1984. ISBN 90-247-2997-1, 0-7923-0950-2 (paperback). xiv + 196 pp. LCCN Q175 .N3885 1984.

Nersessian:1987:BRM

[Ner87] Nancy J. Nersessian. Book review: From Maxwell to Microphysics: Aspects of Electromagnetic Theory in the Last Quarter of the Nineteenth Century by Jed Z. Buchwald. Philosophy of Science, 54 (3):489–490, September 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187592.

Newman:1955:JCM

[New55] James R. Newman. James Clerk Maxwell. Scientific American, 192(6):58–71, June 1955. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v192/n6/pdf/scientificamerican0655-
58.pdf.

Niven:1890:SPJa

[Niv90a] William Davidson Niven, editor. The Scientific papers of James Clerk Maxwell, volume 1. Cambridge University Press, Cambridge, UK,

1890. xxx + 607 pp. URL http://www.archive.org/download/
scientificpapers01maxwuoft/scientificpapers01maxwuoft.pdf;
 http://www.biodiversitylibrary.org/item/59553;   https:/
/hdl.handle.net/2027/umn.31951p011043735.

Niven:1890:SPJb

[Niv90b] William Davidson Niven, editor. The Scientific papers of James Clerk Maxwell, volume 2. Cambridge University Press, Cambridge, UK,

1890. viii + 806 pp. URL http://www.archive.org/download/
scientificpapers02maxwuoft/scientificpapers02maxwuoft.pdf;
 http://www.biodiversitylibrary.org/item/59554;   https:/
/hdl.handle.net/2027/nyp.33433087548644.

Niven:1890:SPJc

[Niv90c] William Davidson Niven, editor. The Scientific papers of James Clerk Maxwell, volume 1. Librairie Scientifique J. Hermann, Paris, France, 1890. xxx + 607 pp.

Niven:1890:SPJd

[Niv90d] William Davidson Niven, editor. The Scientific papers of James Clerk Maxwell, volume 2. Librairie Scientifique J. Hermann, Paris, France,

1890. viii + 806 (vol. 2) pp. URL https://hdl.handle.net/2027/
msu.31293102595331. Two volumes.

Niven:1927:SPJ

[Niv27] William Davidson Niven, editor. The Scientific papers of James Clerk Maxwell. Librairie Scientifique J. Hermann, Paris, France, 1927. xxix + 3 + 607 (vol. 1), vii + 806 (vol. 2) pp. Two volumes.

Niven:1952:SPJ

[Niv52] William Davidson Niven, editor. The Scientific papers of James Clerk Maxwell. Dover Publications, Inc., New York, NY, USA, 1952. xxix + 3 + 607 (vol. 1), vii + 806 (vol. 2) pp. URL https://hdl.handle.

net/2027/mdp.39015040426655. Two volumes.

Niven:1965:SPJ

[Niv65] William Davidson Niven, editor. The Scientific papers of James Clerk Maxwell. Dover Publications, Inc., New York, NY, USA,

1965. URL https://hdl.handle.net/2027/mdp.39076006363597;
 https://hdl.handle.net/2027/mdp.49015000116476.

Niven:2003:SPJ

[Niv03] Sir W. D. (William Davidson) Niven, editor. The scientific papers of James Clerk Maxwell. Dover phoenix editions. Dover Publications, Inc., New York, NY, USA, 2003. ISBN 0-486-49560-4 (vol. 1), 0-486-49561-2 (vol. 2). xxix + 3 + 607 (vol. 1), vii + 806 (vol. 2) pp. LCCN

QC3 .M5 2003. URL http://www.loc.gov/catdir/enhancements/
fy0614/2003055158-d.html.

Niven:2011:SPJ

[Niv11] William Davidson Niven, editor. The scientific papers of James Clerk Maxwell. Cambridge University Press, Cambridge, UK, 2011. ISBN 1-108-01225-6 (vol. 1), 1-108-01538-7 (vol. 2), 1-108-01227-2 (set). 607 pp. Two-volume set. Reprint of the 1890 hardcover edition.

Norton:2005:ELL

[Nor05] John D. Norton. Eaters of the lotus: Landauer’s principle and the return of Maxwell’s Demon. Studies in History and Philosophy of Modern Physics, 36(2):375–411, June 2005. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1355219804000851.

Nugayev:2016:MEG

[Nug16] Rinat Magdievich Nugayev. Maxwellian electrodynamics genesis and development: Intertheoretic context. Spontaneous Generations: A Journal for the History and Philosophy of Science, 8 (1):55–92, ???? 2016. CODEN ???? ISSN 1913-0465. URL

https://spontaneousgenerations.library.utoronto.ca/index.
php/SpontaneousGenerations/article/view/19467.

Navarro:2019:LTN

[NV19] Juan F. Navarro and Juan Vargas. On the long term numerical integration of planetary rings. Journal of Computational and Applied Mathematics, 354(??):390–401, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0377042718304229.

Nye:2002:CHS

[Nye02] Mary Jo Nye, editor. The Cambridge History of Science. Volume 5. The Modern Physical and Mathematical Sciences. The Cambridge History of Science. Cambridge University Press, Cambridge, UK, 2002. ISBN 0-521-57199-5 (hardcover), 1-139-05352-3 (e-book). xxix + 678 pp. LCCN Q125 .C32 2003.

Ogilvie:2008:JCMa

[Ogi08a] Gordon I. Ogilvie. James Clerk Maxwell and the dynamics of astrophysical discs. arXiv.org, ??(??):??, February 10, 2008. URL https:

//arxiv.org/abs/0802.1333.

Ogilvie:2008:JCMb

[Ogi08b] Gordon I. Ogilvie. James Clerk Maxwell and the dynamics of astrophysical discs. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 366 (1871):1707–1715, May 2008. CODEN PTRMAD, PTMSFB. ISSN 1364-503X (print), 1471-2962 (electronic). URL http://rsta.

royalsocietypublishing.org/content/366/1871/1707;   http:
//www.jstor.org/stable/25190779.

OHara:1975:GJS

[O’H75] J. G. O’Hara. George Johnstone Stoney, F.R.S. and the concept of the electron. Notes and Records of the Royal Society of London, 29(2):265–276, March 1975. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/531468.

Opatrny:2011:EMD

[OM11] Tomás Opatrný and Zuzana Misáková. Exorcising Maxwell’s demon from Liboff’s three-channel conundrum. Foundations of Physics, 41 (2):261–269, February 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-010-9505-4.

OHara:1987:HMS

[OP87] J. G. (James G.) O’Hara and W. (Willibald) Pricha. Hertz and the Maxwellians: a study and documentation of the discovery of electromagnetic wave radiation, 1873–1894, volume 8 of IEE History of technology series. P. Peregrinus Ltd. in association with the Science Museum, London, UK, 1987. ISBN 0-86341-101-0. xiv + 154 pp. LCCN QC670 .O47 1987.

OConnor:1997:JCM

[OR97a] J. J. O’Connor and E. F. Robertson. James Clerk Maxwell. Web document., November 1997. URL http://www-groups.dcs.

st-and.ac.uk/~history/Biographies/Maxwell.html;   https:/
/web.archive.org/web/20110128034939/.

OConnor:1997:VJC

[OR97b] J. J. O’Connor and E. F. Robertson. A visit to James Clerk Maxwell’s house, November 1997. Web document., November 1997.

URL http://www-history.mcs.st-andrews.ac.uk/HistTopics/
Maxwell_House.html.

Oziewicz:1994:CFT

[Ozi94] Zbigniew Oziewicz. Classical field theory and analogy between Newton’s and Maxwell’s equations. Foundations of Physics, 24(10):1379– 1402, October 1994. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/
10.1007/BF02283039.

Page:1938:RJL

[Pag38] Leigh Page. Review: Joseph Larmor, Origins of Clerk Maxwell’s Electric Ideas as Described in Familiar Letters to William Thomson. Bulletin of the American Mathematical Society, 44(5):320, 1938. CO-DEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

URL http://projecteuclid.org/euclid.bams/1183500463.

Peierls:1963:FTS

[Pei63] Rudolf Peierls. Field theory since Maxwell. In Clerk Maxwell and modern science; six commemorative lectures [Dom63], pages 26–42. LCCN QC71 .D63.

Pickover:2008:AHL

[Pic08] Clifford A. Pickover. Archimedes to Hawking: laws of science and the great minds behind them. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-533611-9 (hardcover). ix + 514 pp. LCCN Q175.32.R45 P53 2008. URL http://catdir.loc.

gov/catdir/toc/ecip087/2007051167.html;   http://www.gbv.
de/dms/goettingen/555633241.pdf;   http://www.oup.com/us/
catalog/general/subject/Physics/?view=usa&ci=9780195336115.

Pippard:1999:BRG

[Pip99] Brian Pippard, F.R.S. Book review: Genius in context: P. M. Harman, The natural philosophy of James Clerk Maxwell. Cambridge University Press, 1998. Pp. xiv + 232, £35/US$59.95 (hardback). ISBN 0-521-56102-7. Notes and Records of the Royal Society of London, 53(2):279–281, May 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/532212.

Pippard:2004:BRP

[Pip04] Brian Pippard, FRS. Book review: A paragon of physics: Basil Mahon, The man who changed everything — the life of James Clerk Maxwell. Wiley, Chichester, 2003. Pp. xx + 226, £18.99 (hardback). ISBN 0-470-86088-X. Notes and Records of the Royal Society of London, 58(1):103–104, January 22, 2004. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.

org/stable/4142040.

Planck:1894:BMG

[Pla94] Max Planck. Über den Beweis des Maxwellschen Geschwindigkeitsverteilungsgesetzes unter Gasmolekülen. (German) [On the evidence for Maxwell’s law of velocity distribution of gas molecules]. Sitzungsberichte der Mathematisch-Naturwissenschaftlichen Abteilung der Bayerischen Akademie der Wissenschaften zu München, 24(??):391– 394, ???? 1894. CODEN SMNBAC. ISSN 0376-1037.

Planck:1895:BMG

[Pla95] Max Planck. Ueber den Beweis des Maxwell’schen Geschwindigkeitsvertheilungsgesetzes unter Gasmolecülen. (German) [On the evidence for Maxwell’s law of velocity distribution of gas molecules]. Annalen der Physik (Berlin), 291(5):220–222, 1895. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Planck:1898:MTE

[Pla98] Max Planck. Über die Maxwellsche Theorie der Elektrizität. (German) [On Maxwell’s theory of electricity]. Naturwissenschaftliche Rundschau, 13(42):541, October 15, 1898. CODEN NARSAC. ISSN 0028-1050.

Planck:1899:MTE

[Pla99] Max Planck. Die Maxwellsche Theorie der Elektrizität von der mathematischen Seite betrachtet. (German) [Maxwell’s theory of electricity from the mathematical side]. Jahresbericht der Deutschen Mathematiker-Vereinigung, 7(1):77–89, ???? 1899. ISSN 0012-0456 (print), 1869-7135 (electronic).

Planck:1903:MMT

[Pla03] Max Planck. Metalloptik und Maxwell’sche Theorie. (German) [Metal optics and Maxwell’s theory]. Sitzungsberichte der Königlich Preussischen Akademie der Wissenschaften, 26(??):558–559, May 14, 1903. CODEN SPWPAI. ISSN 0371-2435.

Planck:1931:JCM

[Pla31a] Max Planck. James Clerk Maxwell in seiner Bedeutung für die theoretische Physik in Deutschland. (German) [James Clerk Maxwell and his importance to theoretical physics in Germany]. Die Naturwissenschaften, 19(44):889–894, October 30, 1931. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://link.

springer.com/article/10.1007/BF01516331.

Planck:1931:MIT

[Pla31b] Max Planck. Maxwell’s influence on theoretical physics in Germany. In Thomson [Tho31b], page ?? LCCN QC16.M4 J3. Essays by Sir J. J. Thomson, Max Planck, Albert Einstein, Joseph Larmor, James Jeans, William Garnett, Ambrose Fleming, Oliver Lodge, R. T. Glazebrook, and Horace Lamb.

Poincare:1899:TMO

[Poi99] Henri Poincaré. La Théorie de Maxwell et les oscillations Hertziennes. (French) [Maxwell’s theory and Hertzian waves]. Scientia, Chartres, France, 1899. 80 pp. LCCN ????

Poincare:1907:TMO

[Poi07] Henri Poincaré. La Théorie de Maxwell et les oscillations Hertziennes. (French) [Maxwell’s theory and Hertzian waves]. ????, Paris, France, 1907. ???? pp. LCCN ????

Polak:1985:MRF

[Pol85] L. S. (Lev Solomonovich) Polak, editor. Maksvell i razvitie fiziki XIX– XX vekov. (Russian) [Maxwell and the development of physics XIX– XX centuries]. Nauka, Moscow, Russia, 1985. 244 + 1 pp. LCCN QC19.6 .M35 1985. 1.80rub.

Porter:1981:SSG

[Por81] Theodore M. Porter. A statistical survey of gases: Maxwell’s social physics. Historical Studies in the Physical Sciences, 12(1):77–116, ???? 1981. CODEN HSPSAS. ISSN 0073-2672. URL http://www.

jstor.org/stable/27757490.

Porter:1994:QMS

[Por94] Theodore M. Porter. From Quetelet to Maxwell: Social statistics and the origins of statistical physics. In Cohen [Coh94], chapter 11, pages 345–362. ISBN 0-7923-2223-1, 94-017-3391-0 (e-book). ISSN 0068-0346. LCCN Q174 .B67 vol. 150. URL https://link.springer.

com/chapter/10.1007/978-94-017-3391-5_11.

Pitowsky:1996:LFM

[PS96] Itamar Pitowsky and Noam Shoresh. Locality, factorizability, and the Maxwell Boltzmann distribution. Foundations of Physics, 26(9):1231– 1242, September 1996. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/
10.1007/BF02275627.

Puppin:1925:LDH

[Pup25] M. I. Puppin. Law, description and hypothesis in electrical science. Sci., 62(??):17–22, ???? 1925.

Poincare:1904:MTW

[PV04] Henri Poincaré and Frederick K. Vreeland. Maxwell’s theory and wireless telegraphy. Part one. Maxwell’s theory and Hertzian oscillations. Part two. The principles of wireless telegraphy. ????, London, 1904. xiii + 1 + 255 pp. LCCN ???? Translation from French of Part One by Frederick K. Vreeland. Part Two by Frederick K. Vreeland.

Pyle:1996:OMP

[Pyl96] Andrew Pyle, editor. The Origins of Modern Philosophy of Science, 1830–1914. Routledge/Thoemmes Press, London, UK, 1996. ISBN 0-415-13267-3 (set). xvi + vii + 372 (vol. 1), cxx + 523 (vol. 2), iv + 586 (vol. 3), x + vii + 351 (vol. 4), xi–xiv + 348 (vol. 5), 291 (vol. 6), ix + 128 (vol. 7), xi–xvi + 463 (vol. 8), vii + 480 (vol. 9), ix + 271 (vol. 10), xlii + 380 (vol. 11), xi + 288 (vol. 12) pp. LCCN QA9 .C63 1996. 12 volumes.

Qadir:2015:SPJ

[QM15] A. Qadir and D. P. Mason. Sesquicentennial of the presentation by James Clerk Maxwell of his paper “A Dynamical Theory of the Electromagnetic Field” to the Royal Society of London. In International Journal of Modern Physics Conference Series, volume 38 of International Journal of Modern Physics Conference Series, page 1560070. ????, ????, April 2015. URL http://adsabs.harvard.edu/

abs/2015IJMPS..3860070Q.

Randall:1962:ALW

[Ran62] Sir John Randall, F.R.S. Aspects of the life and work of James Clerk Maxwell. Nature, 195(4840):427–434, August 4, 1962. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v195/n4840/pdf/
195427a0.pdf.

Randall:1963:ALW

[Ran63] J. Randall. Aspects of the life and work of James Clerk Maxwell. In Clerk Maxwell and modern science; six commemorative lectures [Dom63], page ?? LCCN QC71 .D63.

Rapoport:2005:CWD

[Rap05] Diego L. Rapoport. Cartan–Weyl Dirac and Laplacian operators, Brownian motions: The quantum potential and scalar curvature, Maxwell’s and Dirac–Hestenes equations, and supersymmetric systems. Foundations of Physics, 35(8):1383–1431, August 2005. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/s10701-005-
6443-7.

Rautio:2014:LRM

[Rau14a] J. C. Rautio. The long road to Maxwell’s equations. IEEE Spectrum, 51(12):36–56, December 2014. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rautio:2014:CSL

[Rau14b] James C. Rautio. Cover story: The long road to Maxwell’s equations: How four enthusiasts helped bring the theory of electromagnetism to light. IEEE Spectrum, 51(12):36–40, 54–56, December 1, 2014. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340

(electronic). URL https://spectrum.ieee.org/magazine/2014/
December;   https://spectrum.ieee.org/tech-history/dawn-
of-electronics/the-long-road-to-maxwells-equations.

Rodrigues:2007:MFM

[RdO07] W. A. Rodrigues and Edmundo Capelas de Oliveira. The many faces of Maxwell, Dirac and Einstein equations: a Clifford bundle approach, volume 722 of Lecture notes in physics. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 3-540-71292-5 (hardcover), 3-540-71293-3 (e-book). ISSN 0075-8450. xiv + 445 pp. LCCN QC173.59.S65 R63 2007. URL

http://www.loc.gov/catdir/enhancements/fy0826/2007923174-
d.html;   http://www.loc.gov/catdir/enhancements/fy0826/
2007923174-t.html.

Rodrigues:2016:MFM

[RdO16] W. A. Rodrigues and Edmundo Capelas de Oliveira. The many faces of Maxwell, Dirac and Einstein equations: a Clifford bundle approach, volume 922 of Lecture notes in physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2016. ISBN 3-319-27636-0 (print), 3-319-27637-9 (e-book). ISSN 0075-8450 (print), 1616-6361 (electronic). xvi + 587 pp. LCCN

QC173.59.S65. URL http://link.springer.com/10.1007/978-3-
319-27637-3.

Reid:2008:JCMb

[Rei08] John S. Reid. James Clerk Maxwell’s Scottish chair. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 366(1871):1661–1684, May 2008. CODEN PTR-MAD, PTMSFB. ISSN 1364-503X (print), 1471-2962 (electronic).

URL http://www.jstor.org/stable/25190776.

Reid:2015:JCM

[Rei15] John S. Reid. James Clerk Maxwell’s class of 1856/57. arXiv.org, ??

(??):??, April 2015. URL https://arxiv.org/abs/1504.07587.

Rex:2017:MDH

[Rex17] Andrew Rex. Maxwell’s Demon — a historical review. Entropy, 19 (6):240:1–240:13, May 2017. CODEN ENTRFG. ISSN 1099-4300.

Rey:1921:CQD

[Rey21] A. Rey. La contribution que les divers pays ont donnée aux progrès de la physique. 1. Physique Newtonnienne et physique de Fresnel, Maxwell, Clausius. (French) [the contribution that the various countries have made in the progress of physics. 1. Newtonian physics and the physics of Fresnel, Maxwell, Clausius]. Scientia, 15(29):345–360, May 1, 1921.

Rice:2014:BRJ

[Ric14] Adrian Rice. Book review: James Clerk Maxwell. Perspectives on his life and work, by Raymond Flood, Mark McCartney and Andrew Whitaker (eds). BSHM Bulletin: Journal of the British Society for the History of Mathematics, 29(3):219–221, 2014. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/17498430.2014.919433.

Riddell:1930:CMG

[Rid30] J. Oliver Riddell. Clerk Maxwell of Glenlair: a centenary booklet. Adam Rae, Bookseller, Castle-Douglas, Scotland, 1930. 47 + 2 pp. LCCN ????

Rinaldi:2008:SWU

[Rin08] Giancarlo Rinaldi. The science world’s unsung hero? BBC News Channel Web site., November 25, 2008. URL http://news.bbc.co.

uk/2/hi/uk_news/scotland/south_of_scotland/7746365.stm.

Roche:1987:BRJ

[Roc87] John Roche. Book review: J. Z. Buchwald. From Maxwell to Microphysics. Chicago: University of Chicago Press, 1985. Pp. xv + 339. ISBN 0-226-07882-5. £59.50. British Journal for the History of Science, 20(2):229–230, April 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.jstor.org/

stable/4026315.

Rodgers:2002:TAA

[Rod02] Glen E. Rodgers. Traveling with the atom: Allegheny College: James Clerk Maxwell, Scottish physicist (1831–1879). Web site, 2002.

Rodgers:2019:TAS

[Rod19] Glen E. Rodgers. Travelling with the Atom: a Scientific Guide to Europe and Beyond. Royal Society of Chemistry, Cambridge, UK, 2019. ISBN 1-78801-528-2 (paperback), 1-78801-702-1 (e-book). xxxii + 551 pp. LCCN QC171.2 .R63 2020.

Rowe:2002:PJC

[Row02] Ray Rowe. The poetry of James Clerk Maxwell. Physics World, 15(5):64, May 2002. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://stacks.iop.org/2058-7058/
15/i=5/a=42.

Rowlinson:2005:BRM

[Row05] J. S. Rowlinson, FRS, FREng. Book review: Mixed mathematics: Andrew Warwick, Masters of theory: Cambridge and the rise of mathematical physics. University of Chicago Press, 2003. Pp. xiv + 572, $29.00 (paperback). ISBN 0-226-87375-7. Notes and Records of the Royal Society of London, 59(1):94–96, January 22, 2005. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL

http://www.jstor.org/stable/30041476.

Rauen:2011:PEM

[RS11] Kenneth M. Rauen and Daniel P. Sheehan. The Proell Effect: A macroscopic Maxwell’s demon. AIP Conference Proceedings, 2011. CODEN APCPCS. ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465.

Ringe:1992:CLC

[RSB^{+}92] Dagmar Ringe, Barry L. Stoddard, John Bruhnke, Paula Koenigs, and Ned Porter. Can Laue catch Maxwell?: Observation of short-lived species by Laue X-ray crystallography. Philosophical Transactions of the Royal Society of London: Physical Sciences and Engineering, 340 (1657):273–284, 1992. ISSN 0962-8428 (print), 2054-0299 (electronic).

URL http://www.jstor.org/stable/53893.

Reid:2008:JCMa

[RWT08] John S. Reid, Charles H.-T. Wang, and J. Michael T. Thompson. James Clerk Maxwell 150 years on. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 366(1871):1651–1659, May 2008. CODEN PTRMAD, PTMSFB. ISSN 1364-503X (print), 1471-2962 (electronic). URL http://www.

jstor.org/stable/25190775.

Rynasiewicz:2002:FUM

[Ryn02] Robert Rynasiewicz. Field unification in the Maxwell–Lorentz theory with absolute space. Philosophy of Science, 70(5):1063–1072, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/10.1086/
377389.

Rynasiewicz:2016:BRH

[Ryn16] Robert Rynasiewicz. Book review: Howard J. Fisher. Maxwell’s “Treatise on Electricity and Magnetism”: The Central Argument. Isis, 107(4):877–878, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Saatsi:2005:RFM

[Saa05] Juha Saatsi. Reconsidering the Fresnel–Maxwell theory shift: how the realist can have her cake and EAT it too. Studies in History and Philosophy of Science Part A, 36(3):509–538, September 2005. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0039368105000506.

Sommerfeld:1950:PMS

[SB50] Arnold Sommerfeld and Fritz Bopp. Zum Problem der Maxwellschen Spannungen. (German) [The problem of Maxwell’s voltages]. Annalen der Physik (1900), 443(1–2):41–45, 1950. CODEN ANPYA2. ISSN 1521-3889.

Sun:2008:UDH

[SBB^{+}08] H. Sun, P. W. Benzie, N. Burns, D. C. Hendry, M. A. Player, and J. Watson. Underwater digital holography for studies of marine plankton. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 366(1871):1789–1806, May 2008. CODEN PTRMAD, PTMSFB. ISSN 1364-503X (print), 1471-2962 (electronic).

Schuster:1926:LR

[Sch26] Sir Arthur Schuster. Lord Rayleigh. Nature, 118(2981 (supplement)):47–49, December 18, 1926. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.

com/nature/journal/v118/n2981supp/pdf/118044a0.pdf.

Schneider:1989:BRM

[Sch89] Ivo Schneider. Book review: Maxwell on Molecules and Gases by Elizabeth Garber; Stephen G. Brush; C. W. F. Everitt. Isis, 80(3):535– 536, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234977.

Schaffer:1994:REE

[Sch94] Simon Schaffer. Rayleigh and the establishment of electrical standards. European Journal of Physics, 15(6):277–285, November 1994. CODEN EJPHD4. ISSN 0143-0807 (print), 1361-6404 (elec-

tronic). URL https://iopscience.iop.org/article/10.1088/
0143-0807/15/6/001.

Schubring:1996:HGG

[Sch96] Gert Schubring, editor. Hermann Günther Graßmann (1809–1877): Visionary Mathematician, Scientist and Neohumanist Scholar: Papers from a Sesquicentennial Conference, volume 187 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1996. ISBN 90-481-4758-1, 94-015-8753-1 (e-book). ISSN 0068-0346. LCCN QA8.9-

10.3. URL https://link.springer.com/book/10.1007/978-94-
015-8753-2.

Schweller:2014:MDG

[Sch14] Randall L. Schweller. Maxwell’s demon and the golden apple: global discord in the new millennium. The Johns Hopkins University Press, Baltimore, MD, USA, 2014. ISBN 1-4214-1277-2 (hardcover), 1-4214-1278-0 (electronic). xiv + 196 pp. LCCN JZ1318 .S366 2014.

Schiesser:2017:SCM

[Sch17] William E. Schiesser. Spline Collocation Methods for Partial Differential Equations: with Applications in R. Wiley, New York, NY, USA, 2017. ISBN 1-119-30103-3 (hardcover), 1-119-30105-X (PDF), 1-119-30104-1 (ePub), 1-119-30106-8 (online). xv + 549 pp. LCCN QA377 .S355 2017.

Scully:2007:DQP

[Scu07] Robert J. Scully. The Demon and the Quantum: from the Pythagorean Mystics to Maxwell’s Demon and Quantum Mystery. Wiley, New York, NY, USA, 2007. ISBN 3-527-40688-3 (paperback). 271 pp. LCCN QC174.12 S38 2007.

Shukla:2008:LIE

[SE08] Padma Kant Shukla and Bengt Eliasson. Localization of intense electromagnetic waves in plasmas. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 366 (1871):1757–1769, May 2008. CODEN PTRMAD, PTMSFB. ISSN 1364-503X (print), 1471-2962 (electronic).

Segala:1992:BRP

[Seg92] Marco Segala. Book review: Peter M. Harman, The scientific letters and papers of James Clerk Maxwell, 1846–1862, vol. I, Cambridge, Cambridge University Press, 1990, xxviii + 748 pp. (ISBN 0-521-25625-9). Nuncius, 7(1):278–281, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic).

URL http://booksandjournals.brillonline.com/content/10.
1163/182539192x00352.

Seitz:2001:JCM

[Sei01] Frederick Seitz. James Clerk Maxwell (1831–1879); member APS 1875. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 145(1):1–44, 2001. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

http://www.jstor.org/stable/1558323.

Selvan:2009:RSP

[Sel09] K. T. Selvan. A revisiting of scientific and philosophical perspectives on Maxwell’s displacement current. IEEE Antennas Propagat. Mag., 51(??):36–46, June 2009.

Sen:2021:EFH

[Sen21] Paul Sen. Einstein’s fridge: how the difference between hot and cold explains the universe. Scribner, New York, NY, USA, 2021. ISBN 1-5011-8130-0 (hardcover), 1-5011-8132-7 (e-book). xii + 305 pp. LCCN QC311 .S5118 2021.

Sheehan:2002:PSP

[SGD^{+}02] D. P. Sheehan, J. Glick, T. Duncan, J. A. Langton, M. J. Gagliardi, and R. Tobe. Phase space portraits of an unresolved gravitational Maxwell demon. Foundations of Physics, 32(3):441–462, March 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/
A%3A1014813413305.

Shapiro:1973:HDM

[Sha73] I. S. Shapiro. On the history of the discovery of the Maxwell equations. Soviet Physics. Uspekhi, 15(5):651–659, May 1973. CODEN SOP-

UAP. ISSN 0038-5670. URL http://ufn.ru/en/articles/1973/5/
h/.

Shenker:1999:MDB

[She99] Orly R. Shenker. Maxwell’s Demon and Baron Munchausen: Free will as a Perpetuum Mobile. Studies in History and Philosophy of Modern Physics, 30(3):347–372, September 1999. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219899000143.

Shenker:2004:BRB

[She04a] Orly R. Shenker. Book review: Maxwell’s Demon 2: Entropy, classical and quantum information, computing: Harvey Leff and Andrew Rex (Eds.); Institute of Physics, Bristol, 2003, 500 pp., US $55, ISBN 0-7503-0759-5. Studies in History and Philosophy of Modern Physics, 35(3):537–540, September 2004. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1355219804000279. See [LR03].

Shenker:2004:BRM

[She04b] Orly R. Shenker. Book review: Maxwell’s Demon 2: Entropy, classical and quantum information, computing: Harvey Leff and Andrew Rex (Eds.); Institute of Physics, Bristol, 2003, 500 pp., US $55, ISBN 0-7503-0759-5. Studies in History and Philosophy of Modern Physics, 35(3):537–540, September 2004. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1355219804000279. See [LR03].

Sherwood:2015:MMD

[She15] Bruce Sherwood. Modeling Maxwell’s Demon. Physics Today, 68(3):9, March 2015. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). See [LMJ14].

Sheehan:2018:MZC

[She18] Daniel Peter Sheehan. Maxwell zombies: Conjuring the thermodynamic undead. American Scientist, 106(4):234–??, July/August 2018. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786

(electronic). URL https://www.americanscientist.org/article/
maxwell-zombies-conjuring-the-thermodynamic-undead.

Shorter:2014:JCM

[Sho14] Richard Shorter. James Clerk Maxwell: Perspectives on His Life and Work, edited by Raymond Flood, Mark McCartney, and Andrew Whitaker, Scope: Review. Level: general readership. Contemporary Physics, 55(4):356, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Siegel:1975:CGM

[Sie75] Daniel M. Siegel. Completeness as a goal in Maxwell’s electromagnetic theory. Isis, 66(3):361–368, September 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/228842.

Siegel:1985:BRT

[Sie85a] Daniel Siegel. Book review: I. Tolstoy, James Clerk Maxwell: A Biography and M. Goldman, The Demon in the Aether: The Life of James Clerk Maxwell. Physics Today, 38(8):66–67, August 1985. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). See comments [TS86].

Siegel:1985:BRM

[Sie85b] Daniel M. Siegel. Book review: Maxwell on Saturn’s Rings by Stephen G. Brush; C. W. F. Everitt; Elizabeth Garber. Isis, 76(1):100, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/232813.

Siegel:1985:MIR

[Sie85c] Daniel M. Siegel. Mechanical image and reality in Maxwell’s electromagnetic theory. In Harman [Har85], pages 180–201. ISBN 0-7190-1756-4. LCCN QC9.G7 W73 1985.

Siegel:1991:IME

[Sie91] Daniel M. Siegel. Innovation in Maxwell’s electromagnetic theory: molecular vortices, displacement current, and light. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-35365-3. x + 225 pp. LCCN QC670 .S48 1991. URL http://www.

loc.gov/catdir/description/cam024/90042511.html;   http:/
/www.loc.gov/catdir/toc/cam029/90042511.html.

Siegel:1993:BRB

[Sie93a] Daniel M. Siegel. Book review: The Scientific Letters and Papers of James Clerk Maxwell. Volume 1: 1846–1862 by James Clerk Maxwell; P. M. Harman. Isis, 84(4):810, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/235158.

Siegel:1993:BRS

[Sie93b] Daniel M. Siegel. Book review: The Scientific Letters and Papers of James Clerk Maxwell. Volume 1: 1846–1862 by James Clerk Maxwell; P. M. Harman. Isis, 84(4):810, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/235158.

Siegel:1995:TCM

[Sie95] Daniel M. Siegel. Text and context in Maxwell’s electromagnetic theory. In No Truth Except in the Details: Essays in Honor of Martin J. Klein [KS95], chapter 12, pages 281–297. ISBN 94-010-4097-4, 94-011-0217-1 (e-book). ISSN 0068-0346. LCCN ???? URL https:/

/link.springer.com/chapter/10.1007/978-94-011-0217-9_12.

Siegel:1996:BRM

[Sie96] Daniel Siegel. Book review: Maxwell on Heat and Statistical Mechanics: On “Avoiding All Personal Enquiries” of Molecules. Isis, 87(3):511–516, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235986.

Simpson:1966:MDE

[Sim66] Thomas K. Simpson. Maxwell and the direct experimental test of his electromagnetic theory. Isis, 57(4):411–432, Winter 1966. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/228514.

Simpson:1968:CSM

[Sim68] Thomas King Simpson. A Critical Study of Maxwell’s Dynamical Theory of the Electromagnetic Field in the “Treatise On Electricity And Magnetism”. Ph.D. dissertations, The Johns Hopkins University, Baltimore, MD, USA, December 1968. xx + 580 pp. URL https://

search.proquest.com/pqdtglobal/docview/302334022/.

Simpson:1970:SOM

[Sim70] Thomas K. Simpson. Some observations on Maxwell’s treatise on electricity and magnetism: On the role of the ‘dynamical theory of the electromagnetic field’ in part IV of the treatise. Studies in History and Philosophy of Science Part A, 1(3):249–263, November 1970. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0039368170900129.

Simon:2006:AEA

[Sim06] Dieter Simon. Albert Einstein: Akademie-Vorträge: Sitzungsberichte der Preußischen Akademie der Wissenschaften, 1914–1932. (German) [Albert Einstein: Academy lectures: Proceedings of the Prussian Academy of Sciences, 1914–1932]. Wiley-VCH, Weinheim, Germany, 2006. ISBN 3-527-40609-3 (hardcover), 3-527-60895-8 (e-book), 3-527-62244-6 (e-book). xiv + 431 pp. LCCN QC71 .E36 2006.

Simpson:2014:JMJ

[Sim14] Richard Simpson. A joint meeting of the James Clerk Maxwell Foundation and the British Society for the History of Mathematics in celebration of the 400th anniversary of the publication of John Napier’s Mirifici Logarithmorum Canonis Descriptio. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 29(3):213– 214, 2014. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.1080/
17498430.2014.948348.

Smith:1963:SER

[Smi63] R. A. Smith. Stimulated emission of radiation and its practical application to masers and lasers. In Clerk Maxwell and modern science; six commemorative lectures [Dom63], page ?? LCCN QC71 .D63.

Smith:1982:BRB

[Smi82a] Crosbie Smith. Book review: James Clerk Maxwell: A Biography by Ivan Tolstoy. Isis, 73(3):480, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/231509.

Smith:1982:BRJ

[Smi82b] Crosbie Smith. Book review: James Clerk Maxwell: A Biography by Ivan Tolstoy. Isis, 73(3):480, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/231509.

Smith:1985:BRM

[Smi85] Crosbie Smith. Book review: Martin Goldman, The demon in the aether. The story of James Clerk Maxwell. Edinburgh: Paul Harris Publishing, 1983. Pp. 224. ISBN 0-86228-026-5. £18. British Journal for the History of Science, 18(1):120–121, March 1985. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://

www.jstor.org/stable/4026300.

Smith:2000:BRB

[Smi00a] C. Smith. Book review: The Natural Philosophy of James Clerk Maxwell. P. M. Harman. The British Journal for the Philosophy of Science, 51(3):531–534, September 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://

bjps.oxfordjournals.org/content/51/3/531.full.pdf+html;
http://www.jstor.org/stable/3541783.

Smith:2000:BRN

[Smi00b] C. Smith. Book review: The Natural Philosophy of James Clerk Maxwell. P. M. Harman. The British Journal for the Philosophy of Science, 51(3):531–534, September 2000. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://

bjps.oxfordjournals.org/content/51/3/531.full.pdf+html;
http://www.jstor.org/stable/3541783.

Smith:2002:FET

[Smi02] Crosbie Smith. Force, energy, and thermodynamics. In Nye [Nye02], pages 289–310. ISBN 0-521-57199-5 (hardcover), 1-139-05352-3 (ebook). LCCN Q125 .C32 2003.

Smil:2017:TMC

[Smi17] V. Smil. Thank Maxwell for cellphones [numbers don’t lie]. IEEE Spectrum, 54(3):24, March 2017. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sarkar:2006:HW

[SMO^{+}06] Tapan (Tapan K.) Sarkar, Robert J. Mailloux, Arthur A. Oliner, M. Salazar-Palma, and Dipak L. Sengupta. History of Wireless. Wiley series in microwave and optical engineering. Wiley-Interscience, New York, NY, USA, 2006. ISBN 0-471-71814-9 (hardcover). xix + 655

pp. LCCN TK6547 .H57 2006. URL http://www.loc.gov/catdir/
enhancements/fy0622/2005022232-d.html;   http://www.loc.
gov/catdir/enhancements/fy0644/2005022232-b.html;   http:
//www.loc.gov/catdir/toc/ecip0517/2005022232.html.

Spencer:1965:BRF

[Spe65] J. Brookes Spencer. Book review: Faraday, Maxwell, and Kelvin by D. K. C. MacDonald. Isis, 56(3):392–393, Autumn 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/228141.

Spranzi:2006:BRF

[Spr06] Marta Spranzi. Book review: Fernand Hallyn, Les Structures rhétoriques de la science. De Kepler à Maxwell. Collection ≪Des travaux≫. Paris: Editions du Seuil, 2004. Pp. 323. ISBN 2-02-063249-7. EUR24.00. British Journal for the History of Science, 39(2):285– 286, June 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-

001X (electronic). URL http://www.jstor.org/stable/4028613.

Sheehan:2002:SSM

[SPW02] D. P. Sheehan, A. R. Putnam, and J. H. Wright. A solid-state Maxwell demon. Foundations of Physics, 32(10):1557–1595, October 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1023/
A%3A1020479302947.

Squires:1989:CMR

[Squ89] Euan J. Squires. A comment on Maxwell’s resolution of the wave/particle dilemma. The British Journal for the Philosophy of Science, 40(3):413–417, September 1989. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://

bjps.oxfordjournals.org/content/40/3/413.full.pdf+html;
http://www.jstor.org/stable/687789.

Smirnov-Rueda:2005:TCT

[SR05] R. Smirnov-Rueda. On two complementary types of total time derivative in classical field theories and Maxwell’s equations. Foundations of Physics, 35(10):1695–1723, October 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-005-6515-8.

Stachel:2009:GC

[Sta09] John J. Stachel. Going Critical: Volume One: The Challenge of Practice, volume 201 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 1-4020-1308-6 (hardcover). 407 (est.) pp. LCCN ????

Stanley:2012:DJC

[Sta12] Matthew Stanley. By design: James Clerk Maxwell and the evangelical unification of science. British Journal for the History of Science, 45(1):57–73, March 2012. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Staley:2014:BRP

[Sta14] Kent W. Staley. Book review: Peter Achinstein. Evidence and Method: Scientific Strategies of Isaac Newton and James Clerk Maxwell. Isis, 105(3):672–673, September 2014. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/679176.

Stanley:2015:HCM

[Sta15] Matthew Stanley. Huxley’s church and Maxwell’s demon: from theistic science to naturalistic science. University of Chicago Press, Chicago, IL, USA, 2015. ISBN 0-226-16487-X (hardcover), 0-226-16490-X (e-book). 364 pp. LCCN Q127.G4 S73 2015.

Stanley:2017:RSV

[Sta17] Matthew Stanley. Religious and scientific virtues: Maxwell, Eddington, and persistence. In van Dongen and Paul [vDP17], pages 49– 61. ISBN 3-319-48892-9 (hardcover), 3-319-48893-7 (e-book). ISSN 0068-0346. LCCN Q174 .B67 v. 321; Q175.32.V57 E65 2017. URL

https://link.springer.com/book/10.1007/978-3-319-48893-6.

Steinle:2005:EEA

[Ste05] Friedrich Steinle. Explorative Experimente: Ampère, Faraday und die Ursprünge der Elektrodynamik. (German) [Exploratory experiments: Ampère, Faraday, and the origins of electrodynamics], volume 50 of Boethius. Steiner, Stuttgart, Germany, 2005. ISBN 3-515-08185-2 (hardcover). 450 pp. LCCN QC630.5 .S74 2005.

Steinle:2016:EEA

[Ste16] Friedrich Steinle. Exploratory experiments: Ampère, Faraday, and the origins of electrodynamics. University of Pittsburgh Press, Pittsburgh, PA, USA, 2016. ISBN 0-8229-4450-2 (hardcover). x + 494 pp. LCCN QC630.5 .S7413 2016. English translation by Alex Levine of [Ste05].

Stork:1982:BRC

[Sto82] David G. Stork. Book reviews: Colour Vision in the Nineteenth Century: The Young–Helmholtz–Maxwell Theory by P. Sherman, 233pp. Hilger, Bristol (US dist. Heyden, Philadelphia), 1981. $77.00; Color Measurement: Theme and Variations, D. L. MacAdam, 228 pp., Springer, New York, 1981. $39.50. Physics Today, 35(5):83–84, May 1982. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Susskind:1968:EHEa

[Sus68a] C. Susskind. The early history of electronics I. Electromagnetics before Hertz. IEEE Spectrum, 5(8):90–98, August 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Susskind:1968:EHEb

[Sus68b] C. Susskind. The early history of electronics II. The experiments of Hertz. IEEE Spectrum, 5(12):57–60, December 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Susskind:1969:EHEa

[Sus69a] C. Susskind. The early history of electronics. III. Prehistory of radiotelegraphy. IEEE Spectrum, 6(4):69–74, April 1969. CODEN IEESAM. ISSN 1939-9340.

Susskind:1969:EHEb

[Sus69b] C. Susskind. The early history of electronics. IV. First radiotelegraphy experiments. IEEE Spectrum, 6(8):66–70, August 1969. CODEN IEESAM. ISSN 1939-9340.

Smith:1989:EEB

[SW89] Crosbie Smith and M. Norton Wise. Energy and Empire: a Biographical Study of Lord Kelvin. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-26173-2. xxvi + 866 pp. LCCN QC16.K3 S65

1989. URL http://www.loc.gov/catdir/description/cam041/
88025685.html;   http://www.loc.gov/catdir/toc/cam041/88025685.
html.

Tait:1873:BRC

[Tai73] Peter Guthrie Tait. Book review: Clerk-Maxwell’s Electricity and magnetism. Nature, 7(182):478–480, April 24, 1873. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https:/

/www.nature.com/nature/journal/v7/n182/pdf/007478a0.pdf.

Tait:1877:BRE

[Tai77] Peter Guthrie Tait. Book review: Elementary physics: Matter and motion, by James Clerk Maxwell. Nature, 16(398):119–122, June 14, 1877. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v16/n398/
pdf/016119a0.pdf.

Tait:1880:CMS

[Tai80a] Peter Guthrie Tait. Clerk-Maxwell’s scientific work. Nature, 21 (536):317–321, February 5, 1880. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v21/n536/pdf/021317a0.pdf.

Tait:1880:OJC

[Tai80b] Peter Guthrie Tait. [obituary]: James Clerk Maxwell. Proceedings of the Royal Society of Edinburgh, 10:331–339, 1880. CODEN PRSEAE. ISSN 0080-4541 (print), 2053-5902 (electronic).

Tait:1886:NTC

[Tai86] Peter Guthrie Tait. 5. Note on a theorem of Clerk-Maxwell. Proceedings of the Royal Society of Edinburgh, 13:21–23, 1886. CODEN PRSEAE. ISSN 0080-4541 (print), 2053-5902 (electronic).

Tait:1895:NAC

[Tai95] Peter Guthrie Tait. Note on the antecedents of Clerk-Maxwell’s electrodynamical-wave-equations. Proceedings of the Royal Society of Edinburgh, 20:213–214, 1895. CODEN PRSEAE. ISSN 0080-4541 (print), 2053-5902 (electronic).

Tait:1897:NCM

[Tai97] Peter Guthrie Tait. Note on Clerk-Maxwell’s law of distribution of velocity in a group of equal colliding spheres. Proceedings of the Royal Society of Edinburgh, 21:123–128, 1897. CODEN PRSEAE. ISSN 0080-4541 (print), 2053-5902 (electronic).

Tait:1911:JCM

[Tai11a] Peter Guthrie Tait. James Clerk Maxwell. In Encyclopaedia Britannica, volume 17, pages 929–930. ????, ????, eleventh edition, 1911. LCCN ????

Tait:1911:OMD

[Tai11b] Peter Guthrie Tait. [Origin’s of Maxwell’s demon]. In Knott

[Kno11], pages 213–216. URL https://archive.org/stream/
lifescientificwo00knotuoft#page/213. From the text: “When

Tait was preparing his Thermodynamics for the press he asked Maxwell for some hints. Maxwell’s reply of date Dec. 11, 1867, was very characteristics and of great interest as begin probably the first occasion on which he put in writing his conception of those fine intelligences — Maxwell’s demons as Kelvin nicknamed them — who operating on the individual molecules of a gas could render nugatory the second law of thermodynamics.” That is followed by three more pages of the Tait–Maxwell correspondence on the matter of demons.

Tait:1874:DR

[TC74] Peter Guthrie Tait and James Clerk Maxwell. Double rainbow. Nature, 10(257):437, October 1, 1874. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v10/n257/pdf/010437b0.pdf.

Theerman:1986:JCM

[The86] Paul Theerman. James Clerk Maxwell and religion. American Journal of Physics, 54(4):312–317, April 1986. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Theerman:1987:BRJ

[The87] Paul Theerman. Book review: John Hendry. James Clerk Maxwell and the Theory of the Electromagnetic Field. Bristol and Boston: Adam Hilger, 1986. Pp. xix + 305. ISBN 0-85274-563-X. £30.00. British Journal for the History of Science, 20(3):365–366, July 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/stable/4026375.

Thomson:1879:SDM

[Tho79] Sir William Thomson. The sorting demon of Maxwell. Nature, 20 (501):126, June 5, 1879. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL https://www.nature.com/articles/

020126a0. Abstract of Lecture at the Royal Institution, Friday, February 28, 1879, by Sir William Thomson, LL.D., F.R.S..

Thomson:1882:SDM

[Tho82] Sir William Thomson. The sorting demon of Maxwell: [abstract of a Friday evening lecture before the Royal Institution of Great Britain, February 28, 1879]. Proceedings of the Royal Institution of Great Britain, 9(??):113–114, ???? 1882. CODEN PIGBAI. ISSN 0035-8959. See also an earlier publication about the same lecture [Tho79].

Thomson:1889:SDM

[Tho89] Sir William Thomson [Lord Kelvin]. The sorting demon of Maxwell. In Popular Lectures and Addresses: Constitution of Matter, volume 1, pages 137–141. MacMillan Publishing Company, New York, NY, USA, 1889. LCCN ???? Three volumes (1: 1889, second edition 1891), 2: 1894, 3: 1891.

ThomsonLordKelvin:1891:STC

[Tho91] Sir William Thomson [Lord Kelvin]. On some test cases for the Maxwell–Boltzmann doctrine regarding distribution of energy. Proceedings of the Royal Society of London, 50(??):79–88, ???? 1891. CODEN PRSLAZ. ISSN 0370-1662 (print), 2053-9126 (electronic).

ThomsonLordKelvin:1892:DTC

[Tho92] Sir William Thomson [Lord Kelvin]. On a decisive test-case disproving the Maxwell–Boltzmann doctrine regarding distribution of kinetic energy. Proceedings of the Royal Society of London, 51(??):397–399, ???? 1892. CODEN PRSLAZ. ISSN 0370-1662 (print), 2053-9126 (electronic).

Thomson:1893:NRR

[Tho93] Sir J. J. (Joseph John) Thomson. Notes on recent researches in electricity and magnetism, intended as a sequel to Professor Clerk-Maxwell’s Treatise on electricity and magnetism. Clarendon press series. Clarendon Press, New York, NY, USA, 1893. xvi + 578 pp. LCCN QC518 .T5.

Thompson:1910:LWT

[Tho10] Silvanus P. (Silvanus Phillips) Thompson. The life of William Thomson, Baron Kelvin of Largs. ????, ????, 1910. ???? pp. LCCN ????

Thomson:1931:JCMa

[Tho31a] J. J. Thomson. James Clerk Maxwell. In Thomson [Tho31b], page ?? LCCN QC16.M4 J3. Essays by Sir J. J. Thomson, Max Planck, Albert Einstein, Joseph Larmor, James Jeans, William Garnett, Ambrose Fleming, Oliver Lodge, R. T. Glazebrook, and Horace Lamb.

Thomson:1931:JCMb

[Tho31b] Sir Joseph John Thomson, editor. James Clerk Maxwell; a commemoration volume, 1831–1931. Cambridge University Press, Cambridge, UK, 1931. 3 + 146 pp. LCCN QC16.M4 J3. Essays by Sir J. J. Thomson, Max Planck, Albert Einstein, Joseph Larmor, James Jeans, William Garnett, Ambrose Fleming, Oliver Lodge, R. T. Glazebrook, and Horace Lamb.

Thomson:1936:RR

[Tho36] J. J. (Joseph John) Thomson. Recollections and Reflections. G. Bell and Sons, Ltd., London, UK, 1936. viii + 451 pp. LCCN QC16.T45

A3 1936. URL https://archive.org/stream/in.ernet.dli.2015.
176233/2015.176233.Recollections-And-Reflections_djvu.txt.

Thomson:1937:RR

[Tho37] J. J. (Joseph John) Thomson. Recollections and Reflections. The Macmillan Company, New York, NY, USA, 1937. viii + 451 + 9 pp. LCCN QC16.T45 A3 1937.

Thomson:1968:NRR

[Tho68] Sir J. J. (Joseph John) Thomson. Notes on recent researches in electricity and magnetism, intended as a sequel to Professor Clerk-Maxwell’s ’Treatise on Electricity and Magnetism’. Dawsons, London, UK, 1968. ISBN 0-7129-0132-9. xvi + 578 pp. LCCN QC518 .T5 1968. Reprint of [Tho93].

Thomson:1975:RR

[Tho75] J. J. (Joseph John) Thomson. Recollections and reflections. History, philosophy, and sociology of science. Arno Press, New York, NY, USA, 1975. ISBN 0-405-06622-8. viii + 451 + 9 pp. LCCN QC16.T45 A3 1975.

Thompson:1976:LLK

[Tho76] Silvanus P. (Silvanus Phillips) Thompson. The life of Lord Kelvin. Chelsea Pub. Co., New York, NY, USA, second edition, 1976. ISBN 0-8284-0292-3. xx + 1297 + 14 pp. LCCN QC16.K3 T7 1976. Two volumes.

Thompson:2005:LLK

[Tho05] Silvanus P. (Silvanus Phillips) Thompson. The life of Lord Kelvin. American Mathematical Society, Providence, RI, USA, second edition, 2005. ISBN 0-8218-3745-1 (set), 0-8218-3743-5 (vol. 1), 0-8218-3744-3 (vol. 2). xx + xi + 1297 + 16 pp. LCCN QC16 .K3 T7 2005.

Thomson:2011:JCM

[Tho11] Joseph John Thomson, editor. James Clerk Maxwell: a commemoration volume 1831–1931. Cambridge University Press, Cambridge, UK, 2011. ISBN 1-107-67095-0 (paperback). 6 + 146 pp. LCCN QC16.M4 J3. Essays by Sir J. J. Thomson, Max Planck, Albert Einstein, Joseph Larmor, James Jeans, William Garnett, Ambrose Fleming, Oliver Lodge, R. T. Glazebrook, and Horace Lamb.

Tiscareno:2017:CRS

[Tis17] Matthew S. Tiscareno. Cassini and the rings of Saturn. American Scientist, 105(6):356–??, 2017. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL https://www.

americanscientist.org/article/cassini-and-the-rings-of-saturn.

Tolstoy:1981:JCM

[Tol81] Ivan Tolstoy. James Clerk Maxwell: a biography. Canongate, Edinburgh, Scotland, 1981. ISBN 0-86241-010-X. viii + 184 + 8 pp. LCCN QC16.M4 T64 1981.

Tolstoy:1982:JCM

[Tol82] Ivan Tolstoy. James Clerk Maxwell: a biography. University of Chicago Press, Chicago, IL, USA, 1982. ISBN 0-226-80785-1, 0-226-80787-8 (paperback). viii + 184 + 8 pp. LCCN QC16.M4 T64 1982.

Topper:1970:JJT

[Top70] David Roy Topper. J. J. Thomson and Maxwell’s Electromagnetic Theory. Ph.D. thesis, Division of Special Interdisciplinary Studies, Case Western Reserve University, Cleveland, OH, USA, September

1970. URL http://search.proquest.com/pqdtglobal/docview/
302507275/.

Torrance:2000:SPF

[Tor00] Thomas F. Torrance. The significance of Philoponos as a forerunner of Clerk Maxwell. IOP History of Physics Group Newsletter, 13: 48–56, Spring 2000. ISSN 1756-168X. URL http://www.iop.org/

activity/groups/subject/hp/newsletter/archive/file_66506.
pdf.

Toyoda:1999:NLM

[Toy99] Toshiyuki Toyoda. Notes on the life of Maxwell: Beyond the classification of physical sciences. Mathematical Analysis Laboratory Record [Kyoto University], 1142(??):137–147, February 18, 1999.

URL http://hdl.handle.net/2433/63886;   https://core.ac.
uk/download/pdf/39197271.pdf.

Trotter:1901:EGS

[Tro01] Alexander Trotter. James Clerk Maxwell, LL.D., F.R.S., Aberdeen, London, and Cambridge professor. In East Galloway sketches: or Biographical, historical, and descriptive notices of Kirkcudbrightshire, chiefly in the Nineteenth Century, pages 253–256. A. Rae, Castle-Douglas, Scotland, 1901. LCCN DA880.K6 T8.

JCMF:2014:FLS

[Tru14] Trustees of the James Clerk Maxwell Foundation. Family, life and scientific achievements of James Clerk Maxwell. Lecture slides., 2014.

URL www.homepages.ed.ac.uk/pmg/temp.ppt. 112 slides.

Tolstoy:1986:MM

[TS86] Ivan Tolstoy and Daniel M. Siegel. More on Maxwell. Physics Today, 39(10):13, 15, October 1986. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). Letters about the book review [Sie85a].

Tedre:2009:CNM

[TS09] Matti Tedre and Erkki Sutinen. Crossing the Newton–Maxwell gap: Convergences and contingencies. Spontaneous Generations: A Journal for the History and Philosophy of Science, 3(1):195– 212, ???? 2009. CODEN ???? ISSN 1913-0465. URL

https://spontaneousgenerations.library.utoronto.ca/index.
php/SpontaneousGenerations/article/view/3388.

Turner:1927:MSE

[Tur27] Dorothy Mabel Turner. Makers of Science: electricity and magnetism. Humphrey Milford, London, UK, 1927. xv + 184 pp. LCCN QC514 .T8. With an introduction by Charles Singer.

Turner:1953:MJC

[Tur53] Joseph Turner. The Methodology of James Clerk Maxwell. Ph.D. thesis, Faculty of Philosophy, Columbia University, New York, NY, USA,

1953. URL http://search.proquest.com/pqdtglobal/docview/
301981496.

Turner:1955:MMP

[Tur55a] Joseph Turner. Maxwell on the method of physical analogy. The British Journal for the Philosophy of Science, VI(23):226– 238, November 1955. CODEN BJPIA5. ISSN 0007-0882 (print),

1464-3537 (electronic). URL http://bjps.oxfordjournals.org/
content/VI/23/226.full.pdf+html;   http://www.jstor.org/
stable/685678.

Turner:1955:NMI

[Tur55b] Joseph Turner, Ph.D. A note on Maxwell’s interpretation of some attempts at dynamical explanation. Annals of Science, 11(3):238– 245, 1955. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Turner:1956:MLD

[Tur56] Joseph Turner. Maxwell on the logic of dynamical explanation. Philosophy of Science, 23(1):36–47, January 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.

jstor.org/stable/184996.

Turner:1972:MSE

[Tur72] D. M. (Dorothy Mabel) Turner. Makers of science: electricity and magnetism. University Microfilms, Ann Arbor, MI, USA, 1972. xv + 184 pp. LCCN QC514 .T8 1971. With an introduction by Charles Singer.

Turner:1983:BRM

[Tur83] R. Steven Turner. Book review: Mathematical Sciences: Paul D. Sherman, Colour vision in the nineteenth century. The Young– Helmholtz–Maxwell theory. Bristol: Adam Hilger Ltd, 1981. Pp. xiii + 233. $77.00/£35.00. British Journal for the History of Science, 16(3):297–298, November 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.jstor.org/

stable/4026167.

Turner:1996:OCW

[Tur96] R. Steven Turner. The origins of colorimetry: What did Helmholtz and Maxwell learn from Grassmann? In Schubring [Sch96], pages 71– 86. ISBN 90-481-4758-1, 94-015-8753-1 (e-book). ISSN 0068-0346.

LCCN QA8.9-10.3. URL https://link.springer.com/chapter/
10.1007/978-94-015-8753-2_8.

Tweney:2011:REF

[Twe11] Ryan D. Tweney. Representing the electromagnetic field: How Maxwell’s mathematics empowered Faraday’s field theory. Science & Education (Springer), 20(7–8):687–700, July 2011. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic).

Vargas:1982:NPM

[Var82] Jose G. Vargas. Nonrelativistic para-Maxwellian electrodynamics with preferred reference frame in the universe. Foundations of Physics, 12(9):889–905, September 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/BF01884999.

vonBaeyer:1998:MDW

[vB98] Hans Christian von Baeyer. Maxwell’s Demon: Why Warmth Disperses and Time Passes. Random House, New York, NY, USA, 1998. ISBN 0-679-43342-2. xxi + 207 pp. LCCN QC318.M35 V66 1998.

vanDongen:2017:EVS

[vDP17] Jeroen van Dongen and Herman Paul, editors. Epistemic Virtues in the Sciences and the Humanities, volume 321 of Boston Studies in the Philosophy and History of Science. Springer International Publishing, Cham, Switzerland, 2017. ISBN 3-319-48892-9 (hardcover), 3-319-48893-7 (e-book). ISSN 0068-0346. vi + 198 pp. LCCN Q174 .B67 v. 321; Q175.32.V57 E65 2017. URL https://link.springer.com/

book/10.1007/978-3-319-48893-6.

Visscher:1989:DFM

[Vis89] P. B. Visscher. Discrete formulation of Maxwell equations. Computers in Physics, 3(2):42–??, March 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168342.

vonMeyenn:1982:RLB

[vM82] Karl von Meyenn. Rezensionen: Ludwig Boltzmann Gesamtausgabe, herausgegeben von Roman U. Sexl. Band 1: Vorlesungen über Gastheorie. I. und II. Teil. Einleitung, Anmerkungen und Bibliographie von Stephen G. Brush. — Band 2: Vorlesungen über Maxwells Theorie der Elektricität und des Lichtes. I. und II. Teil. Einleitung, Anmerkungen und Bibliographie von Walter Kaiser. Akademische Druck — und Verlagsanstalt, Graz; Friedrich Vieweg und Sohn, Braunschweig/ Wiesbaden 1981 und 1982. Berichte zur Wissenschaftsgeschichte, 5(3– 4):263–264, ???? 1982. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Vorgul:2008:MEN

[Vor08] I. Vorgul. On Maxwell’s equations in non-stationary media. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 366(1871):1781–1788, May 2008. CO-DEN PTRMAD, PTMSFB. ISSN 1364-503X (print), 1471-2962 (electronic).

Vargas:1989:EML

[VT89] Jose G. Vargas and Douglas G. Torr. Electrodynamics of the Maxwell– Lorentz type in the ten-dimensional space of the testing of special relativity: A case for Finsler type connections. Foundations of Physics, 19 (3):269–291, March 1989. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/
10.1007/BF00734560.

Walker:1916:JCM

[Wal16] Robert Walker. James Clerk Maxwell. Aberdeen University Review, 3(9):193–213, June 1916. ISSN 0001-320X. URL https://archive.

org/stream/aberdeenuniversi03univuoft#page/n209/mode/2up/
search/James+Clerk+Maxwell;   https://archive.org/stream/
aberdeenuniversi03univuoft/aberdeenuniversi03univuoft_djvu.
txt.

Walker:1985:AUM

[Wal85] G. B. Walker. The axioms underlying Maxwell’s electromagnetic equations. American Journal of Physics, 53(12):1169–1172, December 1985. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Wang:2008:GFG

[Wan08] Charles H.-T Wang. Gauge formulation of general relativity using conformal and spin symmetries. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 366 (1871):1867–1874, May 2008. CODEN PTRMAD, PTMSFB. ISSN 1364-503X (print), 1471-2962 (electronic).

Warwick:2003:MTC

[War03] Andrew Warwick. Masters of theory: Cambridge and the rise of mathematical physics. University of Chicago Press, Chicago, IL, USA, 2003. ISBN 0-226-87374-9 (hardcover), 0-226-87375-7 (paperback). xiv + 572 pp. LCCN QC19.6 .W37 2003. URL http:/

/www.loc.gov/catdir/bios/uchi051/2002153732.html;   http:
//www.loc.gov/catdir/description/uchi051/2002153732.html;
 http://www.loc.gov/catdir/toc/fy037/2002153732.html.

Watanabe:1961:DDT

[Wat61] Masao Watanabe. The development of the dynamic theory of heat in England. Journal of the History of Science, Japan, 58(??):9–15, ???? 1961.

Weatherall:2016:MHN

[Wea16] James Owen Weatherall. Maxwell–Huygens, Newton–Cartan, and Saunders–Knox space–times. Philosophy of Science, 83(1):82–92, January 2016. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/10.1086/
684080.

Weinstein:2002:OIM

[Wei02] Steven Weinstein. Objectivity, information, and Maxwell’s demon. Philosophy of Science, 70(5):1245–1255, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/377404.

West:1999:IRJ

[Wes99] Thomas G. West. Images and reversals: James Clerk Maxwell, working in wet clay. Computer Graphics, 33(1):15–17, February 1999.

CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic).

Winiarski:2008:EIC

[WG08] B. Winiarski and I. A. Guz. The effect of the interaction of cracks in orthotropic layered materials under compressive loading. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 366(1871):1841–1847, May 2008. CO-DEN PTRMAD, PTMSFB. ISSN 1364-503X (print), 1471-2962 (electronic).

Wheen:2011:DDD

[Whe11] Andrew Wheen. Dot-dash to Dot.com: how modern telecommunications evolved from the telegraph to the Internet. Springer-Praxis books in popular science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 1-4419-6759-1, 1-4419-6760-5 (e-book). xiv + 301 pp.

LCCN HE7631 .W5 2011. URL http://www.loc.gov/catdir/
enhancements/fy1303/2010929000-b.html;   http://www.loc.
gov/catdir/enhancements/fy1303/2010929000-d.html;   http:
//www.loc.gov/catdir/enhancements/fy1303/2010929000-t.html.

Whiting:1885:MD

[Whi85] Harold Whiting. Maxwell’s demons. Science, ns-6(130):83, July 1885. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Whittaker:1910:HTA

[Whi10] E. T. (Edmund Taylor) Whittaker. A History of the Theories of Aether and Electricity from the Age of Descartes to the Close of the Nineteenth Century. Dublin University Press series. Longmans, Green and Co., London, UK, 1910. xiii + 475 pp. LCCN QC177 .W6.

Whittaker:1951:HTA

[Whi53] Sir E. T. (Edmund Taylor) Whittaker. A History of the Theories of Aether and Electricity. T. Nelson, London, UK, revised and enlarged edition, 1951–1953. ???? pp. LCCN QC177 .W62.

Whittaker:1954:HTA

[Whi54] Sir E. T. (Edmund Taylor) Whittaker. A history of the theories of aether and electricity. Philosophical Library, New York, NY, USA, 1954. ???? pp. LCCN QC177 .W63.

Whittaker:1987:HTA

[Whi87] Sir E. T. (Edmund Taylor) Whittaker. A history of the theories of aether and electricity, volume 7 of The History of modern physics. Tomash Publishers, Los Angeles, CA, USA, 1987. ISBN 0-88318-523-7. ISSN 1800-1950. 434 (vol. 1), 319 (vol. 2) pp. LCCN QC177 .W63 1987.

Whittaker:1989:HTA

[Whi89] Sir E. T. (Edmund Taylor) Whittaker. A History of the Theories of Aether and Electricity. Dover Publications, Inc., New York, NY, USA, 1989. ISBN 0-486-26126-3. 434 + 319 pp. LCCN QC177 .W63 1989.

Whittaker:2008:HTA

[Whi08] Edmund Taylor Whittaker. A history of the theories of aether and electricity: from the age of Descartes to the close of the nineteenth century. Kessinger Publishing’s rare reprints. Kessinger Publishing, Whitefish, MT, USA, 2008. ISBN 0-548-64276-1. xiii + 475 pp. LCCN ????

Wignall:1990:MET

[Wig90] J. W. G. Wignall. Maxwell electrodynamics from a theory of macroscopically extended particles. Foundations of Physics, 20(2):139–158, February 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL http://link.springer.com/article/10.1007/
BF00731644.

Wilson:1851:LHH

[Wil51] George Wilson. The life of the Honble. Henry Cavendish, including abstracts of his more important scientific papers, etc. Printed for the Cavendish Society, London, UK, 1851. xiv + 478 pp. LCCN QD22.C4 W7.

Wilson:1855:RCB

[Wil55] George Wilson. Researches on colour-blindness. With a supplement on the danger attending the present system of railway and marine coloured signals. Sutherland and Knox, South Bridge, Edinburgh, Scotland, 1855. 4 + xx + 5–180 pp. LCCN ????

Williams:1962:PSF

[Wil62] L. Pearce Williams. The physical sciences in the first half of the Nineteenth Century: Problems and sources. History of Science (UK), 1(1):1–15, March 1962. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/1/
1/1.full.pdf+html.

Williams:1968:BRN

[Wil68] L. Pearce Williams. Book review: Nineteenth Century: The Contributions of Faraday and Maxwell to Electrical Science. By R. A. R. Tricker. Pp. ix + 289. 6 plates. Oxford: Pergamon Press. 1966. 25s. British Journal for the History of Science, 4(1):78–79, June 1968. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

http://www.jstor.org/stable/4024996.

Wilson:1972:LHH

[Wil72] George Wilson. The life of the Honble. Henry Cavendish: including abstracts of his more important scientific papers, and a critical inquiry into the claims of all the alleged discoverers of the composition of water. Works of the Cavendish Society. Readex Microprint Corporation, New York, NY, USA, 1972. xiv + 478 pp. LCCN ????

Wilson:1979:MCG

[Wil79] Mark Wilson. Maxwell’s condition — Goodman’s problem. The British Journal for the Philosophy of Science, 30(2):107–123, June 1979. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.oxfordjournals.org/content/30/2/
107.full.pdf+html;   http://www.jstor.org/stable/686943.

Williams:1997:BRM

[Wil97] L. Pearce Williams. Book review: Maxwell on the Electromagnetic Field: A Guided Study by Thomas K. Simpson. Isis, 88(4):715–716, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237858.

Wilson:2005:BRJ

[Wil05] David B. Wilson. Book review: James Clerk Maxwell popularized: The Man Who Changed Everything: The Life of James Clerk Maxwell. Journal for the History of Astronomy, 36(3):352–354, August 2005. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Wilson:2016:LHH

[Wil16] George Wilson. The life of the Honble. Henry Cavendish: including abstracts of his more important scientific papers, and a critical inquiry into the claims of all the alleged discoverers of the composition of water. Forgotten Books, London, UK, 2016. ISBN 1-330-74657-0. LCCN ????

Wise:1977:FAE

[Wis77] Matthew Norton Wise. The flow analogy to electricity and magnetism: Kelvin and Maxwell. Ph.d. dissertation, Department of History, Princeton University, Princeton, NJ, USA, January 1977. ix + 217

+ n1–n77 pp. URL https://search.proquest.com/pqdtglobal/

docview/302837465. Number AAT7714252.

Wise:1979:MEE

[Wis79] M. N. Wise. The mutual embrace of electricity and magnetism. Science, 203(4387):1310–1318, March 30, 1979. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Wise:1982:MLB

[Wis82] M. Norton Wise. The Maxwell literature and British dynamical theory. Historical Studies in the Physical Sciences, 13(1):175–205, ???? 1982. CODEN HSPSAS. ISSN 0073-2672. URL http://www.jstor.

org/stable/27757511.

Yaghjian:2007:RMT

[Yag07] Arthur D. Yaghjian. Reflections on Maxwell’s Treatise. In ET2007 XXIII Annual Meeting of the Research Groups on Electrical Engineering, Florence, Italy, June 2007, page ?? ????, ????, 2007.

Yaghjian:2014:RMTa

[Yag14a] Arthur D. Yaghjian. Reflections on Maxwell’s Treatise. In Proceedings of PIERS, Guangzhou, China, Aug. 2014, page ?? ????, ????, 2014.

Yaghjian:2014:RMTb

[Yag14b] Arthur D. Yaghjian. Reflections on Maxwell’s Treatise. Progress in Electromagnetics Research (PIER), 149:217–249, 2014. CODEN PEL-REX. ISSN 1070-4698 (print), 1559-8985 (electronic). URL http:

//www.jpier.org/PIER/pier.php?paper=14092503;   https://
pdfs.semanticscholar.org/b395/1ef16ff023237fa3516627c5264fc9fd57bc.
pdf.

Yan:1995:CMM

[Yan95] C. C. Yan. Can the macroscopic Maxwell equations be obtained from the microscopic Maxwell–Lorentz equations by performing averages? Foundations of Physics, 25(3):491–502, March 1995. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02059233.

Yang:2014:COM

[Yan14] Chen Ning Yang. The conceptual origins of Maxwell’s equations and gauge theory. Physics Today, 67(11):45–51, November 2014. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Yost:1999:BRP

[Yos99] Robinson M. Yost. Book review: P. M. Harman, The Natural Philosophy of James Clerk Maxwell. Cambridge: Cambridge University Press, 1998. Pp. xiv + 232. ISBN 0-521-56102-7. £35.00, $59.95. British Journal for the History of Science, 32(3):363–378, September 1999. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X

(electronic). URL http://www.jstor.org/stable/4028003.

Zito:2002:MHM

[Zit02] F. A. Zito. Maxwell, Hertz and Marconi, using the history of science and technology in science education. Ph.D. dissertation, Program in Science Education, Department of Teaching and Learning, New York University, New York, NY, USA, 2002. URL http://ezproxy.lib.

utah.edu/docview/305561906.